A Complete Bibliography of the Journal of Econometrics (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.05

Title word cross-reference

5 [LPS15]. AR(1) [Kru08, Kru14]. b [Sun14, SY19, VW14a]. ϵ [BBCL18]. F [SW16]. g [LS12]. Γ [DGG14]. I(1) [Phi10]. I(2) [GBH15b, JJFG10]. K [Kau15, LLL16]. L^{2} [And17]. l_{1} [MM16, Zhu18]. M [IL10, IL18, Yan18, MV13]. N [Che10b, FVW16, Gay13, AG10a, EJC12]. p [BBKU19]. R [HSVV13, KZ15]. S [Lea16]. σ [KPS19]. ^{3} \sqrt{n} [JPW15]. T [DMR16, FVW16, MMV14, ZG10].

-complete [And17]. -consistent [JPW15, Che10b, Gay13]. -contamination [BBCL18]. -convergence [KPS19]. -dependence [MV13]. -estimation [HSVV13, Yan18]. -estimators [IL18, IL10]. -minute [LPS15]. -moment [DGG14]. -nearest-neighbor [LLL16]. -priors [LS12]. -regularization [MM16]. -regularized [Zhu18]. -squared [KZ15]. -state [Kau15]. -statistics [MMV14]. -test [SW16]. -uniformly [EJC12]. -values [BBKU19, Lea16].

/unit [Kru14, Kru08].

1 [Ano11n, Ano11r]. 134 [HLT12a]. 144 [Kru14]. 145-406 [Ano19n]. 159 [IL18]. 164 [PST22]. 170 [HINR14]. 178 [CW14a]. 182 [GMR14a].

2010 [Ano10t, Ano10q, Ano10s, Ano10p, Ano10m, Ano10r, Ano10o, Ano10n]. 2011 [Ano11n, Ano11o, Ano11p, Ano11s, Ano11q, Ano11r, Ano12b, Ano12p, Ano12o]. 2012 [Ano12t, Ano12u, Ano12r, Ano12s, Ano12q, Ano13a, Ano13o]. 2013 [Ano13v, Ano13x, Ano13r, Ano13w, Ano13t, Ano13q, Ano13s, Ano13p, Ano13y, Ano13u]. 2014 [Ano14x, Ano14v, Ano14w, Ano14u, Ano14s, Ano14r, Ano14q, Ano14t]. 2015 [Ano15t, Ano15s, Ano15m, Ano15r, Ano15p, Ano15o, Ano15n, Ano15q, Ano16a]. 2016 [Ano16t, Ano16p, Ano16r, Ano16q, Ano16s, Ano16n, Ano16o, Ano17a]. 2017 [Ano17t, Ano17v, Ano17r, Ano17u, Ano17s, Ano17o, Ano17q, Ano17p, Ano17n]. 2018 [Ano18v, Ano18x, Ano18w, Ano18t, Ano18o, Ano18u, Ano18q, Ano18p, Ano18r, Ano18s]. 2019 [Ano19o, Ano19q, Ano19m, Ano19r]. 212 [CCCM22]. 2SLS [Zhu18].

319-618 [Ano19p]. 37 [Xie11].

40 [Kra13].

90th [Slo17].

ABC [FN18]. ability [BK10, CM12]. absences [Gil10]. Absolute [BFS19a, Jus11]. accelerated [LLS15]. Accelerating [BGK19]. Accounting [LW10a, Zam10, PT11a]. accumulation [PT11a]. accuracy [BR11, CM15b, CS11b, GMP17, Hay10, LP18a]. accurate [HS18b, SK12]. ACF [ACT13]. achievement [BDV17]. Achieving [BFS19b]. acquisition [LZ12]. across [GK18a, LPS15]. Activity [TT10, HV17c, JKLM12, TTG14]. Adaptive [CN14, CLR10, GKP13, Hor14, KKLW18, KP19, LTT17, Yu15, AC16, CF15a, GD19, GGL13, HOvD12, JPW15, QS16]. Adaptively [ZLZ13]. addition [BD15]. additional [Hay10, PS17]. Additive [CG10, DL18, BMDL17, FMV15, LPLZ18, PHL16]. additively [JCLL10]. adequacy [ZL15a]. ADF [GBH15b]. adjusted [Sai13]. adjustment [FP10]. Advance [CHH18]. Advances [ET14, CS14b, MXY12a, MXY12b]. Adverse [KS16a]. affiliation [HLP12, JPW10]. affine [CDR11, CW15b, FMS13, GZ16, HW12a, HW14a, MPRR17]. after [BI14]. against [CAS15, KL18, Lee14b, SLB18]. age [PSW10]. agents [Poi12]. aggregate [DNR10, Tri11]. aggregated [HLW15]. aggregates [cCTZ14]. Aggregation [PC14, DS15, HV17a, Has11, HS14, IDF11, MR16, Tho14]. aggressive [ILM19]. aging [GK10]. Agricultural [PLJ11]. Aigner [Ano12b, Ano16a]. airline [AH12a]. Akaike [Spa10]. Akaike-type [Spa10].

ALASSO [CGL15]. algebra [MZC19]. algorithm [FGS18, Swe11]. algorithms [GD19, HOvD12]. allocation [HV19, PT11c, PST22]. Alternative [RS19, AG10a, AKM11, CF15b, CAS15, Kim14]. alternatives [Hag12, Lee14b]. ambiguity [AGKZ18]. Amemiya [CHH18]. American [GR13]. among [ASX16, Gil10]. Analysis [ACR19, Bha14a, CF15b, ASX17, ABHP16, BCF18, BI14, BKMZ16, BH16b, CKM16, CM15a, CS14b, sCJP14, CFSHP14, DZ15, EG14, FF14, FHLZ17, GM16, GBH15a, GMR17, GKP12b, HM14, Ho15, ILM19, JJFG10, Kim14, LB19, LMT12, LZ12, LS15b, LL15b, Li17, LLL18, MZ18, MSS11, Nie10, OY19, Ona15, OW19, SX17b, TT10, Tod11, Ton15, Wes12, WP14, ZA10, ZWXW10]. Analyzing [DLM17]. and/or [LZ12]. Anderson [WT18]. Annals [ASLN19, CL10, ET14, HKRV18, ILT11, MT12, GKP12a]. Announcement [Ano14a, Ano17a, Ano18a, Ano18b, Ano19a]. anomalies [CO15]. anonymous [Lam12]. anything [LPS15]. apples [BDV17]. Application [GMR17, ALGQ16, ACG17, ALX11, BSV12, BKY10, BS18, BJW15, CCF18, Cam12, CNT15, Cha13b, CDS13, CR10, CTV19, CKT17, DM13, DDSW18, DJfL15, FGPZ17, Gut16, JWFS16, KP17, KF19, Kue19, LLS15, LT11, LP17, MR16, MPP10, MPRR17, ZZ16, ZG10]. Applications [AG10b, GKP12a, Hu17, CFHZ19, CN14, Fas13, FMV15, HM18, KPS19, LS12, Zha11b]. Applied [LB19, LW14a, MP13a]. approach [ACT13, AHS10, An17, AFPZ16, AC15, ABHP16, BPS19, BPQ16, Car12, CMM18, CFHZ19, Cha12, CJMMPA15, CCP17, CS13, CFW14, CS14a, CH14b, CLL19, CH15, CGP18, CDDH15, CW11a, CPTV17, CGP16, CM12, CM15b, DGG14, DMP11, DSW12, Ell11, FL16, FS12, FiKY16, FL13, GHK18, GLS11, GK18b, HL19b, Hil16, HMT12, HZ15a, Kai16, KP15, Kim14, KS18b, Kol18, KP10b, LR19, Lee16a, LZY14, LZL18, LT14b, LT15b, MSX13, MW16, PWM11, PS12, Pet19, PTV16, Phi91, RL17, RH16, Son11, SHL11, SI19, SY19, XTZ10, ZA10]. approaches [Blu10, CM17a, Kea10, KPT12, Rus10]. approximate [BL16, BN19a, Cho15, DGR11, FWZ19, GOS19, KS17]. approximation [GK17]. approximations [BR11, DT19, FMS13, GP14a, LR19]. April [Ano11n, Ano19n, Ano10t, Ano13v, Ano14x, Ano15t, Ano17t, Ano18v]. AR-GARCH [AW15]. arbitrage [BC10, CG11, CDR11, FH15a, PLJ11]. arbitrage-free [CDR11]. arbitrary [Kru13]. ARCH [CZ15a, HP12a, NSS14]. ARCH-type [CZ15a]. ARCH/GARCH [HP12a]. Archimedean [OOS13]. ArCo [CMM18]. area [Fav13]. ARFIMA [CNT15]. ARMA [ADFK16, ZXC15, ZL15a]. Arnold [Ano13a, Ano17a]. artificial [CMM18]. ascending [FT14]. ask [CLY17, CLSY19]. aspects [AS14a]. Assessing [AG12a]. assessment [CHP18, ERS17]. Asset [CO15, ASX16, AG12a, BH14b, FL19b, GKR13, GK18b, KS18b, LF18, LZK18, LPS15, LYCP19, PT11c, PST22]. Asset-pricing [CO15]. assignment [BD12, CF19, PWM11]. assisted [Sha11]. assuming [HK11]. assumptions [Gha17]. asymmetric [BEE15, EM14b, GKS16, JM14, KLMP15, Lam12, LT15b, PPW19, Par16, Tet12, ZG10]. asymmetry [AM15a, ACM17, Cam12, DS15]. Asymptotic [CS11a, DMN19, FHLZ17, Fre15, GMHS12, HP12a, HS17b, KS19a, Lee14a, Lee16b, LOS18, LLZ15, LB17, LP18a, Ona15, Par19, Rob11, ASP16, AKM10, BR11, BCCK15, CF13, CC15b, sCJP14, CP18b, DH18, Geo10, IL10, IL18, JP12, JJ14, Ket19, KSY17, LKH14, LT14a, WL15]. Asymptotically [Arm15, CK18, CZP15, FGL11]. Asymptotics [AG12b, CM16, DFL18, KGMR12, KP17, LL15a, Ona12, Can14, CGP18, Kap15a, MP14, Sun14, WY16, Zha16a]. asynchronicity [BLL^{+}17, HS18a]. asynchronous [DLX19, SX17b]. asynchronously [PHL16]. ATE [CF19]. atheoretic [Blu10, Kea10, Rus10]. attitudes [FT14]. attrition [HL19b]. auction [HLM^{+}12, HN12, JPW10]. auctions [AHS10, An17, ALGQ16, AGKZ18, BS12, Cam12, FT14, HS13b, HMS13, HLP12, KPT12, Lam12, LZ12, MMS19, MS12, MSX13, Zin18]. Augmented [Han15, CH15, CS14c, DP17, GP14b, GMHS12, KRW17]. August [Ano19p, Ano10q, Ano11o, Ano13x, Ano14v, Ano18x]. autocorrelated [HLT10, MM19]. autocorrelation [KW10a, KS11, PP18, RR15, Sun14, Vog12]. autocovariance [GP12, PL15]. autoregression [Dua16, HT19, KMS16, Mag12, Qia16, ZZL18, ZWWH19]. autoregressions [Bog22, BCRT16, CS11a, CCM19, CCCM22, CRT10, DHY16, GMWY19, Han10, HL14b, JM19, KC15, KKP19, KP19, LMS17, LZZZ19]. Autoregressive [CKM16, Gup18a, AD13, Cha13a, CYZ15, ChKS15, DK18, DPSS11, FP11b, GTY13, GKS16, GGL12, GJ17, GJ18, GS11b, GLL19, GR15, GR18b, HM18, JfL19, JN10, Joh10, JL13, JN18, KG18, KP10a, fLY10, LOS18, LL12a, LLZ15, LfL10a, LfLB10, LT15b, LXC16, MS17, MSW15, Pos16, QfL15, Rob10, SJ10, Su12, Sun16, SM18, SHHW18, XfL15a, XfL15b, XfL18, YfL17b, YL17, ZY18, ZZC18]. auxiliary [FN18]. average [Cha13b, CPP19, DMP18, DK18, ES10, Far15, HS18a, LY16, LB17, MSV19, Pos16, SWC12, SY19]. averages [WU15]. Averaging [Han10, CH15, CIY15, CKP10, DDMP15, DiT16, GZWZ16, HR12a, HP12b, KM16, KLGS12, LS12, LZZZ19, Liu15, LS15c, MPP10, MZ16, WZZ10, ZWZ13, ZY18]. averse [Wil11, Zin18]. aversion [BGZ11, Cam12, Jus11]. Awaji [FH15c]. Award [Ano12b, Ano13a, Ano15u, Ano16a, Ano17a]. Axiomatic [Fat10].

B [MXY12a, MXY12b, MP14]. Bad [BEE15, APT18, KT16]. Bahadur [CF13, Ots10]. band [CV17]. Banded [GMWY19]. bands [FR18, HL12, KS18a, KS19b]. Bandwidth [ASP16, KSY17, LP13]. banking [HL19a]. banks [FGPZ17]. bargaining [FM12, MT19]. Barnett [BI14]. Based [KMN11, ASP12, ASAM15, ABH11, AW15, AS14b, AS17b, AMM11, BCCFV19, BHP12, BvH17, BD15, Bre15, BCDN12, CN12, CGL15, CZ11a, CS13, CZ15a, CL15b, CW11a, sCJP14, COP10, DPvD11, DGR11, FF11, FS10, FS14, FS12, FZ10, GT18, GZWZ16, GLPR11, Geo10, GP14a, GLS11, GBH15b, GBH15a, HvdAW11, HvdAW16, HO12, HLR11, HLPY15, HLW13, Kim14, KKLW18, KR10, Las10, LLL16, LZK18, LYZ18, LZZZ19, LW10b, MS16, MS12, McC17, MMV14, MV13, OP16, PG16, Pat11a, Pel19, PdSSGK12, PKA18, PM11, SG19, Siz11, Son11, TBE14, Tet12, Tri11, WT18, WL14, Xiu14, Yan15b, YL17, ZWZ19, ZL15b, Zu15, dHvdG11]. Basmann [RS17, Slo17]. Bayesian [Bog22, CCCM22, Kar17, ABHP16, BCL14, BCR19, BvH17, CCM19, CEK16, CGP18, CFSHP14, Des11, FJ19, FS12, FL19b, GGR17, GHK18, GT18, Gew96, GKP12a, Gew12, GD19, HPK14, sHDW17, HP12b, ILM19, JWFS16, JM10, JM13, JM16a, KG18, Kas19, KS19c, KFSvD19, Kim14, KLGS11, KLGS12, KKP19, KPT12, LS12, LT11, LMT12, LY12, LZY14, LLY15, LS19b, Mac10, MZ18, MW19a, NP12, Nor15, Pel14, Pet19, ZA10, vH11]. be [Ket18, Poi12]. bearing [AT14]. beat [LPS15]. behavior [CGH^{+}19, GK10, HP14a, SGF11, TPY19]. behaviour [O’D16]. being [MR16]. beliefs [An17]. beneath [CFFL15]. benefits [KW10b]. bequests [GGHM15]. Besov [Kue19]. Best [CL18a, KT12]. Beta [BCL14]. Beta-product [BCL14]. betas [DFL18, TB10]. better [BC11]. Between [MZ16, CFL16, Dez15, HLOW16, HR15b, JKM17]. Beyond [MP12b]. Bias [FP14, FVV11, Lee12, Pak19, WPY11, Yu12a, CF15b, HL19b, Ike16, Sia17, Yan15a]. bias-corrected [Ike16]. Bias-correction [FP14]. bid [CLY17, CLSY19, HLP12, LZ12]. bid-ask [CLY17, CLSY19]. bidder [FT14]. bidders [AHS10, Lam12, Zin18]. bidding [LZ12]. Bierens [Kas10]. Big [ACR19, DGMZ19, CDG14, GG17, HL19a, KFSvD19, HL19b]. Binary [CZ15b, GLL15, Kom13, AG10a, BN12, BC14b, CKT16, CL18a, DGJ19, DL11, DHKS18, EL13, HS15a, Kha13, KSV15a, KT12, LVX17, MM17b, MJ11, Mou15, WX14, WX15]. binomial [LML11, Sho16]. bipower [CPR10]. bivariate [CZ11b, HV17b, HS18a, LPZ19]. black [DMZ18]. black-white [DMZ18]. blame [ASX16]. block [AGS11, CR10, PSU11]. blocks [HL11a]. Board [Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17b, Ano17c, Ano17d, Ano17e]. Board [Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l]. Bond [ACR19, BCN11, Fav13]. Bonferroni [McC17, WT18]. Bonferroni-based [McC17, WT18]. book [BNW19, CP19]. boomerang [HL17b]. Bootstrap [CNT15, GK15, WK16, WT18, BCIT14, Can10a, CNT17, Dav17, DMN19, GH13, HV17c, JfL15, KS12c, Lee14a, Lee16b, LTTC17, LSW10, PSU11, PSS18, PGM15, Ric19, Sha11, Swe11, Yan15b]. bootstrap-assisted [Sha11]. bootstrapped [DMC12, ZL15a]. Bootstrapping [BCMH19, DGM13, DG17, GP14b, Hou17, Phi10, AGS11, HS18a, Tra13]. both [ASX16, CCH13, DL16, DL18, HLT12b, HZ15a, Ike16, LZL18, OÁ19]. bound [AC12, KS12b]. boundary [DGG14, GP12, Ket18, Ped17]. bounded [CX14a, GS13]. boundedly [And17]. boundedly-complete [And17]. Bounding [BKM14, CDD19]. Bounds [KT12, VRW18, BSV12, Chi10, GKP12b, JPX11, Mou15, ST12]. Box [GM11]. break [EM14a, FMV15, GJ11, HLT16, HLT12c, HL15, ILT13, JWY18, Kim11, KS15]. breakdowns [JM16b]. breaking [GS11b]. breaks [Bai10, BFK16, BCLL16, BBKU19, BE11, CFH10, CDH12, CH14a, CDG14, HLT10, HLT13, HS17a, MS18, MSW19, OP18, PPP13, PT11c, PST22, QS16, STZ19, WBH13, YT15]. Breusch [HOY17]. BRIKT [BdCK^{+}14]. budget [BD12, GJZ15]. building [Zha15].

C [MXY12a, MXY12b]. calibrated [HMT12, JK15]. Calibrating [Jus11]. calibration [KS18b]. call [FH15a]. call-option [FH15a]. Can [CF12, CM17b, HU10, LLS17, MW17]. cannot [HU10]. canonical [OW19]. capital [iKPS16]. Carlo [LS15b, NR10, PdSSGK12, ZA10]. CARMA [BL15]. case [ADFK16, BMS18, CGE10, Gug10, HR12b, Myk19, OP15, Sia17, YdJfL12]. cash [BXZ15]. categorical [Xu17]. catheterization [BSV12]. Causal [WXC14, LT11, LZW^{+}18]. Causality [CS14b, RW11, AS14a, BM12, DT10, GHM16, GHS16, NHKJ11, TBE14, WP14]. causation [Slo17]. caused [Tho14]. CCE [KRW17]. CDF [GK18a]. CDS [ASLP14]. celebration [Slo17]. Censored [BFS19a, Che10a, Che18, Cíz12, FL18, LO15]. censoring [Che10b, CFVK15, KPT16]. censorship [CFPY10]. central [CKT17, Xie88, Xie11]. centrality [LfL10b]. CEOs [LM11]. certain [Kas10]. chain [PdSSGK12, dHvdG11]. chains [NNN11]. chance [WW15]. change [ALX11, BCF18, BDR14, BCMH19, Die14, FS10, FJ19, FH19, GJZ15, GKP13, HO12, HLT06, HLT12a, JJ10, Lar11, PLG17, YT15, ZWXW10]. change-point [BCF18, BDR14]. change-points [ZWXW10]. changes [CF12, Che15a, CH18b, HHB12, KT11, LW10a, OQ11, Xu13]. chaos [PW12]. characteristics [AFPZ16]. Characterization [FR18, IL10, IL18, FP11b, SV12]. Chasing [CRdM17]. checking [Zha16b]. checks [LW14a, WWZ18]. Chi [GKR13, LLY15, Xie88, Xie11]. chi-square [Xie88, Xie11]. Chi-squared [GKR13, LLY15]. child [BK10, GKP12b, LMvdK10]. childcare [BK10]. China [AFPZ16, BEK17, DZ15, OP15]. choice [AM10, AG10a, Arm18, BKY10, BDKS18, BHH12, CLLL16, CFVN19, CS12, CHM11, Gew12, HL11b, Hor14, Kom13, LB19, Mat19, MM17b, Sto12, Tet12, Wil11, YY19]. choices [AHK12, BBB^{+}19, BK10, LS15a]. Cholesky [DFL18]. chosen [MO12]. cities [BLO14]. class [AFPZ16, CDH10, CDR11, FGL11, GMO18, HvdAW11, HOvD12, HSS17, KV10a]. classes [ASX16, Kri10, Kue19, LPS15].

Classical [JPW15, HS15b]. classified [DGJ19]. cleaning [MZ16]. Climate [HLX19, ALX11]. clinical [ILM19]. Closed [Cho13, HS15b, LC16]. Closed-form [Cho13, HS15b, LC16]. cluster [BCH11]. clustered [DMN19]. clustering [ZWZ19]. clusters [Hag19]. co [AM15a, AGHP16, BW15, BCRT16, CRT10, Fas13, HLT16]. co-integration [BCRT16, CRT10, Fas13, HLT16]. co-jumps [BW15]. co-movements [AGHP16]. co-volatilities [AM15a]. Coauthorship [ASK16]. codifference [LT14a]. coefficient [CJY15, CFL16, FGPZ17, GLL15, GKY14, GL14, HSU10, HKLP15, HL17b, HT16, HT19, JL13, KKL14, LZL^{+}16, LL12b, MKS16, MS17, SH16, Sun16, SM18, TPY19, WL12, ZLX15]. coefficients [AKM11, CX12, DPY16, DHKS18, FiKRB12, FiKY16, GHPP18, HS15a, HHM17, KW10a, Ket19, Lea16, LT11, MSW18, RV14]. cognitive [BK10]. coherent [GR14]. cointegrated [BBKU19, FP10, GHX14, Geo10, HR10, Joh10, JN18, Phi14, Shi12, SHL11, TY17]. Cointegrating [CP12, CK12, DGTY17, KP12, KU14, MP17, PSW10, VW14a, WWZ18]. Cointegration [BFvD10a, JW10, BFvD10b, BJN15, CN12, ChKS15, CV17, GBH15b, GBH15a, Gra10, GL14, HvdAW16, Hua14, KLGS11, KR10, Las10, MW13, Nie10, OW19, PLG17, Swe11, TU10, WP14, YdJfL12, ZC18]. college [AHK12]. combination [Bha13, HW14c]. combinations [BBG^{+}19, BCRvD13, KMPF15, WY12]. combined [ZLZ13]. Combining [BBKU19, HL14b, Myk19]. COMFORT [PP15]. Comment [Ahn15, Bog22]. Comments [Blu10, Rus10]. Commercial [BCG19, FGPZ17]. commodity [CNT15, cCTZ14, FFG10]. Common [Bai10, CP15, BE10, FP11b, HS13b, JJ14, Kim11, OP18, PP15, PT11b, QS16, Wol16, YT15]. comovements [HS12a]. compactly [HSS17]. Comparing [GK18a, HU10, DPvD11]. Comparison [HMT12, RH16, AS10, GKR13, HS18b, LPS15, MPP10, MP12a, MO12, Pat11b]. Comparisons [Hil16, CM15b]. competing [MP13a]. competition [BE16]. competitive [CHHW11]. complements [HHS12]. Complete [EGT13, And17, DHKS18, Kli15]. complexity [KFSvD19]. component [ASX17, CEG11, GMR17]. components [BN13, MBC15, MRSE11, Ona12, WU15]. composite [LL16]. compressed [KKP19]. computation [BFS19a, LPLZ18, RH16, ZLCL19]. computational [LS15b]. computationally [Qia16, SI19]. computing [FAM13, Xie88, Xie11]. conclusions [Gra12]. condition [FGL11, HNS14, HLT12b, HA13, KRW17]. Conditional [AL14, BCCFV19, RS13, WZ16, AG12b, AS14b, AS17b, Arm14, Arm15, AC16, Arm18, BN12, BJT16, CH14b, CT17, CLLL19, DFL18, DS18, EV10, FL19a, FLM19, FJGM17, GK17, GGL12, GO12, Hon17, HK11, KV17, KV10b, LP13, LCLS15, LT15b, NP12, OSW12, PPW19, PS17, Ped17, QY15, RL17, RS19, SC15a, Son10, Son15, SW14b, TBE14, WT18, Zha10, ZZC18]. conditioning [CLLL16, CLL19]. conditions [CO12, CC15b, Con10, GGR17, HZ18, Kru13, Mat16, MP17, Oku11, PP12]. Confidence [CQYZ18, FP12, HL15, BDK14, BKY10, CK18, FP14, FR18, HL17b, HL12, IK16, KS18a, KS19b, Mik10, Sha15]. Confronting [WZ12]. conjugate [Bog22, CCM19, CCCM22]. connectedness [DY14, HL19a]. conservative [CK18, ILM19]. consistency [DHY16]. Consistent [BNHLS11a, BDH16, BPMSW13, BI18, CSW14, FS19, HP18, LLL16, MSW17, Men14, MHU19, AG10a, Che10b, CT17, DMR16, EJC12, FH19, Gay13, HS14, JPW10, JPW15, KS11, LKH14, LLS14, NHKJ11, PSS18]. constant [BJW15, KL18, Lar11, Ped17]. constrained [LLL18, Par19, Phi91, XTZ10]. constraints [BD12, FH15a, FWZ13, LTU14, SWD11]. Constructing [HK14b]. construction [Cam12, CS11b, IDF11, LW10b]. Constructive [Mat19]. consumer [HL11b]. consumers [Hod11]. consumption [AT14, PSW10]. contagion [CW11a, GMR14a, GMR14b]. contain [AGJ11]. containing [AC12]. contamination [BBCL18, CV17]. contaminations [HP14b]. Contemporaneous [HOY17, DPSS11]. contemporaneous-threshold [DPSS11]. content [CKT16, CP19, vEGM14]. context [Lea16]. context-minimal [Lea16]. contextual [Wil11]. continuous [ASP16, Cha16, CH11, CS13, CPY17, DL17, KW16, LTT17, MP11, PQ12, San11, SL12, TPY19, WY16, Yu12a]. continuous-time [ASP16, CH11, CS13, KW16, LTT17]. contour [Cha12]. Control [AT11, HR19, Ell11, HR17, MW16, iKPS16]. Controlling [PP18]. Conventional [Lea16, HL14b]. convergence [AR12, CF13, CC15b, EJCL14, ID16, JPW15, KPS19]. convex [LS19b]. convexity [SWD11]. Convolution [Sch19]. Convolutional [LXC16]. Copula [ZZ16, CT15b, DT19, MSW19, OP16, ZOZS16]. copula-based [OP16]. copula-like [DT19]. copulas [OOS13]. corner [Jud13]. corporate [BdCK^{+}14, DSW12, XX19]. correct [CX14b, RS19]. corrected [Ike16]. correction [AG10b, CFH10, FP14, HvdAW16, KR10, LS19a, McC17, WT18]. corrections [FVV11, Yan15a]. Correlated [Woo19, CP15, CL18b, GHK18, GHX14, GLL15, Gay13, GHPP18, HSU10, HS15a, Koc16, KLS11, LT11, LR10b, LVX17, WX14]. correlation [ASX16, BPS19, BJS10, CFL16, DR15, GS15, GLH15, HOY17, JKM17, Ped17, PT11b, TH13, Vog12]. correlations [CEG11, OW19, PPW19, SHL11]. Corrigendum [CW14a, CCCM22, HINR14, HLT12a, IL18, IK19, Kar17, Kra13, Kru14, PST22, Xie11]. corrupted [PHL16]. cost [Die14]. costs [HV19]. could [BC11]. counterfactual [CMM18, CDD19, DZ15, FP14]. counterfactuals [AHK12]. Counting [BDV17]. countries [BdCK^{+}14]. country [CGP16, ID16, Sal12]. Covariance [GO11, AM15a, BL17, BCH11, BLL^{+}17, CLL19, CKP10, DLX19, FWZ19, HC16, Hou17, JM16a, JM16b, Kas19, KS11, Kru08, Kru14, LF18, LKH14, LT14b, Por19]. covariances [BPQ18]. covariate [BC14a, HP12a]. covariates [CLLL19, DMP11, DDMP15, Far15, GM17, GKCS18, KLS11, Lee16a, SH13, Sto12, Wol16]. covariation [BNHLS11a, Ike16, LT14a, PHL16, PM17, SX17b, Zha11a]. Cox [GM11]. creative [Slo17]. Credible [Man16]. credit [LW10b, PMM15, TZ11]. crime [DNR10]. crisis [BC11]. criteria [CK12, HINR12, HINR14, KT11, Spa10, Tet12]. criterion [FOZ19, GOS19, Kha13, WZZ10, ZWZ13]. critical [Yan15b]. critique [BI14].

Cross [PSU11, WU15, ZY15, AAMS10, Arb13, BFK12, CGL12, CFVN19, DL11, DGP15, EM14b, EV17, GZWZ16, HV17a, HR15a, HLOW16, HP15, HS17a, ID16, KMS14, Kim11, KS18b, KP13, LR16, LZZZ19, MK14, OÁ19, Pak19, Rob12, Sal12, SJ12, WWYL10]. cross-country [ID16, Sal12]. cross-quantilogram [HLOW16]. cross-racial [AAMS10]. cross-section [CFVN19, EV17, MK14, Pak19]. Cross-sectional [PSU11, WU15, BFK12, CGL12, DGP15, HV17a, HP15, KMS14, LR16, OÁ19, Rob12, WWYL10]. cross-validated [HR15a]. Cross-validation [ZY15, GZWZ16, LZZZ19]. cubic [CG10]. CUE [CY12, HLMN11]. cumulant [EJW14]. Curve [AB18, BC10, LL10a]. Curves [Kra13, ELN11, VA89]. cycle [HP14a, Khw10]. cycles [ADG11]. cyclical [FP11b].

D [Ahn15, ADG11]. Daily [CGH^{+}19, DLM17]. Data [ACR19, Gal17, Pat11a, ALS13, AG10a, ALLM15, ASJL12, ASX17, ASX19, AT14, AH12b, ACG17, ALX11, Bai10, BL17, BJS10, BFK12, BBCL18, BN12, BBP15, BBG^{+}19, BCH11, Bha14a, Bha13, BW15, BKMZ16, BLV15, CCF18, CMM18, Cha16, CCC15, Che10b, CGL12, CM17a, CH18b, CW18, CFVN19, sCJP14, CKP10, CPvdW16, CPTV17, CP15, CL18b, CP18b, DLX19, DGMZ19, DL11, DGP15, DEvG17, DEMY18, FK19, FLWZ15, FGPZ17, FPSY19, FMV15, FVV11, FF14, FAM13, GG17, Gal11, GK16, GLL15, GN13, Gay13, Gha17, GM16, Ghy16, GHM16, GKS16, GK15, GHPP18, GLH15, GV19, Gug10, HL19a, HOY17, HMR13, Hay10, HP15, HP18, HW12b, HLW15, HZ18, IDF11, JvN11, JJ14, JJ11, Kal11, KMS14, KFSvD19, KPT16, Kim14, KS14, KW16, KSY17, KKLW18, KF19, Koc16, KL18]. data [KLZ19, KP13, LF18, Lam10, fLY10, fLY14, Lee14b, LB19, LML11, LL10b, LL15b, LTT17, LB17, Li17, LP18a, LLS14, LZW^{+}18, LT14b, LKJ18, LS16, MKS16, MZ18, MXY12b, MSW15, MM17a, MSW17, MK14, MSS11, MZ16, OP16, OY19, PU11, PHY18, Phi10, QW12, Qia16, QS16, QfLY17, Rob11, RV15, RR15, RV18, Sai13, Sas15, SX17a, SW10, SfL17, SH13, SW14a, SJ12, SL13, SY15, SJZ15, SJ18, SUZ19, SM18, SHHW18, TT10, WX15, WWYL10, WL12, WL14, Xiu10, YY19, YdJfL12, Zad16, ZWZ13, ZZ16, ZWZ19, ZB14, dHvdG11, HL19b]. Data-based [Pat11a, Kim14]. data-driven [FLWZ15]. datasets [DK18]. date [HL15]. deadweight [BS19]. dealing [D’H10]. decades [ASK16, BI14]. December [Ano12o, Ano10s, Ano11p, Ano16t, Ano18w, Ano19o]. decision [ILM19, SL12, WX14]. Decisions [MT12, BPQ18, BSW10, DM13, Zam10]. declines [AX18]. decompose [O’D16]. decomposition [AT13, Egg16]. Decompositions [Die14, DY14]. deconvolution [KS18a]. decoupled [Jus11]. decreasing [Jus11]. default [DSW12, GMR14a, GMR14b]. Defaults [BCG19, KLS11]. defined [BCS15]. definite [BNHLS11a, CP18a]. degree [Sch11]. degrees [PL13]. Delta [GMO18, HL18, LSL14, YCW19]. Delta-method [GMO18]. demand [BKM14, CDD19, Fre15, KS16a, Khw10, MSW18, RV14, SI19]. demographic [AV11]. Dennis [Ano12b, Ano16a]. densities [BCRvD13, CKP16, FAM13, sHDW17, RS19, Wu10]. Density [FMS13, Zha10, BBG^{+}19, Cho15, CS11b, DMR16, DPvD11, EJC12, KMPF15, LSL14, MP14, Pel14, RS13, TBE14, Zu15]. Dependence [BCG19, AGHP16, BFK12, BJW15, CLO18, CHC10, CGL12, CC15b, CL15a, CSS18, DMC12, Dez15, DGP15, EV17, GM15, HLOW16, HS17a, Jen12, JP12, JfL15, KMS14, Kim11, KS13, KLMP15, KP13, LR16, LP18b, Mag12, MV13, Pak19, PSU11, Rob12, RT12, SG19, Sha11, SJ12, TH13, WWYL10, WKM15, Yan15b, ZOZS16]. dependencies [BTL13, LTT16]. dependent [Abr19, ASMZ11, BCL14, BCH11, CN12, CCC15, GS13, Kue19, LL10b, Mat12, Wol16, XfL15b, ZWZ13]. derivative [HY10, LP17, LCLS15]. derivatives [CM15a, CPP19, ES10, HMN15, Par16, ZL15b]. description [DHY16]. Design [ADZ14, CY12, DL17, FFM12, ILM19, OXM15, Xu18]. Design-free [ADZ14]. designs [CHS19, HS19b, PY15]. desparsified [CK18]. destructive [CGH^{+}19]. detail [AV11]. Detecting [CDG14, FF11, HLT10]. Determinacy [Fan12]. Determinants [BEK17, BC10]. Determination [LS19a, KKL14]. determine [MP12b]. Determining [Arb13, LLS17]. deterministic [JN18, Kim11]. detrending [Wes15a]. development [Gra10, LMvdK10]. developments [KFSvD19, LT19, PRSV13]. Deviations [BFS19a, Can10a]. Diagnostic [LS15b, GOS19, Zha16b]. diagnostics [KV17]. diagonal [DPY16, HC16]. Dickey [HLT13]. Difference [CF15b, AL12, CLO18, HP13, ZXC15, SY19]. Difference-in-Difference [CF15b]. Difference-in-Differences [SY19]. differences [AAMS10, CLO18, SY19]. differentiability [HR19]. differential [Gil10]. differentials [CS10, Fla10]. differentiated [Fre15]. diffusion [CS13, CZP15, sCJP14, Cho15, sC16, CKP10, CS11b, FMS13, Hou17, KYS14, KL12, Kri10, Kri11, LSL14, Son11, SC15b]. diffusions [ASP12, Cho13, GS19, HLM13, KP17, LC16, WPY11, YCW19]. diffusive [TTG11]. Dilation [GH13]. dimension [CH10, GR18b]. dimensional [ASX17, BW14, BCF18, BMN19, CK18, CCGR15, CMM18, CCC15, CGY15, CP11, CT15b, FGZ19, FHLZ15, FHLZ17, GM15, GKCS18, HC16, KC15, Koc16, KP19, LZL^{+}16, LLL18, MS18, Mas17, MM16, MHU19, OP16, OW19, Pel19, Shi16, SUZ19, TYZ18, WWZ16, WLC19, Zhu18]. dimensions [CGL15, Kas19, LS19a]. Direct [Kri17, FL16, ZA10]. Directional [AT16, FS14, HLOW16, SV12]. directions [AT16, APT18, CS14b]. Dirichlet [CG10, JM14]. discern [CM17a]. discontinuity [CRY16, CHS19, DL17, FFM12, HS19b, OXM15, PWM11, PY15, Xu17, Xu18]. discontinuous [BNW19, CF19]. discount [PS12, Wes12]. discounts [FS16]. discourse [WXC14]. discovery [FFG10]. discrete [AM10, BKY10, BDKS18, BHH12, CLR10, Dav17, Gew12, HR16, HLW15, KV17, KW16, Kom13, LS15a, LB19, LS15b, Mat19, NR10, Wil11, Xu18, ZLZ13]. discrete-time [KW16]. discretely [FF11, GS19]. discrimination [BCDN12]. Disentangling [BBGV14, FH15c, TB10]. distance [AR12, AT16, EJW14, FS10, FS14, GKN17, HLR11, HMT12, Kol18, LP13, Tor17]. distance-based [HLR11]. distances [SV12]. distortion [RY11]. distortions [Ket19]. distributed [ChKS15]. distributed-lag [ChKS15]. Distribution [DV10, DE12, Kha13, Liu15, MP13b, BTL13, CH14b, CZP15, CTV19, DLM17, DH14, DHKS18, FGZ17, FR19, FiKY16, GT18, GM17, GMHS12, HvdAW11, IL10, IL18, JWY18, KW10a, KS18a, LML11, MRT16, MP14, Myk19, Sho16, ZG10]. Distribution-free [DV10, DE12, CZP15, HvdAW11]. distributional [Rot10, XX19]. distributions [ALW12, And17, Bas19, BDK14, CS11a, DT19, DKK14, Egg16, GLPR11, GRV11, GK18a, Kau15, MM11, MP13a, NP12, Oga13, OP16, PRSV13, PSW10, PQ12, RS17]. disturbances [KP10a]. diurnal [CHP18]. diversification [XX19]. Divided [ST15]. dividend [WW15]. Divisia [FS10, FS14]. Dmk/$ [JJFG10]. Do [AGJ11, MM11, XX19]. doctor [Gil10]. Does [Jus11, Kau15, LPS15, vEGM14]. dollar [ELN11]. domain [CET14, LR19]. domestic [BEK17]. dominance [AT17, BDH16, CJMMPA15, CKT17, LSW10, PKA18, Sch11]. Dominating [FM10]. Double [GG17, WY16, GLL19, LLZ15, ZZL18]. drift [dHvdG11]. driven [BBG^{+}19, BGK19, CHO19, FLWZ15, FZ18, HWY16, KT14, Tod11, BL15]. drug [Dez15]. drugs [Dez15]. DSGE [HINR14, CR10, DSS17, Gal17, GQIK17, HINR12, HS12a, Sai13]. dual [Kai16]. dummy [HV17b]. duration [ACG17, CDS13, Che19, CW10, HW14b, LS15b, LT15b, NR10, Xu18]. durations [Ike16, LKJ18]. during [CFH10, Gil10]. Dynamic [AM10, BE16, BGZ11, BC14b, DHS16, FHLZ15, FHLZ17, HL11a, HHV16, KP12, MW19a, PZC19, SS16, TvD13, AH12a, AK12, AFS13, AC15, ABHP16, BW14, BH17, BN12, BBG^{+}19, BPMSW13, BKMZ16, BE10, BE11, CX12, CN14, CLLL16, CLL19, CPvdW16, CP15, CEG11, CT15b, DGMZ19, DGR11, EMvS11, EW15, EV10, Fan12, FVGQRR15, FGS18, FL11, FL19b, GG19, GHK18, Gal11, GK15, GKN17, GPY10, HMPV11, HL17a, HOY17, HP13, Han18, Hay10, HP15, HR16, HS17a, HLW15, HZ15a, HZ15b, HZ18, HS12b, KG14, KS19c, Khw10, KV10b, KS12d, LW10a, Lee12, fLY14, Lee14b, LMZ17, LOS18, Li17, LR10a, LX19, LS16, MK14, Pak19, PPW19, PC14, QfLY17, RvdHW14, Sas15, SX17a, SfL17, SH13, SL13, SY15, SI19, Tho14, Yan18, YdJfL12]. Dynamics [LPT10, ABTT15, BBGV14, BKLS16, BS18, CDS14, CF12, CNT15, CLR10, GR18a, JvN11, NNN11, OY19, PTV16, Wu11].

earnings [BS18, CF15b, JWFS16]. earthquake [FH15c]. Econom [HINR14, HLT12a, Kra13, Kru14, Xie11]. Econometric [CM15a, Hau19, JJ10, Jud13, KMN11, SX17b, Shi16, Tod11, Wes12, DGMZ19, GM16, HM14, Kai16, KS19a, Li10, LL10b, PL16, TvD13]. Econometrics [ASK16, BW15, CW14a, CCCM22, CL10, ET14, GKP12a, GMR14a, HKRV18, IL18, IK19, LMT15, PST22, BDM16, CHL14, CHH18, CCGR15, DKK14, Gew96, Hu17, Kar17, Kea10, Lam12, Rus10, Slo17, ZZ11, Zha11b, ZG10, Blu10, Ano12a, Ano13b, ET14]. Economic [McF12, BSW10, DHN16, DEvG17, FH15c, HM14, LT19, LS12, OP15, WP14].

Economics [HKRV18, BDM16, Hu17, KFSvD19, LW14a, ZZ11]. economy [JW10]. Edgeworth [Kap15a, Por19, ZMAS11]. Editor [Bha14a]. Editorial [DS10, KFSvD19, LZ19, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17b]. Editorial [Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l]. Editors [ASLN19, CCKR12, DMR17, DKK14, ET14, HKW12, HW17, ILT11, KS12e, KS16b]. educational [BDS14]. effect [BC14a, BSV12, BNW19, CCF18, Che10b, CFW14, Che17, Dez15, DGJ19, DH14, FR19, FM12, Geo10, HS19b, LY16, LT11, LZW^{+}18, MSV19, OP15, SS16, WH10, Wes15a, Zha11a]. effects [Abr19, ALS13, ALLM15, AM15b, BL17, BJS10, BFK12, BR12, BI18, BHP12, BH16a, BLV15, CLO18, Cat10, CSW14, CHS19, Chi10, CP15, CL18b, EV17, FP12, Far15, FGPZ17, FPSY19, FVV11, FVW16, FFM12, FH15c, GG19, Gal11, GN13, Gay13, GK15, GLH15, GV19, Hag19, HL14a, Hay10, HHV16, HR16, HL11b, HW12b, Hsi18, Ike16, JWFS16, JSE14, Jus11, KGMR12, Kau15, KS13, KM16, KZ15, Kle10, KSV15a, Koc16, KTT19, fLY10, jL12, fLY14, fLY15, LR15, Lee18, LL15a, LO15, LB17, Li17, LLS14, LS16, MKS16, MK14, MSW18, Mou15, NDT19, OÁ19, PSW10, Par16, PHY18, RW11, RV15, Rot10, SWC12, SfL17, SJZ15, SJ18, SM18, SY19, VRW18, Vog12, Woo19, WL14, XX19, Yan18, YdJfL12]. Efficiency [GL16, ST12, AC12, AS10, AT17, CFFS16, FLZ11, GKS16, HLX19, Ots10]. Efficient [AR12, AB18, BL16, Cat10, CP18b, Cos13, FR17, FL13, HL10, Han16, HN12, KV10b, fLY14, Lee18, LXZ16, LR10b, LPLZ18, LX19, Poi17, Rob10, RV15, ZLCL19, APT18, Egg16, FGL11, HOvD12, HS17b, LZW^{+}18, LfLB10, McC12, MZ16, Qia16, SK16, SI19]. eigenpairs [HK14b]. eigenvalue [LF18]. eigenvalue-regularized [LF18]. either [HZ15a]. elementary [BDV17]. elicited [AHK12]. Eliminating [dHvdG11]. ellipse [HL17b]. Elliptical [Kra13, VA89]. ellipticity [FJGM17]. emerging [RY11]. Empirical [AGS11, AGKZ18, DS15, HLM^{+}12, Hor14, OXM15, AG12a, CHHW11, Can14, CF15a, CCC15, CGL15, CTV19, HP16, HN12, Hu17, KS12b, KS14, LW10b, MJ11, Oga13, Ots10, OSW12, PKA18, SW14b, YXZ14, Zha16a]. empirics [MPP10]. employment [AV11, LW10a, LMvdK10]. Endogeneity [APS16, KT16, CGY15, CFVK15, FH19, GH16, GKCS18, Gut16, MM17a, SW10, SS16, SVV16, YP18]. Endogenous [APS17, DF12, HLP16, AMM11, AT16, BMS18, Che15a, CDD19, D’H10, DGJ19, FT14, FF14, HHB12, HV17b, HKY19, Ike16, KPT16, KTT19, MKS16, MW16, NDT19, PM17, QfL15, QfLY17, SW16, ST12, Woo14, Zhu18]. endogenously [CFR18, Cos13]. energy [BJW15]. engineering [LZ19]. English [ALGQ16]. entries [CQYZ18]. entropy [RS17]. entry [FT14, LZ12, MSX13]. environmental [APS17, KT16]. environments [BEE15]. episode [Gil10]. epoch [Jen12, JP12]. Epps [Zha11a]. equal [CM15b, GMP17]. equalities [Shi16]. equality [CP18a]. equation [AK12, AKM11, Bas19, DL16, Slo17]. equations [CH14a, DFH^{+}14, EJW14, GP13, HZ15b, KV10a, LP13, LW14b, MP17, OP18, YfL17b]. equilibria [CC15a, Xia18]. equilibrium [An17, CFH10, CPvdW16, FVGQRR15, LW10a, LM12]. equilibrium-correction [CFH10]. equities [CDS13]. equity [ABTT15, BNHLS11a, CDS14, SX17b, WW15]. equity-index [ABTT15]. equivalence [Lav14]. ERA [Sia17]. Erratum [GMR14a]. error [AH12b, An17, ADN^{+}16, BC14a, BvH17, CRdM17, Che17, DL17, FLM19, GKS15, GM17, GLH15, HR17, HvdAW16, Hu17, HW17, JvN11, Joc15, KPY11, KS18a, KSV15b, KR10, LMZ17, LS19a, MSW17, MRSE11, Ona15, PHL16, Pel14, PK17, PSY13, iKPS16]. error-correction [HvdAW16, KR10]. errors [BPQ16, CCJ12, CIY15, CG10, DL16, DMN19, EJW14, FGS17, GKN17, HSVV13, HP15, HW10, HS15b, KS19b, LR10b, MM16, MM19, NOR13, Son15, SY15, VW14b, WH11, ZXC15]. errors-in-variables [EJW14, GKN17, KS19b, WH11]. Essays [CHH18]. estimands [And19]. estimate [ASX17, GK18b, HZ15a]. estimated [GMP17, HL12]. Estimates [FM12, AV11, Gup18a, HR16, Man16]. Estimating [AHS10, CH18a, CS10, CPvdW16, CGE10, DP17, FVGQRR15, HL14a, HW14b, JLZ19, JM14, Khw10, Kim11, KV10a, Lar11, LML11, LC16, LKJ18, OQ11, PHL16, PLG17, QW12, QfL15, Sai13, SW10, SW14a, WH10, Zha11a, ZB14, iKPS16, CCF18, CFW14, CP18b, DMC12, FiKY16, HK14b, HLM13, LP13, LS15b, MW16, Phi91, ST12, TB10, WPY11, XTZ10]. Estimation [AH12a, ALLM15, ACR19, BFK16, BNW19, BH13, CFPY10, CCeC18, CW15b, DK18, DH14, EV17, FZ18, GTY13, GLPR11, GRV11, GN13, HKRV18, Han18, Has11, HLW15, HK11, Hua14, KSV15a, KL12, KS12d, KS12f, fLY10, LTU14, fLY15, LB17, MS18, Mar14, MSW15, MSW18, Oga13, PM17, SX17b, SW17, SM18, YdJfL12, ZLX15, ADZ14, Ahn15, ASMZ11, ASAM15, AFS13, ADS12, ALW12, AC15, AC13, AR12, AB18, AL10, AL12, AMM11, AdCGIV11, BN13, BL16, BN19a, BKW17, BH17, BN12, BPMSW13, BC14a, BC15, BK10, BS12, BGZ11, BMN19, BJR17, BvdW11, BLL^{+}17, BMS18, CX12, Cam12, Cat10, CNT17, CFFS16, Cha17, Cha13a, Cha16, CYZ15, CDH10, Che10b, CZ10, CZ11b, CZ12, CSW14, CZJ18, Che18, CW18, CF19, CLLL19, CLL19, CLSY19, CKK15, sC16, COP10, CV17]. estimation [CP15, Cíz12, CL18b, CLR10, Cos13, DGG14, DDMP15, DMP18, DMR16, Des11, DPY16, DFH^{+}14, EJW14, EJC12, EJCL14, FWZ19, FK19, Fas13, FH15a, FS12, FSV14, FV15, FN18, FLZ11, FZ15, FR17, FNN12, FH15b, FL19b, GR13, GG17, GGR17, GHK18, GJZ15, GP13, Gay13, GP12, GR18a, GO12, GKN17, GJ17, GLL19, GR18b, HL10, HR17, HY10, HINR12, HINR14, HSVV13, HL17a, HLR11, HW12a, HP13, HK14a, HP15, HLPY15, HKLP15, HP16, HP18, HS15a, HN12, HMN15, Hor12, Hor14, HL17c, HV17c, HS15b, HKY19, HLP12, JCLL10, JWY18, JfL18, JfL19, JL13, JPW15, KRW17, Kas19, KS19c, KP10a, Kha13, KS11, KK12, KKLW18, KS18b, KM16, KYS14, KSV15b, KV10b, Kri17, Kri10, Kri11, Kru08, Kru13, Kru14, KPT12, Lam10, LMS17, LP13]. estimation [jL12, fLY14, LKH14, LS15a, LR16, LMZ17, Leu15, LT15a, LL10b, LL12a, LLL16, LXZ16, LTT17, LZL18, LPZ19, LR10a, LR10b, LfL10a, LCLS15, LPLZ18, LZW^{+}18, LX19, LfL10b, LT14b, LS16, MS17, MR11, Mas17, MSW17, Men14, MHU19, Miy16, MW19b, MZC19, NDT19, OOS13, Oku11, Ona15, OP15, PPT11, Phi14, PP12, Poi17, PY15, QS16, QY15, QfLY17, RL17, RS15, Rob10, Rob14, Rot10, RS17, SX17a, Shi16, Shi12, SH13, SWD11, SVV16, Slo17, Son15, SL12, SJ10, SJ12, Su12, SL13, SY15, SC15b, SH16, SI19, TBE14, TTG11, TTG14, Tor17, Tsi12, VW14a, WH11, WIW13, WZ16, Wol16, Woo14, Xia18, Xiu10, XfL15a, XfL18, YY19, YfL17b, Yan18, Yu12a, Yu12c, Yu15, Zha16a, Zha10, ZLCL19, Zin18]. Estimator [Wu10, AG10a, AK12, AKM10, And10, BFS19a, Che10a, DGR11, HLMN11, Hay10, HP14b, Ike16, KS19a, LF18, LfLB10, Liu15, MMS19, Ona12, PSS18, Sun11, Xu18, Yan15a]. estimators [ASP16, AG12a, AKM11, BR11, BNHLS11a, BCH11, BH16a, Bre15, Can14, CF15a, CF13, CC15b, CKP10, DLX19, DFS12, FGL11, FVV11, FS19, FM10, GS12, GS19, GMHS12, GQIK17, HO12, Han10, Han12, Hou17, HLW13, Ibr14, IL10, IL18, JPW15, KS17, Kue12, Lee14a, Lee16b, LOS18, Lee18, MJ11, MZ16, Pat11a, PM11, ZMAS11]. ETFs [CDS14]. euro [ELN11, Fav13]. European [Xiu14]. Eurozone [ASLP14]. Evaluating [Bha13, HS12a, DLX19]. Evaluation [BA11, AT14, BCIT14, DSS17, GKR13, MP11, RS13, RY11]. Event [HWY16, CHO19]. events [GMR14a, GMR14b, GKM14]. Evidence [BEK17, BNW19, BdCK^{+}14, HL19b, MS16, Sia17, DS15, KS14, Wu11, WW15]. Ex [KS18b, ABHP16, CKP10, KS19c]. Ex-post [KS18b, ABHP16, CKP10, KS19c]. Exact [BDK14, BFS19a, GT18, HM18, Shi12, Arm15, GJZ15, GS13]. Examining [AFPZ16, BH14b, AGHP16]. example [DHN16]. Examples [And17]. excess [AGJ11]. exchange [BCN11]. exchangeable [BJR17]. excitation [ASLP14, BLY18]. exciting [DEMY18, GKM14]. excluded [LT15a]. exclusion [KV10a]. Exit [GR18a]. exogeneity [Can14, GMR14a, GMR14b, JW10, KP13, WP14]. exogenous [CFR18, CP15]. expansion [Kap15a, LSL14, Xiu14, YCW19]. expansions [Cho13, LC16, ZMAS11]. expectation [Egg16]. expectations [AHK12, Bia16, Fan12, YfL17a]. expected [BA11, LM11, PZC19]. expenditures [CDD19]. experience [CGE10]. experiment [HL19b, Sia17]. experiments [AH12a, FP12, Gew12, Sto12]. expiry [FH15a]. explain [CF12, CHP18]. explained [CP19]. Explaining [AAMS10]. explanatory [DEvG17, MW16, Woo14]. Explicit [Cho15]. Exploiting [BPQ16]. exploratory [CFSHP14]. explosive [CPY17, Mag12, PL16, WY16]. exponent [PW12]. Exponential [FWZ13, KT14, Wu10, CW10, DJfL15, JfL18]. Exponentially [TYZ18]. exports [BEK17]. Extended [MP13a, Zad16, CLY17, CLSY19, DM13, Des11, Ped17]. Extending [PWM11]. externality [Miy16]. extraction [Mac10]. Extremal [DMZ18, DMC12]. Extreme [CDES14, OW19, AGHP16, BTL13, GKM14, TPY19]. Extreme-quantile [CDES14]. extremogram [DMC12]. Extremum [HMN15].

F [HS17b]. facing [CH14a]. Factor [Han15, KF19, LZK18, PU11, WLC19, Ahn15, ASX17, AFTV19, AM15a, ABHP16, BW14, BL16, BN19a, BKW17, BH17, BCF18, BPMSW13, BKMZ16, BE11, CH18a, CRT15, CDG14, CH15, CP11, CFSHP14, CS14c, DLX19, DL11, DGR11, EMvS11, FXY17, FWZ19, FGS18, FL11, FHLZ15, FHLZ17, GOS19, GP14b, GMHS12, Han18, KRW17, KS19c, KL18, LLL18, MS18, MSW19, Mas17, MHU19, Ona12, Ona15, PP15, Pel19, PS12, RS15, SW17, Tra13, Wes12, YT15]. Factor-Augmented [Han15, CH15, GP14b, KRW17]. factorizations [BLL^{+}17]. factors [BN13, BE10, CCH13, CGY15, DLX19, GMP17, HL11a, HMPV11, KZ15, KLZ19, LL10a, LLS17, LZK18, Ona12, PT11a, PT11b]. failure [AC13, LLS15]. failure-time [LLS15]. families [HY10]. family [MJ11]. fast [ABH14]. Fat [DJS^{+}13, KLMP15, PPW19]. fat-tailed [PPW19]. FAVARs [MS16]. FDA [ILM19]. FDI [CCF18, DJfL15]. feasible [AG12b]. features [FP11b, LML11]. February [Ano10p, Ano11s, Ano12t, Ano14w, Ano15s, Ano17v, Ano18t, Ano19q]. feedback [AM15b]. feedbacks [Hay10]. Fellow [Jud13]. Fellows [Ano12p, Ano13o]. female [CGE10, FM12]. festschrift [MXY12a, MXY12b]. Filter [MRR15, KP15, PdSSGK12]. Filtered [GS18]. filtering [BCRvD13, DGR11, HS19a]. filters [fLY15, MP11]. Finance [HKRV18, BJW15, KFSvD19]. Financial [CMY16, LMT15, AT14, AGHP16, BC11, Bha14a, BKLS16, BPQ18, CM15a, CJMMPA15, CDES14, CPvdW16, DDSW18, DHS16, DY14, DEMY18, FK19, KS14, KW16, KKLW18, LZ19, MZ18, TZ11, ZZ16, Zha11b, ZG10, ZL15b, ZB14]. financing [LPT10]. fine [ABTT15]. Finite [GS13, And10, AKM11, CPT14, GM15, Mar14, Tra16, VKN12]. Finite-sample [GS13]. Firm [Bha14b, BEK17, CO15, CS10, HMR13, Ibr14, KS12f, O’D16, vEGM14]. firm-level [BEK17]. firms [DY14, GR18a]. First [HP13, AHS10, An17, AGKZ18, BS12, DG17, HLM^{+}12, HS13b, HMS13, HLP12, KPT12, MMS19, MS12, MSX13, PTV16, Zin18]. first-order [DG17]. first-price [AHS10, An17, AGKZ18, BS12, HLM^{+}12, HS13b, HMS13, HLP12, KPT12, MMS19, MS12, MSX13, Zin18]. fiscal [LPT10, MS16]. fish [HHS12]. fit [BDK14, Bre15, CTV19, KV17, MO12, MP13a, ZOZS16]. Fitting [EMvS11, PK17]. fix [Sun14]. Fixed [Li17, Sun14, Zha16a, Abr19, ALLM15, AFTV19, BFK12, BI18, BH16a, BLV15, Che10b, CSW14, EV17, FGPZ17, FPSY19, FVV11, Gal11, GK15, GLH15, GV19, HW12b, JSE14, Kap15a, KS13, KSY17, fLY10, fLY14, fLY15, LR15, LL15a, LO15, LLS14, LS16, MKS16, MP14, MSW18, PHY18, RV15, SfL17, SJZ15, SJ18, SM18, SI19, SY19, Vog12, VW14a, WH10, Yan18, YdJfL12]. Fixed- [Sun14, SY19, VW14a]. fixed-b [MP14]. fixed-bandwidth [KSY17]. fixed-effect [Che10b, WH10]. Fixed-effects [Li17, Abr19, BI18, Vog12, Yan18]. Fixed-smoothing [Zha16a, Kap15a]. fixity [Hil16]. flexibility [FM12]. flexible [CFW14, KP10b, LLL15a]. flights [DEMY18]. flow [BXZ15]. flows [AV11, DLM17]. focus [ES10]. Focused [DiT16]. followers [BE16]. follows [CS13]. food [HL11b]. force [AAMS10, DP17]. Forecast [BBG^{+}19, Arb13, Cha13b, CM15b, DSS17, Ell11, HW14c, KMPF15, MP12a, Pat11b, WY12]. Forecasting [AM15a, BV11, BN19b, CFH10, CCH13, cCTZ14, CH15, KS14, TY17, WBH13, ABM11, AdCGIV11, BH17, BKMZ16, BPQ16, BPQ18, BCN11, CG11, CGP16, CS14c, CPR10, DHS16, DK18, FXY17, Fav13, GS11a, GR14, GKP13, HM14, IJR17, ILT11, KP15, KS15, LRV13, LTU14, LLL15a, MW19a, PPT11, RS11, Siz11, TvD13, ZLZ13]. forecasts [DPvD11, HS12a, MM11, MMFG12, PPP13]. foreign [BCN11]. form [ABH11, BJT16, CNT17, Cho13, Cho15, HS15b, LC16, LLS14, Sal12, Siz11]. formal [PW12]. formation [Leu15]. forms [PG17]. forward [DSW12]. Four [ASK16]. Fourier [Tsi12]. Fourth [HMR11]. Fractional [GK17, CNT15, CNT17, CHL18, CAS15, CV17, FNN12, GP14a, HV17a, Has11, JN10, Las10, LT14a, Nie10]. fractionally [EV17, HR10, PGM15, RV15, Shi12]. Frailty [BCG19, KLS11]. frailty-correlated [KLS11]. framework [ABH11, BFS19b, ChKS15, HS18b, Lav14, MBC15, TB10, Zha16a]. France [DM13]. Fréchet [ZZC18]. free [ADZ14, CZP15, CM17a, CDR11, DV10, DE12, FH19, HvdAW11, Kha13]. frequencies [AGK10, LS15a]. frequency [ASMZ11, ASJL12, ASX17, ASX19, ADFK16, And16, BW15, BKMZ16, Cha17, CET14, Cha16, CM17a, sCJP14, CPvdW16, CPTV17, CV17, CP18b, DEMY18, FK19, FP11a, GM16, Ghy16, GHM16, GHS16, HP14b, Kal11, KW16, KP17, KKLW18, KF19, KL18, KLZ19, LR19, LTT17, LP18a, LT14b, LKJ18, MM11, MZ18, MS16, MW13, MZC19, OP16, PHL16, Pel19, Qia16, TT10, Wu11, Xiu10, Zad16, ZZ16, ZB14]. frequentist [CH15, CGP18]. frictionless [Cha17]. frictions [DLM17, DHS16]. Frontier [FSV14, KSV15b, APS16, APS17, BI18, CFFS16, CSW14, DG11, DFS12, DP17, GKS16, KPT13, OÁ19, SVV16, Tsi12, WH10]. Frontiers [LMT15, GH16, MR16, PS12]. Fuller [HLT13]. fully [CP19]. function [CZP15, CTV19, Ell11, FS10, FS14, FOZ19, GP12, Gut16, HINR12, HINR14, LCLS15, LPLZ18, LCY10, MW16, Pel14, PL15, Sch11, SWD11]. Functional [BCF17, CJY15, CFHZ19, LL12b, MSS11, PQ12, Sun16, CGL15, FV15, HKR14, HR15b, LSW13, LTT16, LLS14, LXC16, Sal12, SG19, SH16, SM18, Zha16b]. Functional-coefficient [Sun16, SM18]. functionals [FGZ17, FR19, San11, ST12]. functions [AC12, AT16, BKM14, CS11a, CFGV13, CFPY10, CLLL19, Cho15, Die14, DH14, GP14a, HL12, HLP16, IK13, IK19, Kha13, KT12, Kri17, LRV13, MJ11, TT10, Wol16, YXZ14, ZLX15, ZLCL19, iKPS16]. fund [BCIT14]. fundamentals [Bha14b]. further [HS18b]. fused [QS16]. future [BCN11, CS14b]. futures [CNT15, ST15].

game [AH12a, AL10, HHS12, WXC14]. games [AL12, DHKS18, KT12, Kli15, LT15a, LVX17, Mar14, WX14, Xia18]. Ganz [BSV12]. gap [DMZ18, MW17]. GARCH [AW15, BDR14, BvdW11, BHK^{+}13, CFHZ19, CDI12, Con10, CM16, DFL18, DMR16, FLZ11, FWZ13, FJGM17, HP12a, HP16, JM13, JM16b, KF19, LLL15b, LZZL18, Ped17]. GARCH-in-Mean [CM16, CDI12]. GARCH-Itô [KF19]. Gaussian [AC15, BCC^{+}15, BEE15, BHP12, CCJ12, CFL16, FS19, FLZ11, FSW10, HvdAW16, HW12a, KMS16, LMS17, Mac10, MM16, PP15, TH13]. GDP [ADN^{+}16, BCLL16]. GEL [CPP19, GRS12, HP16, JfL19, Lee16b]. gender [Fla10]. general [ACM17, CFPY10, FL11, JN18, MBC15, Yan15a]. Generalised [KMPF15, LPLZ18, PL15, ZLX15]. generalization [BRG14, CP18a, HV17b]. Generalized [BH17, BDKS18, CFGV13, CH11, CW11b, DPY16, FGZ19, MK14, Tri11, VKN12, ZXC15, BBP15, Can14, CF15a, CCC15, Che10a, DMP11, DDMP15, DMP18, Han12, HW12b, HS18b, Lee14a, LLS15, Oga13, WWYL10, Zha16a, ZLCL19, ZG10]. generate [CM17b, MP10]. generated [LP18b]. Generating [CHL18]. generation [ERS17]. geo [Fat10]. geo-logarithmic [Fat10]. Geometric [AS14a, RS17]. Gibbs [FAM13]. GLAD [GLL19]. glass [CC15a]. Global [JJ11]. GLS [AG12b]. GMM [Tri11, Can10b, CZ11a, CL15a, CL15b, DiT16, DG17, DH18, GO12, HZ15a, HS17b, Kue12, fLY14, LfL10a, LfLB10, LfL10b, Su12, SK12, dCGKL19]. go [HS18b, BvdW11]. GO-GARCH [BvdW11]. good [BEE15, KT16]. Goodness [Bre15, ZOZS16, BDK14, CTV19, KV17, MP13a]. Goodness-of-fit [Bre15, ZOZS16, BDK14, CTV19, KV17, MP13a]. governance [BdCK^{+}14]. government [Fav13, SGF11]. governments [ST15]. Gradient [HLPY15]. Gradient-based [HLPY15]. Granger [AS14a, BM12, GHM16, GHS16, TBE14, WP14]. Graphical [BBCG19]. great [FH15c]. gross [AV11]. group [QS16]. Grouped [BH16a, GV19, SJ18]. growing [TYZ18]. growth [CCF18, CGP16, Ibr14, Kau15, MPP10, Sal12]. Guerre [MMS19]. Guest [ET14, ILT11]. GVAR [Fav13].

Hahn [LO12]. Hahn-Hausman [LO12]. Hanshin [FH15c]. Harris [LL16]. Hastings [WWZ16]. Hausman [LO12, ASLN19, ASX19, Gug10, HHM11, Jos19, Kue19]. hazard [KS16a, Wol16]. hazards [HL11b]. Health [BSW10, GK10, GKP12b, HL11b, Khw10]. heaped [ACG17]. Heavy [DKK14, MdV13, CZ15a, HA13, HP16, Ibr14, MP13a, PRSV13, YL17]. heavy-tailed [CZ15a, HP16, MP13a, PRSV13, YL17]. heavy-tailedness [Ibr14]. Hermite [Xiu14]. HESSIAN [McC12]. Heterogeneity [LL15b, Sas15, BBP15, BC14b, CDD19, CW10, DP17, Dez15, HNS14, HW14b, HS19b, HMS13, Kaw17, LT11, Mat12, NR10, SVV16]. Heterogeneous [Kle10, BFK16, BLV15, CRT15, CP15, GHK18, Lee15, LT11, MSW15, OY19, TU10, YfL17a, Zam10]. heteroscedastic [Pel14, ZWXW10]. heteroscedasticity [LZZL18, Zha10]. heteroskedastic [BCRT16, CNT15, CKT16, GM17, KP10a, Kha13, MM16, MM19]. Heteroskedasticity [KS13, Vog12, AG12b, BJS10, BC15, BS18, BJT16, CNT17, CHN^{+}14, FL19a, HOY17, HP15, JSE14, KS11, LfL10a, LX19, MRSE11, Nor15, Sun14]. hidden [Zha11b]. hierarchical [KP19, OOS13]. High [BMN19, CCGR15, CCC15, CGY15, CT15b, FP11a, MZC19, OP16, ASMZ11, ASJL12, ASX17, ASX19, And16, BW14, BCF18, BW15, CK18, CMM18, Cha17, CGL15, CM17a, sCJP14, CPTV17, CP18b, DEMY18, FGZ19, FK19, GKCS18, HC16, Kal11, KW16, KP17, KKLW18, KF19, KC15, Koc16, KL18, KLZ19, KP19, LZL^{+}16, LTT17, LP18a, LLL18, LS19a, LT14b, LKJ18, MM11, MZ18, MM16, MHU19, OW19, PHL16, Pel19, Shi16, SUZ19, TT10, WWZ16, WLC19, Wu11, Xiu10, ZZ16, Zhu18, ZB14]. High-dimensional [BMN19, OP16, BCF18, CMM18, FGZ19, Koc16, KP19, MM16, MHU19, OW19, Shi16, SUZ19, WWZ16, WLC19, Zhu18]. High-frequency [FP11a, ASX17, BW15, CM17a, CPTV17, FK19, KW16, KKLW18, KF19, KL18, KLZ19, LTT17, LKJ18, MM11, Pel19, TT10, Wu11, ZZ16, ZB14]. Higher [KS12c, KS17, PGM15, GR15, HM18, MP17]. Higher-order [KS17, PGM15, GR15, HM18]. Highly [McC12]. historical [BFvD10a]. hitting [RvdHW14]. hitting-time [RvdHW14]. Hodges [CO12]. Hodges-Lehmann [CO12]. Home [DZ15, BLO14]. Home-purchase [DZ15]. homoskedasticity [WT18]. honest [CK18]. Honor [ASLN19, CHH18, MXY12a, MXY12b]. horizon [PL13]. house [HPY10]. household [DF12]. housing [DZ15]. Hybrid [CF15a, AG10b]. hyperbolic [Con10, LLL15b]. hypotheses [CS14a, JJFG10]. Hypothesis [Cal11, CW11b, DLM17, FP11a, GGL13, Gug10, LY12, LZY14, LLY15, LWY16].

ICP [RH16]. ID [Ano15l, Ano16m]. identically [BS12]. identifiability [HK11, WH11]. Identification [An17, BW14, BKW17, BMDL17, BKY10, Cam12, CF19, CL18b, DF15, FH15b, GG19, Gay13, HW12a, HV17b, HS15a, HMS13, HKY19, JCLL10, KU14, KPT16, Kli15, Kom13, LMS17, LT15a, LR10a, Mat12, MT12, MP17, SX17a, SH13, Xia18, YfL17b, AD13, AC13, AL14, BN13, BMM12, CRY16, Can14, CLY17, CZJ18, Che15b, CFVH^{+}15, CKK15, DL11, DH18, DHKS18, FGZ17, FL18, FR19, GJ17, GRS12, HL14b, HW12b, Hon17, HS12b, IR11, Joh10, JJ11, KS19c, LO15, LPZ19, LVX17, Mat19, MT19, Pos16, Ros12, Slo17, SK12, WX14, Zad16]. identified [BCS15, Can10a, CZ11a, FiKRB12, GMO18, KW14, Kai16, LS19b]. identify [HU10]. Identifying [DGJ19, LY16, SJ18, Gha17, Han18, LKH14]. idiosyncratic [BBGV14]. if [CP19]. IFC [Ano15l, Ano16m]. ignorability [Cat10]. IID [CW11b]. illicit [Dez15]. illness [Gil10]. immigrants [CGE10]. Impact [AT14, CPR10, FH15c, Gug10, GKP12b]. impacts [Lee15, XX19]. imperfect [Pat11b]. implementation [HLM^{+}12, LB17]. Implications [KS15, BCDN12, HW14a]. implied [BGZ11, BCN11, CX14b, HK11]. Importance [DT19, DS15, HOvD12, SK16]. imposing [Hil16]. improve [MM11]. Improved [BCIT14, BJN15, JPW15, Kap15a, BPQ16, LSW10, Seo18]. improvements [PGM15]. Improving [ADN^{+}16, RV14]. Impulse [GQIK17, CS11a, HINR12, HINR14, IK13, IK16, IK19, Li17]. imputation [MM17a]. imputed [DMP11, DDMP15]. In-sample [CM12]. incentive [CF19]. Incidental [HZ18, LP15]. income [BS19, LB19, RS17]. incomplete [AL10, AL12, Leu15, LT15a, Tho14, WX14, WXC14, Xia18, YfL17a]. Inconsistency [FZ10, WT18]. incorporating [LXZ16]. increase [LLS17]. Increased [ASX16]. Increasing [WH19, GR18b]. increasingly [GR15]. increment [Lar11]. independence [HR15b, Mat16, Poi17, Sch19, SW14b, Tor17]. independent [GMR17, KS19c]. independently [BS12]. indeterminacy [Fan12]. Index [CKM16, ABTT15, CJY15, CL18b, DGP15, ES10, FS10, FS14, FPSY19, HS13b, HV17c, JKLM12, Lee15, MZ18, O’D16, Son14]. indexed [Kue19]. indexes [IDF11, dHvdG11]. indicator [CS14a, DMP11]. Indicators [AV11]. indices [Fat10]. Indirect [CC15a, CFR18, FOZ19, GPY10, HKRV18, Li10, BD18, BDKS18, CH18a, DH18, GRV11, GR18a, GK18b, HL17b, SWC12]. Individual [FVW16, ALS13, BHP12, GN13, GK15, KGMR12, KTT19, MP12b, WL14]. industrial [Hu17]. industry [AV11, AH12a, ID16]. inefficiencies [KS12f]. inefficiency [KPT13, OÁ19]. inequalities [AS14b, AS17b, Arm14, AC16, Arm18, BKM14, BCS15, KC15, Koc16, LSW13, Men14]. inequality [AFPZ16, ALGQ16, Arm15, CS14a, Shi15]. Inference [Ana12, AS17b, BCH11, BCRT16, BJT16, CRT15, CFR18, CFL16, CPY17, CPTV17, DM13, DDSW18, FT14, GJ11, GKY14, GR15, HHB12, HS17a, Hon17, IK13, IK19, LTT16, MMS19, Ped17, RV18, SY19, WX15, AFTV19, ALW12, AC13, AS14b, Arm14, Arm15, AC16, APT18, BCL14, BI18, BCIT14, BD18, BH13, BKY10, BvH17, BD15, BDKS18, CC15a, CH18a, Can10a, CCJ12, CNT17, CZ11a, CPT14, CLS14, CL14, Che15b, CHS19, CCeC18, CDI12, DG11, DMN19, DH14, DH18, DT10, EM14a, FP14, FL16, FHLV16, FL18, Far15, FJ19, FZ10, FOZ19, GMO18, GT18, GRV11, Geo10, GP14a, GD19, GK17, GR18a, GK15, GPY10, GMR17, GK18b, Hag19, HR19, HPK14, Han18, HS14, HP16, HM18, HL17c, HT16, HZ15b, HZ18, HR10, JP12, JN10, JPW15]. inference [Kai16, Kap15a, Ket18, KU14, KS13, Kim14, KSY17, Koc16, Kol18, KP10b, KLGS11, KR10, Kru08, Kru14, LOS18, Li10, LT11, LLZ15, LTTC17, LP18a, LS19b, LZW^{+}18, MS12, MP14, MZ16, Par19, PPT11, PL16, PG17, QY15, RT12, RV15, RR15, SC15a, Spa10, Sun11, Sun14, SM18, TBE14, TYZ18, Tri11, Vog12, VW14a, WK16, WZ16, Xia12, YL17, Yu12c, vH11]. Inferences [BL17, KW16]. Inferring [BD12]. Infinite [CP11, HR15a, AD13, BCC^{+}15, FHLZ15, FHLZ17, KKL14, LOS18]. Infinite-dimensional [CP11, FHLZ15, FHLZ17]. infinite-state [BCC^{+}15]. infinitesimal [CS13, Son11]. inflation [CF12, Cha13b, cCTZ14, FS16, Kau15, NNN11]. influence [vEGM14]. influential [Ona12]. Information [sC16, HINR12, HINR14, LZ12, AGJ11, AB18, AL10, AL12, CL15b, DLM17, DHKS18, HL11b, Jud13, Kli15, Leu15, LT15a, LXZ16, O’D16, RY11, Tho14, WX14, WXC14, Xia18, YfL17a, vEGM14]. information-based [CL15b]. Informational [CKT16, CP19]. informative [FL19b]. infrequent [Geo10]. inhomogeneous [Cho13]. initial [HLT12b, HZ18, Kru13]. initiations [vEGM14]. Injectivity [HSS17]. innovation [BE10, BE16]. Innovations [BDM16, BH16b, CZ15a]. inputs [APT18]. instabilities [RS13]. instability [BKW17, BPMSW13, GHLT18, TvD13]. institutions [AT14]. Instrument [CF15a, SW16]. Instrumental [BD15, FV15, HR17, HK14a, Kle10, MSV19, Oku11, San11, BC15, Bre15, CCJ12, CFFS16, D’H10, DFH^{+}14, FS12, FH15b, GS12, GG17, Hor12, HL12, Hor14, HKY19, KLGS12, Kri17, MM19, SH16, Tor17, WK16]. instruments [AKM10, AKM11, BMDL17, BCF12, CY12, Car12, CT15a, CHN^{+}14, DiT16, HHM11, HK14a, Kol18, Mat16, Mik10, MMV14, Phi14, PG17, RV14, WK16, Zhu18]. insufficiently [HKY19]. insurance [KS16a, Khw10]. integer [CDDH15, CAS15]. integral [HSS17]. Integrated [Siz11, VW14a, BR11, BS12, BLL^{+}17, Che10a, COP10, EV17, GLPR11, Hou17, JLZ19, LF18, LXZ16, LT14b, LKJ18, MP17, PL16, PGM15, PM17, RV15, SHL11]. integration [BRG14, BCRT16, CRT10, CNT15, CHL18, CAS15, Fas13, HV17a, HLT16, HL15, Has11, ILT13, Nie10, QW12]. Intelligible [LL10a]. intensity [BE10, CSS18, DSW12]. inter [CDS13]. inter-trade [CDS13]. interact [MP10]. interaction [DF12, GG17, LfL10b]. interactions [YfL17a]. interactive [BL17, FPSY19, HL14a, Hsi18, LS16, MSW18, SfL17, SJZ15, SJ18]. interest [AGJ11, BC13, CG11, Dua16, ELN11, ERS17, MR11]. Intergenerational [HR16]. internally [HS14]. International [CDF12, Hil16, RH16]. interval [LB19, Man16, SHHW18, WX15]. interval-data [LB19]. interval-valued [SHHW18]. intervals [FP12, HL17b, Myk19]. intra [CS10]. intra-firm [CS10]. Intraday [LT15b, CHP18]. Introduction [ASLN19, BDZ11, GKP12a, HS14, HKW12, Bha14a, CCKR12, DMR17, DKK14, DS10, ET14, HW17, ILT11, KFSvD19, KS12e, KS16b, ZZ11]. invalid [DiT16]. Invariance [Kue19, JJ14]. invariant [BBP15, LR10a]. inverse [Kri17, Xie88, Xie11]. Inverting [HL17b]. Investigating [PSW10]. investigation [DHS16, KS12b]. investment [AT14]. investor [BGZ11]. Irregular [JfL18, CL14]. irrelevant [CLLL19, Phi14]. Issue [ASLN19, HKRV18, MT12, ET14, GKP12a, ILT11, LZ19, ZZ11]. Itô [GT18, KF19, YCW19]. Italian [BDV17]. iteration [Dav17]. Iterative [DFH^{+}14]. IV [And19, CPP19, CS12, FP14, HLR11, HU10, HN12, HZ15a, MMV14, RS15, SW16, WWYL10, Ahn15]. IVs [LMZ17]. IVX [Lee16a]. IVX-QR [Lee16a].

J [Ano12b, Ano16a, CW14a, CCCM22, GMR14a, HINR14, HLT12a, IL18, IK19, Kar17, Kra13, Kru14, PST22, Xie11]. Jackknife [BC15, Cha13a, HR12a, LS15c, ZWZ13]. January [Ano10m, Ano12p, Ano12o, Ano13o, Ano13r, Ano14u, Ano15m, Ano16p, Ano17r, Ano18o]. Japanese [KS12b]. JE [Ano12p, Ano13o]. Jerry [ASLN19, Jos19]. JIVE [HK14a]. job [AV11, FM12]. JoE [LZ19]. joiners [DP17]. Joint [IK16, Wu11, ELN11, FGZ17, NP12]. Journal [ASK16, CL10, GKP12a, Ano12a, Ano13b, ET14]. July [Ano13t, Ano15r, Ano16r, Ano17s]. Jump [ADS12, BTL13, BT14, CSS18, FMS13, GS19, Hou17, HV17c, JKLM12, LC16, LTTC17, Tod11, TTG11]. jump-diffusion [FMS13, Hou17]. jump-diffusions [GS19, LC16]. jump-driven [Tod11]. Jump-robust [ADS12]. Jumps [LM12, TB10, ASJL12, ASX16, BW15, BLY18, BCN11, CRdM17, CSS18, CPR10, DEMY18, DS15, FF11, FP11a, JKM17, MR11, Par16, TZ11]. June [Ano10r, Ano11q, Ano12u, Ano13w, Ano14s, Ano17u, Ano18u]. juries [McF12].

k/n [Cal11]. Kalman [DGR11, MRR15]. Keane [Blu10, Rus10]. Kernel [Kue12, SL13, Wol16, Can10b, KSY17, RL17]. Kernel-weighted [Kue12]. kernels [BNHLS11a, BNHLS11b, HSS17, SX16]. Keynesian [AB18, BI14]. Kingdom [ALX11]. kink [CF19]. Knowing [DLX19]. Knowledge [BE10]. known [Tri11]. KS [Arm14].

labor [AAMS10, DP17, Fla10, FM12, GJZ15, GK10, Hu17, LW10a]. Labour [BBB^{+}19]. LAD [YL17]. LAD-based [YL17]. lag [BM12, ChKS15]. Lagrange [BFK12]. lags [CK12]. Laplace [GG19, JPW15, TTG11]. LARCH [FZ10]. Large [Bog22, Can10a, CCM19, CCCM22, CEK16, DJfL15, HV19, KK13, Pel19, PT11b, AFTV19, BL17, BPS19, BKW17, BHP12, Cal11, CQYZ18, CDG14, CLL19, CHL18, DLX19, DGR11, FLS15, FHLV16, FVW16, GHK18, Geo10, GHS16, Han12, HS17a, HZ15a, Kim11, KKLW18, KS18b, KF19, KL18, MS18, MRT16, Mas17, Ona12, PC14, Sia17, SW14a, ZCLW19]. Large-dimensional [Pel19]. large-sample [MRT16]. Large-scale [HV19, Sia17]. Lasso [CK18, QS16, CF15a, CYZ15, CL15b, JfL18]. last [Fla10]. latent [APT18, BJR17, CFFL15, CH18a, CGY15, GHK18, LY12, SJ18]. latent-variable [BJR17]. Latest [PRSV13]. leaders [BE16]. leads [CK12]. leads-and-lags [CK12]. learn [MW17]. Learning [CM17b, Sha15, FL13, GD19]. Least [BFS19a, JL13, Mas17, WZZ10, BCCK15, BH13, CV17, DMP18, HY10, Hil16, HL17b, LL12a, Liu15, RW17, SWC12]. leave [GZWZ16, JWFS16, LZZZ19]. leave-subject-out [GZWZ16, LZZZ19]. leavers [DP17]. Lehmann [CO12]. length [DHY16]. lens [BH14b]. Let [Sun14]. level [BEK17, CO15, HLT10, HL15, KS12f]. Leverage [AM15b, BR12, BNW19, LWY16, MBC15]. Lévy [CDH10, BL15]. Lévy-driven [BL15]. Life [Hau19, HP14a, Khw10]. life-cycle [HP14a, Khw10]. like [DT19]. Likelihood [CPT14, DPvD11, GP14a, JN10, KR10, Las10, Yu12c, AGS11, AG12a, AC13, BL16, BN12, BDR14, BKMZ16, BJN15, Can14, CF15a, CNT17, CFHZ19, CCC15, Cho13, sCJP14, sC16, Cos13, FS19, GS18, GS19, GR18b, HP13, HP15, HP16, HMR11, HZ15a, HLM13, KS19a, Kim14, KYS14, Kri10, KS12d, Kru08, Kru14, LS15a, LC16, LLL18, LPZ19, LT14b, LfL19, MP11, MJ11, Oga13, Ots10, OSW12, OXM15, Pet19, Phi91, PKA18, RR15, SJ10, SW14b, SC15b, TYZ18, Tsi12, WIW13, WW10, Woo14, Xiu10, XTZ10, XfL15a, XfL18, YXZ14, Zha16a]. Likelihood-based [DPvD11, GP14a, KR10, sCJP14]. Limit [KP13, CP19]. limited [Mat12, Sto12]. limiting [HK14b]. limits [CGH^{+}19]. LIML [AK12, AKM10, And10, CT15a]. Linear [NOR13, PM11, ZZL18, AC15, BBCL18, BBP15, BCF17, BRG14, Cal11, DDMP15, DMP18, Ell11, EW15, EJW14, EG14, Fan12, FV15, GJZ15, GK15, GS13, HHB12, HSVV13, HP18, JJFG10, JJ14, KS13, Kue12, LZL^{+}16, Lee14b, MSW17, Ric19, SW16, San11, ST12, SJ10, Vog12, Zhu18]. linearity [BCP15, CH10, Kue19]. link [LPLZ18, ZLX15]. links [CDF12]. Linux [HR12b]. liquidity [Cha17, CM17a, DLM17, HMPV11]. List [Ano12p, Ano12o, Ano13o, Ano14p]. LM [Yan15b]. loading [KL18, YT15]. loadings [DLX19, MHU19]. loan [Kau15]. loans [LMT12]. Local [Dua16, FNN12, GO12, LL16, SWC12, TH13, ASP16, CZ15b, Des11, HR15a, HLT12c, HN12, HP14b, HV17c, Pet19, PMG10, Shi12, WL15]. Local-momentum [Dua16]. local-to-moderate [PMG10]. locally [KL12, PHY18, RL17]. location [FSV14, HL17a, Ibr14]. location-scale [FSV14, HL17a]. Lock [HR12b]. Lock-in [HR12b]. log [FWZ13, LT14a]. log-fractional [LT14a]. logarithmic [Fat10]. logit [BN12, FiKRB12, Ket19, MSW18, Kha13]. Long [GZ16, HV17a, AS14a, AM15a, ACM17, AdCGIV11, BGMT15, CM17b, DT10, ERS17, GK18b, HR16, Ho15, HP14b, Hua14, JW10, JJFG10, LR19, MP13b, MP14, PL13, Rob14, SLB18, WBH13]. long-horizon [PL13]. long-memory [Ho15]. long-run [AdCGIV11, GK18b, Hua14, JW10]. long-term [HR16]. Longevity [HP14a]. looking [CC15a]. Lorenz [Kra13, VA89]. loss [BS19, LRV13]. Low [MW13, CH10, CV17, HP14b, KW16]. low-dimension [CH10]. Low-frequency [MW13, CV17, KW16]. LS [AG12b, FZ10, Hay10]. Lunch [GKP12b]. Lyapunov [PW12].

M [CL14]. M4 [ZZ16]. macro [CPvdW16, TU10]. macroeconomic [Bha14b, BC10, BH14b, KS14, KLS11, Pos16]. Macroeconomics [Ghy16, WZ12]. major [AHK12]. Mallows [WZZ10]. Management [CMY16, CJMMPA15, LZ19]. manipulation [BDV17]. Many [Can14, LMZ17, Ana12, AKM10, AKM11, AS17b, BCCFV19, CY12, Car12, CT15a, CHN^{+}14, CL15b, Fre15, GR15, HK14a, Hod11, HKY19, Kas19, KP15, Kol18, KLS11, Men14, Oku11, PG17, Ric19, WK16]. March [Ano10o, Ano12r, Ano13q, Ano14r, Ano15p, Ano16q, Ano17o, Ano18q, Ano19m]. Marginal [BDR14, BS19, FAM13, KSV15a, QW12, SWC12]. marginalized [FL13]. Mark [CWZ19]. Market [ASAM15, BR11, HMPV11, ASX19, ABM11, BV11, BH14a, BE16, Bon14, CGH^{+}19, CWZ19, CP19, CDF12, Fla10, FM12, GGL12, GO11, HLX19, HS18a, KS12f, LW10a, LM12, LZL18, Myk19, PP15]. Market-based [ASAM15]. markets [BCN11, CNT15, FFG10, Fre15, GL16, O’D16, SX16, TZ11]. Markov [BDR14, Bia16, BBCG19, CDS13, HL14b, LL16, MR11, PdSSGK12, SL12, Zha11b]. Markov-switching [BDR14, Bia16, CDS13]. Markowitz [AT17]. martingale [Son10, Son11, ZXC15]. matched [HP18]. Matching [CF15b, HLR11, AAMS10, GQIK17, HINR12, HINR14]. Maternal [LMvdK10]. maternity [JWFS16]. matrices [ADZ14, BL17, BPS19, CLL19, CP18a, DLX19, HC16, JM16a, QfLY17]. matrix [CQYZ18, CKP10, DJfL15, FK19, Hou17, HS18a, Ike16, JfL18, KS11, KKLW18, KF19, KL18, LF18, LKH14, LT14b, PHL16, QfL15, WLC19, ZY18]. matrix-valued [WLC19]. matter [Jus11]. maxima [ZZC18]. maximal [BC14b]. maximisation [MP11]. maximization [LW10b]. maximizers [LM11]. maximizing [BD12]. Maximum [AC13, KYS14, Kru08, Kru14, RS17, Tsi12, XfL15a, BL16, BN12, BKMZ16, CNT17, Che10a, sC16, Egg16, FS19, GR18b, HP13, HMR11, HLM13, KS19a, Kri10, KS12d, LS15a, LLL18, LPZ19, LT14b, MRT16, PSS18, Phi91, RR15, SJ10, WIW13, Woo14, Xiu10, XTZ10, XfL18]. maximum-likelihood [Phi91]. may [Ket18, Ano13s, Ano14q, Ano15o, Ano16s, Ano17q, Ano18p]. MCMC [CR10, LYZ18]. Mean [GP12, AS10, AHH12, CDI12, FLM19, JWY18, MP13b, Pel14, PS12, RS17, XX19, Yu12a, CM16]. mean-nonstationary [AHH12]. mean-variance [AS10, PS12]. means [Bai10, FP14]. measure [AGHP16, HLR11, MO12, RS17, TH13]. measured [iKPS16]. Measurement [FGS17, HW17, JvN11, AH12b, An17, ADN^{+}16, BDZ11, BDM16, BC14a, BKLS16, Che17, DL17, DL16, FS16, GO11, HR17, HW10, HS15b, Hu17, KSV15b, LMZ17, MSW17, NNN11, PHL16, Son15, VW14b]. measurement-error [ADN^{+}16]. measures [ABH11, And16, AHK12, BBGV14, BDS14, DGM13, DT10, Lea16, LPS15, MM11, RW11, TBE14]. Measuring [DY14, FH15c, HLOW16, ID16, SHL11, WKM15, AFPZ16, Bia16, FFG10]. mechanisms [CF19]. median [Bas19]. Medicare [Khw10]. Medigap [KS16a]. Medium [CV17]. membership [AFPZ16]. Memorium [Ano11m]. memory [AM15a, ACM17, CM17b, GZ16, HV17a, Ho15, HP14b, LR19, MP13b, MP14, MP10, PS13, Rob14, SLB18, WBH13]. men [BSW10, Gil10, HLR11]. merged [KW16]. Method [BvdW11, CKT17, WH11, AT11, ABH14, BS12, CZ11a, D’H10, DV13, GMO18, Han12, HL18, HLM13, HS18b, Lee14a, LLL16, McC12, Qia16, Tri11, Yan15a]. methodology [Pos16]. Methods [Bia16, BdCK^{+}14, CL10, FAM13, GKP12a, HKRV18, KMN11, AG10b, BDK14, BCIT14, BKY10, BN19b, CHL14, CR10, HLT10, HMR11, KS14, Kru08, Kru14, MV13, PdSSGK12, San11]. Metropolis [WWZ16]. Michael [Rus10, Blu10]. Micro [TU10]. microstructure [ASMZ11, ASX19, ABM11, BR11, CM17a, CP19, GS11a, GO11, HS18a, Ike16, LM12, LZL18, Zha11a]. MIDAS [And16, PTV16]. migration [HLR11, KW10b, LMvdK10]. Mildly [Mag12, CPY17, PL16]. mind [Slo17]. minimal [Lea16, MW19b]. Minimax [Sto12, Tet12]. Minimum [EJW14, Kol18, Tor17, AR12, DHY16, FM10, GKN17, HLT13, HMT12, LP13]. minimum-variance [FM10]. minute [LPS15]. mis [DGJ19, KS15, NR10]. mis-classified [DGJ19]. mis-specification [KS15]. Misclassification [MM17b, AHS10, BvH17]. mismeasured [BMDL17, GM17]. Misreported [BDS14]. misreporting [NDT19]. miss [XX19]. Missing [Abr19, MM17a, CFR18, DMP11, LZW^{+}18]. missing-indicator [DMP11]. Misspecification [CHL14, GJ18, Fan12, GKS15, KP12, KS12c, Kom13, Kri11, Lee12, Lee14a, Lee16b, PG16, WZ12]. misspecification-robust [Lee14a, Lee16b]. misspecified [AG12a, CLS14, CS11b, GH19, HMT12, MO12, Ona15, Rob14]. Mixed [LTTC17, ADFK16, AGK10, BKMZ16, Cha16, Che15b, CPvdW16, GM16, Ghy16, GHM16, GHS16, MS16, OP16, PL16, Qia16, RvdHW14, Zad16]. mixed-frequency [GHS16, Zad16]. Mixed-scale [LTTC17]. Mixture [CHM11, DLM17, BHK^{+}13, BHH12, CPT14, CG10, DF15, FP11a, JM14, KMS16]. Mixtures [LS12, sHDW17, VKN12]. ML [Kru13]. MLE [FZ10, HP12a, HM18]. MM [Hay10]. mode [sHDW17, KS19c, KS12a]. Model [AD13, AdCGIV11, CDH12, CH14a, CM17a, CK12, DDMP15, GHX14, GZWZ16, Geo10, Kim12, KM16, KT11, Lav14, LP15, LZZZ19, Shi15, Siz11, Spa10, WWZ18, ZWZ13, ZL15b, ADG11, AAMS10, ASX17, AG12b, ALX11, BJS10, BFK12, BKW17, BN12, BEE15, BK10, BFS19b, BE10, BvH17, BLV15, BJN15, BHH12, CHHW11, CCF18, CRdM17, CCP17, CFPY10, CS10, Che10a, CZ11b, CZ12, CDS13, CN14, CSW14, CFL16, CH15, CIY15, sCJP14, CDI12, CM15b, CEG11, CLR10, Con10, DM13, Dav17, DHY16, Des11, EW15, EJW14, Fav13, FGPZ17, FJ19, FM12, FL11, FiKRB12, FSW10, FH19, GM17, GN13, Gay13, GJ11, GKS16, GGL12, GHPP18, GK18b, GL14, HL10, HR12a, HL14a, HW14b, HV19, HSU10, HR16, HS12a, HS15a, HHM17, HPY10, HP12b, HN12, HT16, HT19, HKY19, JCLL10]. model [JM14, JfL18, JN10, Joh10, JJFG10, JL13, JN18, KMS14, Ket19, Khw10, Kim14, KLGS11, KLGS12, Kru08, Kru13, Kru14, KPT13, KTT19, LW10a, Lee14b, Lee15, LS12, LT11, LLL15a, LLL15b, LO15, LLL16, LZZL18, LPZ19, LP17, LR10a, LfL19, LS15c, MPP10, MKS16, MZ18, MSX13, MJ11, MRSE11, OÁ19, PP15, PHY18, PK17, Por19, QfL15, RvdHW14, Rob10, RS11, WZ12, WZZ10, WH10, WL12, XfL15a, XfL15b, XfL18, YY19, Yu12b, ZA10, ZY15, Zha15, Zha16b, ZY18, Zha11b, ZZC18, vH11]. Model-based [Geo10, ZL15b]. Model-free [CM17a, FH19]. Modeling [AHK12, ACG17, BBCG19, BPQ18, Che15a, ELN11, GKM14, JvN11, JM16b, KLS11, LT11, ZZC18, ABH11, ChKS15, DG11, DDSW18, DGMZ19, Dua16, HS14, HM14, JM10, JM13, JJ10, JM16a, MR11, NP12, Pel19, PTV16, Pos16, TvD13]. Modelling [AT13, Fav13, FFG10, Gew12, JW10, MPRR17]. Models [GKP12a, Abr19, AM10, ALS13, AG10a, AC12, ASAM15, ASP16, AG12a, AFS13, APS16, APS17, AH12b, Ana12, AFTV19, AC15, AGK10, AW15, AL14, ALGQ16, AMM11, ABH14, Arm15, AM15a, ABHP16, AdCGIV11, BW14, BL16, BN19a, BBCL18, BH17, BBP15, BBG^{+}19, BPMSW13, BDR14, BI18, BMDL17, BH16a, Bia16, BBCG19, BCR19, BKLS16, BGK19, BH13, BKY10, BJR17, BH14b, BvdW11, BE11, BHK^{+}13, BC14b, BDKS18, BCS15, BA11, CRY16, CX12, CW14a, CW14b, CJY15, CLO18, CH18a, Can10a, CKM16, CFH10, CNT15, CNT17, CFHZ19, Cha13a, Cha16, Cha13b, CYZ15, Che10b, CZ10, CH11, CGL12, CDG14, CFW14, CPT14, CLS14, CH14b, CZ15a, CKT16, CT17, CLY17, CZJ18, CH18b, Che18, CW18, Che19, CLSY19, CP12, Che15b, CIY15, CFVN19, CS12, CKK15, CR10, CW10, sC16]. models [CKP10, CDR11, CPvdW16, CP11, CP15, Cíz12, CL18b, CM16, CHM11, CS11b, CS14c, CW15b, CT15b, DF15, DMZ18, DDMP15, DFL18, DMP18, DL16, DMR16, DV10, DK18, DSS17, DH14, DGP15, DL18, DPY16, DGR11, DEvG17, DPSS11, EMvS11, Ell11, EJC12, EG14, FGL11, FL16, FXY17, FWZ19, Fan12, FPSY19, FMV15, FVW16, FVGQRR15, FF14, FGS18, FGS17, FSV14, FV15, FHLZ15, FHLZ17, FiKY16, FZ10, FLZ11, FWZ13, FZ15, FJGM17, FZ18, Fre15, FSW10, FL19b, GG17, GG19, GGR17, GHK18, Gal17, GTY13, GLL15, GKS15, GM17, GHLT18, Gha17, GK10, GKY14, GZ16, GP14b, GK15, GMP17, GO12, GKR13, GKN17, GPY10, GMR17, GHM14, GLH15, GH16, GQIK17, GLL19, GR15, GR18b, HNS14, HR17, HINR12, HHB12, HPK14, HINR14, HSVV13]. models [HvdAW16, HL17a, HOY17, HW12a, HW14a, Han15, HV17b, Han18, Han16, HLT16, Hay10, HP15, HU10, HLM^{+}12, HP16, HM18, HP18, HMT12, Ho15, HW10, HW12b, HLW15, Hon17, Hou17, HZ15a, HZ15b, HZ18, Hsi18, HS12b, HS15b, Hu17, HW17, HLP12, Ibr14, IR11, IK13, IK19, JWFS16, JW10, JK15, JfL19, Joc15, JSE14, JPW10, KW14, Kai16, KG14, KG18, KS19a, Kas10, KGMR12, Kau15, KS19c, KP10a, Kha13, KPT16, KV17, KS12b, KS13, KW16, KF19, KV10a, KSV15a, KYS14, Koc16, Kom13, KV10b, KL12, KS15, KP10b, Kri10, Kri11, KS12d, KT14, Kue12, KP13, KT16, KT11, KKL14, LZL^{+}16, LRV13, fLY10, Lee12, jL12, LSL14, fLY14, fLY15, LMZ17, Leu15, LLS15, LY16, Li10, LL10b, LY12, LL12a, LLZ15, LL15a, LS15b, LTT17, Li17, LZK18, LLL18, LL12b, LS19a]. models [LS19b, LR10b, LfL10a, LLS14, LPLZ18, LfLB10, LfL10b, LCY10, LXC16, LS16, MS18, MS17, MR11, MSW19, MXY12b, MO12, Mas17, Mat12, Mat16, Mat19, MM16, MSW17, MT19, MK14, MM17b, MHU19, MSW18, MW16, NR10, NSS14, Ona12, Ona15, PZC19, Ped17, PMM15, Pet19, PLG17, Poi17, QW12, QS16, QfLY17, RV14, Ric19, Rob14, RV15, Sai13, SW16, SW10, ST12, Shi15, SfL17, SH13, STZ19, SC15a, Son11, Son14, Son15, SJ10, SJ12, Su12, SL13, SY15, SJZ15, SH16, SW17, SJ18, SUZ19, Sun16, SM18, SHHW18, Swe11, TYZ18, Tod11, TTG11, Ton15, Tor17, Tsi12, Vog12, WWZ16, WX15, WH10, WH11, WIW13, WK16, WY16, WLC19, Wes12, Woo14, Woo19, WL14, Xia12, YfL17b, YL17, Yan18, Yu12a, Zad16, ZWZ13, ZLX15, Zha15, ZOZS16, ZY18, Zha10, Zha11b, ZXC15, ZLCL19]. models [ZL15a, Zhu18, Zu15, dCGKL19]. moderate [PMG10]. modification [HLMN11]. Modified [HLT06, HLT12a, HP14b, VW14a]. Moment [AMM11, HA13, KMN11, WL14, AC12, AS14b, AS17b, AL14, ALGQ16, Arm14, Arm15, AC16, Arm18, BvH17, BCS15, CO12, Can14, CCC15, DGG14, DiT16, FGL11, GGR17, GT18, GM17, GO12, HNS14, Hon17, HK11, MO12, Men14, Oku11, OSW12, PS17, PTV16, SG19, Shi15, Shi16, Son10, Son15, TYZ18]. Moment-based [AMM11, WL14, BvH17, SG19]. moments [And10, BS12, BvdW11, Can10b, CL15b, GLPR11, Han12, HLMN11, HK11, HS18b, Lee14a, MS19, Tra16, Tri11, WH11, Xie88, Xie11]. momentum [Dua16]. Monetary [BC13, MS16, BC11, CF12]. Monitoring [HL19a]. monotone [Gew12]. monotonic [DGG14]. monotonicity [BB14, CW18, DE12, Gut16, HSWY16, Kle10, LTU14, Seo18, SWD11]. Monte [LS15b, NR10, PdSSGK12, ZA10]. moral [KS16a]. Moran [JfL15]. mortality [ACG17, GGHM15]. Mortgage [BCG19]. move [FL13]. movements [AGHP16, GT18]. Moving [Cha13b, DK18, Pos16]. MSE [Arb13]. MSE-optimal [Arb13]. Multi [GS15, Cat10, CGP16, PPT11]. multi-country [CGP16]. multi-period [PPT11]. Multi-scale [GS15]. multi-valued [Cat10]. multicountry [BdCK^{+}14]. multidimensional [MR16]. multifactor [AM15b, KPY11, PSY13]. multifractal [CDS13]. multifrequency [CFFL15]. multinomial [CFVN19, LR10a]. Multiperiod [DSW12]. multiple [ALS13, AR12, BPS19, BCF18, BDS14, BBKU19, BH16b, CDH12, CS14a, CLLL16, CLL19, CCeC18, CR16, CS11b, FJ19, FH15c, GJ11, GK18a, HHB12, HLT10, HLT13, HKY19, Ike16, KT11, LL12a, LLZ15, LPS15, MBC15, MP12a, MP12b, OP18, PWM11, RvdHW14, SW16, Xia18]. multiple-change-point [FJ19]. multiple-regime [LL12a]. multiple-structural-break [GJ11]. multiple-threshold [LLZ15]. Multiplicative [Joc15, CRdM17, GKS15, HL10, PK17]. Multiplicative-error [Joc15]. Multiplier [BFK12]. Multiscale [AC16]. multivalued [Lee18]. Multivariate [BNHLS11a, CKM16, CR16, DPSS11, NSS14, AFS13, ADFK16, BV11, BMN19, CFPY10, CH11, CH14b, Cho13, DHN16, DGTY17, DGM13, Egg16, Fas13, FMS13, FJGM17, FZ18, GGL12, GKM14, HL10, HR10, HLM13, JM13, JR15, JM16b, KT11, LRV13, LP17, Oga13, PU11, SX17b, SLB18, WPY11, WKM15, Wu10, Xu12, YfL17b, YCW19]. Mutual [ASLP14, BCIT14]. mutually [DEMY18].

N [HZ15a, BHP12, Cal11]. N2SLS [JfL18]. NASDAQ [BNW19]. National [AV11, GKP12b]. Near [Can14, GP12, Han10, Jen12, JP12, Ket18, Wes12]. near-epoch [Jen12, JP12]. nearest [ADS12, LLL16]. Nearly [MW19b, CY12]. negative [DHN16, GKS15, MP13b, Sho16]. negativity [Con10]. Neglected [HNS14]. neighbor [ADS12, LLL16]. neither [CCH13, DLX19]. Nelson [ACT13, CN14, CDR11]. Nelson-Plosser [ACT13]. neonatal [ACG17]. Nested [CM15b, DR15, GMP17, GHM14, Hag12, OSW12]. net [FH15c]. Network [ZWWH19, DY14, HLP16, Leu15, LP18b, HWY16]. network-formation [Leu15]. networks [ASK16, CHO19, Miy16, ZCLW19]. neutral [GLPR11]. news [LT19]. Neyman [MP13a]. No [BC10, CG11, FH15a, Las10]. No-arbitrage [BC10, CG11, FH15a]. nodes [ZCLW19]. noise [ASMZ11, ASX19, ABM11, BR11, BNHLS11a, BW15, CM17a, CP19, GS11a, GO11, GGL13, HS18a, LM12, LT14a, LZL18, OÁ19, Sha11, Zha11a, Zha16b]. noisy [ASJL12, CKP10, DLX19, DFH^{+}14, Hou17]. nominal [BCLL16]. Non [Con10, DR15, FF14, SUZ19, An17, AC15, BCC^{+}15, BNHLS11a, BEE15, BRG14, Bog22, BJR17, CCM19, CCCM22, CH10, CCJ12, CRT10, CM17a, CKP10, EMvS11, EW15, EJCL14, FK19, FS19, FLZ11, FOZ19, FSW10, GKS15, GO11, Hag12, Hou17, HMS13, HS15b, JKM17, KPY11, LMS17, Mac10, MM16, OSW12, PP15, PVWW15, Xie88, Xie11, ZWXW10]. non-central [Xie88, Xie11]. non-classical [HS15b]. non-conjugate [Bog22, CCM19, CCCM22]. non-correlation [JKM17]. non-equilibrium [An17]. non-exchangeable [BJR17]. non-Gaussian [AC15, BCC^{+}15, BEE15, CCJ12, FS19, FSW10, LMS17, MM16, PP15]. non-linear [AC15, BRG14, EW15]. non-linearity [CH10]. non-negative [GKS15]. Non-negativity [Con10]. Non-nested [DR15, Hag12, OSW12]. non-parametric [Mac10, ZWXW10, ZLCL19]. Non-separable [SUZ19, HMS13]. non-smooth [FOZ19]. non-stationary [CRT10, CM17a, EMvS11, KPY11, PVWW15]. non-synchronized [FK19]. non-synchronous [BNHLS11a, CKP10, GO11, Hou17]. Noncausal [GJ17, GJ18]. nonclassical [HR17, Son15]. nonconvex [GJZ15]. Nonlinear [CW15a, CL10, AFTV19, BCRvD13, BS19, BH13, CDS14, Che15b, DGTY17, DFH^{+}14, FVW16, GG19, GN13, Gay13, HR17, HPK14, IR11, KMS14, KK12, KP10b, KR10, Lee14b, MXY12b, NHKJ11, NOR13, Pak19, SH13, SC15a, WWYL10, WH11, WWZ18, Woo14, XfL15b, Yan15a, Zha10]. Nonlinearity [CHC10, MP10, CGY15]. nonnegative [Miy16]. nonnested [HP12b]. Nonparametric [ALW12, AS14b, BB14, BJR17, BS18, BMS18, BJW15, CFFS16, CL10, CDH10, CZJ18, CH18b, CLLL19, CFVH^{+}15, CKK15, CCeC18, DL16, DS18, DL11, DHKS18, FP14, FS12, GJZ15, Gew12, Gup18b, HY10, HW12b, HL17c, HS12b, Jen12, KAP15b, LTU14, LL10b, LCY10, Nie10, PVWW15, PHY18, QY15, RL17, Rob12, Rot10, SWD11, SL13, TBE14, Zha11b, Zin18, Zu15, ASP16, BS12, BCR19, Bre15, CZ10, CLS14, CPP19, CHP18, DG11, DFS12, DL18, DFH^{+}14, EJC12, FL16, FH15a, FJ19, FSV14, FiKY16, FH15b, GS12, GK17, GLH15, GKP12b, HLM^{+}12, HLPY15, HS13b, Hor12, Hor14, HS15b, JCLL10, JPW10, KG18, KP12, KS19b, Kri17, Kri11, KS12d, LF18, jL12, LR15, LL15a, LL15b, LLS14, LCLS15, LWY16, LX19, MXY12b, MS12, MSX13, Mat12, NHKJ11, Nor15, Rob11, Sch11]. nonparametric [ST12, SVV16, Sun16, Xu18, ZLCL19]. Nonparametric/semiparametric [LL10b]. nonparametrically [HL12]. nonresponse [Man16]. Nonseparable [CFVN19, CRY16, CL18b, Gha17, HW10, HW12b, fLY15, Mat16, Mat19, SWC12, Tor17]. nonstandard [McC17, PP12]. Nonstationarity [CKP16, GP13, Kim12, MP10, MP12b]. nonstationary [ASP12, ASP16, AHH12, CW14a, CW14b, CW15a, DL18, JN10, Kas10, KK12, MXY12a, TY17, Xia12, Xu12]. norm [FGZ19]. normal [Egg16]. normality [CF13, CC15b]. normals [FP11a]. note [Ano10u, XTZ10]. notion [SG19]. novel [Cha12]. November [Ano10n, Ano12s, Ano13p, Ano15n, Ano16n, Ano17p, Ano18r]. nowcasting [BCLL16]. nuclear [FGZ19]. nuisance [CZ11a, Son14]. null [CAS15]. number [AHS10, GHK18, GHM14, Hag19, KLZ19, LLS17, MS18]. numerical [HMN15, HL18]. nutshell [McC12].

Obituary [Aig16]. objective [GLPR11]. Observable [Mat16]. observables [JM19]. observational [BCDN12]. observations [Far15, FL19b, JLZ19, Pel19, Wol16]. observed [Cha17, FF11, GS19, KYS14, PHL16]. occupational [CGE10]. October [Ano11r, Ano13y, Ano14t, Ano19r]. off [BGMT15]. official [Man16]. oil [MS16]. older [BSW10]. oligopoly [AH12a]. OLS [MdV13, VW14a]. omissions [vEGM14]. One [ACR19, FL11, FHLZ15, HSVV13, HS18b, MRSE11, Wes12]. One-sided [FL11, FHLZ15]. One-Step [ACR19, HSVV13, HS18b]. one-way [MRSE11]. ongoing [GKP13]. only [CLO18]. operating [HR12b]. operator [CS13, DFH^{+}14, Son11]. operator-based [CS13]. operators [HSS17]. opinion [Jud13]. Optimal [AT16, CCJ12, CF13, CC15b, GA11, LCLS15, MO12, MM19, PPP13, Phi14, PP12, Arb13, APT18, BHP12, CIY15, GGL13, HvdAW16, HW14c, RV14, Sun11]. optimality [AKM10, Can10a, CO12, FR18]. optimization [GD19, GK10, PKA18]. Optimum [FLM19]. Option [BCC^{+}15, ABTT15, AM15b, BGZ11, FH15a, GR13, JK15, KLMP15, MBC15, Sha15, SX16, Xiu14, CFFL15]. option-implied [BGZ11]. options [AGJ11, CX14b, Wu11]. Oracle [KC15, Koc16, CGL15]. order [BNW19, CP19, DG17, DH18, GK17, GJ18, GR15, HR15a, HL15, HM18, HMR11, ILT13, JR15, KS19c, KS12c, KS17, LOS18, MP17, PGM15, Yan15a]. order-independent [KS19c]. ordered [CS12, LY16, YY19]. orders [Nie10]. organization [Hu17]. Ornstein [Fas13]. other [HP14b, SWD11]. out-of-sample [IJR17]. outcome [BC14b]. outcomes [CFL16, EL13, FGZ17, FM12, JWFS16, KSV15a, Xu17, Xu18]. output [CGP16, FS10, FS14, FAM13, LYZ18]. outputs [APT18, FS14]. over-identification [SK12]. over-identifying [Han18, LKH14]. overidentification [DG17]. overidentifying [CHN^{+}14, Han15]. overlap [ALW12]. Overview [CMY16, BDM16, CM15a, KMN11, LMT15].

P [Blu10]. package [OP15]. Pagan [HOY17]. Pages [Ano10t, Ano10q, Ano10s, Ano10p, Ano10m, Ano10r, Ano10o, Ano10n, Ano11n, Ano11o, Ano11p, Ano11s, Ano11q, Ano11r, Ano12t, Ano12u, Ano12r, Ano12s, Ano12q, Ano13v, Ano13x, Ano13r, Ano13w, Ano13t, Ano13q, Ano13s, Ano13p, Ano13y, Ano13u, Ano14x, Ano14v, Ano14w, Ano14u, Ano14s, Ano14r, Ano14q, Ano14t, Ano15t, Ano15s, Ano15m, Ano15r, Ano15p, Ano15o, Ano15n, Ano15q, Ano16t, Ano16p, Ano16r, Ano16q, Ano16s, Ano16n, Ano16o, Ano17t, Ano17v, Ano17r, Ano17u, Ano17s, Ano17o, Ano17q, Ano17p, Ano17n, Ano18v, Ano18x, Ano18w, Ano18t, Ano18o, Ano18u, Ano18q, Ano18p, Ano18r, Ano18s, Ano19o, Ano19q, Ano19m, Ano19r, Ano19n, Ano19p]. Pairwise [AL12, Miy16]. Pairwise-difference [AL12]. Panel [ALS13, GV19, Hsi18, LR15, OY19, PSY13, AG10a, ALLM15, AK12, AT14, ALX11, Bai10, BL17, BJS10, BFK12, BBCL18, BN12, BBP15, BCIT14, BLV15, BS18, CCF18, CS11a, CMM18, Cha12, Che10b, CGL12, CH18b, CW18, CFVN19, CP15, CL18b, DGP15, FGPZ17, FPSY19, FVV11, FVW16, FF14, FJ19, GG19, Gal11, GK16, GLL15, GN13, Gay13, Gha17, GKS16, GK15, GPY10, GHPP18, GLH15, GMHS12, Gug10, HOY17, HP13, HMR13, HL19b, Hay10, HP15, HW12b, HT16, HZ15a, HZ15b, HZ18, JWFS16, KMS14, KRW17, KGMR12, KPT16, KS13, Koc16, KL18, Kru08, Kru13, Kru14, KP13, Lam10, fLY10, Lee12, fLY14, Lee14b, fLY15, LMZ17, LOS18, LL15a, LL15b, LB17, Li17, LLS14, LS16, MKS16, MXY12b, MSW15, MSW17, MK14, MP12b, PSU11, PU11, PHY18, QW12, QS16, QfLY17]. panel [RV15, RR15, RV18, Sas15, SX17a, SW10, SfL17, SH13, STZ19, SJ12, SL13, SY15, SJZ15, SJ18, SM18, Tra13, Vog12, WH10, WWYL10, WL12, Wes15a, WL15, WL14, YdJfL12, ZWZ19]. Panels [KPY11, Ahn15, AFTV19, BFK16, BBGV14, BHP12, CRT15, CN12, CFVH^{+}15, EV17, HS17a, Kim11, LO15, Pak19, PVWW15, PT11b, PC14, RS15, Ros12, SS16, TU10, Woo19, Yan18]. PANIC [Wes15b]. parameter [BFS19b, BKY10, EM14a, FZ15, GHLT18, HLPY15, Hor14, HK11, Ket18, KK13, Pet19, Sun11, Yu12a]. parameters [AFS13, BCRT16, CK18, CZ11a, CFW14, CL14, GR15, HZ18, Hua14, HLM13, IJR17, LP15, LL10b, LPZ19, MSW15, Son14, Zha15, ZLCL19]. parametric [ABH14, CFW14, CM16, EV10, FPSY19, FF14, FZ18, GR13, GJ17, HR19, HLR11, Han16, HW14b, KK12, KP10b, Lav14, LLS14, Mac10, Pel14, Zha10, ZWXW10, ZLCL19]. Paretian [DKK14]. parities [RH16]. Partial [BMM12, FGZ17, FL18, FR19, KK12, WIW13, DG11, FSW10, Hon17, LPZ19]. partially [BCS15, CX12, Can10a, KW14, Kai16, KYS14, LS19b, SJ10]. participation [AAMS10]. Particle [MP11, SK16, HS19a, PdSSGK12]. partitioning [CF13]. partly [ALLM15]. pass [KP15, KZ15]. past [MBC15]. Patent [BE16, BE10]. patents [BE10]. path [CSS18, TPY19]. pattern [CHP18]. patterns [ASK16, SJ18]. pay [Khw10]. payday [LMT12]. payments [Jus11]. payout [vEGM14]. Penalized [BD18, CPP19, BL16, Lam10]. pension [HP14a]. performance [BCIT14, DHS16, HLW13, RV14, RS11]. performances [LRV13]. period [PPT11]. periods [CLO18]. permutation [CR16]. Perrigne [MMS19]. Persistence [BM12, FL19a, HO12, HLT06, HLT12a, MP10, PL13]. Persistence-robust [BM12]. persistent [BS18, CRdM17, HP12a, Ike16, JJFG10, Lee16a]. perspective [ADN^{+}16, BFvD10a, KFSvD19]. perturbed [FNN12]. Peter [MXY12a, MXY12b]. phase [PK17]. Phillips [AB18, MXY12a, MXY12b, XTZ10]. piecewise [DHY16, GJZ15, JJFG10]. piecewise-linear [GJZ15]. Pitman [BCL14]. Placebo [Hag19]. players [GHK18]. Plosser [ACT13]. PMIs [CGP16]. point [BCF18, BDR14, FJ19, GS18, GKM14, HO12, HLT16, JWY18, SI19, Yu15]. points [PY15, SW14a, ZWXW10]. Pointwise [BCCK15]. Poisson [BHH12]. polarization [AFPZ16]. policies [Gil10]. policy [AH12a, AX18, BC13, CF12, HS14, Rot10, WP14]. polynomial [BMDL17, CZ15b, FNN12, GP14a, HR15a, Xiu14]. pools [DHS16, GA11]. Portal [ZCLW19]. Portfolio [PKA18, BBG^{+}19, CLLL16, HV19]. portfolios [FLS15, FHLV16, FM10, FZ18]. Portmanteau [ZC18, AH15, CH10, CZ15a]. Portmanteau-type [ZC18]. posedness [AH12b]. Positive [BLL^{+}17, BNHLS11a, CP18a]. positivity [FWZ13]. possible [HLT16, HLT13]. possibly [ASP16, AKM10, CLS14, CS11b, Far15, FH19, IR11]. post [ABHP16, CKP10, EM14a, KS19c, KS18b]. posterior [DT19, FS12, HOvD12]. potential [CFL16, FGZ17]. poverty [AFPZ16, BDH16]. power [BCP15, DEvG17, HV17c, KS12f, RH16, Seo18, WL15, Wes15b, WH19]. Powerful [Xu13, SK12]. practice [FPSY19]. Pre [CKP10, EM14a, MZ16, KSY17]. pre-asymptotic [KSY17]. Pre-averaging [CKP10, MZ16]. precision [CQYZ18, DEvG17]. Predictability [PT11c, PST22, BPTT19, BXZ15, CJP16, GHM14, HLOW16, LYCP19]. Predicting [EL13, LYCP19, TZ11]. Prediction [BL15, CS14b, CL18a, DHS16, DSW12, DS15, FJ19, GJ11, GA11, GLS11, HP12b, KF19]. Predictive [CS11b, FL19a, Lee16a, PL13, BCRvD13, BD15, CW14a, CW14b, CM12, DGMZ19, GHLT18, GH19, HOvD12, KAP15b, LP18a, MW19a, RS13, RS19]. predictors [KP15]. preference [BKM14, CDDH15, JJ11]. preferences [Gew12, Gil10, HR12b, JJ11, PLJ11]. Prejudice [Fla10]. premia [BGZ11, KS18b, MMFG12]. premium [BH14a, EW15, WW15]. preparation [LZ12]. preschool [HR16]. preschool-structural [HR16]. presence [ASX19, CLLL19, CV17, GKP13, GO11, HL11a, HLT16, HLT12b, HLT13, HP14b, Kim12, KLMP15, KSV15b, KS12f, LP15, LZL18, LX19, MW17, MKS16, Mik10, Oku11, Pak19, PPP13, PSY13, RS13, SW10, SGF11, Son15, Swe11, TvD13]. present [Wes12]. pretest [Gug10]. pretesting [LR10a]. Price [FS16, NNN11, AHS10, An17, AGKZ18, BS12, Bon14, CDS14, CNT15, Cha17, cCTZ14, CGH^{+}19, DZ15, Fat10, FH15a, FFG10, HS13b, HMS13, HLP12, IDF11, JKM17, KPT12, MMS19, MS12, MSX13, MZ16, Myk19, WW15, Zin18, dHvdG11]. Priced [EM14b]. Prices [ACR19, ABH11, APT18, BLO14, BNHLS11a, Bha14b, GR13, HPY10, Ike16, Lar11, LM12, Myk19, RvdHW14, Sha15, SX17b, ST15, Xiu14]. Pricing [GMR14a, GMR14b, GM15, SX16, AG12a, AM15b, BCC^{+}15, BH14b, CFFL15, CO15, FL19b, GKR13, GK18b, HMR13, JK15, KLMP15, KS18b, LP17, MBC15, Par16, ZL15b]. primal [FS10, FS14]. Principal [BN13, ASX17, Ona12, WU15]. principles [Kue19, LW10b]. Priors [JM19, Bog22, CCM19, CCCM22, FJ19, KP19, LS12]. private [WX14]. Probabilistic [MMFG12, SV12]. probability [Cho15]. probit [HV17b, LPZ19, LR10a, LR10b, WIW13, Kha13]. Probit/Logit [Kha13]. problem [ABHP16, Car12]. problems [CCGR15, Joh10, LZW^{+}18, McC17, PPT11]. procedure [AC15, KW14, PSS18, ZY15]. procedures [HS18b]. process [ASP12, CGL15, CS13, CG10, DDSW18, FF11, GJ17, HL11b, HK14b, JM14, Par19, SC15b, WBH13]. processes [AT11, AD13, BCL14, BCCFV19, BRG14, BS18, BL15, CF15b, CDH10, Cho15, Fas13, FMS13, FP11b, FNN12, GKS15, GS18, GJ18, GS11b, GKM14, HP14b, JP12, KT14, Kue19, LOS18, LL16, MPRR17, PMM15, PGM15, QY15, SL12, ZZ16]. procurement [Cam12]. procurements [Cam12]. product [BCL14]. production [APT18, BK10, Gil10, HLP16, KT16, iKPS16]. productivity [FS14, ID16, O’D16]. products [Fre15]. Profile [SJ10]. profiling [BCDN12]. profitability [Die14]. Program [GKP12b, Hil16, RH16]. programming [CDDH15, HR16, PM11]. programming-based [PM11]. programs [SGF11]. project [McF12]. projection [CZ11a]. projection-based [CZ11a]. Proofs [Han12]. propensity [BE16, HLW13, Lee18]. proper [O’D16]. Properties [HLMN11, ASP16, AK12, AKM11, DJfL15, DH18, Fat10, Han12, HM18, KS19a, KS12c, KS17, KKL14, LYCP19, MP12b, PdSSGK12]. Property [BLO14, CGL15, DZ15, JJ14]. prospect [AT17]. protocols [Bha13]. provision [McF12]. proxies [Pat11b]. Pseudo [BN12, BHP12, GR18b, HvdAW16, Kri10, FS19, HMR11, LPZ19]. Pseudo-Gaussian [BHP12, HvdAW16]. Pseudo-maximum [Kri10]. public [McF12]. Publisher [Ano10u]. purchase [DZ15]. purchasing [RH16]. pure [HV17c]. pure-jump [HV17c]. purpose [DiT16]. Pythagorean [CP18a].

QML [FLZ11, QfLY17, SX17b, SY15, YfL17b]. QMLE [FLZ11]. QR [Lee16a]. Quadratic [MRR15, CFGV13, Lar11, PHL16, PM17]. qualifications [BDS14]. Quantile [CLO18, Che19, CFVK15, CGHX19, ChKS15, FFM12, Gal11, MS12, ZWZ19, BCCFV19, BKM14, CX12, CCF18, CDES14, Che10a, CZ15b, Che18, CLLL19, CPP19, Che17, CHS19, COP10, DMZ18, DH14, EG14, FP12, FL16, FL18, FL19a, FGS17, GK16, GH19, GK17, GHPP18, GV19, HLOW16, HL14a, HL19b, Ho15, Kap15a, KGMR12, KV10b, Lam10, Lee16a, LO15, LL16, LCLS15, LS15c, Par19, QY15, RL17, Ros12, SH16, ZWWH19, dCGKL19]. Quantile-based [MS12, COP10]. Quantile-regression-based [ZWZ19]. Quantiles [MS19, CFVH^{+}15, DHN16, DV13, EV10, GK18a, OQ11, SC15a, WKM15]. quantilogram [HLOW16]. Quanto [KLMP15]. quarterly [Bha14b, AV11]. Quasi [BK10, CNT17, Kru13, LLL18, SC15b, Woo14, Xiu10, CFHZ19, FS12, HLM13, Kim14, LT14b, Pet19, SJ10]. quasi-Bayesian [FS12, Pet19]. Quasi-likelihood [SC15b, CFHZ19]. Quasi-maximum [CNT17, Woo14, Xiu10, HLM13, LT14b, SJ10]. Quasi-structural [BK10]. quoted [Lar11].

R [HL17a]. R&D [BE16]. R-estimation [HL17a]. racial [AAMS10, BCDN12, MW17]. Random [PW12, TPY19, BJS10, BHP12, BMM12, BN19b, CL18b, DL11, DHKS18, FiKRB12, FiKY16, GLL15, GHPP18, HSU10, HS15a, HHM17, HT16, HT19, JL13, Ket19, Koc16, LT11, LZL18, MK14, MSW18, RW11, RV14, WL12, Woo19, Xie88, Xie11, YY19]. randomisation [Sia17]. randomized [CR10, FP12, HL19b]. randomness [HT19]. range [CW11a]. range-based [CW11a]. Rank [BN19a, PS13, AS17a, AW15, BHP12, BJN15, CP12, Gew96, HvdAW11, HvdAW16, HLT16, KRW17, Kar17, KLZ19, PVWW15, Swe11, YY19]. rank-based [BHP12, HvdAW11, HvdAW16]. rank-ordered [YY19]. ranking [LRV13, Pat11a]. rate [AGJ11, Dua16, ERS17, GGL13, JJFG10, JPW15, MR11]. rate-adaptive [JPW15]. rate-optimal [GGL13]. rates [AR12, BC13, CG11, CF13, CC15b, ELN11, JPW15]. rating [VW14b, WW10]. Ratio [HO12, BJN15, LfL19, MRT16, WW15, WW10]. Ratio-based [HO12]. rational [Fan12, Hod11]. rationality [Kaw17]. Rationalization [LVX17]. ratios [BKY10]. Real [BCLL16, DSS17]. Real-time [BCLL16, DSS17]. Realised [COP10, BNHLS11a, BNHLS11b, Pat11a, SX17b]. reality [Ghy16]. Realized [ABM11, ACM17, MMS11, TZ11, TTG11, ABH11, AM15a, BV11, BGZ11, BPQ18, BMN19, BCN11, CTV19, DGM13, FMV15, GM11, HS18a, JM16a, LPS15, ZMAS11]. reasons [CX14b]. recovering [San11]. recovery [Jud13]. recurrent [KP17, LL16]. recursive [Wes15a]. reduced [ABH11, Gew96, Kar17, PG17, Siz11]. reduction [CP18b, KS14, Pak19]. reexamination [CJP16]. Referees [Ano12o, Ano14p]. Refinements [RR15, Lee14a, Lee16b]. reflections [Ton15]. reform [HP14a]. regarding [HHB12]. regardless [LYCP19]. Regime [PPW19, CCP17, CFW14, CFL16, GTY13, KS19a, LL12a]. regimes [BC13]. region [Hil16]. regions [CK18, CQYZ18]. Regression [AGK10, DMP11, DL17, GBH15a, KS12a, PY15, Xu17, ALLM15, Ana12, AMM11, BCF17, BvH17, Bre15, CX12, CW14a, CW14b, Cal11, CCJ12, CET14, CGY15, Che10a, CZ12, CZ15b, Che18, Che19, CH15, CIY15, CFVK15, CGHX19, CHS19, CG10, CCeC18, CK12, Cíz12, Cos13, DFH^{+}14, Ell11, EJC12, FL18, FGZ19, FGS17, FS12, FFM12, FH19, Gal11, GK16, GHLT18, Gew96, GH19, GP14b, GMHS12, GV19, Hag12, HR15a, HL14a, HLPY15, HKLP15, HP18, sHDW17, HS19b, Jen12, KG14, Kar17, KP12, KAP15b, KGMR12, KS19b, KP15, KP17, KS18b, KLGS12, Kri17, KKL14, Lam10, LZL^{+}16, Lea16, fLY15, LR15, Lee16a, LMZ17, Lee18, LO15, LL16, LTT17, LL12b, LLS14, LX19, MMV14, MM19, MW16, NHKJ11, NOR13, Nor15, OQ11, OXM15, PWM11, PSW10, PQ12, Par19, Pel14, PL13, PM11]. regression [Rob11, Rob12, RT12, ST12, SH16, WWZ18, WKM15, Xu18, Yu12c, Yu15, YP18, ZA10, Zha15, ZWZ19, ZWXW10, ZLCL19]. Regression-based [GBH15a]. regression-calibration [KS18b]. regression-discontinuity [PWM11]. regressions [And16, AHH12, BBKU19, BD15, CW15a, CPP19, Che17, DMZ18, DGTY17, DNR10, EGT13, FL19a, GS13, KRW17, KU14, KK12, LTU14, LTTC17, LS15c, Phi91, Sal12, SYN12, VW14a, WW15, XTZ10]. regressor [DGJ19]. regressors [Ana12, BCP15, BCCFV19, BMDL17, CW14a, CW14b, Che15a, CKT16, CPY17, Che19, CP15, DL18, HHB12, HV17b, Kom13, KTT19, LT15a, PHY18, PL16, Ric19, ST12, Zhu18]. regret [Sto12, Tet12]. Regular [ADFK16, GP14a]. regularisation [BPS19]. Regularization [DFS12, Car12, FGZ19, GS12, Hor14, MM16]. Regularized [CT15a, BN19a, FS12, HK14a, LF18, Zhu18]. regulation [LMT12]. Reinforced [PMM15]. related [Egg16, ZMAS11]. relations [BDH16]. relationship [GL14]. relationships [Tho14]. relevance [AGKZ18, CHHW11]. relevant [CL15b]. reliability [Spa10]. reliable [BPQ18]. repeated [BS12]. reported [AH12b]. Representation [GJ17, CF13, FL11]. representations [FHLZ15]. resample [ABH14, FL13]. resample-move [FL13]. reservation [Gut16]. Residential [BCG19]. Residual [AW15, CN12, Ric19, Ahn15, CGL15, GBH15b]. Residual-based [AW15, GBH15b]. residuals [EJCL14, RS15]. Resolution [AX18]. resources [BSW10]. respect [Sch11]. response [BCF17, CS11a, CKT16, DHKS18, GLL15, GQIK17, HINR12, HINR14, HS15a, IK13, IK19, Kha13, KV17, KT12, Li17, MJ11, WX15, ZLZ13]. responses [IK16]. restricted [BBB^{+}19, GM17]. restricting [Kaw17]. Restriction [KMN11, DZ15, MO12]. Restriction-Based [KMN11]. restrictions [AC12, AdCGIV11, CLO18, CG11, CCC15, CHN^{+}14, DHN16, ES10, FH15b, GO12, Han15, Han18, HL17c, HK11, KV10a, LKH14, OSW12, PS17, Poi17, Ros12, Son10, Son15]. Results [AAMS10, BCCK15, FL11, Gra12, Hay10, MT19]. Resurrecting [RW17]. retail [NNN11]. return [BXZ15, BGMT15, CJP16, CGE10, LF18, MM11, PS12, RW11]. returns [AGJ11, BDS14, BTL13, FP11a, LT11, LZK18, LYCP19, PP15, PPW19, PHL16, PT11c, PST22, Wu11, Zam10]. Revealed [CDDH15, BKM14, JJ11]. revenue [CFGV13]. reversion [Yu12a]. Review [Bas19]. revisions [Gal17, JvN11]. revisited [ACT13, BCF12]. rich [DK18]. Risk [Cam12, CMY16, FZ15, LT15b, SGF11, WZ16, AT14, BBCG19, BKLS16, BGZ11, BXZ15, Bon14, BGMT15, CJMMPA15, CWZ19, CHM11, EM14b, ERS17, EW15, FT14, FZ18, GGHM15, GLPR11, GH19, GK18b, HWY16, Jus11, JJ11, KS18b, LZ19, MMFG12, MW19b, PMM15, PLJ11, RY11, SX16, Wil11, ZZ11, Zin18, PZC19]. risk-averse [Zin18]. risk-bearing [AT14]. risk-neutral [GLPR11]. Risk-parameter [FZ15]. risk-return [BGMT15]. Risks [FLS15, FHLV16, HLX19, TB10, vEGM14]. Robert [Slo17]. Robertson [Ahn15]. Robust [AH15, BBCL18, CST12, Che15b, CHS19, FHLV16, FWZ19, Far15, GGL13, HMR13, HLT10, HP15, Lam10, LZW^{+}18, Mik10, MRSE11, PL16, Sun11, WY12, Xia12, ADS12, BM12, BKY10, Cíz12, HOY17, HP16, HOvD12, KU14, KS13, Lee14a, Lee16b, LP18b, MW13, MZ16, MZC19, PSU11, PL13, PP18, SV12, Sun14, Vog12]. Robustifying [Xu12]. Robustness [DG11, LW14a, Ibr14]. role [BXZ15, BCN11, Gil10, KRW17, RV14]. Roll [CLY17, CLSY19]. Rolling [IJR17]. Root [Che10b, DMR16, AG10a, Cha12, GTY13, HvdAW11, Han10, HLT12c, Kru08, Kru14, LP17, MP12b, PSU11, PSY13, PMG10, WWYL10, WL12, Wes15a, WL15, ZC18]. Root- [Che10b, DMR16, AG10a]. roots [CX14a, CET14, HLT12b, HLT13]. rotated [NSS14]. rotation [ABHP16]. rotten [BDV17]. rounding [LZL18]. Roy [DM13]. Rubin [WT18]. rules [DPvD11, LW10b]. run [AS14a, AdCGIV11, DT10, ERS17, GK18b, Hua14, JW10]. running [DL17]. runs [CW11b]. RV [LPS15].

S [MZC19]. S-TSRV [MZC19]. Saddlepoint [LR19]. sale [LZ12]. sales [BEK17, MRT16]. same [Por19]. Sample [Ho15, AKM11, BPS19, CZ10, CZJ18, CM12, DJfL15, GS13, Han12, HS13a, HZ18, IJR17, Joc15, KW10a, jL12, LLS17, MW17, MRT16, MV13, TPY19, vH11]. sampler [WWZ16]. samples [Tri11]. sampling [AGK10, DT19, GM16, HOvD12, Ike16, MZC19, PM17, SK16]. Sarafidis [Ahn15]. saving [PSW10]. scale [BPTT19, Des11, FSV14, GS15, HL17a, HV19, LTTC17, Sia17, ZCLW19]. scales [MZC19]. scaling [BCC^{+}15]. Scanner [IDF11, dHvdG11]. Scenario [ERS17]. School [GKP12b]. schooling [BDS14, DM13, LT11, Zam10]. Schools [BDV17]. Schröder’s [Bas19]. score [AH15, BDV17, BGK19, CNT15, Che10a, HLW13, Lee18, PSS18, VW14b]. score-driven [BGK19]. score-test [VW14b]. scoring [DPvD11, LW10b]. screening [ZCLW19]. sealed [HLP12]. sealed-bid [HLP12]. search [FM12, FSW10]. seasonal [CET14]. seasonally [Sai13]. second [DH18, JR15, KZ15, PTV16, Sch11]. second-order [DH18]. secret [BE10]. section [Arb13, CFVN19, DL11, EM14b, EV17, MK14, Pak19, SJ12]. sectional [BFK12, CGL12, DGP15, HV17a, HP15, HS17a, KMS14, Kim11, KP13, LR16, OÁ19, PSU11, Rob12, WWYL10, WU15]. sections [KS18b]. security [Ike16]. seemingly [HKLP15, ZA10]. Segmenting [AHH12]. Select [CL15b]. selecting [ZY15]. selection [ASP16, AdCGIV11, BMS18, CJY15, Can10b, CF15a, CDH12, CH14a, CF15b, CFPY10, CZ10, CZJ18, CP12, CK12, sCJP14, D’H10, DMZ18, DHY16, DiT16, HLPY15, HP12b, IJR17, JWFS16, Joc15, KS16a, Kim12, KSV15a, Koc16, KT11, jL12, LP15, LfL19, MW17, PPT11, Sal12, Sas15, SW10, Shi15, STZ19, Spa10, ZY15, ZY18, vH11]. selectivity [HLP16, MKS16]. Self [YL17, AH12b, BLY18, GKM14]. self-excitation [BLY18]. self-exciting [GKM14]. self-reported [AH12b]. Self-weighted [YL17]. Semi [BS12, FPSY19, FH15a, GR13, Kri11, Sch11, AL14, BNHLS11a, CLS14, FZ18, GJ17, HR19, HLR11, HW14b, Xu18, ZLCL19]. semi-definite [BNHLS11a]. Semi-nonparametric [BS12, FH15a, Kri11, Sch11, CLS14, Xu18]. Semi-parametric [FPSY19, GR13, FZ18, GJ17, HR19, HLR11, HW14b]. semi-strong [AL14]. semi/non [ZLCL19]. semi/non-parametric [ZLCL19]. semidefinite [BLL^{+}17]. semidefiniteness [DHN16]. semimartingale [GT18]. semimartingales [HV17c, JKLM12]. Semiparametric [AL10, CX12, CZ10, CZ11b, CZ12, CGL12, CLLL16, CLY17, CW18, CLSY19, Chi10, CDI12, Cíz12, DGP15, GP13, HvdAW16, HR10, HLP12, MS17, Son14, Son15, SL12, Su12, WX14, WZ16, YY19, Zha15, AC12, ABH14, ALX11, BLV15, CCF18, Cat10, CFPY10, CL15a, CLL19, Cos13, EJCL14, FL16, GN13, GL14, HL17a, Hon17, IL10, IL18, JM10, JM13, JM14, JM16a, KM16, KSV15a, KL12, Kri10, LTU14, Lee15, LL10b, LL15a, LLL15a, PSW10, PZC19, QW12, Rob10, Son10, SHL11, WX15, WH11, Yu12b, ZOZS16]. sensible [Poi12]. separability [CW18, CDDH15, LW14b]. separable [HMS13, JCLL10, fLY15, SUZ19]. Separating [Dez15]. September [Ano12q, Ano13u, Ano15q, Ano16o, Ano17n, Ano18s]. Sequential [AFS13, Che18, AC12, KP13, Mar14, Zam10]. Sequentially [GD19]. serial [BJS10, DMC12, GS15, GLH15]. serially [GHK18]. Series [ET14, LR16, LMT15, Wu10, AHH12, BCF18, BCCK15, BCCFV19, BKLS16, BGK19, BH16b, Bre15, CMM18, CX14a, CNT17, CW15a, CKP16, CDES14, CLS14, CH14b, CC15b, CIY15, DDSW18, DV10, EMvS11, GG19, GKS15, GO12, HPK14, HLOW16, HLT10, HA13, HKR14, HR15b, JR15, KSY17, KP10b, Kue12, LR19, Lee12, LLL15a, LXC16, Mac10, MXY12a, MW19a, MP13b, MP14, MM16, MP12b, NHKJ11, PU11, Por19, SG19, SC15a, Sun11, SHHW18, Ton15, Tra13, TY17, WLC19, Xia12, ZLZ13, Zha16b]. server [HR12b]. Set [LO15, Ros12, BMM12, BKY10, GMO18]. set-identified [GMO18]. sets [BDK14, FP14, GJZ15, HL15, IK16, Mik10, SW14a]. setting [HS17b]. settings [GBH15b]. shape [AFS13, FH15b, HL17c, SWD11]. shares [Tri11]. Sharp [Mou15]. Sharpe [MRT16]. shocks [Geo10, HL14b, MS16]. Short [DT10, AdCGIV11, BGMT15, CS11a, LR19, LO15, MRT16, MSW15, MP13b, PS13, Ros12, Yan18]. short-run [AdCGIV11]. shortfall [PZC19]. Should [HS18b, Poi12]. Shrinkage [LS16, QS16, BFS19b, Han16]. sided [AAMS10, FL11, FHLZ15, MM19]. sides [DL16]. Siegel [CN14, CDR11]. Sieve [CLS14, CL14, CL15a, SJ12, SH16, XfL18, CPP19, PGM15]. SIFIs [CHO19]. Sign [CZ15a, MSV19]. Sign-based [CZ15a]. signal [Kau15, Mac10]. signaled [BC11]. signals [WX14]. signature [TT10]. similarity [BA11, GLS11]. similarity-based [GLS11]. Simple [SK12, ALGQ16, BPQ16, CC15a, CT17, FiKY16, Hag12, HvdAW11, HU10, JL13, PM11, Xie88, Xie11]. Simulated [GS19, GKN17, LS15a, BS12, DV13, KS12d]. simulation [FN18, FL13, HOvD12, McC12, PdSSGK12]. simultaneity [Mat12]. Simultaneous [BCF18, Can10b, GM17, AL10, HZ15b, KV10a, MP17, YfL17b]. single [ADFK16, CL18b, DGP15, ES10, FPSY19, LZL^{+}16, Lee15, Son14, Zad16]. single- [Zad16]. single-index [CL18b, DGP15, ES10, FPSY19, Son14]. singular [ADFK16, CY12]. size [AG10b, Gug10, HZ18, Ket19, LLS17, McC17, PP18, SK12, WT18]. size-correction [AG10b, McC17, WT18]. skewness [EM14b]. Slutsky [DHN16]. small [BHP12, GHK18, GT18, GHM14, Hag19, Sun14]. small- [Sun14]. small-time [GT18]. Smile [MBC15, CX14b]. Smooth [HKLP15, LP13, AMM11, Che15a, CH18b, FLWZ15, FOZ19, HK14b, LS19b, MS17, MP13a, PLG17, VKN12]. Smoothed [GK16, dCGKL19, CS14a, Cos13]. Smoothing [PMG10, CZ15b, HLPY15, Kap15a, KSY17, LL10b, LL16, LCLS15, McC12, Sun11, Zha16a]. Social [YfL17a, LfL10b, Sia17, ZCLW19]. solution [MR16]. solutions [DFH^{+}14]. Some [AK12, BCCK15, CX14b, Gra10, GH16, Hay10, Joh10, Ton15, CPT14, CP19, Mat19, PdSSGK12]. sovereign [ASLP14]. space [FHLZ17, FSW10, GGR17, fLY15, SL12]. space-time [fLY15]. spaces [FHLZ15]. spacing [SX17a]. span [AFTV19]. spanned [CW15b]. Spanning [PS12]. Sparse [Kas19, Zhu18, BCR19, KS19c]. sparsity [KG14]. Spatial [BCG19, KS11, SfL17, ZY18, BKLS16, DJfL15, DR15, GKS16, GR15, Gup18a, GR18b, HM18, Jen12, JP12, JfL15, JfL18, JfL19, KP10a, fLY10, fLY14, Li17, LfL10a, LfLB10, LfL19, MS17, PT11b, QfL15, QfLY17, Rob10, Rob11, RT12, RR15, SJ10, Su12, SY15, Sun16, SM18, Vog12, WIW13, XfL15a, XfL15b, XfL18, Yan15b, YfL17b, Yan18, YdJfL12]. spatio [DPY16, GMWY19, HPY10]. spatio-temporal [DPY16, GMWY19, HPY10]. spatiotemporal [KS13]. special [CKT16, LZ19, ZZ11]. Specification [BCS15, DGTY17, EV10, EG14, FMV15, Hor12, JK15, KP10a, LLS15, LYZ18, NHKJ11, SJZ15, TTG14, ASP12, AW15, CS14b, DJfL15, DV10, FSW10, GHX14, Gup18b, Hag12, JfL18, Kim14, KS15, LO12, LLL16, NR10, RS19, Spa10, SL13, WH19, ZOZS16, Zu15]. specified [ALLM15, CH14a]. specifying [HW14b]. Spectral [MP14, CH11, FGS18, Gup18a, Sha11, ZL15a]. speculative [ABH11]. Speed [FP10]. Spillover [BKLS16, AH15]. spillovers [BEK17, BE10, BE16, GKS16]. spline [CG10]. splitting [MV13]. sporadic [AC13]. spot [CNT15, ZB14]. spread [CLY17, CLSY19]. spreads [Fav13, TZ11]. Spurious [SYN12, SLB18]. square [Xie88, Xie11]. squared [FLM19, GKR13, KZ15, LLY15, Ona15]. squares [BCCK15, BH13, CV17, DMP18, HY10, Hil16, HL17b, JL13, LL12a, Liu15, Mas17, RW17, SWC12, WZZ10]. SR [AC15]. stability [CS14c, HS13a]. Stable [BHK^{+}13, Bas19, BDK14, CH18a, DKK14, GRV11, HSVV13, HV17c, KT14, LT14a, Miy16, NOR13, Oga13]. stage [FLZ11, HR19, HL17b, HS18a, KW14]. standard [HP15]. start [GR18a]. start-up [GR18a]. state [BCC^{+}15, CKP16, Dez15, FSW10, GGR17, HS12b, Kau15, Myk19, PQ12, SL12]. statement [Ano16m]. states [CFFL15, GHK18, LPT10]. static [ABHP16, BN13, BW14, BBCL18, DHKS18]. Stationarity [ASP12, GP12, GLL19, HKR14, JR15, PS13, WW10]. Stationarity-based [ASP12]. stationary [AT11, CRT10, Cha13a, CM17a, DL18, EMvS11, GTY13, GBH15b, HS18a, KPY11, KL12, Kru08, Kru14, PVWW15, PHY18, Swe11]. statistic [Arm18, GK17]. Statistical [APT18, Fas13, GMR17, HT16, HZ15b, MP12a, RT12, SC15a, Tet12, HL14b, HZ18, KW16, Myk19, Spa10, YXZ14]. Statistics [CMY16, AG12b, Arm14, BC11, FN18, GRS12, HLT13, Man16, MMV14]. Statment [Ano15l]. Staying [MPRR17]. Step [ACR19, CL15a, DGR11, FVV11, FR17, GK18b, HSVV13, HS18b, Leu15]. stepping [Dez15]. Stick [GS11b]. Stick-breaking [GS11b]. stimulus [OP15]. Stochastic [FSW10, MZ18, RH16, ASAM15, APS16, APS17, AT17, AM15b, ACM17, BI18, BRG14, Bog22, CCM19, CCCM22, Cha13b, CGY15, CJMMPA15, CSW14, CTV19, CLR10, CW15b, CT15b, DP17, DE12, Des11, DSS17, FVGQRR15, GG19, GKY14, GKS16, GH16, Ho15, JM10, JM14, KPT13, LP17, LSW10, MT19, OÁ19, PS12, PKA18, Sch11, Seo18, Tod11, Tsi12, WH10, Wil11, Yu12b, Zu15]. Stochastically [Wil11]. Stock [BXZ15, BV11, BH14a, Bha14b, BTL13, BCN11, CJP16, GGL12, KT14, PHL16, PT11c, PST22]. stone [Dez15]. storage [CHHW11]. Strategic [HHS12]. strategies [BBG^{+}19, CJMMPA15, LS15b]. stratified [Cos13, Tri11]. strength [Che15b]. Striated [WWZ16]. Strict [GLL19]. strike [FH15a]. strong [AL14, HS17a, LWY16, Mag12]. stronger [Kau15]. Structural [Blu10, CKM16, GK10, GR18a, HL14b, HW10, Kea10, KS15, Miy16, PG17, Rus10, AM10, AK12, AKM11, BFK16, BKW17, BCLL16, BPMSW13, BBKU19, BK10, BCMH19, BE11, CFH10, CDG14, Che15a, CH18b, CS14c, FMV15, FiKY16, FH19, GJZ15, GJ11, GKP13, GMR17, HHB12, HPK14, Han18, HU10, HR16, HS13a, HL17b, IK13, IK16, IK19, JW10, JWY18, KT11, LMS17, Li10, LZ12, LW14b, MSW19, MS16, OQ11, PPP13, PT11c, PST22, PLG17, SWC12, Slo17, SI19, WBH13, Xu13, YT15]. Structural-break [KS15]. Structure [ACR19, ABTT15, ADFK16, AC15, And19, BC13, CG11, CN14, CDR11, CW15b, Dua16, ELN11, GOS19, GZ16, HW12a, HW14a, HC16, KS12b, MPRR17, OOS13, PSY13, Por19].

Structured [FK19, ZZ16]. structures [CRT15, KPY11, PU11]. Student [BDV17, ZG10]. Student- [ZG10]. Studentized [MP13b]. Studies [Slo17, BdCK^{+}14]. study [LS15b, LZZL18, NR10, OP15]. studying [AFPZ16]. sturdiness [Lea16]. subadditivity [DJS^{+}13]. subject [GZWZ16, LTU14, LZZZ19, SWD11, WBH13]. Subjective [GGHM15, DEvG17]. Subsampling [BNHLS11b, Kal11, AG10b, CST12, CPTV17]. subset [CL18a, EGT13, WT18]. subspace [BN19b]. substitutes [HHS12]. Subvector [Ket18]. success [PT11a]. sudden [AX18]. Sufficient [FXY17, CHP18]. Summability [BRG14]. sums [EJCL14]. super [HC16]. super-diagonal [HC16]. supply [BBB^{+}19, GJZ15]. support [DF15, RY11]. supported [HSS17]. surface [CFFL15, FH15a]. surplus [BM12]. surplus-lag [BM12]. Surprise [GMR14a, GMR14b]. surprising [Gra12]. survey [AM10, Man16]. survival [CFPY10]. SVARs [GMO18]. Swan [BSV12]. swings [JJFG10]. Switching [BBCG19, BCC^{+}15, BDR14, Bia16, CCP17, CDS13, CFW14, CFL16, HL14b, KS19a, Kau15, MW16, PPW19, Phi91, XTZ10]. symmetric [CP18a]. symmetrically [CT17]. symmetry [CZ10, CT17, CZJ18, DS18, FLWZ15]. synchronized [FK19]. synchronous [BNHLS11a, CKP10, GO11, Hou17]. synthesis [MW19a]. system [GP13, HL19a, Mou15, OP18]. systematic [BBGV14, TB10]. systemic [BBCG19, BKLS16]. systems [ADFK16, CPY17, FP10, GBH15a, HR12b, HR10, JPX11, MP17, Nie10, Phi14, SWC12, Shi12, TPY19].

T [BHP12, HZ15a, HS17b]. Tail [AGHP16, CHO19, HWY16, BB14, BJW15, DDSW18, HS13b, MZ18, WKM15]. Tail-Event [HWY16]. tail-index [HS13b]. tailed [CZ15a, HA13, HP16, MP13a, PRSV13, PPW19, YL17]. tailedness [Ibr14]. Tailored [CR10]. tails [BTL13, BT14, DJS^{+}13, DPvD11, DKK14, KLMP15, MdV13, MP10]. Takeshi [CHH18]. Taking [Cha12]. tale [BLO14, ELN11, SX16]. targeting [FR17]. taste [BCDN12]. taste-based [BCDN12]. tax [BS19, DZ15]. taxation [BBB^{+}19]. taxes [BLO14]. Taylor [YCW19]. technical [FS10, JJ10, SYN12]. techniques [MPP10]. technologies [O’D16]. technology [KT16]. teenage [AAMS10]. Tempered [HS19a, KT14]. temporal [CHC10, DPY16, GMWY19, Has11, HPY10]. TENET [HWY16]. Term [ACR19, KS12b, AC15, BC13, CG11, CN14, CDR11, CW15b, Dua16, ELN11, GZ16, HW12a, HW14a, HR16, MPRR17]. terms [JN18, OÁ19]. test [ASX19, ALGQ16, Arm18, BFK12, BBKU19, CRY16, CHL14, CH10, CZ15a, CT17, CAS15, DG17, FLWZ15, FH19, GHX14, GKS15, GBH15b, GHM14, Gug10, Hag12, HHM11, HOY17, HS13b, JR15, JfL15, JPW10, Kas10, KLZ19, Kue19, LO12, LLY15, LLS14, LSW10, LWY16, LP18b, LfL19, LYCP19, NHKJ11, PW12, SW16, Sch11, Sha11, SLB18, SJZ15, VW14b, WWYL10, WT18, WW10, ZOZS16, ZL15a]. Testable [HW14a]. Testing [ASJL12, AT17, BCP15, BJS10, BBP15, BLY18, BE11, CW14a, CW14b, CRT10, CX14a, CET14, CHN^{+}14, CS13, CS14a, CW10, CKT17, CP19, CS14c, CSS18, DHN16, DEMY18, ES10, FF11, GHLT18, Gha17, GHM16, GHS16, GLH15, GL14, GKCS18, Gut16, HLT12b, HLT13, HC16, HSU10, HS13a, HSWY16, HKR14, HR15b, HT19, HS19b, ILT13, IR11, JKM17, JJFG10, JSE14, Kaw17, KL18, LZL^{+}16, LSW13, LKH14, Lee14b, LW14b, MSW19, OP18, OSW12, PY15, Son10, SW14b, TPY19, Tra16, WL12, YT15, BPS19, BM12, Cal11, CO12, CH11, Che15a, CH18b, CP18a, CTV19, CM15b, CS11b, DR15, DGTY17, Ell11, EJCL14, FLZ11, GK18a, GGL13, GLL19, Gup18b, HLOW16, HL14a, HLT12c, HS17a, Hor12, Kri11, Las10, LP13, LLS15, LL10b, LY12, LZY14, LLY15, MS17, McC17, MP12b, MW13, Ped17]. testing [Son11, SL13, SW17, Sun11, Swe11, Woo14, Zha16b]. testing-optimal [Sun11]. Tests [FJGM17, GMP17, Han15, HLT16, MMV14, MV13, PS17, Seo18, AH15, ASP12, AS17a, AS10, AW15, BDH16, BHP12, BB14, BCMH19, BJN15, Bre15, BJW15, BCS15, CNT15, CN12, Cha12, CZP15, CDDH15, CW11b, CR16, CM12, DV10, DE12, DS18, EV10, FMV15, GS15, GKR13, GS13, Gup18b, HvdAW11, HvdAW16, HLT06, HLT12a, HA13, HK14b, HS17b, JK15, JfL19, KS18b, Lav14, LLL16, LYZ18, LW14a, MP12a, MP13a, MRSE11, MP12b, MM19, PSU11, PG16, PVWW15, PS13, PS12, PSY13, PP18, Ric19, RS19, Shi15, SK12, Wes15a, WL15, WH19, WL14, Xu12, Xu13, Yan15b, ZC18, Zu15]. their [KZ15, jL12, LMT12, SV12]. theoretical [CHH18]. Theory [DT10, FGPZ17, FPSY19, GR14, KPS19, AS17a, AT14, BW14, BDZ11, BCCK15, BMM12, sC16, DMN19, Fre15, HP12a, JWY18, KP13, LP13, LTT16, LB17, Liu15, MP13b, PMG10, Rob11, Wil11]. Theory-coherent [GR14]. there [CDH12, Dez15, HS13b, HW14c]. thick [GL16, MP10]. thin [GL16]. third [Yan15a]. third-order [Yan15a]. thoughts [Gra10]. Three [HR19, BI14, KP15]. three-pass [KP15]. Three-stage [HR19]. Threshold [CPR10, MR11, SHHW18, Ton15, YP18, CYZ15, DPSS11, GTY13, LL12a, LLZ15, Mas17, PK17, SS16, SC15b, YL17, Yu12c, Yu15]. thresholding [FLM19, KKLW18]. thresholds [CCeC18]. tick [JLZ19]. tier [DP17]. Tighter [JPX11]. Tikhonov [GS12]. tilted [TYZ18]. timber [LZ12]. Time [BR12, BCRvD13, BT14, ET14, KG14, LMT15, ALS13, ASP16, AFTV19, AB18, AHH12, BCF18, BCLL16, BBP15, BFS19b, BKLS16, BGK19, BH16b, CLO18, CMM18, CX14a, CNT17, Cha16, CW15a, CKP16, CDES14, CH11, CS13, CLS14, CH14b, CM17a, Che19, CP12, CIY15, Cho13, CLR10, DFL18, DDSW18, DV10, DSS17, DL18, EMvS11, FH15a, FVW16, GG19, GT18, GKS15, GKY14, GO12, HPK14, HLOW16, HLT10, HA13, HLW15, HKR14, HR15b, IJR17, IDF11, JR15, Kas19, Kau15, Kim11, KW16, KSY17, KP10b, KLGS11, KK13, Kue12, KTT19, LR19, Lee12, fLY15, LLS15, LS15b, LLL15a, LTT17, LL12b, LXC16, Mac10, MXY12a, MW19a, MP13b, MP14, MM16, MK14, MHU19, NHKJ11, PU11, Pet19, Por19, QfLY17, RvdHW14, SG19, SC15a, SW17, SHHW18, Ton15, TY17, WW15, WY16, WLC19]. time [WL14, Xia12, Yu12a, ZLZ13, Zha15, Zha16b]. time-inhomogeneous [Cho13]. time-invariant [BBP15]. time-series [CMM18, MM16]. time-to-expiry [FH15a]. Time-varying [BR12, BCRvD13, BT14, KG14, ALS13, AB18, BFS19b, CM17a, Che19, CP12, DFL18, GKY14, IJR17, Kas19, Kau15, KK13, KTT19, MHU19, SW17, TY17, Zha15]. times [MZC19, PM17, RW11, RvdHW14, VW14b]. tobit [CT17, ALLM15, CZ11b, XfL15a, XfL18]. too [ILM19]. tools [WL15]. topology [DY14]. tournament [CS10]. trace [FGZ19]. tracking [CDES14]. trade [BGMT15, CDS13]. trade-off [BGMT15]. trading [BNHLS11a, Bon14, GO11, Hou17, LXZ16, SX17b, SYN12]. training [CGE10]. transaction [HV19, RvdHW14]. transfer [LCY10]. transform [Tsi12]. transformation [Che10b, CKK15, LLS15, WH10, XfL15b]. transformations [JJ14]. transformed [HP15, JCLL10, LS15a, LZK18]. transforms [BCP15, GG19, GM11, Son10, TTG11]. transition [AMM11, Cho15, Kau15, LSL14, VW14b]. transitions [VRW18, WW10]. transmission [CDF12]. trapezoidal [Can10b]. treated [Lee18]. Treatment [BC14a, BSV12, jL12, BD12, Bha13, CLO18, Cat10, CFW14, CHS19, Chi10, DH14, FP12, Far15, FR19, FFM12, GM17, Hag19, HHV16, HS19b, JWFS16, KM16, Kle10, Lee18, LY16, LB17, LZW^{+}18, MSV19, Mou15, NDT19, OP15, Sto12, SY19, Tet12, VRW18]. treatment-effect [OP15]. treatments [FH15c]. trend [ADG11, HLT16, HLT12b, HLT12c, HLT13, HL15, ILT13, Kim11, LL12b, Sun11, Xu12]. trending [AHH12, CGL12, Ell11, Rob12, RV18]. trends [HvdAW16, JJFG10, JJ14]. trial [ILM19]. triangular [CRY16, GP13, sHDW17, HHM17, HKY19, JPX11, KV10a, Mou15]. trimmed [CT17]. trinity [Gup18b]. true [JvN11, Ket18, KTT19, LPZ19]. truncated [CZ12, Cíz12, Sho16]. truncation [ADS12]. TSRV [MZC19]. turning [SW14a]. Twenty [BFvD10b, Fla10]. Two [FLZ11, HS18a, JW10, Leu15, AAMS10, ALW12, BLO14, CLO18, CL15a, CP18a, DP17, DGR11, ELN11, FVV11, FR17, GK18b, HL17b, HS18b, KW14, Kri10, MPP10, MM19, MZC19, PK17, SX16]. two-sided [AAMS10, MM19]. Two-stage [FLZ11, HS18a, HL17b, KW14]. Two-step [Leu15, CL15a, DGR11, FVV11, FR17, GK18b, HS18b]. two-tier [DP17]. type [AW15, CZ15a, Spa10, ZC18]. types [LVX17].

U.S. [CDS14]. Uhlenbeck [Fas13]. UK [JW10]. Ultra [ASMZ11]. unadjusted [Sai13]. unbalanced [ALX11, Hua14, Woo19]. unbiased [MW19b]. uncertainty [AX18, Bia16, HLT12b, HV19, Sal12]. unconditional [HK11, OSW12, TYZ18]. unconfoundedness [HSWY16, LY16]. under-specified [CH14a]. Underidentification [AHS12, DG17]. underlying [CS13]. Understanding [Che17, DNR10, RS11, WL15]. Undesirable [FS14]. Unequal [SX17a]. Unexplained [KZ15]. Unified [AFTV19, KW16, Yan18, CH14b, LZL18, LYCP19]. Uniform [EJCL14, FR18, HL12, KS18a, KS19b, LP13, AC13, BCCK15, CC15b, CHS19, Koc16]. uniformly [EJC12]. unifying [AS17a, PS12]. Unit [HLT12c, CX14a, CET14, Cha12, GTY13, HvdAW11, Han10, HLT12b, HLT13, Kru08, Kru14, LP17, MP12b, PSU11, PSY13, PMG10, WWYL10, WL12, Wes15a, WL15, ZC18]. unit-root [ZC18]. United [ALX11, LPT10]. univariate [CH14b, CHL18, WPY11]. unknown [AC12, AHS10, BJT16, CNT17, DPY16, HLT16, HL15, ILT13, KS18a, KSV15b, LfL10a, LPLZ18, MS18, Nie10, PY15, Sha11]. unobservable [Mat16]. unobservables [HSWY16, Hu17]. Unobserved [SVV16, BBP15, CDD19, CW10, Dez15, GG19, Gay13, HR12b, HS12b, HMS13, Lee15, Mat12, NR10, SH13]. Unpredictability [HM14]. unrelated [HKLP15, ZA10]. unspanned [CW15b]. unspecified [GM17, ZLX15]. unstable [BH13]. urban [AFPZ16]. urn [PMM15]. USA [HPY10]. use [And16, Cha17, Dez15]. Useful [Gra12, CG11]. usefulness [Ell11]. Using [ASX17, ACR19, DiT16, MP12b, O’D16, AT14, ADS12, AHK12, APT18, BCP15, BL17, BBCL18, BMM12, BCH11, Bha13, BCRvD13, BCIT14, BKLS16, BKM14, BN19b, Can10b, cCTZ14, CFVH^{+}15, sCJP14, CPvdW16, CGP16, DF15, DP17, DT19, DHN16, Die14, EJW14, FXY17, FLM19, GGR17, GR18a, GKM14, HR17, HK14a, HLT13, HP18, sHDW17, KP15, Kha13, Khw10, KS14, KS18b, KLS11, LS15a, LT15a, LC16, LTT17, LP18a, LKJ18, OSW12, Pat11b, Sai13, Son10, SW14a, WKM15]. utility [BA11, LM11, YY19].

V [Ahn15]. valid [CL15b]. validated [HR15a]. validating [CH14b]. validation [GZWZ16, LZZZ19, ZY15]. validations [Zha11b]. validity [BCF12, Can10a, Sto12]. Value [LT15b, WZ16, CWZ19, LT19, Wes12, PZC19]. Value-at-Risk [LT15b, WZ16, PZC19]. valued [Cat10, SHHW18, WLC19]. values [BBKU19, GK18a, HS13b, JvN11, Lea16, Yan15b]. VAR [IK19, AdCGIV11, BCR19, BJN15, CHL18, GMR17, Han15, HLT16, IK13, Pet19, Swe11, WKM15, DJS^{+}13, FZ18]. Variable [PPT11, Sal12, STZ19, BC15, BJR17, BD15, CJY15, DL17, DL18, FV15, GS12, GG17, HR17, Hor12, JWFS16, Koc16, KLGS12, Lee15, LY12, Mat12, Oku11, San11, SH16, WK16, Xie88, Xie11, XfL15b]. variables [Abr19, APS17, AMM11, Bha14b, CCH13, CCJ12, CFFS16, Cha17, CFR18, CLLL16, CLL19, CP19, DL16, Ell11, EJW14, FF14, FH15b, GHX14, Geo10, GKN17, GS13, HR17, HR19, HK14a, HL12, Hor14, HS12b, HKY19, KS19b, Kau15, KS14, Kle10, KKL14, LYCP19, MSV19, MM19, MP17, MW16, PWM11, SW16, SHL11, Swe11, Tor17, WH11, Woo14]. Variance [Bon14, Sho16, Wu11, ADZ14, AT13, AS10, AD13, BR11, BH14a, CP12, COP10, CTV19, CP18b, DY14, EW15, FMV15, FM10, KSV15b, KKL14, PS12, Siz11, Sun11, TY17, Yan15a, ZWXW10]. variances [Bai10]. variate [AT11]. Variation [PG16, ABH11, CHP18, CPR10, Lar11]. Variation-based [PG16]. variations [DF15, HV17c]. varies [WW15]. various [PL13]. VARMA [Zad16]. VARMAs [CEK16]. VARs [BJT16, CP11, GHS16, KK13]. Varying [MKS16, ALS13, AB18, BR12, BCRvD13, BFS19b, BT14, CX12, CM17a, Che19, CP12, DFL18, FGPZ17, GKY14, GL14, IJR17, KG14, Kas19, Kau15, KP10b, KLGS11, KK13, KTT19, MHU19, Pet19, QfLY17, SW17, TY17, ZLX15, Zha15]. varying-coefficient [FGPZ17, ZLX15]. Vector [Pos16, Bog22, BCRT16, CS11a, CCM19, CCCM22, CRT10, DK18, FP11b, GJ17, HL14b, JM19, Joh10, JN18, KG18, KMS16, Ket18, KC15, KKP19, KP19, LMS17, LS19a, LZZZ19, Qia16]. versions [SV12]. versus [Spa10, Sun14, TU10, WU15]. via [AM15a, AHH12, BL16, BS12, CC15a, Cos13, DMC12, FGZ19, FR17, FL13, Gup18b, Kap15a, MS19, MP14, QS16, Ros12, SW14b, YCW19]. view [KSY17]. violations [BA11]. visits [Gil10]. Visualizing [HL17b]. VIX [BH14a, Par16]. volatilities [AM15a, BH17, BGZ11]. Volatility [CW11a, GS11a, MMS11, Pat11b, TTG14, AH15, ASMZ11, ASAM15, ASX16, AT13, AX18, ABH11, ABM11, ADS12, And16, AM15b, ACM17, BCC^{+}15, BBGV14, BV11, BH14a, BEE15, Bog22, BGZ11, BXZ15, BPQ16, BMN19, BCN11, CRdM17, CCM19, CCCM22, CRT10, Cha13b, CX14b, CZP15, CHP18, CP18b, CLR10, CDF12, CPR10, CW15b, Des11, DSS17, DGM13, DS15, EM14b, FK19, FVGQRR15, FZ15, GLPR11, GGL12, GM11, HL10, Ho15, HS18a, JKM17, JLZ19, JM10, JM14, KKLW18, KF19, LRV13, LXZ16, LTT16, LZL18, LKJ18, MM11, MBC15, MMFG12, MSS11, PG16, Par16, Pat11a, RW11, SX16, Tod11, TTG11, Xiu10, Xu12, Xu13, Yu12b, ZMAS11, ZB14, Zu15]. vs [Blu10, HR12b, Kea10, Rus10, Siz11]. Vuong [MMS19].

wage [CS10, DMZ18, Gut16, LW10a]. Wages [KW10b, HLR11]. walk [PW12]. Walker [DPY16, Zad16]. Wavelet [ZWXW10, CF19]. way [LZZL18, MRSE11, Xie88, Xie11]. Weak [KPS19, CY12, Can14, CC15b, CL15a, CDDH15, FP14, GRS12, HHM11, HK14a, IR11, JW10, Mag12, Mik10, MMV14, SW16, WK16, ZL15a]. weakly [CZ11a, CP15, Ona12]. Wealth [PT11a, Jus11]. Weibull [CW10]. weight [QfL15]. Weighted [Arm14, BKMZ16, DMP18, CF19, CPP19, EJCL14, FZ10, HOvD12, Kue12, MW19b, RL17, RW17, RS17, YL17]. Weighted-average [DMP18]. weighting [Can10b]. weights [QfLY17, Sun16, ZY18]. welfare [BD12, GJZ15, KW10b, LB19]. Well [AH12b, MR16]. well-being [MR16]. Well-posedness [AH12b]. where [HHS12]. whether [CS13]. White [Zha16b, DMZ18, GGL13, Sha11]. Whittle [FNN12, HP14b, Shi12]. whole [HS13a]. wild [DMN19, KS12c]. willingness [Khw10]. window [IJR17]. Windows [HR12b]. Wishart [AM15b, GGL12]. within [CHL18, Dez15, Hil16]. within-region [Hil16]. without [BMDL17, CW18, FWZ13, HW14b, HK11, HK14b, Kaw17, KV10a, Kle10, LKH14, LY16, MRT16, Sch19]. women [Gil10]. Work [Gil10, BSW10]. Workforce [AV11]. worst [Myk19].

year [Slo17]. years [BFvD10b, Fla10]. yield [BC10, ELN11, LL10a]. yields [KS12b]. Yor [BCL14]. young [HLR11]. yourself [Poi12]. Yule [DPY16, Zad16].

ZD [LZZL18]. ZD-GARCH [LZZL18]. Zellner [Ano13a, Ano17a, Ano15u]. zero [KS12b, KPT13, LKJ18, MPRR17].

References

Ahn:2010:ECR

[AAMS10] Tom Ahn, Peter Arcidiacono, Alvin Murphy, and Omari Swinton. Explaining cross-racial differences in teenage labor force participation: Results from a two-sided matching model. Journal of Econometrics, 156(1):201–211, May 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002164.

Antoine:2018:EET

[AB18] Bertille Antoine and Otilia Boldea. Efficient estimation with time-varying information and the new Keynesian Phillips curve. Journal of Econometrics, 204(2):268–300, June 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300320.

Andersen:2011:RFF

[ABH11] Torben G. Andersen, Tim Bollerslev, and Xin Huang. A reduced form framework for modeling volatility of speculative prices based on realized variation measures. Journal of Econometrics, 160(1):176–189, January 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000710.

Armstrong:2014:FRM

[ABH14] Timothy B. Armstrong, Marinho Bertanha, and Han Hong. A fast resample method for parametric and semiparametric models. Journal of Econometrics, 179(2):128–133, April 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000025.

Assmann:2016:BAS

[ABHP16] Christian Aßmann, Jens Boysen-Hogrefe, and Markus Pape. Bayesian analysis of static and dynamic factor models: an expost approach towards the rotation problem. Journal of Econometrics, 192(1):190–206, May 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615002626.

Andersen:2011:RVF

[ABM11] Torben G. Andersen, Tim Bollerslev, and Nour Meddahi. Realized volatility forecasting and market microstructure noise. Journal of Econometrics, 160(1):220–234, January 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000746.

Abrevaya:2019:MDV

[Abr19] Jason Abrevaya. Missing dependent variables in fixed-effects models. Journal of Econometrics, 211(1):151–165, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302422.

Andersen:2015:FSE

[ABTT15] Torben G. Andersen, Oleg Bondarenko, Viktor Todorov, and George Tauchen. The fine structure of equity-index option dynamics. Journal of Econometrics, 187(2):532–546, August 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000627.

Ai:2012:SEB

[AC12] Chunrong Ai and Xiaohong Chen. The semiparametric efficiency bound for models of sequential moment restrictions containing unknown functions. Journal of Econometrics, 170(2):442–457, October 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001273.

Andrews:2013:MLE

[AC13] Donald W. K. Andrews and Xu Cheng. Maximum likelihood estimation and uniform inference with sporadic identification failure. Journal of Econometrics, 173(1):36–56, March 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612002357.

Andreasen:2015:SAN

[AC15] Martin M. Andreasen and Bent Jesper Christensen. The SR approach: a new estimation procedure for non-linear and non-Gaussian dynamic term structure models. Journal of Econometrics, 184(2):420–451, February 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614002292.

Armstrong:2016:MAI

[AC16] Timothy B. Armstrong and Hock Peng Chan. Multiscale adaptive inference on conditional moment inequalities. Journal of Econometrics, 194(1):24–43, September 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300574.

Arulampalam:2017:MHD

[ACG17] Wiji Arulampalam, Valentina Corradi, and Daniel Gutknecht. Modeling heaped duration data: an application to neonatal mortality. Journal of Econometrics, 200(2):363–377, October 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301021.

Asai:2017:RSV

[ACM17] Manabu Asai, Chia-Lin Chang, and Michael McAleer. Realized stochastic volatility with general asymmetry and long memory. Journal of Econometrics, 199(2):202–212, August 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300726.

Andreasen:2019:TSA

[ACR19] Martin M. Andreasen, Jens H. E. Christensen, and Glenn D. Rudebusch. Term structure analysis with big data: One-step estimation using bond prices. Journal of Econometrics, 212(1):26– 46, September 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300740.

Abadir:2013:NPR

[ACT13] Karim M. Abadir, Giovanni Caggiano, and Gabriel Talmain. Nelson-plosser revisited: The ACF approach. Journal of Econometrics, 175(1):22–34, July 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613000419.

Andrews:2013:MII

[AD13] Beth Andrews and Richard A. Davis. Model identification for infinite variance autoregressive processes. Journal of Econometrics, 172(2):222–234, February 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001935.

Athanasopoulos:2011:MSE

[AdCGIV11] George Athanasopoulos, Osmani Teixeira de Carvalho Guillén, João Victor Issler, and Farshid Vahid. Model selection, estimation and forecasting in VAR models with short-run and long-run restrictions. Journal of Econometrics, 164(1):116–129, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000364.

Anderson:2016:SMA

[ADFK16] Brian D. O. Anderson, Manfred Deistler, Elisabeth Felsenstein, and Lukas Koelbl. The structure of multivariate AR and ARMA systems: Regular and singular systems; the single and the mixed frequency case. Journal of Econometrics, 192(2):366–373, June 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300070.

Abadir:2011:DMT

[ADG11] Karim M. Abadir, Walter Distaso, and Liudas Giraitis. An i( d ) model with trend and cycles. Journal of Econometrics, 163(2):186–199, August 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611000686.

Aruoba:2016:IGM

[ADN^{+}16] S. Boragan Aruoba, Francis X. Diebold, Jeremy Nalewaik, Frank Schorfheide, and Dongho Song. Improving GDP measurement: a measurement-error perspective. Journal of Econometrics, 191(2): 384–397, April 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002857.

Andersen:2012:JRV

[ADS12] Torben G. Andersen, Dobrislav Dobrev, and Ernst Schaumburg. Jump-robust volatility estimation using nearest neighbor truncation. Journal of Econometrics, 169(1):75–93, July 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000127.

Abadir:2014:DFE

[ADZ14] Karim M. Abadir, Walter Distaso, and Filip Zikes. Designfree estimation of variance matrices. Journal of Econometrics, 181(2):165–180, August 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614000591.

Anderson:2016:NAM

[AFPZ16] Gordon Anderson, Alessio Farcomeni, Maria Grazia Pittau, and Roberto Zelli. A new approach to measuring and studying the characteristics of class membership: Examining poverty, inequality and polarization in urban China. Journal of Econometrics, 191(2):348–359, April 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002821.

Amengual:2013:SES

[AFS13] Dante Amengual, Gabriele Fiorentini, and Enrique Sentana. Sequential estimation of shape parameters in multivariate dynamic models. Journal of Econometrics, 177(2):233–249, December 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000870.

Andersen:2019:UIN

[AFTV19] Torben G. Andersen, Nicola Fusari, Viktor Todorov, and Rasmus T. Varneskov. Unified inference for nonlinear factor models from panels with fixed and large time span. Journal of Econometrics, 212(1):4–25, September 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300739.

Ai:2010:ARC

[AG10a] Chunrong Ai and Li Gan. An alternative root-n consistent estimator for panel data binary choice models. Journal of Econometrics, 157(1):93–100, July 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002723.

Andrews:2010:ASH

[AG10b] Donald W. K. Andrews and Patrik Guggenberger. Applications of subsampling, hybrid, and size-correction methods. Journal of Econometrics, 158(2):285–305, October 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000059.

Almeida:2012:AMA

[AG12a] Caio Almeida and René Garcia. Assessing misspecified asset pricing models with empirical likelihood estimators. Journal of Econometrics, 170(2):519–537, October 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001327.

Andrews:2012:ALG

[AG12b] Donald W. K. Andrews and Patrik Guggenberger. Asymptotics for LS, GLS, and feasible GLS statistics in an AR(1) model with conditional heteroskedasticity. Journal of Econometrics, 169(2):196–210, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612000279.

Asimit:2016:TDM

[AGHP16] Alexandru V. Asimit, Russell Gerrard, Yanxi Hou, and Liang Peng. Tail dependence measure for examining financial extreme co-movements. Journal of Econometrics, 194(2):330–348, October 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301026.

Almeida:2011:DIR

[AGJ11] Caio Almeida, Jeremy J. Graveline, and Scott Joslin. Do interest rate options contain information about excess returns? Journal of Econometrics, 164(1):35–44, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000340.

Andreou:2010:RMM

[AGK10] Elena Andreou, Eric Ghysels, and Andros Kourtellos. Regression models with mixed sampling frequencies. Journal of Econometrics, 158(2):246–261, October 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000072.

Aryal:2018:ERA

[AGKZ18] Gaurab Aryal, Serafin Grundl, Dong-Hyuk Kim, and Yu Zhu. Empirical relevance of ambiguity in first-price auctions. Journal of Econometrics, 204(2):189–206, June 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300204.

Allen:2011:ELB

[AGS11] Jason Allen, Allan W. Gregory, and Katsumi Shimotsu. Empirical likelihood block bootstrapping. Journal of Econometrics, 161(2):110–121, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610002046.

Aguirregabiria:2012:DOG

[AH12a] Victor Aguirregabiria and Chun-Yu Ho. A dynamic oligopoly game of the US airline industry: Estimation and policy experiments. Journal of Econometrics, 168(1):156–173, May 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761100176X.

An:2012:WPM

[AH12b] Yonghong An and Yingyao Hu. Well-posedness of measurement error models for self-reported data. Journal of Econometrics, 168(2):259–269, June 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612000462.

Aguilar:2015:RSP

[AH15] Mike Aguilar and Jonathan B. Hill. Robust score and portmanteau tests of volatility spillover. Journal of Econometrics, 184(1): 37–61, January 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001821.

Aue:2012:SMN

[AHH12] Alexander Aue, Lajos Horváth, and Marie Husková. Segmenting mean-nonstationary time series via trending regressions. Journal of Econometrics, 168(2):367–381, June 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000590.

Arcidiacono:2012:MCM

[AHK12] Peter Arcidiacono, V. Joseph Hotz, and Songman Kang. Modeling college major choices using elicited measures of expectations and counterfactuals. Journal of Econometrics, 166(1):3–16, January 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001151.

Ahn:2015:CIE

[Ahn15] Seung C. Ahn. Comment on ‘IV estimation of panels with factor residuals’ by D. Robertson and V. Sarafidis. Journal of Econometrics, 185(2):542–544, April 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614002905. See [RS15].

An:2010:EFP

[AHS10] Yonghong An, Yingyao Hu, and Matthew Shum. Estimating first-price auctions with an unknown number of bidders: a misclassification approach. Journal of Econometrics, 157(2):328– 341, August 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000308.

Arellano:2012:U

[AHS12] Manuel Arellano, Lars Peter Hansen, and Enrique Sentana. Underidentification? Journal of Econometrics, 170(2):256–280, October 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001170.

Aigner:2016:O

[Aig16] Dennis J. Aigner. Obituary. Journal of Econometrics, 194(1): iv, September 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616301300.

Akashi:2012:SPL

[AK12] Kentaro Akashi and Naoto Kunitomo. Some properties of the LIML estimator in a dynamic panel structural equation. Journal of Econometrics, 166(2):167–183, February 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001631.

Anderson:2010:AOL

[AKM10] T. W. Anderson, Naoto Kunitomo, and Yukitoshi Matsushita. On the asymptotic optimality of the LIML estimator with possibly many instruments. Journal of Econometrics, 157(2):191– 204, August 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609002991.

Anderson:2011:FSP

[AKM11] T. W. Anderson, Naoto Kunitomo, and Yukitoshi Matsushita. On finite sample properties of alternative estimators of coefficients in a structural equation with many instruments. Journal of Econometrics, 165(1):58–69, November 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000984.

Aradillas-Lopez:2010:SES

[AL10] Andres Aradillas-Lopez. Semiparametric estimation of a simultaneous game with incomplete information. Journal of Econometrics, 157(2):409–431, August 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000953.

Aradillas-Lopez:2012:PDE

[AL12] Andres Aradillas-Lopez. Pairwise-difference estimation of incomplete information games. Journal of Econometrics, 168(1):120– 140, May 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001734.

Antoine:2014:CMM

[AL14] Bertille Antoine and Pascal Lavergne. Conditional moment models under semi-strong identification. Journal of Econometrics, 182(1):59–69, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614000670.

Aradillas-Lopez:2016:STM

[ALGQ16] Andrés Aradillas-López, Amit Gandhi, and Daniel Quint. A simple test for moment inequality models with an application to English auctions. Journal of Econometrics, 194(1):96–115, September 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300628.

Ai:2015:EPD

[ALLM15] Chunrong Ai, Hongjun Li, Zhongjian Lin, and Meixia Meng. Estimation of panel data partly specified Tobit regression with fixed effects. Journal of Econometrics, 188(2):316–326, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000718.

Ahn:2013:PDM

[ALS13] Seung C. Ahn, Young H. Lee, and Peter Schmidt. Panel data models with multiple time-varying individual effects.

Journal of Econometrics, 174(1):1–14, May 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761300002X.

Anderson:2012:NEI

[ALW12] Gordon Anderson, Oliver Linton, and Yoon-Jae Whang. Nonparametric estimation and inference about the overlap of two distributions. Journal of Econometrics, 171(1):1–23, November 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001108.

Atak:2011:SPM

[ALX11] Alev Atak, Oliver Linton, and Zhijie Xiao. A semiparametric panel model for unbalanced data with application to climate change in the United Kingdom. Journal of Econometrics, 164(1):92–115, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611000352.

Aguirregabiria:2010:DDC

[AM10] Victor Aguirregabiria and Pedro Mira. Dynamic discrete choice structural models: a survey. Journal of Econometrics, 156(1): 38–67, May 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001985.

Asai:2015:FCV

[AM15a] Manabu Asai and Michael McAleer. Forecasting co-volatilities via factor models with asymmetry and long memory in realized covariance. Journal of Econometrics, 189(2):251–262, December 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000986.

Asai:2015:LFE

[AM15b] Manabu Asai and Michael McAleer. Leverage and feedback effects on multifactor Wishart stochastic volatility for option pricing. Journal of Econometrics, 187(2):436–446, August 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000548.

Areosa:2011:MBE

[AMM11] Waldyr Dutra Areosa, Michael McAleer, and Marcelo C. Medeiros. Moment-based estimation of smooth transition regression models with endogenous variables. Journal of Econometrics, 165(1):100–111, November 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611001011.

An:2017:IFP

[An17] Yonghong An. Identification of first-price auctions with nonequilibrium beliefs: a measurement error approach. Journal of Econometrics, 200(2):326–343, October 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617301008.

Anatolyev:2012:IRM

[Ana12] Stanislav Anatolyev. Inference in regression models with many regressors. Journal of Econometrics, 170(2):368–382, October 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001236.

Anderson:2010:LEF

[And10] T. W. Anderson. The LIML estimator has finite moments! Journal of Econometrics, 157(2):359–361, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000801.

Andreou:2016:UHF

[And16] Elena Andreou. On the use of high frequency measures of volatility in MIDAS regressions. Journal of Econometrics, 193(2):367– 389, August 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300719.

Andrews:2017:ECB

[And17] Donald W. K. Andrews. Examples of L^{2}-complete and boundedly-complete distributions. Journal of Econometrics, 199(2):213–220, August 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617300738.

Andrews:2019:SIE

[And19] Isaiah Andrews. On the structure of IV estimands. Journal of Econometrics, 211(1):294–307, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407618302537.

Anonymous:2010:EBa

[Ano10a] Anonymous. Editorial Board. Journal of Econometrics, 154(1): ifc, January 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609002401.

Anonymous:2010:EBb

[Ano10b] Anonymous. Editorial Board. Journal of Econometrics, 154(2): ifc, February 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609002498.

Anonymous:2010:EBc

[Ano10c] Anonymous. Editorial Board. Journal of Econometrics, 155(1): ifc, March 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000175.

Anonymous:2010:EBd

[Ano10d] Anonymous. Editorial Board. Journal of Econometrics, 155(2): ifc, April 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000321.

Anonymous:2010:EBe

[Ano10e] Anonymous. Editorial Board. Journal of Econometrics, 156 (1):ifc, May 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000874.

Anonymous:2010:EBf

[Ano10f] Anonymous. Editorial Board. Journal of Econometrics, 156 (2):ifc, June 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610001004.

Anonymous:2010:EBg

[Ano10g] Anonymous. Editorial Board. Journal of Econometrics, 157 (1):ifc, July 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610001168.

Anonymous:2010:EBh

[Ano10h] Anonymous. Editorial Board. Journal of Econometrics, 157(2): ifc, August 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610001326.

Anonymous:2010:EBi

[Ano10i] Anonymous. Editorial Board. Journal of Econometrics, 158(1): ifc, September 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610001491.

Anonymous:2010:EBj

[Ano10j] Anonymous. Editorial Board. Journal of Econometrics, 158(2): ifc, October 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610001624.

Anonymous:2010:EBk

[Ano10k] Anonymous. Editorial Board. Journal of Econometrics, 159(1): ifc, November 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610001740.

Anonymous:2010:EBl

[Ano10l] Anonymous. Editorial Board. Journal of Econometrics, 159(2): ifc, December 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610001958.

Anonymous:2010:PJa

[Ano10m] Anonymous. Pages 1–100 (January 2010). Journal of Econometrics, 154(1):??, January 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2010:PNa

[Ano10n] Anonymous. Pages 1–250 (November 2010). Journal of Econometrics, 159(1):??, November 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2010:PM

[Ano10o] Anonymous. Pages 1–98 (March 2010). Journal of Econometrics, 155(1):??, March 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2010:PF

[Ano10p] Anonymous. Pages 101–202 (February 2010). Journal of Econometrics, 154(2):??, February 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2010:PAb

[Ano10q] Anonymous. Pages 191–512 (August 2010). Journal of Econometrics, 157(2):??, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2010:PJb

[Ano10r] Anonymous. Pages 239–426 (June 2010). Journal of Econometrics, 156(2):??, June 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2010:PD

[Ano10s] Anonymous. Pages 251–330 (December 2010). Journal of Econometrics, 159(2):??, December 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2010:PAa

[Ano10t] Anonymous. Pages 99–194 (April 2010). Journal of Econometrics, 155(2):??, April 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2010:PNb

[Ano10u] Anonymous. Publisher’s note. Journal of Econometrics, 159(2): 251, December 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610001442.

Anonymous:2011:EBa

[Ano11a] Anonymous. Editorial Board. Journal of Econometrics, 160(1): ifc, January 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610002174.

Anonymous:2011:EBb

[Ano11b] Anonymous. Editorial Board. Journal of Econometrics, 160(2): ifc, February 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610002368.

Anonymous:2011:EBc

[Ano11c] Anonymous. Editorial Board. Journal of Econometrics, 161(1): ifc, March 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611000133.

Anonymous:2011:EBd

[Ano11d] Anonymous. Editorial Board. Journal of Econometrics, 161(2): ifc, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611000169.

Anonymous:2011:EBe

[Ano11e] Anonymous. Editorial Board. Journal of Econometrics, 162 (1):ifc, May 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611000522.

Anonymous:2011:EBf

[Ano11f] Anonymous. Editorial Board. Journal of Econometrics, 162 (2):ifc, June 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611000716.

Anonymous:2011:EBg

[Ano11g] Anonymous. Editorial Board. Journal of Econometrics, 163 (1):ifc, July 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611000844.

Anonymous:2011:EBh

[Ano11h] Anonymous. Editorial Board. Journal of Econometrics, 163(2): ifc, August 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001060.

Anonymous:2011:EBi

[Ano11i] Anonymous. Editorial Board. Journal of Econometrics, 164(1): ifc, September 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001333.

Anonymous:2011:EBj

[Ano11j] Anonymous. Editorial Board. Journal of Econometrics, 164(2): ifc, October 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001497.

Anonymous:2011:EBk

[Ano11k] Anonymous. Editorial Board. Journal of Econometrics, 165(1): ifc, November 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001886.

Anonymous:2011:EBl

[Ano11l] Anonymous. Editorial Board. Journal of Econometrics, 165(2): ifc, December 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002247.

Anonymous:2011:M

[Ano11m] Anonymous. In memorium. Journal of Econometrics, 160(2):iv– v, February 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610002423.

Anonymous:2011:PAa

[Ano11n] Anonymous. Pages 101–338 (1 April 2011). Journal of Econometrics, 161(2):??, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2011:PAb

[Ano11o] Anonymous. Pages 127–230 (August 2011). Journal of Econometrics, 163(2):??, August 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2011:PD

[Ano11p] Anonymous. Pages 137–274 (December 2011). Journal of Econometrics, 165(2):??, December 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2011:PJ

[Ano11q] Anonymous. Pages 149–396 (June 2011). Journal of Econometrics, 162(2):??, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2011:PO

[Ano11r] Anonymous. Pages 207–404 (1 October 2011). Journal of Econometrics, 164(2):??, October 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2011:PF

[Ano11s] Anonymous. Pages 289–372 (February 2011). Journal of Econometrics, 160(2):??, February 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2012:JE

[Ano12a] Anonymous. 2011 Journal of Econometrics. Journal of Econometrics, 168(2):vi, June 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612000917.

Anonymous:2012:DJA

[Ano12b] Anonymous. 2011 Dennis J. Aigner Award. Journal of Econometrics, 168(2):v, June 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612000905.

Anonymous:2012:EBa

[Ano12c] Anonymous. Editorial Board. Journal of Econometrics, 166(1): ifc, January 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002466.

Anonymous:2012:EBb

[Ano12d] Anonymous. Editorial Board. Journal of Econometrics, 166(2): ifc, February 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002612.

Anonymous:2012:EBc

[Ano12e] Anonymous. Editorial Board. Journal of Econometrics, 167(1): ifc, March 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000176.

Anonymous:2012:EBd

[Ano12f] Anonymous. Editorial Board. Journal of Econometrics, 167(2): ifc, April 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000498.

Anonymous:2012:EBe

[Ano12g] Anonymous. Editorial Board. Journal of Econometrics, 168 (1):ifc, May 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000620.

Anonymous:2012:EBf

[Ano12h] Anonymous. Editorial Board. Journal of Econometrics, 168 (2):ifc, June 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000814.

Anonymous:2012:EBg

[Ano12i] Anonymous. Editorial Board. Journal of Econometrics, 169 (1):ifc, July 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001005.

Anonymous:2012:EBh

[Ano12j] Anonymous. Editorial Board. Journal of Econometrics, 169(2): ifc, August 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001376.

Anonymous:2012:EBi

[Ano12k] Anonymous. Editorial Board. Journal of Econometrics, 170(1): ifc, September 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001637.

Anonymous:2012:EBj

[Ano12l] Anonymous. Editorial Board. Journal of Econometrics, 170(2): ifc, October 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001790.

Anonymous:2012:EBk

[Ano12m] Anonymous. Editorial Board. Journal of Econometrics, 171(1): ifc, November 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612002060.

Anonymous:2012:EBl

[Ano12n] Anonymous. Editorial Board. Journal of Econometrics, 171(2): ifc, December 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612002230.

Anonymous:2012:LRJ

[Ano12o] Anonymous. List of referees from January 1, 2011 to December 31, 2011. Journal of Econometrics, 170(1):I–V, September 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001704.

Anonymous:2012:LJF

[Ano12p] Anonymous. List of the JE Fellows as of January 2011. Journal of Econometrics, 168(2):vii–xix, June 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000929.

Anonymous:2012:PS

[Ano12q] Anonymous. Pages 1–250 (September 2012). Journal of Econometrics, 170(1):??, September 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2012:PM

[Ano12r] Anonymous. Pages 1–294 (March 2012). Journal of Econometrics, 167(1):??, March 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2012:PN

[Ano12s] Anonymous. Pages 1–98 (November 2012). Journal of Econometrics, 171(1):??, November 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2012:PF

[Ano12t] Anonymous. Pages 167–354 (February 2012). Journal of Econometrics, 166(2):??, February 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2012:PJ

[Ano12u] Anonymous. Pages 175–408 (June 2012). Journal of Econometrics, 168(2):??, June 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2013:AZA

[Ano13a] Anonymous. 2012 Arnold Zellner Award. Journal of Econometrics, 173(1):v, March 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613000122.

Anonymous:2013:JE

[Ano13b] Anonymous. 2012 Journal of Econometrics. Journal of Econometrics, 173(1):vi, March 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613000134.

Anonymous:2013:EBa

[Ano13c] Anonymous. Editorial Board. Journal of Econometrics, 172(1): ifc, January 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612002370.

Anonymous:2013:EBb

[Ano13d] Anonymous. Editorial Board. Journal of Econometrics, 172(2): ifc, February 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612002515.

Anonymous:2013:EBc

[Ano13e] Anonymous. Editorial Board. Journal of Econometrics, 173(1): ifc, March 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613000195.

Anonymous:2013:EBd

[Ano13f] Anonymous. Editorial Board. Journal of Econometrics, 173(2): ifc, April 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613000237.

Anonymous:2013:EBe

[Ano13g] Anonymous. Editorial Board. Journal of Econometrics, 174 (1):ifc, May 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613000444.

Anonymous:2013:EBf

[Ano13h] Anonymous. Editorial Board. Journal of Econometrics, 174 (2):ifc, June 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613000717.

Anonymous:2013:EBg

[Ano13i] Anonymous. Editorial Board. Journal of Econometrics, 175 (1):ifc, July 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613000973.

Anonymous:2013:EBh

[Ano13j] Anonymous. Editorial Board. Journal of Econometrics, 175(2): ifc, August 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613001061.

Anonymous:2013:EBi

[Ano13k] Anonymous. Editorial Board. Journal of Econometrics, 176(1): ifc, September 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613001188.

Anonymous:2013:EBj

[Ano13l] Anonymous. Editorial Board. Journal of Econometrics, 176(2): ifc, October 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761300136X.

Anonymous:2013:EBk

[Ano13m] Anonymous. Editorial Board. Journal of Econometrics, 177(1): ifc, November 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613001711.

Anonymous:2013:EBl

[Ano13n] Anonymous. Editorial Board. Journal of Econometrics, 177(2): ifc, December 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761300211X.

Anonymous:2013:LJF

[Ano13o] Anonymous. List of the JE Fellows as of January 2012. Journal of Econometrics, 173(1):vii–xx, March 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613000146.

Anonymous:2013:PN

[Ano13p] Anonymous. Pages 1–130 (November 2013). Journal of Econometrics, 177(1):??, November 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2013:PMa

[Ano13q] Anonymous. Pages 1–142 (March 2013). Journal of Econometrics, 173(1):??, March 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2013:PJa

[Ano13r] Anonymous. Pages 1–182 (January 2013). Journal of Econometrics, 172(1):??, January 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2013:PMb

[Ano13s] Anonymous. Pages 1–44 (May 2013). Journal of Econometrics, 174(1):??, May 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2013:PJc

[Ano13t] Anonymous. Pages 1–60 (July 2013). Journal of Econometrics, 175(1):??, July 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2013:PS

[Ano13u] Anonymous. Pages 1–92 (September 2013). Journal of Econometrics, 176(1):??, September 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2013:PAa

[Ano13v] Anonymous. Pages 143–204 (April 2013). Journal of Econometrics, 173(2):??, April 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2013:PJb

[Ano13w] Anonymous. Pages 45–194 (June 2013). Journal of Econometrics, 174(2):??, June 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2013:PAb

[Ano13x] Anonymous. Pages 61–154 (August 2013). Journal of Econometrics, 175(2):??, August 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2013:PO

[Ano13y] Anonymous. Pages 93–172 (October 2013). Journal of Econometrics, 176(2):??, October 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2014:A

[Ano14a] Anonymous. Announcement. Journal of Econometrics, 178(2): 829, February 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613002662.

Anonymous:2014:EBa

[Ano14b] Anonymous. Editorial Board. Journal of Econometrics, 178(Part 1):ifc, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613002376.

Anonymous:2014:EBb

[Ano14c] Anonymous. Editorial Board. Journal of Econometrics, 178(Part 2):ifc, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613002455.

Anonymous:2014:EBc

[Ano14d] Anonymous. Editorial Board. Journal of Econometrics, 178(Part 3):ifc, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613002558.

Anonymous:2014:EBd

[Ano14e] Anonymous. Editorial Board. Journal of Econometrics, 178(2): ifc, February 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613002698.

Anonymous:2014:EBe

[Ano14f] Anonymous. Editorial Board. Journal of Econometrics, 179(1): ifc, March 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614000050.

Anonymous:2014:EBf

[Ano14g] Anonymous. Editorial Board. Journal of Econometrics, 179(2): ifc, April 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614000220.

Anonymous:2014:EBg

[Ano14h] Anonymous. Editorial Board. Journal of Econometrics, 180 (1):ifc, May 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614000517.

Anonymous:2014:EBh

[Ano14i] Anonymous. Editorial Board. Journal of Econometrics, 180 (2):ifc, June 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614000815.

Anonymous:2014:EBi

[Ano14j] Anonymous. Editorial Board. Journal of Econometrics, 181 (1):ifc, July 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614000931.

Anonymous:2014:EBj

[Ano14k] Anonymous. Editorial Board. Journal of Econometrics, 181(2): ifc, August 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001195.

Anonymous:2014:EBk

[Ano14l] Anonymous. Editorial Board. Journal of Econometrics, 182(1): ifc, September 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001298.

Anonymous:2014:EBl

[Ano14m] Anonymous. Editorial Board. Journal of Econometrics, 182(2): ifc, October 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001614.

Anonymous:2014:EBm

[Ano14n] Anonymous. Editorial Board. Journal of Econometrics, 183(1): ifc, November 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761400219X.

Anonymous:2014:EBn

[Ano14o] Anonymous. Editorial Board. Journal of Econometrics, 183(2): ifc, December 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614002371.

Anonymous:2014:LR

[Ano14p] Anonymous. List of referees for 2013. Journal of Econometrics, 178(2):825–828, February 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613002650.

Anonymous:2014:PMb

[Ano14q] Anonymous. Pages 1–116 (May 2014). Journal of Econometrics, 180(1):??, May 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2014:PMa

[Ano14r] Anonymous. Pages 1–98 (March 2014). Journal of Econometrics, 179(1):??, March 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2014:PJb

[Ano14s] Anonymous. Pages 117–264 (June 2014). Journal of Econometrics, 180(2):??, June 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2014:PO

[Ano14t] Anonymous. Pages 235–412 (October 2014). Journal of Econometrics, 182(2):??, October 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2014:PJa

[Ano14u] Anonymous. Pages 383–706 (January 2014). Journal of Econometrics, 178(Part 3):383–706, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2014:PAb

[Ano14v] Anonymous. Pages 65–194 (August 2014). Journal of Econometrics, 181(2):??, August 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2014:PF

[Ano14w] Anonymous. Pages 707–830 (February 2014). Journal of Econometrics, 178(2):??, February 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2014:PAa

[Ano14x] Anonymous. Pages 99–178 (April 2014). Journal of Econometrics, 179(2):??, April 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2015:EBa

[Ano15a] Anonymous. Editorial Board. Journal of Econometrics, 184(1): ifc, January 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614002590.

Anonymous:2015:EBb

[Ano15b] Anonymous. Editorial Board. Journal of Econometrics, 184(2): ifc, February 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614002711.

Anonymous:2015:EBc

[Ano15c] Anonymous. Editorial Board. Journal of Econometrics, 185(1): ifc, March 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000032.

Anonymous:2015:EBd

[Ano15d] Anonymous. Editorial Board. Journal of Econometrics, 185(2): ifc, April 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000160.

Anonymous:2015:EBe

[Ano15e] Anonymous. Editorial Board. Journal of Econometrics, 186 (1):ifc, May 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000895.

Anonymous:2015:EBf

[Ano15f] Anonymous. Editorial Board. Journal of Econometrics, 186 (2):ifc, June 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615001402.

Anonymous:2015:EBg

[Ano15g] Anonymous. Editorial Board. Journal of Econometrics, 187 (1):ifc, July 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615001451.

Anonymous:2015:EBh

[Ano15h] Anonymous. Editorial Board. Journal of Econometrics, 187(2): ifc, August 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615001670.

Anonymous:2015:EBi

[Ano15i] Anonymous. Editorial Board. Journal of Econometrics, 188(2): ifc, October 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002171.

Anonymous:2015:EBj

[Ano15j] Anonymous. Editorial Board. Journal of Econometrics, 189(1): ifc, November 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002262.

Anonymous:2015:EBk

[Ano15k] Anonymous. Editorial Board. Journal of Econometrics, 189(2): ifc, December 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002419.

Anonymous:2015:IIS

[Ano15l] Anonymous. IFC: ID statment. Journal of Econometrics, 188(1): ifc, September 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615001979.

Anonymous:2015:PJa

[Ano15m] Anonymous. Pages 1–208 (January 2015). Journal of Econometrics, 184(1):??, January 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2015:PN

[Ano15n] Anonymous. Pages 1–244 (November 2015). Journal of Econometrics, 189(1):??, November 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2015:PMb

[Ano15o] Anonymous. Pages 1–276 (May 2015). Journal of Econometrics, 186(1):??, May 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2015:PMa

[Ano15p] Anonymous. Pages 1–304 (March 2015). Journal of Econometrics, 185(1):??, March 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2015:PS

[Ano15q] Anonymous. Pages 1–312 (September 2015). Journal of Econometrics, 188(1):??, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2015:PJb

[Ano15r] Anonymous. Pages 1–402 (July 2015). Journal of Econometrics, 187(1):??, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2015:PF

[Ano15s] Anonymous. Pages 209–464 (February 2015). Journal of Econometrics, 184(2):??, February 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2015:PA

[Ano15t] Anonymous. Pages 305–544 (April 2015). Journal of Econometrics, 185(2):??, April 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2015:ZA

[Ano15u] Anonymous. Zellner Award. Journal of Econometrics, 184(2): v, February 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614002887.

Anonymous:2016:DJA

[Ano16a] Anonymous. 2015 Dennis J. Aigner Award. Journal of Econometrics, 190(1):iv, January 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615002717.

Anonymous:2016:EBa

[Ano16b] Anonymous. Editorial Board. Journal of Econometrics, 190(2): ifc, February 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615003073.

Anonymous:2016:EBb

[Ano16c] Anonymous. Editorial Board. Journal of Econometrics, 191(1): ifc, March 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002924.

Anonymous:2016:EBc

[Ano16d] Anonymous. Editorial Board. Journal of Econometrics, 191(2): ifc, April 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300276.

Anonymous:2016:EBd

[Ano16e] Anonymous. Editorial Board. Journal of Econometrics, 192 (1):ifc, May 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300367.

Anonymous:2016:EBe

[Ano16f] Anonymous. Editorial Board. Journal of Econometrics, 192 (2):ifc, June 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300483.

Anonymous:2016:EBf

[Ano16g] Anonymous. Editorial Board. Journal of Econometrics, 193 (1):ifc, July 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300811.

Anonymous:2016:EBg

[Ano16h] Anonymous. Editorial Board. Journal of Econometrics, 193(2): ifc, August 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616301117.

Anonymous:2016:EBh

[Ano16i] Anonymous. Editorial Board. Journal of Econometrics, 194(1): ifc, September 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616301245.

Anonymous:2016:EBi

[Ano16j] Anonymous. Editorial Board. Journal of Econometrics, 194(2): ifc, October 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616301397.

Anonymous:2016:EBj

[Ano16k] Anonymous. Editorial Board. Journal of Econometrics, 195(1): ifc, November 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616301531.

Anonymous:2016:EBk

[Ano16l] Anonymous. Editorial Board. Journal of Econometrics, 195(2): ifc, December 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616301841.

Anonymous:2016:IIS

[Ano16m] Anonymous. IFC: ID statement. Journal of Econometrics, 190(1):ifc, January 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615002651.

Anonymous:2016:PN

[Ano16n] Anonymous. Pages 1–168 (November 2016). Journal of Econometrics, 195(1):??, November 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2016:PS

[Ano16o] Anonymous. Pages 1–202 (September 2016). Journal of Econometrics, 194(1):??, September 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2016:PJa

[Ano16p] Anonymous. Pages 1–208 (January 2016). Journal of Econometrics, 190(1):??, January 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2016:PMa

[Ano16q] Anonymous. Pages 1–272 (March 2016). Journal of Econometrics, 191(1):??, March 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2016:PJb

[Ano16r] Anonymous. Pages 1–290 (July 2016). Journal of Econometrics, 193(1):??, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2016:PMb

[Ano16s] Anonymous. Pages 1–328 (May 2016). Journal of Econometrics, 192(1):??, May 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2016:PD

[Ano16t] Anonymous. Pages 169–270 (December 2016). Journal of Econometrics, 195(2):??, December 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2017:AAZ

[Ano17a] Anonymous. Announcement: 2016 Arnold Zellner Award. Journal of Econometrics, 196(2):iv, February 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616302275.

Anonymous:2017:EBa

[Ano17b] Anonymous. Editorial Board. Journal of Econometrics, 196(1): ifc, January 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407616302056.

Anonymous:2017:EBb

[Ano17c] Anonymous. Editorial Board. Journal of Econometrics, 196(2): ifc, February 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616302214.

Anonymous:2017:EBc

[Ano17d] Anonymous. Editorial Board. Journal of Econometrics, 197(1): ifc, March 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300052.

Anonymous:2017:EBd

[Ano17e] Anonymous. Editorial Board. Journal of Econometrics, 197(2): ifc, April 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761730026X.

Anonymous:2017:EBe

[Ano17f] Anonymous. Editorial Board. Journal of Econometrics, 198 (1):ifc, May 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300362.

Anonymous:2017:EBf

[Ano17g] Anonymous. Editorial Board. Journal of Econometrics, 198 (2):ifc, June 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300490.

Anonymous:2017:EBg

[Ano17h] Anonymous. Editorial Board. Journal of Econometrics, 199 (1):ifc, July 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761730057X.

Anonymous:2017:EBh

[Ano17i] Anonymous. Editorial Board. Journal of Econometrics, 199(2): ifc, August 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617301070.

Anonymous:2017:EBi

[Ano17j] Anonymous. Editorial Board. Journal of Econometrics, 200(1): ifc, September 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617301215.

Anonymous:2017:EBj

[Ano17k] Anonymous. Editorial Board. Journal of Econometrics, 200(2): ifc, October 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761730132X.

Anonymous:2017:EBk

[Ano17l] Anonymous. Editorial Board. Journal of Econometrics, 201(1): ifc, November 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617301823.

Anonymous:2017:EBl

[Ano17m] Anonymous. Editorial Board. Journal of Econometrics, 201(2): ifc, December 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617301963.

Anonymous:2017:PS

[Ano17n] Anonymous. Pages 1–150 (September 2017). Journal of Econometrics, 200(1):??, September 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2017:PMa

[Ano17o] Anonymous. Pages 1–172 (March 2017). Journal of Econometrics, 197(1):??, March 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2017:PN

[Ano17p] Anonymous. Pages 1–172 (November 2017). Journal of Econometrics, 201(1):??, November 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2017:PMb

[Ano17q] Anonymous. Pages 1–188 (May 2017). Journal of Econometrics, 198(1):??, May 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2017:PJa

[Ano17r] Anonymous. Pages 1–232 (January 2017). Journal of Econometrics, 196(1):??, January 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2017:PJc

[Ano17s] Anonymous. Pages 1–92 (July 2017). Journal of Econometrics, 199(1):??, July 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2017:PA

[Ano17t] Anonymous. Pages 173–382 (April 2017). Journal of Econometrics, 197(2):??, April 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2017:PJb

[Ano17u] Anonymous. Pages 189–296 (June 2017). Journal of Econometrics, 198(2):??, June 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2017:PF

[Ano17v] Anonymous. Pages 233–378 (February 2017). Journal of Econometrics, 196(2):??, February 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2018:Aa

[Ano18a] Anonymous. Announcement. Journal of Econometrics, 202(2): 308, February 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617302300.

Anonymous:2018:Ab

[Ano18b] Anonymous. Announcement. Journal of Econometrics, 202(2): 309, February 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617302312.

Anonymous:2018:EBa

[Ano18c] Anonymous. Editorial Board. Journal of Econometrics, 202(1): ifc, January 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407617302130.

Anonymous:2018:EBb

[Ano18d] Anonymous. Editorial Board. Journal of Econometrics, 202(2): ii, February 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617302506.

Anonymous:2018:EBc

[Ano18e] Anonymous. Editorial Board. Journal of Econometrics, 203(1): ii, March 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618300137.

Anonymous:2018:EBd

[Ano18f] Anonymous. Editorial Board. Journal of Econometrics, 203 (2):ii, April 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618300241.

Anonymous:2018:EBe

[Ano18g] Anonymous. Editorial Board. Journal of Econometrics, 204(1):ii, May 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300368.

Anonymous:2018:EBf

[Ano18h] Anonymous. Editorial Board. Journal of Econometrics, 204 (2):ii, June 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618300642.

Anonymous:2018:EBg

[Ano18i] Anonymous. Editorial Board. Journal of Econometrics, 205(1):ii, July 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300745.

Anonymous:2018:EBh

[Ano18j] Anonymous. Editorial Board. Journal of Econometrics, 205(2): ii, August 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618300861.

Anonymous:2018:EBi

[Ano18k] Anonymous. Editorial Board. Journal of Econometrics, 206(1): ii, September 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301374.

Anonymous:2018:EBj

[Ano18l] Anonymous. Editorial Board. Journal of Econometrics, 206(2): ii, October 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301477.

Anonymous:2018:EBk

[Ano18m] Anonymous. Editorial Board. Journal of Econometrics, 207(1): ii, November 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761830160X.

Anonymous:2018:EBl

[Ano18n] Anonymous. Editorial Board. Journal of Econometrics, 207(2): ii, December 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301891.

Anonymous:2018:PJa

[Ano18o] Anonymous. Pages 1–124 (January 2018). Journal of Econometrics, 202(1):??, January 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2018:PMb

[Ano18p] Anonymous. Pages 1–130 (May 2018). Journal of Econometrics, 204(1):??, May 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2018:PMa

[Ano18q] Anonymous. Pages 1–186 (March 2018). Journal of Econometrics, 203(1):??, March 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2018:PN

[Ano18r] Anonymous. Pages 1–260 (November 2018). Journal of Econometrics, 207(1):??, November 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2018:PS

[Ano18s] Anonymous. Pages 1–278 (September 2018). Journal of Econometrics, 206(1):??, September 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2018:PF

[Ano18t] Anonymous. Pages 125–310 (February 2018). Journal of Econometrics, 202(2):??, February 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2018:PJb

[Ano18u] Anonymous. Pages 131–320 (June 2018). Journal of Econometrics, 204(2):??, June 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2018:PAa

[Ano18v] Anonymous. Pages 187–378 (April 2018). Journal of Econometrics, 203(2):??, April 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2018:PD

[Ano18w] Anonymous. Pages 261–448 (December 2018). Journal of Econometrics, 207(2):??, December 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2018:PAb

[Ano18x] Anonymous. Pages 303–526 (August 2018). Journal of Econometrics, 205(2):??, August 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2019:A

[Ano19a] Anonymous. Announcement. Journal of Econometrics, 208(2): 323, February 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302616.

Anonymous:2019:EBa

[Ano19b] Anonymous. Editorial Board. Journal of Econometrics, 208(1): ii, January 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302240.

Anonymous:2019:EBb

[Ano19c] Anonymous. Editorial Board. Journal of Econometrics, 208(2): ii, February 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302653.

Anonymous:2019:EBc

[Ano19d] Anonymous. Editorial Board. Journal of Econometrics, 209(1): ii, March 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300132.

Anonymous:2019:EBd

[Ano19e] Anonymous. Editorial Board. Journal of Econometrics, 209 (2):ii, April 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300326.

Anonymous:2019:EBe

[Ano19f] Anonymous. Editorial Board. Journal of Econometrics, 210(1):ii, May 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300430.

Anonymous:2019:EBf

[Ano19g] Anonymous. Editorial Board. Journal of Econometrics, 211(1):ii, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301253.

Anonymous:2019:EBg

[Ano19h] Anonymous. Editorial Board. Journal of Econometrics, 211(2): ii, August 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619301320.

Anonymous:2019:EBh

[Ano19i] Anonymous. Editorial Board. Journal of Econometrics, 212(1): ii, September 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619301538.

Anonymous:2019:EBi

[Ano19j] Anonymous. Editorial Board. Journal of Econometrics, 212(2): ii, October 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619301769.

Anonymous:2019:EBj

[Ano19k] Anonymous. Editorial Board. Journal of Econometrics, 213(1): ii, November 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619301939.

Anonymous:2019:EBk

[Ano19l] Anonymous. Editorial Board. Journal of Econometrics, 213(2): ii, December 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619302167.

Anonymous:2019:PM

[Ano19m] Anonymous. Pages 1–144 (March 2019). Journal of Econometrics, 209(1):??, March 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2019:PAa

[Ano19n] Anonymous. Pages 145-406 (April 2019). Journal of Econometrics, 209(2):??, April 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2019:PD

[Ano19o] Anonymous. Pages 297–632 (December 2019). Journal of Econometrics, 213(2):??, December 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2019:PAb

[Ano19p] Anonymous. Pages 319-618 (August 2019). Journal of Econometrics, 211(2):??, August 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2019:PF

[Ano19q] Anonymous. Pages 323–658 (February 2019). Journal of Econometrics, 208(2):??, February 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Anonymous:2019:PO

[Ano19r] Anonymous. Pages 359–678 (October 2019). Journal of Econometrics, 212(2):??, October 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

Amsler:2016:ESF

[APS16] Christine Amsler, Artem Prokhorov, and Peter Schmidt. Endogeneity in stochastic frontier models. Journal of Econometrics, 190(2):280–288, February 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615001827.

Amsler:2017:EEV

[APS17] Christine Amsler, Artem Prokhorov, and Peter Schmidt. Endogenous environmental variables in stochastic frontier models. Journal of Econometrics, 199(2):131–140, August 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300672.

Atkinson:2018:SIE

[APT18] Scott E. Atkinson, Daniel Primont, and Mike G. Tsionas. Statistical inference in efficient production with bad inputs and outputs using latent prices and optimal directions. Journal of Econometrics, 204(2):131–146, June 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300162.

Antoine:2012:EMD

[AR12] Bertille Antoine and Eric Renault. Efficient minimum distance estimation with multiple rates of convergence. Journal of Econometrics, 170(2):350–367, October 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001224.

Arbues:2013:DMO

[Arb13] Ignacio Arbués. Determining the MSE-optimal cross section to forecast. Journal of Econometrics, 175(2):61–70, August 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000602.

Armstrong:2014:WKS

[Arm14] Timothy B. Armstrong. Weighted KS statistics for inference on conditional moment inequalities. Journal of Econometrics, 181(2):92–116, August 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614000888.

Armstrong:2015:AEI

[Arm15] Timothy B. Armstrong. Asymptotically exact inference in conditional moment inequality models. Journal of Econometrics, 186 (1):51–65, May 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000111.

Armstrong:2018:CTS

[Arm18] Timothy B. Armstrong. On the choice of test statistic for conditional moment inequalities. Journal of Econometrics, 203(2): 241–255, April 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617302373.

Amengual:2010:CMV

[AS10] Dante Amengual and Enrique Sentana. A comparison of meanvariance efficiency tests. Journal of Econometrics, 154(1):16– 34, January 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760900147X.

Al-Sadoon:2014:GLR

[AS14a] Majid M. Al-Sadoon. Geometric and long run aspects of Granger causality. Journal of Econometrics, 178(Part 3):558–568, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001693.

Andrews:2014:NIB

[AS14b] Donald W. K. Andrews and Xiaoxia Shi. Nonparametric inference based on conditional moment inequalities. Journal of Econometrics, 179(1):31–45, March 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613002091.

Al-Sadoon:2017:UTT

[AS17a] Majid M. Al-Sadoon. A unifying theory of tests of rank. Journal of Econometrics, 199(1):49–62, July 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300453.

Andrews:2017:IBM

[AS17b] Donald W. K. Andrews and Xiaoxia Shi. Inference based on many conditional moment inequalities. Journal of Econometrics, 196(2):275–287, February 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616301774.

Ait-Sahalia:2015:MBE

[ASAM15] Yacine Aït-Sahalia, Dante Amengual, and Elena Manresa. Market-based estimation of stochastic volatility models. Journal of Econometrics, 187(2):418–435, August 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000536.

Ait-Sahalia:2012:TJN

[ASJL12] Yacine Aït-Sahalia, Jean Jacod, and Jia Li. Testing for jumps in noisy high frequency data. Journal of Econometrics, 168(2): 207–222, June 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002764.

Andrikopoulos:2016:FDJ

[ASK16] Andreas Andrikopoulos, Aristeidis Samitas, and Konstantinos Kostaris. Four decades of the journal of econometrics: Coauthorship patterns and networks. Journal of Econometrics, 195(1):23– 32, November 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761630121X.

Ait-Sahalia:2019:AIH

[ASLN19] Yacine Aït-Sahalia, Andrew W. Lo, and Whitney K. Newey. Annals issue in honor of Jerry A. Hausman: Editors’ introduction. Journal of Econometrics, 211(1):1–3, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761830229X.

Ait-Sahalia:2014:MEE

[ASLP14] Yacine Aït-Sahalia, Roger J. A. Laeven, and Loriana Pelizzon. Mutual excitation in Eurozone sovereign CDS. Journal of Econometrics, 183(2):151–167, December 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001092.

Ait-Sahalia:2011:UHF

[ASMZ11] Yacine Aït-Sahalia, Per A. Mykland, and Lan Zhang. Ultra high frequency volatility estimation with dependent microstructure noise. Journal of Econometrics, 160(1):160–175, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000709.

Ait-Sahalia:2012:SBS

[ASP12] Yacine Aït-Sahalia and Joon Y. Park. Stationarity-based specification tests for diffusions when the process is nonstationary. Journal of Econometrics, 169(2):279–292, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000401.

Ait-Sahalia:2016:BSA

[ASP16] Yacine Aït-Sahalia and Joon Y. Park. Bandwidth selection and asymptotic properties of local nonparametric estimators in possibly nonstationary continuous-time models. Journal of Econometrics, 192(1):119–138, May 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002730.

Ait-Sahalia:2016:ICA

[ASX16] Yacine Aït-Sahalia and Dacheng Xiu. Increased correlation among asset classes: Are volatility or jumps to blame, or both? Journal of Econometrics, 194(2):205–219, October 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300902.

Ait-Sahalia:2017:UPC

[ASX17] Yacine Aït-Sahalia and Dacheng Xiu. Using principal component analysis to estimate a high dimensional factor model with high-frequency data. Journal of Econometrics, 201(2):384–399, December 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617301677.

Ait-Sahalia:2019:HTP

[ASX19] Yacine Aït-Sahalia and Dacheng Xiu. A Hausman test for the presence of market microstructure noise in high frequency data. Journal of Econometrics, 211(1):176–205, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761830246X.

Amano:2011:CVM

[AT11] Tomoyuki Amano and Masanobu Taniguchi. Control variate method for stationary processes. Journal of Econometrics, 165(1):20–29, November 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611000959.

Amado:2013:MVV

[AT13] Cristina Amado and Timo Teräsvirta. Modelling volatility by variance decomposition. Journal of Econometrics, 175(2):142– 153, August 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761300064X.

Alem:2014:EFI

[AT14] Mauro Alem and Robert M. Townsend. An evaluation of financial institutions: Impact on consumption and investment using panel data and the theory of risk-bearing. Journal of Econometrics, 183(1):91–103, November 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001523.

Atkinson:2016:DDF

[AT16] Scott E. Atkinson and Mike G. Tsionas. Directional distance functions: Optimal endogenous directions. Journal of Econometrics, 190(2):301–314, February 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761500175X.

Arvanitis:2017:TPM

[AT17] Stelios Arvanitis and Nikolas Topaloglou. Testing for prospect and Markowitz stochastic dominance efficiency. Journal of Econometrics, 198(2):253–270, June 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407617300210.

Abowd:2011:NEG

[AV11] John M. Abowd and Lars Vilhuber. National estimates of gross employment and job flows from the Quarterly Workforce Indicators with demographic and industry detail. Journal of Econometrics, 161(1):82–99, March 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001880.

Andreou:2015:RBR

[AW15] Elena Andreou and Bas J. M. Werker. Residual-based rank specification tests for AR-GARCH type models. Journal of Econometrics, 185(2):305–331, April 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614002656.

Amengual:2018:RPU

[AX18] Dante Amengual and Dacheng Xiu. Resolution of policy uncertainty and sudden declines in volatility. Journal of Econometrics, 203(2):297–315, April 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617302439.

Buschena:2011:ESM

[BA11] David E. Buschena and Joseph A. Atwood. Evaluation of similarity models for expected utility violations. Journal of Econometrics, 162(1):105–113, May 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002668.

Bai:2010:CBM

[Bai10] Jushan Bai. Common breaks in means and variances for panel data. Journal of Econometrics, 157(1):78–92, July 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002735.

Bassett:2019:RMS

[Bas19] Gib Bassett. Review of median stable distributions and Schröder’s equation. Journal of Econometrics, 213(1):289–295, November 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300715.

Berghaus:2014:NTT

[BB14] Betina Berghaus and Axel Bücher. Nonparametric tests for tail monotonicity. Journal of Econometrics, 180(2):117–126, June 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000451.

Beffy:2019:LST

[BBB^{+}19] Magali Beffy, Richard Blundell, Antoine Bozio, Guy Laroque, and Maxime Tô. Labour supply and taxation with restricted choices. Journal of Econometrics, 211(1):16–46, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302331.

Bianchi:2019:MSR

[BBCG19] Daniele Bianchi, Monica Billio, Roberto Casarin, and Massimo Guidolin. Modeling systemic risk with Markov Switching Graphical SUR models. Journal of Econometrics, 210(1):58–74, May 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302069.

Baltagi:2018:RLS

[BBCL18] Badi H. Baltagi, Georges Bresson, Anoop Chaturvedi, and Guy Lacroix. Robust linear static panel data models using ϵcontamination. Journal of Econometrics, 202(1):108–123, January 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407617301446.

Basturk:2019:FDC

[BBG^{+}19] N. Bastürk, A. Borowska, S. Grassi, L. Hoogerheide, and H. K. van Dijk. Forecast density combinations of dynamic models and data driven portfolio strategies. Journal of Econometrics, 210(1): 170–186, May 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302124.

Barigozzi:2014:DSI

[BBGV14] Matteo Barigozzi, Christian Brownlees, Giampiero M. Gallo, and David Veredas. Disentangling systematic and idiosyncratic dynamics in panels of volatility measures. Journal of Econometrics, 182(2):364–384, October 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001390.

Bergamelli:2019:CPV

[BBKU19] Michele Bergamelli, Annamaria Bianchi, Lynda Khalaf, and Giovanni Urga. Combining p-values to test for multiple structural breaks in cointegrated regressions. Journal of Econometrics, 211(2):461–482, August 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761930051X.

Bartolucci:2015:TTI

[BBP15] Francesco Bartolucci, Federico Belotti, and Franco Peracchi. Testing for time-invariant unobserved heterogeneity in generalized linear models for panel data. Journal of Econometrics, 184(1):111–123, January 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001833.

Bikbov:2010:NAM

[BC10] Ruslan Bikbov and Mikhail Chernov. No-arbitrage macroeconomic determinants of the yield curve. Journal of Econometrics, 159(1):166–182, November 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610001296.

Barnett:2011:HBM

[BC11] William A. Barnett and Marcelle Chauvet. How better monetary statistics could have signaled the financial crisis. Journal of Econometrics, 161(1):6–23, March 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761000182X.

Bikbov:2013:MPR

[BC13] Ruslan Bikbov and Mikhail Chernov. Monetary policy regimes and the term structure of interest rates. Journal of Econometrics, 174(1):27–43, May 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761300016X.

Battistin:2014:TEE

[BC14a] Erich Battistin and Andrew Chesher. Treatment effect estimation with covariate measurement error. Journal of Econometrics, 178(2):707–715, February 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761300225X.

Browning:2014:DBO

[BC14b] Martin Browning and Jesus M. Carro. Dynamic binary outcome models with maximal heterogeneity. Journal of Econometrics, 178(2):805–823, February 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613002352.

Bekker:2015:JIV

[BC15] Paul A. Bekker and Federico Crudu. Jackknife instrumental variable estimation with heteroskedasticity. Journal of Econometrics, 185(2):332–342, April 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761400267X.

Baldovin:2015:OPN

[BCC^{+}15] Fulvio Baldovin, Massimiliano Caporin, Michele Caraglio, Attilio L. Stella, and Marco Zamparo. Option pricing with non-Gaussian scaling and infinite-state switching volatility. Journal of Econometrics, 187(2):486–497, August 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000585.

Belloni:2019:CQP

[BCCFV19] Alexandre Belloni, Victor Chernozhukov, Denis Chetverikov, and Iván Fernández-Val. Conditional quantile processes based on series or many regressors. Journal of Econometrics, 213(1):4–29, November 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300582.

Belloni:2015:SNA

[BCCK15] Alexandre Belloni, Victor Chernozhukov, Denis Chetverikov, and Kengo Kato. Some new asymptotic theory for least squares series: Pointwise and uniform results. Journal of Econometrics, 186(2): 345–366, June 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761500038X.

Brock:2012:OIT

[BCDN12] William A. Brock, Jane Cooley, Steven N. Durlauf, and Salvador Navarro. On the observational implications of taste-based discrimination in racial profiling. Journal of Econometrics, 166(1): 66–78, January 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001199.

Berkowitz:2012:VIR

[BCF12] Daniel Berkowitz, Mehmet Caner, and Ying Fang. The validity of instruments revisited. Journal of Econometrics, 166(2):255– 266, February 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002193.

Benatia:2017:FLR

[BCF17] David Benatia, Marine Carrasco, and Jean-Pierre Florens. Functional linear regression with functional response. Journal of Econometrics, 201(2):269–291, December 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617301586.

Barigozzi:2018:SMC

[BCF18] Matteo Barigozzi, Haeran Cho, and Piotr Fryzlewicz. Simultaneous multiple change-point and factor analysis for highdimensional time series. Journal of Econometrics, 206(1): 187–225, September 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618300915.

Babii:2019:CRM

[BCG19] Andrii Babii, Xi Chen, and Eric Ghysels. Commercial and residential mortgage defaults: Spatial dependence with frailty. Journal of Econometrics, 212(1):47–77, September 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300752.

Bester:2011:IDD

[BCH11] C. Alan Bester, Timothy G. Conley, and Christian B. Hansen. Inference with dependent data using cluster covariance estimators. Journal of Econometrics, 165(2):137–151, December 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000431.

Blake:2014:IIE

[BCIT14] David Blake, Tristan Caulfield, Christos Ioannidis, and Ian Tonks. Improved inference in the evaluation of mutual fund performance using panel bootstrap methods. Journal of Econometrics, 183(2):202–210, December 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001134.

Bassetti:2014:BPD

[BCL14] Federico Bassetti, Roberto Casarin, and Fabrizio Leisen. Betaproduct dependent Pitman–Yor processes for Bayesian inference. Journal of Econometrics, 180(1):49–72, May 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000190.

Barnett:2016:RTN

[BCLL16] William A. Barnett, Marcelle Chauvet, and Danilo Leiva-Leon. Real-time nowcasting of nominal GDP with structural breaks. Journal of Econometrics, 191(2):312–324, April 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002808.

Boldea:2019:BSC

[BCMH19] Otilia Boldea, Adriana Cornea-Madeira, and Alastair R. Hall. Bootstrapping structural change tests. Journal of Econometrics, 213(2):359–397, December 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619301472.

Busch:2011:RIV

[BCN11] Thomas Busch, Bent Jesper Christensen, and Morten Ørregaard Nielsen. The role of implied volatility in forecasting future realized volatility and jumps in foreign exchange, stock, and bond markets. Journal of Econometrics, 160(1):48–57, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000564.

Baek:2015:TLU

[BCP15] Yae In Baek, Jin Seo Cho, and Peter C. B. Phillips. Testing linearity using power transforms of regressors. Journal of Econometrics, 187(1):376–384, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615001220.

Billio:2019:BNS

[BCR19] Monica Billio, Roberto Casarin, and Luca Rossini. Bayesian nonparametric sparse VAR models. Journal of Econometrics, 212(1):97–115, September 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619300776.

Boswijk:2016:ICI

[BCRT16] H. Peter Boswijk, Giuseppe Cavaliere, Anders Rahbek, and A. M. Robert Taylor. Inference on co-integration parameters in heteroskedastic vector autoregressions. Journal of Econometrics, 192(1):64–85, May 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615002766.

Billio:2013:TVC

[BCRvD13] Monica Billio, Roberto Casarin, Francesco Ravazzolo, and Herman K. van Dijk. Time-varying combinations of predictive densities using nonlinear filtering. Journal of Econometrics, 177(2):213–232, December 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613000869.

Bugni:2015:STP

[BCS15] Federico A. Bugni, Ivan A. Canay, and Xiaoxia Shi. Specification tests for partially identified models defined by moment inequalities. Journal of Econometrics, 185(1):259–282, March 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614002577.

Bhattacharya:2012:IWM

[BD12] Debopam Bhattacharya and Pascaline Dupas. Inferring welfare maximizing treatment assignment under budget constraints. Journal of Econometrics, 167(1):168–196, March 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002697.

Breitung:2015:IVV

[BD15] Jörg Breitung and Matei Demetrescu. Instrumental variable and variable addition based inference in predictive regressions. Journal of Econometrics, 187(1):358–375, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000457.

Blasques:2018:PII

[BD18] Francisco Blasques and Artem Duplinskiy. Penalized indirect inference. Journal of Econometrics, 205(1):34–54, July 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300435.

Black:2014:MMS

[BdCK^{+}14] Bernard Black, Antonio Gledson de Carvalho, Vikramaditya Khanna, Woochan Kim, and Burcin Yurtoglu. Methods for multicountry studies of corporate governance: Evidence from the BRIKT countries. Journal of Econometrics, 183(2):230–240, December 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761400116X.

Barrett:2016:CTP

[BDH16] Garry F. Barrett, Stephen G. Donald, and Yu-Chin Hsu. Consistent tests for poverty dominance relations. Journal of Econometrics, 191(2):360–373, April 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002833.

Beaulieu:2014:ECS

[BDK14] Marie-Claude Beaulieu, Jean-Marie Dufour, and Lynda Khalaf. Exact confidence sets and goodness-of-fit methods for stable distributions. Journal of Econometrics, 181(1):3–14, July 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000323.

Bruins:2018:GII

[BDKS18] Marianne Bruins, James A. Duffy, Michael P. Keane, and Anthony A. Smith. Generalized indirect inference for discrete choice models. Journal of Econometrics, 205(1):177–203, July 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300496.

Barnett:2016:IME

[BDM16] W. A. Barnett, W. E. Diewert, and E. Maasoumi. Innovations in measurement in economics and econometrics: an overview. Journal of Econometrics, 191(2):273–275, April 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615003048.

Bauwens:2014:MLM

[BDR14] Luc Bauwens, Arnaud Dufays, and Jeroen V. K. Rombouts. Marginal likelihood for Markov-switching and change-point GARCH models. Journal of Econometrics, 178(Part 3):508– 522, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S030440761300167X.

Battistin:2014:MSM

[BDS14] Erich Battistin, Michele De Nadai, and Barbara Sianesi. Misreported schooling, multiple measures and returns to educational qualifications. Journal of Econometrics, 181(2):136–150, August 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000414.

Battistin:2017:CRA

[BDV17] Erich Battistin, Michele De Nadai, and Daniela Vuri. Counting rotten apples: Student achievement and score manipulation in Italian elementary schools. Journal of Econometrics, 200(2):344– 362, October 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761730101X.

Barnett:2011:IMT

[BDZ11] William A. Barnett, W. Erwin Diewert, and Arnold Zellner. Introduction to measurement with theory. Journal of Econometrics, 161(1):1–5, March 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610001818.

Blazsek:2010:KSU

[BE10] Szabolcs Blazsek and Alvaro Escribano. Knowledge spillovers in US patents: a dynamic patent intensity model with secret common innovation factors. Journal of Econometrics, 159(1):14– 32, November 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610001107.

Breitung:2011:TSB

[BE11] Jörg Breitung and Sandra Eickmeier. Testing for structural breaks in dynamic factor models. Journal of Econometrics, 163 (1):71–84, July 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610002125.

Blazsek:2016:PPR

[BE16] Szabolcs Blazsek and Alvaro Escribano. Patent propensity, R&D and market competition: Dynamic spillovers of innovation leaders and followers. Journal of Econometrics, 191(1):145– 163, March 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002559.

Bekaert:2015:BEG

[BEE15] Geert Bekaert, Eric Engstrom, and Andrey Ermolov. Bad environments, good environments: a non-Gaussian asymmetric volatility model. Journal of Econometrics, 186(1):258–275, May 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761400178X.

Baltagi:2017:DFL

[BEK17] Badi H. Baltagi, Peter H. Egger, and Michaela Kesina. Determinants of firm-level domestic sales and exports with spillovers: Evidence from China. Journal of Econometrics, 199(2):184– 201, August 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300714.

Baltagi:2012:LMT

[BFK12] Badi H. Baltagi, Qu Feng, and Chihwa Kao. A Lagrange multiplier test for cross-sectional dependence in a fixed effects panel data model. Journal of Econometrics, 170(1):164–177, September 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761200098X.

Baltagi:2016:EHP

[BFK16] Badi H. Baltagi, Qu Feng, and Chihwa Kao. Estimation of heterogeneous panels with structural breaks. Journal of Econometrics, 191(1):176–195, March 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002353.

Bilias:2019:ECC

[BFS19a] Yannis Bilias, Kostas Florios, and Spyros Skouras. Exact computation of censored least absolute deviations estimator. Journal of Econometrics, 212(2):584–606, October 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761930140X.

Bitto:2019:AST

[BFS19b] Angela Bitto and Sylvia Frühwirth-Schnatter. Achieving shrinkage in a time-varying parameter model framework. Journal of Econometrics, 210(1):75–97, May 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618302070.

Boswijk:2010:CHP

[BFvD10a] H. Peter Boswijk, Philip Hans Franses, and Dick van Dijk. Cointegration in a historical perspective. Journal of Econometrics, 158(1):156–159, September 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000679.

Boswijk:2010:TYC

[BFvD10b] H. Peter Boswijk, Philip Hans Franses, and Dick van Dijk. Twenty years of cointegration. Journal of Econometrics, 158 (1):1–2, September 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000436.

Blasques:2019:ASD

[BGK19] F. Blasques, P. Gorgi, and S. J. Koopman. Accelerating scoredriven time series models. Journal of Econometrics, 212(2):359– 376, October 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300557.

Bonomo:2015:LSR

[BGMT15] Marco Bonomo, René Garcia, Nour Meddahi, and Roméo Tédongap. The long and the short of the risk-return trade-off. Journal of Econometrics, 187(2):580–592, August 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000652.

Bollerslev:2011:DEV

[BGZ11] Tim Bollerslev, Michael Gibson, and Hao Zhou. Dynamic estimation of volatility risk premia and investor risk aversion from option-implied and realized volatilities. Journal of Econometrics, 160(1):235–245, January 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000758.

Boldea:2013:EIU

[BH13] Otilia Boldea and Alastair R. Hall. Estimation and inference in unstable nonlinear least squares models. Journal of Econometrics, 172(1):158–167, January 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612002205.

Bekaert:2014:VVP

[BH14a] Geert Bekaert and Marie Hoerova. The VIX, the variance premium and stock market volatility. Journal of Econometrics, 183(2):181–192, December 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001110.

Borovicka:2014:EMM

[BH14b] Jaroslav Borovicka and Lars Peter Hansen. Examining macroeconomic models through the lens of asset pricing. Journal of Econometrics, 183(1):67–90, November 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001511.

Bester:2016:GEE

[BH16a] C. Alan Bester and Christian B. Hansen. Grouped effects estimators in fixed effects models. Journal of Econometrics, 190(1):197– 208, January 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613002030.

Breitung:2016:IMT

[BH16b] Jörg Breitung and Helmut Herwartz. Innovations in multiple time series analysis. Journal of Econometrics, 192(2):329– 331, June 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300045.

Barigozzi:2017:GDF

[BH17] Matteo Barigozzi and Marc Hallin. Generalized dynamic factor models and volatilities: estimation and forecasting. Journal of Econometrics, 201(2):307–321, December 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617301628.

Bhattacharya:2013:ETP

[Bha13] Debopam Bhattacharya. Evaluating treatment protocols using data combination. Journal of Econometrics, 173(2):160– 174, April 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612002497.

Bhargava:2014:EIA

[Bha14a] Alok Bhargava. Editor’s introduction: Analysis of financial data. Journal of Econometrics, 183(2):147–149, December 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001055.

Bhargava:2014:FFM

[Bha14b] Alok Bhargava. Firms’ fundamentals, macroeconomic variables and quarterly stock prices in the US. Journal of Econometrics, 183(2):241–250, December 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001171.

Burda:2012:PMM

[BHH12] Martin Burda, Matthew Harding, and Jerry Hausman. A Poisson mixture model of discrete choice. Journal of Econometrics, 166(2):184–203, February 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611001643.

Broda:2013:SMG

[BHK^{+}13] Simon A. Broda, Markus Haas, Jochen Krause, Marc S. Paolella, and Sven C. Steude. Stable mixture GARCH models. Journal of Econometrics, 172(2):292–306, February 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001960.

Bennala:2012:PGR

[BHP12] Nezar Bennala, Marc Hallin, and Davy Paindaveine. Pseudo-Gaussian and rank-based optimal tests for random individual effects in large n small t panels. Journal of Econometrics, 170(1):50–67, September 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612000772.

Belongia:2014:BCA

[BI14] Michael T. Belongia and Peter N. Ireland. The Barnett critique after three decades: a new Keynesian analysis. Journal of Econometrics, 183(1):5–21, November 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761400147X.

Belotti:2018:CIF

[BI18] Federico Belotti and Giuseppe Ilardi. Consistent inference in fixed-effects stochastic frontier models. Journal of Econometrics, 202(2):161–177, February 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761730194X.

Bianchi:2016:MME

[Bia16] Francesco Bianchi. Methods for measuring expectations and uncertainty in Markov-switching models. Journal of Econometrics, 190(1):79–99, January 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615002249.

Boswijk:2015:ILR

[BJN15] H. Peter Boswijk, Michael Jansson, and Morten Ørregaard Nielsen. Improved likelihood ratio tests for cointegration rank in the VAR model. Journal of Econometrics, 184(1):97–110, January 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001869.

Bonhomme:2017:NEN

[BJR17] Stéphane Bonhomme, Koen Jochmans, and Jean-Marc Robin. Nonparametric estimation of non-exchangeable latent-variable models. Journal of Econometrics, 201(2):237–248, December 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301562.

Baltagi:2010:THS

[BJS10] Badi H. Baltagi, Byoung Cheol Jung, and Seuck Heun Song. Testing for heteroskedasticity and serial correlation in a random effects panel data model. Journal of Econometrics, 154(2):122– 124, February 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440760900164X.

Bruggemann:2016:IVC

[BJT16] Ralf Brüggemann, Carsten Jentsch, and Carsten Trenkler. Inference in VARs with conditional heteroskedasticity of unknown form. Journal of Econometrics, 191(1):69–85, March 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002547.

Bucher:2015:NTC

[BJW15] Axel Bücher, Stefan Jäschke, and Dominik Wied. Nonparametric tests for constant tail dependence with an application to energy and finance. Journal of Econometrics, 187(1):154–168, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000263.

Bernal:2010:QSE

[BK10] Raquel Bernal and Michael P. Keane. Quasi-structural estimation of a model of childcare choices and child cognitive ability production. Journal of Econometrics, 156(1):164–189, May 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002140.

Blasques:2016:SDS

[BKLS16] Francisco Blasques, Siem Jan Koopman, Andre Lucas, and Julia Schaumburg. Spillover dynamics for systemic risk measurement using spatial financial time series models. Journal of Econometrics, 195(2):211–223, December 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616301609.

Blundell:2014:BQD

[BKM14] Richard Blundell, Dennis Kristensen, and Rosa Matzkin. Bounding quantile demand functions using revealed preference inequalities. Journal of Econometrics, 179(2):112–127, April 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000177.

Blasques:2016:WML

[BKMZ16] F. Blasques, S. J. Koopman, M. Mallee, and Z. Zhang. Weighted maximum likelihood for dynamic factor analysis and forecasting with mixed frequency data. Journal of Econometrics, 193(2):405–417, August 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300732.

Baltagi:2017:IEL

[BKW17] Badi H. Baltagi, Chihwa Kao, and Fa Wang. Identification and estimation of a large factor model with structural instability. Journal of Econometrics, 197(1):87–100, March 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616302032.

Bolduc:2010:IRC

[BKY10] Denis Bolduc, Lynda Khalaf, and Clément Yélou. Identification robust confidence set methods for inference on parameter ratios with application to discrete choice models. Journal of Econometrics, 157(2):317–327, August 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761000028X.

Brockwell:2015:PLD

[BL15] Peter J. Brockwell and Alexander Lindner. Prediction of Lévydriven CARMA processes. Journal of Econometrics, 189(2):263– 271, December 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000998.

Bai:2016:EEA

[BL16] Jushan Bai and Yuan Liao. Efficient estimation of approximate factor models via penalized maximum likelihood. Journal of Econometrics, 191(1):1–18, March 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002535.

Bai:2017:IPD

[BL17] Jushan Bai and Yuan Liao. Inferences in panel data with interactive effects using large covariance matrices. Journal of Econometrics, 200(1):59–78, September 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300763.

Boudt:2017:PSI

[BLL^{+}17] Kris Boudt, Sébastien Laurent, Asger Lunde, Rogier Quaedvlieg, and Orimar Sauri. Positive semidefinite integrated covariance estimation, factorizations and asynchronicity. Journal of Econometrics, 196(2):347–367, February 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616301956.

Bai:2014:PTH

[BLO14] ChongEn Bai, Qi Li, and Min Ouyang. Property taxes and home prices: a tale of two cities. Journal of Econometrics, 180(1): 1–15, May 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613002674.

Blundell:2010:CMP

[Blu10] Richard Blundell. Comments on: Michael P. Keane ‘Structural vs. atheoretic approaches to econometrics’. Journal of Econometrics, 156(1):25–26, May 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609001961.

See [Kea10].

Boneva:2015:SMH

[BLV15] Lena Boneva, Oliver Linton, and Michael Vogt. A semiparametric model for heterogeneous panel data with fixed effects. Journal of Econometrics, 188(2):327–345, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761500072X.

Boswijk:2018:TSE

[BLY18] H. Peter Boswijk, Roger J. A. Laeven, and Xiye Yang. Testing for self-excitation in jumps. Journal of Econometrics, 203(2): 256–266, April 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617302397.

Bauer:2012:PRS

[BM12] Dietmar Bauer and Alex Maynard. Persistence-robust surpluslag Granger causality testing. Journal of Econometrics, 169 (2):293–300, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612000334.

Ben-Moshe:2017:IAP

[BMDL17] Dan Ben-Moshe, Xavier D’Haultfœuille, and Arthur Lewbel. Identification of additive and polynomial models of mismeasured regressors without instruments. Journal of Econometrics, 200(2):207–222, October 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617300921.

Beresteanu:2012:PIU

[BMM12] Arie Beresteanu, Ilya Molchanov, and Francesca Molinari. Partial identification using random set theory. Journal of Econometrics, 166(1):17–32, January 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611001163.

Bollerslev:2019:HDM

[BMN19] Tim Bollerslev, Nour Meddahi, and Serge Nyawa. Highdimensional multivariate realized volatility estimation. Journal of Econometrics, 212(1):116–136, September 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300788.

Breunig:2018:NEC

[BMS18] Christoph Breunig, Enno Mammen, and Anna Simoni. Nonparametric estimation in case of endogenous selection. Journal of Econometrics, 202(2):268–285, February 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617302221.

Bartolucci:2012:PCM

[BN12] Francesco Bartolucci and Valentina Nigro. Pseudo conditional maximum likelihood estimation of the dynamic logit model for binary panel data. Journal of Econometrics, 170(1):102–116, September 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000954.

Bai:2013:PCE

[BN13] Jushan Bai and Serena Ng. Principal components estimation and identification of static factors. Journal of Econometrics, 176(1):18–29, September 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613000651.

Bai:2019:RRE

[BN19a] Jushan Bai and Serena Ng. Rank regularized estimation of approximate factor models. Journal of Econometrics, 212(1):78–96, September 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300764.

Boot:2019:FUR

[BN19b] Tom Boot and Didier Nibbering. Forecasting using random subspace methods. Journal of Econometrics, 209(2):391–406, April 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300235.

Barndorff-Nielsen:2011:MRK

[BNHLS11a] Ole E. Barndorff-Nielsen, Peter Reinhard Hansen, Asger Lunde, and Neil Shephard. Multivariate realised kernels: Consistent positive semi-definite estimators of the covariation of equity prices with noise and non-synchronous trading. Journal of Econometrics, 162(2):149–169, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611000029.

Barndorff-Nielsen:2011:SRK

[BNHLS11b] Ole E. Barndorff-Nielsen, Peter Reinhard Hansen, Asger Lunde, and Neil Shephard. Subsampling realised kernels. Journal of Econometrics, 160(1):204–219, January 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000734.

Bibinger:2019:EDL

[BNW19] Markus Bibinger, Christopher Neely, and Lars Winkelmann. Estimation of the discontinuous leverage effect: Evidence from the NASDAQ order book. Journal of Econometrics, 209(2):158– 184, April 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300016.

Bognanni:2022:CLB

[Bog22] Mark Bognanni. Comment on “Large Bayesian vector autoregressions with stochastic volatility and non-conjugate priors”. Journal of Econometrics, 227(2):498–505, April 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407621002554. See [CCM19].

Bondarenko:2014:VTM

[Bon14] Oleg Bondarenko. Variance trading and market price of variance risk. Journal of Econometrics, 180(1):81–97, May 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000207.

Bates:2013:CFE

[BPMSW13] Brandon J. Bates, Mikkel Plagborg-Møller, James H. Stock, and Mark W. Watson. Consistent factor estimation in dynamic factor models with structural instability. Journal of Econometrics, 177(2):289–304, December 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613000912.

Bollerslev:2016:EES

[BPQ16] Tim Bollerslev, Andrew J. Patton, and Rogier Quaedvlieg. Exploiting the errors: a simple approach for improved volatility forecasting. Journal of Econometrics, 192(1):1–18, May 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002584.

Bollerslev:2018:MFR

[BPQ18] Tim Bollerslev, Andrew J. Patton, and Rogier Quaedvlieg. Modeling and forecasting (un)reliable realized covariances for more reliable financial decisions. Journal of Econometrics, 207(1):71– 91, November 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301180.

Bailey:2019:MTA

[BPS19] Natalia Bailey, M. Hashem Pesaran, and L. Vanessa Smith. A multiple testing approach to the regularisation of large sample correlation matrices. Journal of Econometrics, 208(2):507–534, February 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302008.

Bandi:2019:SP

[BPTT19] F. M. Bandi, B. Perron, A. Tamoni, and C. Tebaldi. The scale of predictability. Journal of Econometrics, 208(1):120–140, January 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301738.

Bandi:2011:MMN

[BR11] Federico M. Bandi and Jeffrey R. Russell. Market microstructure noise, integrated variance estimators, and the accuracy of asymptotic approximations. Journal of Econometrics, 160(1):145–159, January 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000692.

Bandi:2012:TVL

[BR12] Federico M. Bandi and Roberto Renò. Time-varying leverage effects. Journal of Econometrics, 169(1):94–113, July 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000115.

Breunig:2015:GFT

[Bre15] Christoph Breunig. Goodness-of-fit tests based on series estimators in nonparametric instrumental regression. Journal of Econometrics, 184(2):328–346, February 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001912.

Berenguer-Rico:2014:SSP

[BRG14] Vanessa Berenguer-Rico and Jesús Gonzalo. Summability of stochastic processes — a generalization of integration for nonlinear processes. Journal of Econometrics, 178(Part 2):331–341, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613001954.

Bierens:2012:SNE

[BS12] Herman J. Bierens and Hosin Song. Semi-nonparametric estimation of independently and identically repeated firstprice auctions via an integrated simulated moments method. Journal of Econometrics, 168(1):108–119, May 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611001758.

Botosaru:2018:NHP

[BS18] Irene Botosaru and Yuya Sasaki. Nonparametric heteroskedasticity in persistent panel processes: an application to earnings dynamics. Journal of Econometrics, 203(2):283–296, April 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617302427.

Blomquist:2019:MDL

[BS19] Sören Blomquist and Laurent Simula. Marginal deadweight loss when the income tax is nonlinear. Journal of Econometrics, 211 (1):47–60, July 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302343.

Bhattacharya:2012:TEB

[BSV12] Jay Bhattacharya, Azeem M. Shaikh, and Edward Vytlacil. Treatment effect bounds: an application to Swan–Ganz catheterization. Journal of Econometrics, 168(2):223–243, June 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000024.

Bound:2010:HER

[BSW10] John Bound, Todd Stinebrickner, and Timothy Waidmann. Health, economic resources and the work decisions of older men. Journal of Econometrics, 156(1):106–129, May 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002097.

Bollerslev:2014:TVJ

[BT14] Tim Bollerslev and Viktor Todorov. Time-varying jump tails. Journal of Econometrics, 183(2):168–180, December 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001109.

Bollerslev:2013:JTE

[BTL13] Tim Bollerslev, Viktor Todorov, and Sophia Zhengzi Li. Jump tails, extreme dependencies, and the distribution of stock returns. Journal of Econometrics, 172(2):307–324, February 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001984.

Bauer:2011:FMR

[BV11] Gregory H. Bauer and Keith Vorkink. Forecasting multivariate realized stock market volatility. Journal of Econometrics, 160(1):93–101, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000631.

Boswijk:2011:MME

[BvdW11] H. Peter Boswijk and Roy van der Weide. Method of moments estimation of GO-GARCH models. Journal of Econometrics, 163(1):118–126, July 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610002150.

Bollinger:2017:BMB

[BvH17] Christopher R. Bollinger and Martijn van Hasselt. Bayesian moment-based inference in a regression model with misclassification error. Journal of Econometrics, 200(2):282–294, October 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300970.

Bai:2014:ITH

[BW14]  Jushan Bai and Peng Wang.  Identification theory for

high dimensional static and dynamic factor models. Journal of Econometrics, 178(2):794–804, February 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613002315.

Bibinger:2015:ECJ

[BW15] Markus Bibinger and Lars Winkelmann. Econometrics of cojumps in high-frequency data with noise. Journal of Econometrics, 184(2):361–378, February 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614002322.

Bollerslev:2015:SRC

[BXZ15] Tim Bollerslev, Lai Xu, and Hao Zhou. Stock return and cash flow predictability: The role of volatility risk. Journal of Econometrics, 187(2):458–471, August 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000561.

Calhoun:2011:HTL

[Cal11] Gray Calhoun. Hypothesis testing in linear regression when k/n is large. Journal of Econometrics, 165(2):163–174, December 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001448.

Campo:2012:RAA

[Cam12] Sandra Campo. Risk aversion and asymmetry in procurement auctions: Identification, estimation and application to construction procurements. Journal of Econometrics, 168(1):96–107, May 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001746.

Canay:2010:IPI

[Can10a] Ivan A. Canay. EL inference for partially identified models: Large deviations optimality and bootstrap validity. Journal of Econometrics, 156(2):408–425, June 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440760900298X.

Canay:2010:SSW

[Can10b] Ivan A. Canay. Simultaneous selection and weighting of moments in GMM using a trapezoidal kernel. Journal of Econometrics, 156(2):284–303, June 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002899.

Caner:2014:NEW

[Can14] Mehmet Caner. Near exogeneity and weak identification in generalized empirical likelihood estimators: Many moment asymptotics. Journal of Econometrics, 182(2):247–268, October 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001031.

Carrasco:2012:RAM

[Car12] Marine Carrasco. A regularization approach to the many instruments problem. Journal of Econometrics, 170(2):383–398, October 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001248.

Cho:2015:TNI

[CAS15] Cheol-Keun Cho, Christine Amsler, and Peter Schmidt. A test of the null of integer integration against the alternative of fractional integration. Journal of Econometrics, 187(1):217–237, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000482.

Cattaneo:2010:ESE

[Cat10] Matias D. Cattaneo. Efficient semiparametric estimation of multi-valued treatment effects under ignorability. Journal of Econometrics, 155(2):138–154, April 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440760900236X.

Calvet:2015:TLG

[CC15a] Laurent E. Calvet and Veronika Czellar. Through the looking glass: Indirect inference via simple equilibria. Journal of Econometrics, 185(2):343–358, April 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614002681.

Chen:2015:OUC

[CC15b] Xiaohong Chen and Timothy M. Christensen. Optimal uniform convergence rates and asymptotic normality for series estimators under weak dependence and weak conditions. Journal of Econometrics, 188(2):447–465, October 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000792.

Chang:2015:HDG

[CCC15] Jinyuan Chang, Song Xi Chen, and Xiaohong Chen. High dimensional generalized empirical likelihood for moment restrictions with dependent data. Journal of Econometrics, 185(1):283– 304, March 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614002553.

Carriero:2022:CLB

[CCCM22] Andrea Carriero, Joshua Chan, Todd E. Clark, and Massimiliano Marcellino. Corrigendum to “Large Bayesian vector autoregressions with stochastic volatility and non-conjugate priors” [j. econometrics 212 (1) (2019) 137–154]. Journal of Econometrics, 227(2):506–512, April 2022. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407621002773. See [CCM19].

Chiou:2018:NRM

[CCeC18] Yan-Yu Chiou, Mei-Yuan Chen, and Jau er Chen. Nonparametric regression with multiple thresholds: Estimation and inference. Journal of Econometrics, 206(2):472–514, October 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301052.

Cai:2018:SQP

[CCF18] Zongwu Cai, Linna Chen, and Ying Fang. A semiparametric quantile panel data model with an application to estimating the growth effect of FDI. Journal of Econometrics, 206(2):531– 553, October 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761830109X.

Carrasco:2015:HDP

[CCGR15] Marine Carrasco, Victor Chernozhukov, Silvia Gonçalves, and Eric Renault. High dimensional problems in econometrics. Journal of Econometrics, 186(2):277–279, June 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000330.

Castle:2013:FFV

[CCH13] Jennifer L. Castle, Michael P. Clements, and David F. Hendry. Forecasting by factors, by variables, by both or neither? Journal of Econometrics, 177(2):305–319, December 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000924.

Cattaneo:2012:OII

[CCJ12] Matias D. Cattaneo, Richard K. Crump, and Michael Jansson. Optimal inference for instrumental variables regression with non-Gaussian errors. Journal of Econometrics, 167(1):1–15, March 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002429.

Carrasco:2012:EI

[CCKR12] Marine Carrasco, Mehmet Caner, Yuichi Kitamura, and Eric Renault. Editors’ introduction. Journal of Econometrics, 170(2):251–255, October 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612001169.

Carriero:2019:LBV

[CCM19] Andrea Carriero, Todd E. Clark, and Massimiliano Marcellino. Large Bayesian vector autoregressions with stochastic volatility and non-conjugate priors. Journal of Econometrics, 212 (1):137–154, September 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761930079X.

See comment [Bog22] and corrigendum [CCCM22].

Chang:2017:NAM

[CCP17] Yoosoon Chang, Yongok Choi, and Joon Y. Park. A new approach to model regime switching. Journal of Econometrics, 196(1):127–143, January 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407616301671.

Chen:2014:FIU

[cCTZ14] Yu chin Chen, Stephen J. Turnovsky, and Eric Zivot. Forecasting inflation using commodity price aggregates. Journal of Econometrics, 183(1):117–134, November 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001547.

Cherchye:2019:BCD

[CDD19] Laurens Cherchye, Thomas Demuynck, and Bram De Rock. Bounding counterfactual demand with unobserved heterogeneity and endogenous expenditures. Journal of Econometrics, 211(2):483–506, August 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619300491.

Cherchye:2015:RPT

[CDDH15] Laurens Cherchye, Thomas Demuynck, Bram De Rock, and Per Hjertstrand. Revealed preference tests for weak separability: an integer programming approach. Journal of Econometrics, 186(1): 129–141, May 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001584.

Chavez-Demoulin:2014:EQT

[CDES14] V. Chavez-Demoulin, P. Embrechts, and S. Sardy. Extremequantile tracking for financial time series. Journal of Econometrics, 181(1):44–52, July 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614000360.

Corradi:2012:IML

[CDF12] Valentina Corradi, Walter Distaso, and Marcelo Fernandes. International market links and volatility transmission. Journal of Econometrics, 170(1):117–141, September 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000759.

Chen:2014:DBS

[CDG14] Liang Chen, Juan J. Dolado, and Jesús Gonzalo. Detecting big structural breaks in large factor models. Journal of Econometrics, 180(1):30–48, May 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614000189.

Chen:2010:NEC

[CDH10] Song X. Chen, Aurore Delaigle, and Peter Hall. Nonparametric estimation for a class of Lévy processes. Journal of Econometrics, 157(2):257–271, August 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000023.

Castle:2012:MSW

[CDH12] Jennifer L. Castle, Jurgen A. Doornik, and David F. Hendry. Model selection when there are multiple breaks. Journal of Econometrics, 169(2):239–246, August 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761200036X.

Christensen:2012:SIG

[CDI12] Bent Jesper Christensen, Christian M. Dahl, and Emma M. Iglesias. Semiparametric inference in a GARCH-in-mean model. Journal of Econometrics, 167(2):458–472, April 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611002090.

Christensen:2011:AAF

[CDR11] Jens H. E. Christensen, Francis X. Diebold, and Glenn D. Rudebusch. The affine arbitrage-free class of Nelson–Siegel term structure models. Journal of Econometrics, 164(1):4–20, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000388.

Chen:2013:MSM

[CDS13] Fei Chen, Francis X. Diebold, and Frank Schorfheide. A Markovswitching multifractal inter-trade duration model, with application to US equities. Journal of Econometrics, 177(2):320–342, December 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613000936.

Caginalp:2014:NPD

[CDS14] Gunduz Caginalp, Mark DeSantis, and Akin Sayrak. The nonlinear price dynamics of U.S. equity ETFs. Journal of Econometrics, 183(2):193–201, December 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001122.

Colacito:2011:CMD

[CEG11] Riccardo Colacito, Robert F. Engle, and Eric Ghysels. A component model for dynamic correlations. Journal of Econometrics, 164(1):45–59, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611000406.

Chan:2016:LBV

[CEK16] Joshua C. C. Chan, Eric Eisenstat, and Gary Koop. Large Bayesian VARMAs. Journal of Econometrics, 192(2):374– 390, June 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300082.

Chambers:2014:TSU

[CET14] Marcus J. Chambers, Joanne S. Ercolani, and A. M. Robert Taylor. Testing for seasonal unit roots by frequency domain regression. Journal of Econometrics, 178(Part 2):243–258, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001899.

Canova:2012:DUI

[CF12] Fabio Canova and Filippo Ferroni. The dynamics of US inflation: Can monetary policy explain the changes? Journal of Econometrics, 167(1):47–60, March 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611002399.

Cattaneo:2013:OCR

[CF13] Matias D. Cattaneo and Max H. Farrell. Optimal convergence rates, Bahadur representation, and asymptotic normality of partitioning estimators. Journal of Econometrics, 174(2):127– 143, June 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613000365.

Caner:2015:HGE

[CF15a] Mehmet Caner and Qingliang Fan. Hybrid generalized empirical likelihood estimators: Instrument selection with adaptive lasso. Journal of Econometrics, 187(1):256–274, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761500069X.

Chabe-Ferret:2015:ABM

[CF15b] Sylvain Chabé-Ferret. Analysis of the bias of matching and difference-in-difference under alternative earnings and selection processes. Journal of Econometrics, 185(1):110–123, March 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614002437.

Chen:2019:IWE

[CF19] Heng Chen and Yanqin Fan. Identification and wavelet estimation of weighted ATE under discontinuous and kink incentive assignment mechanisms. Journal of Econometrics, 212(2):476– 502, October 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619301289.

Calvet:2015:WBS

[CFFL15] Laurent E. Calvet, Marcus Fearnley, Adlai J. Fisher, and Markus Leippold. What is beneath the surface? Option pricing with multifrequency latent states. Journal of Econometrics, 187(2):498– 511, August 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000597.

Cazals:2016:NIV

[CFFS16] Catherine Cazals, Frédérique Fève, Jean-Pierre Florens, and Léopold Simar. Nonparametric instrumental variables estimation for efficiency frontier. Journal of Econometrics, 190(2):349– 359, February 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615001797.

Chambers:2013:GQR

[CFGV13] Robert Chambers, Rolf Färe, Shawna Grosskopf, and Michael Vardanyan. Generalized quadratic revenue functions. Journal of Econometrics, 173(1):11–21, March 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407612002217.

Castle:2010:FEC

[CFH10] Jennifer L. Castle, Nicholas W. P. Fawcett, and David F. Hendry. Forecasting with equilibrium-correction models during structural breaks. Journal of Econometrics, 158(1):25–36, September 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000461.

Cerovecki:2019:FGM

[CFHZ19] Clément Cerovecki, Christian Francq, Siegfried Hörmann, and Jean-Michel Zakoïan. Functional GARCH models: the quasilikelihood approach and its applications. Journal of Econometrics, 209(2):353–375, April 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761930017X.

Chen:2016:ICC

[CFL16] Heng Chen, Yanqin Fan, and Ruixuan Liu. Inference for the correlation coefficient between potential outcomes in the Gaussian switching regime model. Journal of Econometrics, 195(2):255– 270, December 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616301634.

Chen:2010:EMS

[CFPY10] Xiaohong Chen, Yanqin Fan, Demian Pouzo, and Zhiliang Ying. Estimation and model selection of semiparametric multivariate survival functions under general censorship. Journal of Econometrics, 157(1):129–142, July 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002747.

Chaudhuri:2018:IIE

[CFR18] Saraswata Chaudhuri, David T. Frazier, and Eric Renault. Indirect inference with endogenously missing exogenous variables. Journal of Econometrics, 205(1):55–75, July 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300447.

Conti:2014:BEF

[CFSHP14] Gabriella Conti, Sylvia Frühwirth-Schnatter, James J. Heckman, and Rémi Piatek. Bayesian exploratory factor analysis. Journal of Econometrics, 183(1):31–57, November 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001493.

Chernozhukov:2015:NIP

[CFVH^{+}15] Victor Chernozhukov, Iván Fernández-Val, Stefan Hoderlein, Hajo Holzmann, and Whitney Newey. Nonparametric identification in panels using quantiles. Journal of Econometrics, 188(2):378–392, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000755.

Chernozhukov:2015:QRC

[CFVK15] Victor Chernozhukov, Iván Fernández-Val, and Amanda E. Kowalski. Quantile regression with censoring and endogeneity. Journal of Econometrics, 186(1):201–221, May 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001717.

Chernozhukov:2019:NMC

[CFVN19] Victor Chernozhukov, Iván Fernández-Val, and Whitney K. Newey. Nonseparable multinomial choice models in cross-section and panel data. Journal of Econometrics, 211(1):104–116, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302379.

Chen:2014:FPA

[CFW14] Heng Chen, Yanqin Fan, and Jisong Wu. A flexible parametric approach for estimating switching regime models and treatment effect parameters. Journal of Econometrics, 181(2):77–91, August 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000906.

Chib:2010:ACS

[CG10] Siddhartha Chib and Edward Greenberg. Additive cubic spline regression with Dirichlet process mixture errors. Journal of Econometrics, 156(2):322–336, June 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002917.

Carriero:2011:HUN

[CG11] Andrea Carriero and Raffaella Giacomini. How useful are noarbitrage restrictions for forecasting the term structure of interest rates? Journal of Econometrics, 164(1):21–34, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000376.

Cohen-Goldner:2010:ERT

[CGE10] Sarit Cohen-Goldner and Zvi Eckstein. Estimating the return to training and occupational experience: The case of female immigrants. Journal of Econometrics, 156(1):86–105, May 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002085.

Chen:2019:DPL

[CGH^{+}19] Ting Chen, Zhenyu Gao, Jibao He, Wenxi Jiang, and Wei Xiong. Daily price limits and destructive market behavior. Journal of Econometrics, 208(1):249–264, January 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301799.

Chernozhukov:2019:QR

[CGHX19] Victor Chernozhukov, Antonio F. Galvao, Xuming He, and Zhijie Xiao. Quantile regression. Journal of Econometrics, 213(1):1– 3, November 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300570.

Chen:2012:STP

[CGL12] Jia Chen, Jiti Gao, and Degui Li. Semiparametric trending panel data models with cross-sectional dependence. Journal of Econometrics, 171(1):71–85, November 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001613.

Chatterjee:2015:REP

[CGL15] A. Chatterjee, S. Gupta, and S. N. Lahiri. On the residual empirical process based on the ALASSO in high dimensions and its functional oracle property. Journal of Econometrics, 186(2): 317–324, June 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000366.

Chudik:2016:MCA

[CGP16] Alexander Chudik, Valerie Grossman, and M. Hashem Pesaran. A multi-country approach to forecasting output growth using PMIs. Journal of Econometrics, 192(2):349–365, June 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300069.

Cheng:2018:FAB

[CGP18] Tingting Cheng, Jiti Gao, and Peter C. B. Phillips. A frequentist approach to Bayesian asymptotics. Journal of Econometrics, 206(2):359–378, October 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618300976.

Chang:2015:HDS

[CGY15] Jinyuan Chang, Bin Guo, and Qiwei Yao. High dimensional stochastic regression with latent factors, endogeneity and nonlinearity. Journal of Econometrics, 189(2):297–312, December 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001025.

Castle:2010:LDP

[CH10] Jennifer L. Castle and David F. Hendry. A low-dimension portmanteau test for non-linearity. Journal of Econometrics, 158(2):231–245, October 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000096.

Chen:2011:GST

[CH11] Bin Chen and Yongmiao Hong. Generalized spectral testing for multivariate continuous-time models. Journal of Econometrics, 164(2):268–293, October 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761100114X.

Castle:2014:MSU

[CH14a] Jennifer L. Castle and David F. Hendry. Model selection in under-specified equations facing breaks. Journal of Econometrics, 178(Part 2):286–293, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001929.

Chen:2014:UAV

[CH14b] Bin Chen and Yongmiao Hong. A unified approach to validating univariate and multivariate conditional distribution models in time series. Journal of Econometrics, 178(Part 1):22–44, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001541.

Cheng:2015:FFA

[CH15] Xu Cheng and Bruce E. Hansen. Forecasting with factoraugmented regression: a frequentist model averaging approach. Journal of Econometrics, 186(2):280–293, June 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000342.

Calzolari:2018:ESL

[CH18a] Giorgio Calzolari and Roxana Halbleib. Estimating stable latent factor models by indirect inference. Journal of Econometrics, 205(1):280–301, July 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618300538.

Chen:2018:NTS

[CH18b] Bin Chen and Liquan Huang. Nonparametric testing for smooth structural changes in panel data models. Journal of Econometrics, 202(2):245–267, February 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761730221X.

Chang:2012:TNC

[Cha12] Yoosoon Chang. Taking a new contour: a novel approach to panel unit root tests. Journal of Econometrics, 169(1):15–28, July 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000140.

Chambers:2013:JES

[Cha13a] Marcus J. Chambers. Jackknife estimation of stationary autoregressive models. Journal of Econometrics, 172(1):142–157, January 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612002199.

Chan:2013:MAS

[Cha13b] Joshua C. C. Chan. Moving average stochastic volatility models with application to inflation forecast. Journal of Econometrics, 176(2):162–172, October 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613001255.

Chambers:2016:ECT

[Cha16] Marcus J. Chambers. The estimation of continuous time models with mixed frequency data. Journal of Econometrics, 193(2):390–404, August 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616300720.

Chaker:2017:HFE

[Cha17] Selma Chaker. On high frequency estimation of the frictionless price: The use of observed liquidity variables. Journal of Econometrics, 201(1):127–143, November 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617301045.

Chen:2010:NTD

[CHC10] Xiaohong Chen, Lars Peter Hansen, and Marine Carrasco. Nonlinearity and temporal dependence. Journal of Econometrics, 155(2):155–169, April 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002371.

Chen:2010:IMS

[Che10a] Songnian Chen. An integrated maximum score estimator for a generalized censored quantile regression model. Journal of Econometrics, 155(1):90–98, March 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002322.

Chen:2010:RCE

[Che10b] Songnian Chen. Root-N-consistent estimation of fixed-effect panel data transformation models with censoring. Journal of Econometrics, 159(1):222–234, November 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001430.

Chen:2015:MTS

[Che15a] Bin Chen. Modeling and testing smooth structural changes with endogenous regressors. Journal of Econometrics, 185(1):196– 215, March 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614002565.

Cheng:2015:RIN

[Che15b] Xu Cheng. Robust inference in nonlinear models with mixed identification strength. Journal of Econometrics, 189(1):207–228, November 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002055.

Chesher:2017:UEM

[Che17] Andrew Chesher. Understanding the effect of measurement error on quantile regressions. Journal of Econometrics, 200(2):223– 237, October 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300933.

Chen:2018:SEC

[Che18] Songnian Chen. Sequential estimation of censored quantile regression models. Journal of Econometrics, 207(1):30–52, November 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301167.

Chen:2019:QRD

[Che19] Songnian Chen. Quantile regression for duration models with time-varying regressors. Journal of Econometrics, 209(1):1– 17, March 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302574.

Cai:2018:ATE

[CHH18] Zongwu Cai, Yongmiao Hong, and Cheng Hsiao. Advance in theoretical econometrics — essays in honor of Takeshi Amemiya. Journal of Econometrics, 206(2):279–281, October 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300939.

Cafiero:2011:ERC

[CHHW11] Carlo Cafiero, Eugenio S. A. Bobenrieth H., Juan R. A. Bobenrieth H., and Brian D. Wright. The empirical relevance of the competitive storage model. Journal of Econometrics, 162(1):44–54, May 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002619.

Chiburis:2010:SBT

[Chi10] Richard C. Chiburis. Semiparametric bounds on treatment effects. Journal of Econometrics, 159(2):267–275, December 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001582.

Cho:2015:QCA

[ChKS15] Jin Seo Cho, Tae hwan Kim, and Yongcheol Shin. Quantile cointegration in the autoregressive distributed-lag modeling framework. Journal of Econometrics, 188(1):281–300, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001645.

Cai:2014:MTM

[CHL14] Zongwu Cai, Yongmiao Hong, and Qi Li. Misspecification test methods in econometrics. Journal of Econometrics, 178(Part 1): 1–3, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613001516.

Chevillon:2018:GUF

[CHL18] Guillaume Chevillon, Alain Hecq, and Sébastien Laurent. Generating univariate fractional integration within a large VAR(1). Journal of Econometrics, 204(1):54–65, May 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761830006X.

Conte:2011:MMC

[CHM11] Anna Conte, John D. Hey, and Peter G. Moffatt. Mixture models of choice under risk. Journal of Econometrics, 162(1):79–88, May 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002644.

Chao:2014:TOR

[CHN^{+}14] John C. Chao, Jerry A. Hausman, Whitney K. Newey, Norman R. Swanson, and Tiemen Woutersen. Testing overidentifying restrictions with many instruments and heteroskedasticity. Journal of Econometrics, 178(Part 1):15–21, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S030440761300153X.

Choi:2013:CFL

[Cho13] Seungmoon Choi. Closed-form likelihood expansions for multivariate time-inhomogeneous diffusions. Journal of Econometrics, 174(2):45–65, June 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613000341.

Choi:2015:EFA

[Cho15] Seungmoon Choi. Explicit form of approximate transition probability density functions of diffusion processes. Journal of Econometrics, 187(1):57–73, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000275.

Chen:2019:TED

[CHO19] Cathy Yi-Hsuan Chen, Wolfgang Karl Härdle, and Yarema Okhrin. Tail event driven networks of SIFIs. Journal of Econometrics, 208(1):282–298, January 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301829.

Christensen:2018:DPS

[CHP18] Kim Christensen, Ulrich Hounyo, and Mark Podolskij. Is the diurnal pattern sufficient to explain intraday variation in volatility? A nonparametric assessment. Journal of Econometrics, 205(2):336–362, August 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300551.

Chiang:2019:RUI

[CHS19] Harold D. Chiang, Yu-Chin Hsu, and Yuya Sasaki. Robust uniform inference for quantile treatment effects in regression discontinuity designs. Journal of Econometrics, 211(2):589–618, August 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300569.

Cheng:2015:TOM

[CIY15] Tzu-Chang F. Cheng, Ching-Kang Ing, and Shu-Hui Yu. Toward optimal model averaging in regression models with time series errors. Journal of Econometrics, 189(2):321–334, December 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001049.

Cizek:2012:SRE

[Cíz12] Pavel Cízek. Semiparametric robust estimation of truncated and censored regression models. Journal of Econometrics, 168(2): 347–366, June 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000589.

Chang:2015:SDA

[CJMMPA15] Chia-Lin Chang, Juan-Ángel Jiménez-Martín, Esfandiar Maasoumi, and Teodosio Pérez-Amaral. A stochastic dominance approach to financial risk management strategies. Journal of Econometrics, 187(2):472–485, August 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000573.

Choi:2016:RSR

[CJP16] Yongok Choi, Stefan Jacewitz, and Joon Y. Park. A reexamination of stock return predictability. Journal of Econometrics, 192(1):168–189, May 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761500305X.

Cai:2015:FIC

[CJY15] Zongwu Cai, Ted Juhl, and Bingduo Yang. Functional index coefficient models with variable selection. Journal of Econometrics, 189(2):272–284, December 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001001.

Choi:2012:MSC

[CK12] In Choi and Eiji Kurozumi. Model selection criteria for the leads-and-lags cointegrating regression. Journal of Econometrics, 169(2):224–238, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612000310.

Caner:2018:AHC

[CK18] Mehmet Caner and Anders Bredahl Kock. Asymptotically honest confidence regions for high dimensional parameters by the desparsified conservative lasso. Journal of Econometrics, 203(1):143–168, March 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617302282.

Chiappori:2015:NIE

[CKK15] Pierre-André Chiappori, Ivana Komunjer, and Dennis Kristensen. Nonparametric identification and estimation of transformation models. Journal of Econometrics, 188(1):22–39, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761500010X.

Carriero:2016:SAM

[CKM16] Andrea Carriero, George Kapetanios, and Massimiliano Marcellino. Structural analysis with multivariate autoregressive index models. Journal of Econometrics, 192(2):332–348, June 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300057.

Christensen:2010:PAE

[CKP10] Kim Christensen, Silja Kinnebrock, and Mark Podolskij. Preaveraging estimators of the ex-post covariance matrix in noisy diffusion models with non-synchronous data. Journal of Econometrics, 159(1):116–133, November 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001260.

Chang:2016:NTS

[CKP16] Yoosoon Chang, Chang Sik Kim, and Joon Y. Park. Nonstationarity in time series of state densities. Journal of Econometrics, 192(1):152–167, May 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615003036.

Chen:2016:ICS

[CKT16] Songnian Chen, Shakeeb Khan, and Xun Tang. Informational content of special regressors in heteroskedastic binary response models. Journal of Econometrics, 193(1):162–182, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300227.

Chuang:2017:TCD

[CKT17] O-Chia Chuang, Chung-Ming Kuan, and Larry Y. Tzeng. Testing for central dominance: Method and application. Journal of Econometrics, 196(2):368–378, February 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616301968.

Chen:2010:AJE

[CL10] Songnian Chen and Qi Li. Annals journal of econometrics: Nonlinear and nonparametric methods in econometrics. Journal of Econometrics, 157(1):3–5, July 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002760.

Chen:2014:SMI

[CL14] Xiaohong Chen and Zhipeng Liao. Sieve m inference on irregular parameters. Journal of Econometrics, 182(1):70–86, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000682.

Chen:2015:SST

[CL15a] Xiaohong Chen and Zhipeng Liao. Sieve semiparametric twostep GMM under weak dependence. Journal of Econometrics, 189(1):163–186, November 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615002031.

Cheng:2015:SVR

[CL15b] Xu Cheng and Zhipeng Liao. Select the valid and relevant moments: an information-based LASSO for GMM with many moments. Journal of Econometrics, 186(2):443–464, June 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000433.

Chen:2018:BSB

[CL18a] Le-Yu Chen and Sokbae Lee. Best subset binary prediction. Journal of Econometrics, 206(1):39–56, September 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300770.

Cizek:2018:IEN

[CL18b] Pavel Cízek and Jinghua Lei. Identification and estimation of nonseparable single-index models in panel data with correlated random effects. Journal of Econometrics, 203(1):113–128, March 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617302257.

Chen:2019:NSE

[CLL19] Jia Chen, Degui Li, and Oliver Linton. A new semiparametric estimation approach for large dynamic covariance matrices with multiple conditioning variables. Journal of Econometrics, 212(1):155–176, September 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300806.

Chen:2016:SDP

[CLLL16] Jia Chen, Degui Li, Oliver Linton, and Zudi Lu. Semiparametric dynamic portfolio choice with multiple conditioning variables. Journal of Econometrics, 194(2):309–318, October 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301002.

Chen:2019:NEC

[CLLL19] Xirong Chen, Degui Li, Qi Li, and Zheng Li. Nonparametric estimation of conditional quantile functions in the presence of irrelevant covariates. Journal of Econometrics, 212(2):433–450, October 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301034.

Callaway:2018:QTE

[CLO18] Brantly Callaway, Tong Li, and Tatsushi Oka. Quantile treatment effects in difference in differences models under dependence restrictions and with only two time periods. Journal of Econometrics, 206(2):395–413, October 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301027.

Comte:2010:AED

[CLR10] F. Comte, C. Lacour, and Y. Rozenholc. Adaptive estimation of the dynamics of a discrete time stochastic volatility model. Journal of Econometrics, 154(1):59–73, January 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760900150X.

Chen:2014:SIP

[CLS14] Xiaohong Chen, Zhipeng Liao, and Yixiao Sun. Sieve inference on possibly misspecified semi-nonparametric time series models. Journal of Econometrics, 178(Part 3):639–658, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613002066.

Chen:2019:SEB

[CLSY19] Xiaohong Chen, Oliver Linton, Stefan Schneeberger, and Yanping Yi. Semiparametric estimation of the bid-ask spread in extended roll models. Journal of Econometrics, 208(1):160–178, January 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301751.

Chen:2017:SIB

[CLY17] Xiaohong Chen, Oliver Linton, and Yanping Yi. Semiparametric identification of the bid-ask spread in extended roll models. Journal of Econometrics, 200(2):312–325, October 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300994.

Clark:2012:STP

[CM12] Todd E. Clark and Michael W. McCracken. In-sample tests of predictive ability: a new approach. Journal of Econometrics, 170(1):1–14, September 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761200111X.

Chang:2015:EAF

[CM15a] Chia-Lin Chang and Michael McAleer. Econometric analysis of financial derivatives: an overview. Journal of Econometrics, 187(2):403–407, August 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000512.

Clark:2015:NFM

[CM15b] Todd E. Clark and Michael W. McCracken. Nested forecast model comparisons: a new approach to testing equal accuracy. Journal of Econometrics, 186(1):160–177, May 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001699.

Conrad:2016:APG

[CM16] Christian Conrad and Enno Mammen. Asymptotics for parametric GARCH-in-Mean models. Journal of Econometrics, 194(2):319–329, October 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616301014.

Chen:2017:MFA

[CM17a] Richard Y. Chen and Per A. Mykland. Model-free approaches to discern non-stationary microstructure noise and time-varying liquidity in high-frequency data. Journal of Econometrics, 200(1):79–103, September 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617300775.

Chevillon:2017:LCG

[CM17b] Guillaume Chevillon and Sophocles Mavroeidis. Learning can generate long memory. Journal of Econometrics, 198(1):1–9, May 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300027.

Carvalho:2018:AAC

[CMM18] Carlos Carvalho, Ricardo Masini, and Marcelo C. Medeiros. ArCo: an artificial counterfactual approach for high-dimensional panel time-series data. Journal of Econometrics, 207(2):352–380, December 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301544.

Chen:2016:FSR

[CMY16] Rong Chen, Per Mykland, and Qiwei Yao. Financial statistics and risk management: an overview. Journal of Econometrics, 194(2):203–204, October 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616300896.

Chang:2012:RBT

[CN12] Yoosoon Chang and Chi Mai Nguyen. Residual based tests for cointegration in dependent panels. Journal of Econometrics, 167(2):504–520, April 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611002132.

Chen:2014:ADN

[CN14] Ying Chen and Linlin Niu. Adaptive dynamic Nelson–Siegel term structure model with applications. Journal of Econometrics, 180(1):98–115, May 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614000384.

Cavaliere:2015:BST

[CNT15] Giuseppe Cavaliere, Morten Ørregaard Nielsen, and A. M. Robert Taylor. Bootstrap score tests for fractional integration in heteroskedastic ARFIMA models, with an application to price dynamics in commodity spot and futures markets. Journal of Econometrics, 187(2):557–579, August 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000640.

Cavaliere:2017:QML

[CNT17] Giuseppe Cavaliere, Morten Ørregaard Nielsen, and A. M. Robert Taylor. Quasi-maximum likelihood estimation and bootstrap inference in fractional time series models with heteroskedasticity of unknown form. Journal of Econometrics, 198(1):165–188, May 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300234.

Canay:2012:HLO

[CO12] Ivan A. Canay and Taisuke Otsu. Hodges-Lehmann optimality for testing moment conditions. Journal of Econometrics, 171(1):45–53, November 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612001728.

Cederburg:2015:APA

[CO15] Scott Cederburg and Michael S. O’Doherty. Asset-pricing anomalies at the firm level. Journal of Econometrics, 186(1):113–128, May 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001456.

Conrad:2010:NNC

[Con10] Christian Conrad. Non-negativity conditions for the hyperbolic GARCH model. Journal of Econometrics, 157(2):441–457, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000977.

Christensen:2010:RQB

[COP10] Kim Christensen, Roel Oomen, and Mark Podolskij. Realised quantile-based estimation of the integrated variance. Journal of Econometrics, 159(1):74–98, November 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001144.

Cosslett:2013:ESE

[Cos13] Stephen R. Cosslett. Efficient semiparametric estimation for endogenously stratified regression via smoothed likelihood. Journal of Econometrics, 177(1):116–129, November 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613001474.

Chudik:2011:IDV

[CP11] Alexander Chudik and M. Hashem Pesaran. Infinite-dimensional VARs and factor models. Journal of Econometrics, 163(1):4–22, July 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761000206X.

Cheng:2012:CRS

[CP12] Xu Cheng and Peter C. B. Phillips. Cointegrating rank selection in models with time-varying variance. Journal of Econometrics, 169(2):155–165, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612000322.

Chudik:2015:CCE

[CP15] Alexander Chudik and M. Hashem Pesaran. Common correlated effects estimation of heterogeneous dynamic panel data models with weakly exogenous regressors. Journal of Econometrics, 188(2):393–420, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000767.

Cho:2018:PGT

[CP18a] Jin Seo Cho and Peter C. B. Phillips. Pythagorean generalization of testing the equality of two symmetric positive definite matrices. Journal of Econometrics, 202(1):45–56, January 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407617301902.

Clinet:2018:EAV

[CP18b] Simon Clinet and Yoann Potiron. Efficient asymptotic variance reduction when estimating volatility in high frequency data. Journal of Econometrics, 206(1):103–142, September 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300800.

Clinet:2019:TIM

[CP19] Simon Clinet and Yoann Potiron. Testing if the market microstructure noise is fully explained by the informational content of some variables from the limit order book. Journal of Econometrics, 209(2):289–337, April 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300089.

Chen:2019:PSG

[CPP19] Xiaohong Chen, Demian Pouzo, and James L. Powell. Penalized sieve GEL for weighted average derivatives of nonparametric quantile IV regressions. Journal of Econometrics, 213(1):30–53, November 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300594.

Corsi:2010:TBV

[CPR10] Fulvio Corsi, Davide Pirino, and Roberto Renò. Threshold bipower variation and the impact of jumps on volatility forecasting. Journal of Econometrics, 159(2):276–288, December 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001600.

Chen:2014:LIS

[CPT14] Xiaohong Chen, Maria Ponomareva, and Elie Tamer. Likelihood inference in some finite mixture models. Journal of Econometrics, 182(1):87–99, September 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000694.

Christensen:2017:IHF

[CPTV17] K. Christensen, M. Podolskij, N. Thamrongrat, and B. Veliyev. Inference from high-frequency data: a subsampling approach. Journal of Econometrics, 197(2):245–272, April 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616302172.

Christensen:2016:EDE

[CPvdW16] Bent Jesper Christensen, Olaf Posch, and Michel van der Wel. Estimating dynamic equilibrium models using mixed frequency macro and financial data. Journal of Econometrics, 194(1):116–137, September 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616300616.

Chen:2017:ICS

[CPY17] Ye Chen, Peter C. B. Phillips, and Jun Yu. Inference in continuous systems with mildly explosive regressors. Journal of Econometrics, 201(2):400–416, December 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617301689.

Chang:2018:CRE

[CQYZ18] Jinyuan Chang, Yumou Qiu, Qiwei Yao, and Tao Zou. Confidence regions for entries of a large precision matrix. Journal of Econometrics, 206(1):57–82, September 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300782.

Chib:2010:TRB

[CR10] Siddhartha Chib and Srikanth Ramamurthy. Tailored randomized block MCMC methods with application to DSGE models. Journal of Econometrics, 155(1):19–38, March 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609001900.

Chung:2016:MMP

[CR16] EunYi Chung and Joseph P. Romano. Multivariate and multiple permutation tests. Journal of Econometrics, 193(1):76–91, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300021.

Caporin:2017:CVP

[CRdM17] Massimiliano Caporin, Eduardo Rossi, and Paolo Santucci de Magistris. Chasing volatility: a persistent multiplicative error model with jumps. Journal of Econometrics, 198(1):122–145, May 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300192.

Cavaliere:2010:TCI

[CRT10] Giuseppe Cavaliere, Anders Rahbek, and A. M. Robert Taylor. Testing for co-integration in vector autoregressions with non-stationary volatility. Journal of Econometrics, 158(1):7–24, September 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761000045X.

Castagnetti:2015:IFS

[CRT15] Carolina Castagnetti, Eduardo Rossi, and Lorenzo Trapani. Inference on factor structures in heterogeneous panels. Journal of Econometrics, 184(1):145–157, January 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001808.

Caetano:2016:DTI

[CRY16] Carolina Caetano, Christoph Rothe, and Nese Yildiz. A discontinuity test for identification in triangular nonseparable models. Journal of Econometrics, 193(1):113–122, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300215.

Chen:2010:ETM

[CS10] Jiawei Chen and Matthew Shum. Estimating a tournament model of intra-firm wage differentials. Journal of Econometrics, 155(1):39–55, March 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609001912.

Cao:2011:ADI

[CS11a] Bolong Cao and Yixiao Sun. Asymptotic distributions of impulse response functions in short panel vector autoregressions. Journal of Econometrics, 163(2):127–143, August 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000662.

Corradi:2011:PDC

[CS11b] Valentina Corradi and Norman R. Swanson. Predictive density construction and accuracy testing with multiple possibly misspecified diffusion models. Journal of Econometrics, 161(2):304– 324, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761000254X.

Chesher:2012:IMO

[CS12] Andrew Chesher and Konrad Smolinski. IV models of ordered choice. Journal of Econometrics, 166(1):33–48, January 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001175.

Chen:2013:TWU

[CS13] Bin Chen and Zhaogang Song. Testing whether the underlying continuous-time process follows a diffusion: an infinitesimal operator-based approach. Journal of Econometrics, 173(1):83– 107, March 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612002333.

Chen:2014:TMI

[CS14a] Le-Yu Chen and Jerzy Szroeter. Testing multiple inequality hypotheses: a smoothed indicator approach. Journal of Econometrics, 178(Part 3):678–693, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S030440761300208X.

Chen:2014:CPS

[CS14b] Xiaohong Chen and Norman R. Swanson. Causality, prediction, and specification analysis: Recent advances and future directions. Journal of Econometrics, 182(1):1–4, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000621.

Corradi:2014:TSS

[CS14c] Valentina Corradi and Norman R. Swanson. Testing for structural stability of factor augmented forecasting models. Journal of Econometrics, 182(1):100–118, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000700.

Corradi:2018:TJJ

[CSS18] Valentina Corradi, Mervyn J. Silvapulle, and Norman R. Swanson. Testing for jumps and jump intensity path dependence. Journal of Econometrics, 204(2):248–267, June 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300319.

Camponovo:2012:RS

[CST12] Lorenzo Camponovo, Olivier Scaillet, and Fabio Trojani. Robust subsampling. Journal of Econometrics, 167(1):197–210, March 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002594.

Chen:2014:CEF

[CSW14] Yi-Yi Chen, Peter Schmidt, and Hung-Jen Wang. Consistent estimation of the fixed effects stochastic frontier model. Journal of Econometrics, 181(2):65–76, August 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761400058X.

Carrasco:2015:RLM

[CT15a] Marine Carrasco and Guy Tchuente. Regularized LIML for many instruments. Journal of Econometrics, 186(2):427–442, June 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000421.

Creal:2015:HDD

[CT15b] Drew D. Creal and Ruey S. Tsay. High dimensional dynamic stochastic copula models. Journal of Econometrics, 189(2):335– 345, December 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615001050.

Chen:2017:SCT

[CT17] Tao Chen and Gautam Tripathi. A simple consistent test of conditional symmetry in symmetrically trimmed tobit models. Journal of Econometrics, 198(1):29–40, May 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300118.

Christensen:2019:RED

[CTV19] Kim Christensen, Martin Thyrsgaard, and Bezirgen Veliyev. The realized empirical distribution function of stochastic variance with application to goodness-of-fit testing. Journal of Econometrics, 212(2):556–583, October 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619301411.

Christensen:2017:MBL

[CV17] Bent Jesper Christensen and Rasmus Tangsgaard Varneskov. Medium band least squares estimation of fractional cointegration in the presence of low-frequency contamination. Journal of Econometrics, 197(2):218–244, April 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616302160.

Cho:2010:TUH

[CW10] Jin Seo Cho and Halbert White. Testing for unobserved heterogeneity in exponential and Weibull duration models. Journal of Econometrics, 157(2):458–480, August 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000989.

Chiang:2011:VCR

[CW11a] Min-Hsien Chiang and Li-Min Wang. Volatility contagion: a range-based volatility approach. Journal of Econometrics, 165(2):175–189, December 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761100145X.

Cho:2011:GRT

[CW11b] Jin Seo Cho and Halbert White. Generalized runs tests for the IID hypothesis. Journal of Econometrics, 162(2):326–344, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000285.

Cai:2014:CTP

[CW14a] Zongwu Cai and Yunfei Wang. Corrigendum to “Testing predictive regression models with nonstationary regressors” [J. Econometrics 178 (2014) 4–14]. Journal of Econometrics, 181(2): 194, August 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614000487.

Cai:2014:TPR

[CW14b]  Zongwu Cai and Yunfei Wang.  Testing predictive re-

gression models with nonstationary regressors. Journal of Econometrics, 178(Part 1):4–14, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001528.

Chan:2015:NRN

[CW15a] Nigel Chan and Qiying Wang. Nonlinear regressions with nonstationary time series. Journal of Econometrics, 185(1):182– 195, March 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761400253X.

Creal:2015:EAT

[CW15b] Drew D. Creal and Jing Cynthia Wu. Estimation of affine term structure models with spanned or unspanned stochastic volatility. Journal of Econometrics, 185(1):60–81, March 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614002309.

Chen:2018:SEP

[CW18] Songnian Chen and Xi Wang. Semiparametric estimation of panel data models without monotonicity or separability. Journal of Econometrics, 206(2):515–530, October 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301064.

Chen:2019:MMV

[CWZ19] Yu Chen, Zhicheng Wang, and Zhengjun Zhang. Mark to market value at risk. Journal of Econometrics, 208(1):299–321, January 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301830.

Cai:2012:SQR

[CX12] Zongwu Cai and Zhijie Xiao. Semiparametric quantile regression estimation in dynamic models with partially varying coefficients. Journal of Econometrics, 167(2):413–425, April 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002065.

Cavaliere:2014:TUR

[CX14a] Giuseppe Cavaliere and Fang Xu. Testing for unit roots in bounded time series. Journal of Econometrics, 178(Part 2):259– 272, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613001905.

Chen:2014:IVO

[CX14b] Song Xi Chen and Zheng Xu. On implied volatility for options — some reasons to smile and more to correct. Journal of Econometrics, 179(1):1–15, March 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613002200.

Caner:2012:CMW

[CY12] Mehmet Caner and Nese Yildiz. CUE with many weak instruments and nearly singular design. Journal of Econometrics, 170(2):422–441, October 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612001261.

Chan:2015:LET

[CYZ15] Ngai Hang Chan, Chun Yip Yau, and Rong-Mao Zhang. LASSO estimation of threshold autoregressive models. Journal of Econometrics, 189(2):285–296, December 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001013.

Chen:2010:SNE

[CZ10] Songnian Chen and Yahong Zhou. Semiparametric and nonparametric estimation of sample selection models under symmetry. Journal of Econometrics, 157(1):143–150, July 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002759.

Chaudhuri:2011:NMP

[CZ11a] Saraswata Chaudhuri and Eric Zivot. A new method of projection-based inference in GMM with weakly identified nuisance parameters. Journal of Econometrics, 164(2):239–251, October 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001047.

Chen:2011:SEB

[CZ11b] Songnian Chen and Xianbo Zhou. Semiparametric estimation of a bivariate Tobit model. Journal of Econometrics, 165(2):266– 274, December 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001461.

Chen:2012:SET

[CZ12] Songnian Chen and Xianbo Zhou. Semiparametric estimation of a truncated regression model. Journal of Econometrics, 167(2): 297–304, April 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001977.

Chen:2015:SBP

[CZ15a] Min Chen and Ke Zhu. Sign-based portmanteau test for ARCH-type models with heavy-tailed innovations. Journal of Econometrics, 189(2):313–320, December 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001037.

Chen:2015:BQR

[CZ15b] Songnian Chen and Hanghui Zhang. Binary quantile regression with local polynomial smoothing. Journal of Econometrics, 189(1):24–40, November 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615001906.

Chen:2018:NIE

[CZJ18] Songnian Chen, Yahong Zhou, and Yuanyuan Ji. Nonparametric identification and estimation of sample selection models under symmetry. Journal of Econometrics, 202(2):148–160, February 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301938.

Chen:2015:ADF

[CZP15] Qiang Chen, Xu Zheng, and Zhiyuan Pan. Asymptotically distribution-free tests for the volatility function of a diffusion. Journal of Econometrics, 184(1):124–144, January 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001778.

Davidson:2017:DMB

[Dav17] Russell Davidson. A discrete model for bootstrap iteration. Journal of Econometrics, 201(2):228–236, December 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301550.

deCastro:2019:SGQ

[dCGKL19] Luciano de Castro, Antonio F. Galvao, David M. Kaplan, and Xin Liu. Smoothed GMM for quantile models. Journal of Econometrics, 213(1):121–144, November 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300636.

Dardanoni:2015:MAE

[DDMP15] Valentino Dardanoni, Giuseppe De Luca, Salvatore Modica, and Franco Peracchi. Model averaging estimation of generalized linear models with imputed covariates. Journal of Econometrics, 184(2):452–463, February 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001420.

Davis:2018:ITP

[DDSW18] Richard A. Davis, Holger Drees, Johan Segers, and Michal Warchol. Inference on the tail process with application to financial time series modeling. Journal of Econometrics, 205(2):508– 525, August 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618300691.

Delgado:2012:DFT

[DE12] Miguel A. Delgado and Juan Carlos Escanciano. Distributionfree tests of stochastic monotonicity. Journal of Econometrics, 170(1):68–75, September 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612000723.

Dungey:2018:TME

[DEMY18] Mardi Dungey, Deniz Erdemlioglu, Marius Matei, and Xiye Yang. Testing for mutually exciting jumps and financial flights in high frequency data. Journal of Econometrics, 202(1):18–44, January 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407617301914.

Deschamps:2011:BEE

[Des11] Philippe J. Deschamps. Bayesian estimation of an extended local scale stochastic volatility model. Journal of Econometrics, 162(2):369–382, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611000509.

Drerup:2017:PSD

[DEvG17] Tilman Drerup, Benjamin Enke, and Hans-Martin von Gaudecker. The precision of subjective data and the explanatory power of economic models. Journal of Econometrics, 200(2):378–389, October 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617301033.

Deza:2015:TSS

[Dez15] Monica Deza. Is there a stepping stone effect in drug use? Separating state dependence from unobserved heterogeneity within and between illicit drugs. Journal of Econometrics, 184(1):193– 207, January 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761400181X.

DelBoca:2012:EHI

[DF12] Daniela Del Boca and Christopher Flinn. Endogenous household interaction. Journal of Econometrics, 166(1):49–65, January 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001187.

DHaultfoeuille:2015:IMM

[DF15]  Xavier D’Haultfœuille and Philippe Février.  Identifica-

tion of mixture models using support variations. Journal of Econometrics, 189(1):70–82, November 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001931.

Dunker:2014:IES

[DFH^{+}14] Fabian Dunker, Jean-Pierre Florens, Thorsten Hohage, Jan Johannes, and Enno Mammen. Iterative estimation of solutions to noisy nonlinear operator equations in nonparametric instrumental regression. Journal of Econometrics, 178(Part 3):444– 455, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613001322.

Darolles:2018:ACG

[DFL18] Serge Darolles, Christian Francq, and Sébastien Laurent. Asymptotics of Cholesky GARCH models and time-varying conditional betas. Journal of Econometrics, 204(2):223–247, June 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300307.

Daouia:2012:RNF

[DFS12] Abdelaati Daouia, Jean-Pierre Florens, and Léopold Simar. Regularization of nonparametric frontier estimators. Journal of Econometrics, 168(2):285–299, June 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000425.

Daouia:2011:RIN

[DG11] Abdelaati Daouia and Irène Gijbels. Robustness and inference in nonparametric partial frontier modeling. Journal of Econometrics, 161(2):147–165, April 1, 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610002447.

Dovonon:2017:BGO

[DG17] Prosper Dovonon and Sílvia Gonçalves. Bootstrapping the GMM overidentification test under first-order underidentification. Journal of Econometrics, 201(1):43–71, November 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301197.

Daouia:2014:MAM

[DGG14] Abdelaati Daouia, Stéphane Girard, and Armelle Guillou. A Γ-moment approach to monotonic boundary estimation. Journal of Econometrics, 178(2):727–740, February 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613002285.

DiTraglia:2019:IEM

[DGJ19] Francis J. DiTraglia and Camilo García-Jimeno. Identifying the effect of a mis-classified, binary, endogenous regressor. Journal of Econometrics, 209(2):376–390, April 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300181.

Dovonon:2013:BRM

[DGM13] Prosper Dovonon, Sílvia Gonçalves, and Nour Meddahi. Bootstrapping realized multivariate volatility measures. Journal of Econometrics, 172(1):49–65, January 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001765.

Diebold:2019:BDD

[DGMZ19] Francis X. Diebold, Eric Ghysels, Per Mykland, and Lan Zhang. Big data in dynamic predictive econometric modeling. Journal of Econometrics, 212(1):1–3, September 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300727.

Dong:2015:SSI

[DGP15] Chaohua Dong, Jiti Gao, and Bin Peng. Semiparametric single-index panel data models with cross-sectional dependence. Journal of Econometrics, 188(1):301–312, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001700.

Doz:2011:TSE

[DGR11] Catherine Doz, Domenico Giannone, and Lucrezia Reichlin. A two-step estimator for large approximate dynamic factor models based on Kalman filtering. Journal of Econometrics, 164(1):188–205, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761100039X.

Dong:2017:STN

[DGTY17] Chaohua Dong, Jiti Gao, Dag Tjøstheim, and Jiying Yin. Specification testing for nonlinear multivariate cointegrating regressions. Journal of Econometrics, 200(1):104–117, September 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300787.

DHaultfoeuille:2010:NIM

[D’H10] Xavier D’Haultfœuille. A new instrumental method for dealing with endogenous selection. Journal of Econometrics, 154(1):1– 15, January 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001468.

Donald:2014:EID

[DH14] Stephen G. Donald and Yu-Chin Hsu. Estimation and inference for distribution functions and quantile functions in treatment effect models. Journal of Econometrics, 178(Part 3):383–397, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001826.

Dovonon:2018:APG

[DH18] Prosper Dovonon and Alastair R. Hall. The asymptotic properties of GMM and indirect inference under second-order identification. Journal of Econometrics, 205(1):76–111, July 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300459.

Dunker:2018:NID

[DHKS18] Fabian Dunker, Stefan Hoderlein, Hiroaki Kaido, and Robert Sherman. Nonparametric identification of the distribution of random coefficients in binary response static games of complete information. Journal of Econometrics, 206(1):83–102, September 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300794.

Dette:2016:TME

[DHN16] Holger Dette, Stefan Hoderlein, and Natalie Neumeyer. Testing multivariate economic restrictions using quantiles: The example of Slutsky negative semidefiniteness. Journal of Econometrics, 191(1):129–144, March 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761500250X.

DelNegro:2016:DPP

[DHS16] Marco Del Negro, Raiden B. Hasegawa, and Frank Schorfheide. Dynamic prediction pools: an investigation of financial frictions and forecasting performance. Journal of Econometrics, 192(2): 391–405, June 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300094.

deHaan:2011:ECD

[dHvdG11] Jan de Haan and Heymerik A. van der Grient. Eliminating chain drift in price indexes based on scanner data. Journal of Econometrics, 161(1):36–46, March 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001843.

Davis:2016:CMD

[DHY16] Richard A. Davis, Stacey A. Hancock, and Yi-Ching Yao. On consistency of minimum description length model selection for piecewise autoregressions. Journal of Econometrics, 194(2):360– 368, October 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761630104X.

Diewert:2014:DPC

[Die14] W. Erwin Diewert. Decompositions of profitability change using cost functions. Journal of Econometrics, 183(1):58–66, November 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761400150X.

DiTraglia:2016:UII

[DiT16] Francis J. DiTraglia. Using invalid instruments on purpose: Focused moment selection and averaging for GMM. Journal of Econometrics, 195(2):187–208, December 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616301518.

Debarsy:2015:LSP

[DJfL15] Nicolas Debarsy, Fei Jin, and Lung fei Lee. Large sample properties of the matrix exponential spatial specification with an application to FDI. Journal of Econometrics, 188(1):1–21, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001219.

Danielsson:2013:FTV

[DJS^{+}13] Jón Daníelsson, Bjørn N. Jorgensen, Gennady Samorodnitsky, Mandira Sarma, and Casper G. de Vries. Fat tails, VaR and subadditivity. Journal of Econometrics, 172(2):283–291, February 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001959.

Dias:2018:EFV

[DK18] Gustavo Fruet Dias and George Kapetanios. Estimation and forecasting in vector autoregressive moving average models for rich datasets. Journal of Econometrics, 202(1):75–91, January 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407617301707.

Dufour:2014:EIH

[DKK14] Jean-Marie Dufour and Jeong-Ryeol Kurz-Kim. Editors’ introduction: Heavy tails and stable Paretian distributions in econometrics. Journal of Econometrics, 181(1):1–2, July 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613002613.

Dong:2011:NIB

[DL11] Yingying Dong and Arthur Lewbel. Nonparametric identification of a binary random factor in cross section data. Journal of Econometrics, 163(2):163–171, August 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611000650.

DeNadai:2016:NEV

[DL16] Michele De Nadai and Arthur Lewbel. Nonparametric errors in variables models with measurement errors on both sides of the equation. Journal of Econometrics, 191(1):19–32, March 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002390.

Davezies:2017:RDD

[DL17] Laurent Davezies and Thomas Le Barbanchon. Regression discontinuity design with continuous measurement error in the running variable. Journal of Econometrics, 200(2):260–281, October 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300969.

Dong:2018:ANM

[DL18] Chaohua Dong and Oliver Linton. Additive nonparametric models with time variable and both stationary and nonstationary regressors. Journal of Econometrics, 207(1):212–236, November 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301404.

Darolles:2017:MDH

[DLM17] Serge Darolles, Gaëlle Le Fol, and Gulten Mero. Mixture of distribution hypothesis: Analyzing daily liquidity frictions and information flows. Journal of Econometrics, 201(2):367–383, December 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301665.

Dai:2019:KFF

[DLX19] Chaoxing Dai, Kun Lu, and Dacheng Xiu. Knowing factors or factor loadings, or neither? Evaluating estimators of large covariance matrices with noisy and asynchronous data. Journal of Econometrics, 208(1):43–79, January 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301702.

DHaultfoeuille:2013:IER

[DM13] Xavier D’Haultfœuille and Arnaud Maurel. Inference on an extended Roy model, with an application to schooling decisions in France. Journal of Econometrics, 174(2):95–106, June 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000328.

Davis:2012:TEE

[DMC12] Richard A. Davis, Thomas Mikosch, and Ivor Cribben. Towards estimating extremal serial dependence via the bootstrapped extremogram. Journal of Econometrics, 170(1):142–152, September 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000978.

Djogbenou:2019:ATW

[DMN19] Antoine A. Djogbenou, James G. MacKinnon, and Morten Ørregaard Nielsen. Asymptotic theory and wild bootstrap inference with clustered errors. Journal of Econometrics, 212(2):393– 412, October 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300909.

Dardanoni:2011:RIC

[DMP11] Valentino Dardanoni, Salvatore Modica, and Franco Peracchi. Regression with imputed covariates: a generalized missingindicator approach. Journal of Econometrics, 162(2):362–368, June 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611000327.

DeLuca:2018:WAL

[DMP18] Giuseppe De Luca, Jan R. Magnus, and Franco Peracchi. Weighted-average least squares estimation of generalized linear models. Journal of Econometrics, 204(1):1–17, May 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300034.

Delaigle:2016:RCD

[DMR16] Aurore Delaigle, Alexander Meister, and Jeroen Rombouts. Root-T consistent density estimation in GARCH models. Journal of Econometrics, 192(1):55–63, May 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002614.

Darolles:2017:EI

[DMR17] S. Darolles, Alain Monfort, and Eric Renault. Editors’ introduction. Journal of Econometrics, 201(2):173–175, December 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301483.

DHaultfoeuille:2018:EQR

[DMZ18] Xavier D’Haultfœuille, Arnaud Maurel, and Yichong Zhang. Extremal quantile regressions for selection models and the blackwhite wage gap. Journal of Econometrics, 203(1):129–142, March 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617302269.

Durlauf:2010:UAC

[DNR10] Steven N. Durlauf, Salvador Navarro, and David A. Rivers. Understanding aggregate crime regressions. Journal of Econometrics, 158(2):306–317, October 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000060.

Das:2017:ELF

[DP17] Tirthatanmoy Das and Solomon W. Polachek. Estimating labor force joiners and leavers using a heterogeneity augmented twotier stochastic frontier. Journal of Econometrics, 199(2):156– 172, August 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300696.

Dueker:2011:MCT

[DPSS11] Michael J. Dueker, Zacharias Psaradakis, Martin Sola, and Fabio Spagnolo. Multivariate contemporaneous-threshold autoregressive models. Journal of Econometrics, 160(2):311–325, February 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001910.

Diks:2011:LBS

[DPvD11] Cees Diks, Valentyn Panchenko, and Dick van Dijk. Likelihoodbased scoring rules for comparing density forecasts in tails. Journal of Econometrics, 163(2):215–230, August 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611000807.

Dou:2016:GYW

[DPY16] Baojun Dou, Maria Lucia Parrella, and Qiwei Yao. Generalized Yule–Walker estimation for spatio-temporal models with unknown diagonal coefficients. Journal of Econometrics, 194(2):369–382, October 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616301051.

Delgado:2015:NNT

[DR15] Miguel A. Delgado and Peter M. Robinson. Non-nested testing of spatial correlation. Journal of Econometrics, 187(1):385–401, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000950.

Durlauf:2010:EI

[DS10] Steven Durlauf and Aris Spanos. Editorial introduction. Journal of Econometrics, 158(2):175–176, October 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000151.

Duong:2015:EEI

[DS15] Diep Duong and Norman R. Swanson. Empirical evidence on the importance of aggregation, asymmetry, and jumps for volatility prediction. Journal of Econometrics, 187(2):606–621, August 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000676.

Delgado:2018:NTC

[DS18] Miguel A. Delgado and Xiaojun Song. Nonparametric tests for conditional symmetry. Journal of Econometrics, 206(2):447– 471, October 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301040.

Diebold:2017:RTF

[DSS17] Francis X. Diebold, Frank Schorfheide, and Minchul Shin. Real-time forecast evaluation of DSGE models with stochastic volatility. Journal of Econometrics, 201(2):322–332, December 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761730163X.

Duan:2012:MCD

[DSW12] Jin-Chuan Duan, Jie Sun, and Tao Wang. Multiperiod corporate default prediction — a forward intensity approach. Journal of Econometrics, 170(1):191–209, September 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001145.

Dufour:2010:SLR

[DT10] Jean-Marie Dufour and Abderrahim Taamouti. Short and long run causality measures: Theory and inference. Journal of Econometrics, 154(1):42–58, January 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609001493.

Dellaportas:2019:ISP

[DT19] Petros Dellaportas and Mike G. Tsionas. Importance sampling from posterior distributions using copula-like approximations. Journal of Econometrics, 210(1):45–57, May 2019.

CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618302057.

Duan:2016:LMA

[Dua16] Jin-Chuan Duan. Local-momentum autoregression and the modeling of interest rate term structure. Journal of Econometrics, 194(2):349–359, October 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616301038.

Delgado:2010:DFT

[DV10] Miguel A. Delgado and Carlos Velasco. Distribution-free tests for time series models specification. Journal of Econometrics, 155(2): 128–137, April 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609002358.

Dominicy:2013:MSQ

[DV13] Yves Dominicy and David Veredas. The method of simulated quantiles. Journal of Econometrics, 172(2):235–247, February 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001947.

Diebold:2014:NTV

[DY14] Francis X. Diebold and Kamil Yilmaz. On the network topology of variance decompositions: Measuring the connectedness of financial firms. Journal of Econometrics, 182(1):119–134, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000712.

Du:2015:HPR

[DZ15] Zaichao Du and Lin Zhang. Home-purchase restriction, property tax and housing price in China: a counterfactual analysis. Journal of Econometrics, 188(2):558–568, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000871.

Escanciano:2014:SAL

[EG14] J. C. Escanciano and S. C. Goh. Specification analysis of linear quantile models. Journal of Econometrics, 178(Part 3):495– 507, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S030440761300184X.

Eggleston:2016:EDE

[Egg16] Jonathan Eggleston. An efficient decomposition of the expectation of the maximum for the multivariate normal and related distributions. Journal of Econometrics, 195(1):120–133, November 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301361.

Elliott:2013:CSR

[EGT13] Graham Elliott, Antonio Gargano, and Allan Timmermann. Complete subset regressions. Journal of Econometrics, 177 (2):357–373, December 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613000948.

Escanciano:2012:UCD

[EJC12] Juan Carlos Escanciano and David T. Jacho-Chávez. nuniformly consistent density estimation in nonparametric regression models. Journal of Econometrics, 167(2):305–316, April 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001989.

Escanciano:2014:UCW

[EJCL14] Juan Carlos Escanciano, David T. Jacho-Chávez, and Arthur Lewbel. Uniform convergence of weighted sums of non and semiparametric residuals for estimation and testing. Journal of Econometrics, 178(Part 3):426–443, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001462.

Erickson:2014:MDE

[EJW14] Timothy Erickson, Colin Huan Jiang, and Toni M. Whited. Minimum distance estimation of the errors-in-variables model using linear cumulant equations. Journal of Econometrics, 183(2):211– 221, December 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001146.

Elliott:2013:PBO

[EL13] Graham Elliott and Robert P. Lieli. Predicting binary outcomes. Journal of Econometrics, 174(1):15–26, May 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613000171.

Elliott:2011:CFA

[Ell11] Graham Elliott. A control function approach for testing the usefulness of trending variables in forecast models and linear regression. Journal of Econometrics, 164(1):79–91, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000418.

Egorov:2011:TTY

[ELN11] Alexei V. Egorov, Haitao Li, and David Ng. A tale of two yield curves: Modeling the joint term structure of dollar and euro interest rates. Journal of Econometrics, 162(1):55–70, May 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002632.

Elliott:2014:PPB

[EM14a] Graham Elliott and Ulrich K. Müller. Pre and post break parameter inference. Journal of Econometrics, 180(2):141–157, June 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000475.

Engle:2014:PRA

[EM14b] Robert Engle and Abhishek Mistry. Priced risk and asymmetric volatility in the cross section of skewness. Journal of Econometrics, 182(1):135–144, September 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000724.

Eichler:2011:FDF

[EMvS11] Michael Eichler, Giovanni Motta, and Rainer von Sachs. Fitting dynamic factor models to non-stationary time series. Journal of Econometrics, 163(1):51–70, July 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610002113.

Engle:2017:SGL

[ERS17] Robert Engle, Guillaume Roussellet, and Emil Siriwardane. Scenario generation for long run interest rate risk assessment. Journal of Econometrics, 201(2):333–347, December 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301641.

Escanciano:2010:TSI

[ES10]  Juan Carlos Escanciano and Kyungchul Song.  Testing

single-index restrictions with a focus on average derivatives. Journal of Econometrics, 156(2):377–391, June 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002966.

Elliott:2014:AIJ

[ET14] Graham Elliott and A. M. Robert Taylor. Annals issue of Journal of Econometrics “recent advances in time series econometrics”: Guest Editors’ introduction. Journal of Econometrics, 178(Part 2):207–209, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001851.

Escanciano:2010:STP

[EV10] Juan Carlos Escanciano and Carlos Velasco. Specification tests of parametric dynamic conditional quantiles. Journal of Econometrics, 159(1):209–221, November 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001429.

Ergemen:2017:EFI

[EV17] Yunus Emre Ergemen and Carlos Velasco. Estimation of fractionally integrated panels with fixed effects and cross-section dependence. Journal of Econometrics, 196(2):248–258, February 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301737.

Eraker:2015:NLD

[EW15] Bjørn Eraker and Jiakou Wang. A non-linear dynamic model of the variance risk premium. Journal of Econometrics, 187(2):547– 556, August 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000639.

Fuentes-Albero:2013:MCM

[FAM13] Cristina Fuentes-Albero and Leonardo Melosi. Methods for computing marginal data densities from the Gibbs output. Journal of Econometrics, 175(2):132–141, August 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613000560.

Fanelli:2012:DID

[Fan12] Luca Fanelli. Determinacy, indeterminacy and dynamic misspecification in linear rational expectations models. Journal of Econometrics, 170(1):153–163, September 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000966.

Farrell:2015:RIA

[Far15] Max H. Farrell. Robust inference on average treatment effects with possibly more covariates than observations. Journal of Econometrics, 189(1):1–23, November 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001864.

Fasen:2013:SEM

[Fas13] Vicky Fasen. Statistical estimation of multivariate Ornstein– Uhlenbeck processes and applications to co-integration. Journal of Econometrics, 172(2):325–337, February 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612002035.

Fattore:2010:APG

[Fat10] Marco Fattore. Axiomatic properties of geo-logarithmic price indices. Journal of Econometrics, 156(2):344–353, June 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002930.

Favero:2013:MFG

[Fav13] Carlo A. Favero. Modelling and forecasting government bond spreads in the euro area: a GVAR model. Journal of Econometrics, 177(2):343–356, December 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761300081X.

Fan:2011:TDJ

[FF11] Yingying Fan and Jianqing Fan. Testing and detecting jumps based on a discretely observed process. Journal of Econometrics, 164(2):331–344, October 1, 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611001278.

Feve:2014:NPA

[FF14] Frédérique Fève and Jean-Pierre Florens. Non parametric analysis of panel data models with endogenous variables. Journal of Econometrics, 181(2):151–164, August 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000499.

Figuerola-Ferretti:2010:MMP

[FFG10] Isabel Figuerola-Ferretti and Jesús Gonzalo. Modelling and measuring price discovery in commodity markets. Journal of Econometrics, 158(1):95–107, September 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000552.

Frandsen:2012:QTE

[FFM12] Brigham R. Frandsen, Markus Frölich, and Blaise Melly. Quantile treatment effects in the regression discontinuity design. Journal of Econometrics, 168(2):382–395, June 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000607.

Fan:2011:NCA

[FGL11] Yanqin Fan, Matthew Gentry, and Tong Li. A new class of asymptotically efficient estimators for moment condition models. Journal of Econometrics, 162(2):268–277, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000273.

Feng:2017:VCP

[FGPZ17] Guohua Feng, Jiti Gao, Bin Peng, and Xiaohui Zhang. A varying-coefficient panel data model with fixed effects: Theory and an application to US commercial banks. Journal of Econometrics, 196(1):68–82, January 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407616301786.

Firpo:2017:MEQ

[FGS17] Sergio Firpo, Antonio F. Galvao, and Suyong Song. Measurement errors in quantile regression models. Journal of Econometrics, 198(1):146–164, May 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617300209.

Fiorentini:2018:SEA

[FGS18] Gabriele Fiorentini, Alessandro Galesi, and Enrique Sentana. A spectral EM algorithm for dynamic factor models. Journal of Econometrics, 205(1):249–279, July 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300526.

Fan:2017:PIF

[FGZ17] Yanqin Fan, Emmanuel Guerre, and Dongming Zhu. Partial identification of functionals of the joint distribution of “potential outcomes”. Journal of Econometrics, 197(1):42–59, March 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301993.

Fan:2019:GHD

[FGZ19] Jianqing Fan, Wenyan Gong, and Ziwei Zhu. Generalized high-dimensional trace regression via nuclear norm regularization. Journal of Econometrics, 212(1):177–202, September 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300818.

Fengler:2015:SNE

[FH15a] Matthias R. Fengler and Lin-Yee Hin. Semi-nonparametric estimation of the call-option price surface under strike and time-to-expiry no-arbitrage constraints. Journal of Econometrics, 184(2):242–261, February 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001845.

Freyberger:2015:ISR

[FH15b] Joachim Freyberger and Joel L. Horowitz. Identification and shape restrictions in nonparametric instrumental variables estimation. Journal of Econometrics, 189(1):41–53, November 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001918.

Fujiki:2015:DEM

[FH15c] Hiroshi Fujiki and Cheng Hsiao. Disentangling the effects of multiple treatments — measuring the net economic impact of the 1995 great Hanshin–Awaji earthquake. Journal of Econometrics, 186(1):66–73, May 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614002541.

Fu:2019:MFC

[FH19] Zhonghao Fu and Yongmiao Hong. A model-free consistent test for structural change in regression possibly with endogeneity. Journal of Econometrics, 211(1):206–242, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302471.

Fan:2016:RIR

[FHLV16] Jianqing Fan, Fang Han, Han Liu, and Byron Vickers. Robust inference of risks of large portfolios. Journal of Econometrics, 194(2):298–308, October 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616300987.

Forni:2015:DFM

[FHLZ15] Mario Forni, Marc Hallin, Marco Lippi, and Paolo Zaffaroni. Dynamic factor models with infinite-dimensional factor spaces: One-sided representations. Journal of Econometrics, 185(2):359– 371, April 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614002693.

Forni:2017:DFM

[FHLZ17] Mario Forni, Marc Hallin, Marco Lippi, and Paolo Zaffaroni. Dynamic factor models with infinite-dimensional factor space: Asymptotic analysis. Journal of Econometrics, 199(1):74–92, July 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300477.

Fox:2012:RCL

[FiKRB12] Jeremy T. Fox, Kyoo il Kim, Stephen P. Ryan, and Patrick Bajari. The random coefficients logit model is identified. Journal of Econometrics, 166(2):204–212, February 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611001655.

Fox:2016:SNA

[FiKY16] Jeremy T. Fox, Kyoo il Kim, and Chenyu Yang. A simple nonparametric approach to estimating the distribution of random coefficients in structural models. Journal of Econometrics, 195(2):236–254, December 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616301622.

Fisher:2019:BIP

[FJ19] Mark Fisher and Mark J. Jensen. Bayesian inference and prediction of a multiple-change-point panel model with nonparametric priors. Journal of Econometrics, 210(1):187–202, May 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302136.

Francq:2017:TCE

[FJGM17] C. Francq, M. D. Jiménez-Gamero, and S. G. Meintanis. Tests for conditional ellipticity in multivariate GARCH models. Journal of Econometrics, 196(2):305–319, February 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301920.

Fan:2019:SVM

[FK19] Jianqing Fan and Donggyu Kim. Structured volatility matrix estimation for non-synchronized high-frequency financial data. Journal of Econometrics, 209(1):61–78, March 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302604.

Forni:2011:GDF

[FL11] Mario Forni and Marco Lippi. The general dynamic factor model: One-sided representation results. Journal of Econometrics, 163 (1):23–28, July 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610002071.

Fulop:2013:ELS

[FL13] Andras Fulop and Junye Li. Efficient learning via simulation: a marginalized resample-move approach. Journal of Econometrics, 176(2):146–161, October 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613001164.

Fan:2016:DAI

[FL16] Yanqin Fan and Ruixuan Liu. A direct approach to inference in nonparametric and semiparametric quantile models.

Journal of Econometrics, 191(1):196–216, March 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002560.

Fan:2018:PII

[FL18] Yanqin Fan and Ruixuan Liu. Partial identification and inference in censored quantile regression. Journal of Econometrics, 206(1):1–38, September 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618300708.

Fan:2019:PQR

[FL19a] Rui Fan and Ji Hyung Lee. Predictive quantile regressions under persistence and conditional heteroskedasticity. Journal of Econometrics, 213(1):261–280, November 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300697.

Fulop:2019:BED

[FL19b] Andras Fulop and Junye Li. Bayesian estimation of dynamic asset pricing models with informative observations. Journal of Econometrics, 209(1):114–138, March 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618302276.

Flabbi:2010:PGD

[Fla10] Luca Flabbi. Prejudice and gender differentials in the US labor market in the last twenty years. Journal of Econometrics, 156(1): 190–200, May 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609002152.

Figueroa-Lopez:2019:OTU

[FLM19] José E. Figueroa-López and Cecilia Mancini. Optimum thresholding using mean and conditional mean squared error. Journal of Econometrics, 208(1):179–210, January 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301763.

Fan:2015:RLP

[FLS15] Jianqing Fan, Yuan Liao, and Xiaofeng Shi. Risks of large portfolios. Journal of Econometrics, 186(2):367–387, June 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000391.

Fang:2015:DDS

[FLWZ15] Ying Fang, Qi Li, Ximing Wu, and Daiqiang Zhang. A data-driven smooth test of symmetry. Journal of Econometrics, 188(2):490–501, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000822.

Lee:2010:ESA

[fLY10] Lung fei Lee and Jihai Yu. Estimation of spatial autoregressive panel data models with fixed effects. Journal of Econometrics, 154(2):165–185, February 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440760900178X.

Lee:2014:EGE

[fLY14] Lung fei Lee and Jihai Yu. Efficient GMM estimation of spatial dynamic panel data models with fixed effects. Journal of Econometrics, 180(2):174–197, June 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000438.

Lee:2015:EFE

[fLY15] Lung fei Lee and Jihai Yu. Estimation of fixed effects panel regression models with separable and nonseparable space-time filters. Journal of Econometrics, 184(1):174–192, January 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001857.

Francq:2011:TSN

[FLZ11] Christian Francq, Guillaume Lepage, and Jean-Michel Zakoïan. Two-stage non Gaussian QML estimation of GARCH models and testing the efficiency of the Gaussian QMLE. Journal of Econometrics, 165(2):246–257, December 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761100159X.

Frahm:2010:DEM

[FM10] Gabriel Frahm and Christoph Memmel. Dominating estimators for minimum-variance portfolios. Journal of Econometrics, 159(2):289–302, December 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610001594.

Flabbi:2012:EJF

[FM12] Luca Flabbi and Andrea Moro. The effect of job flexibility on female labor market outcomes: Estimates from a search and bargaining model. Journal of Econometrics, 168(1):81–95, May 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001667.

Filipovic:2013:DAM

[FMS13] Damir Filipović, Eberhard Mayerhofer, and Paul Schneider. Density approximations for multivariate affine jump-diffusion processes. Journal of Econometrics, 176(2):93–111, October 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000596.

Fengler:2015:SSB

[FMV15] M. R. Fengler, E. Mammen, and M. Vogt. Specification and structural break tests for additive models with applications to realized variance data. Journal of Econometrics, 188(1):196–218, September 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615001438.

Forneron:2018:ASE

[FN18] Jean-Jacques Forneron and Serena Ng. The ABC of simulation estimation with auxiliary statistics. Journal of Econometrics, 205(1):112–139, July 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618300460.

Frederiksen:2012:LPW

[FNN12] Per Frederiksen, Frank S. Nielsen, and Morten Ørregaard Nielsen. Local polynomial Whittle estimation of perturbed fractional processes. Journal of Econometrics, 167(2):426–447, April 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002077.

Frazier:2019:IIN

[FOZ19] David T. Frazier, Tatsushi Oka, and Dan Zhu. Indirect inference with a non-smooth criterion function. Journal of Econometrics, 212(2):623–645, October 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619301435.

Fanelli:2010:SAC

[FP10] Luca Fanelli and Paolo Paruolo. Speed of adjustment in cointegrated systems. Journal of Econometrics, 158(1):130–141, September 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761000062X.

Fleming:2011:HFR

[FP11a] Jeff Fleming and Bradley S. Paye. High-frequency returns, jumps and the mixture of normals hypothesis. Journal of Econometrics, 160(1):119–128, January 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000667.

Franchi:2011:CVA

[FP11b] Massimo Franchi and Paolo Paruolo. A characterization of vector autoregressive processes with common cyclical features. Journal of Econometrics, 163(1):105–117, July 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610002137.

Fan:2012:CIQ

[FP12] Yanqin Fan and Sang Soo Park. Confidence intervals for the quantile of treatment effects in randomized experiments.

Journal of Econometrics, 167(2):330–344, April 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611002004.

Fan:2014:NIC

[FP14] Yanqin Fan and Sang Soo Park. Nonparametric inference for counterfactual means: Bias-correction, confidence sets, and weak IV. Journal of Econometrics, 178(Part 1):45–56, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001553.

Feng:2019:SPS

[FPSY19] Guohua Feng, Bin Peng, Liangjun Su, and Thomas Tao Yang. Semi-parametric single-index panel data models with interactive fixed effects: Theory and practice. Journal of Econometrics, 212(2):607–622, October 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619301459.

Frazier:2017:ETS

[FR17] David T. Frazier and Eric Renault. Efficient two-step estimation via targeting. Journal of Econometrics, 201(2):212–227, December 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301549.

Freyberger:2018:UCB

[FR18] Joachim Freyberger and Yoshiyasu Rai. Uniform confidence bands: Characterization and optimality. Journal of Econometrics, 204(1):119–130, May 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618300174.

Firpo:2019:PIT

[FR19] Sergio Firpo and Geert Ridder. Partial identification of the treatment effect distribution and its functionals. Journal of Econometrics, 213(1):210–234, November 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300673.

Freyberger:2015:ATD

[Fre15] Joachim Freyberger. Asymptotic theory for differentiated products demand models with many markets. Journal of Econometrics, 185(1):162–181, March 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614002474.

Feng:2010:PDT

[FS10] Guohua Feng and Apostolos Serletis. A primal divisia technical change index based on the output distance function. Journal of Econometrics, 159(2):320–330, December 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001867.

Florens:2012:NEI

[FS12] Jean-Pierre Florens and Anna Simoni. Nonparametric estimation of an instrumental regression: a quasi-Bayesian approach based on regularized posterior. Journal of Econometrics, 170(2):458– 475, October 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001285.

Feng:2014:UOP

[FS14] Guohua Feng and Apostolos Serletis. Undesirable outputs and a primal divisia productivity index based on the directional output distance function. Journal of Econometrics, 183(1):135–146, November 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001559.

Fox:2016:PDM

[FS16] Kevin J. Fox and Iqbal A. Syed. Price discounts and the measurement of inflation. Journal of Econometrics, 191(2):398–406, April 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002869.

Fiorentini:2019:CNG

[FS19] Gabriele Fiorentini and Enrique Sentana. Consistent non-Gaussian pseudo maximum likelihood estimators. Journal of Econometrics, 213(2):321–358, December 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619301423.

Florens:2014:FEN

[FSV14] Jean-Pierre Florens, Léopold Simar, and Ingrid Van Keilegom. Frontier estimation in nonparametric location-scale models. Journal of Econometrics, 178(Part 3):456–470, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001504.

Fruhwirth-Schnatter:2010:SMS

[FSW10] Sylvia Frühwirth-Schnatter and Helga Wagner. Stochastic model specification search for Gaussian and partial non-Gaussian state space models. Journal of Econometrics, 154(1):85–100, January 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609001614.

Fang:2014:IBR

[FT14] Hanming Fang and Xun Tang. Inference of bidders’ risk attitudes in ascending auctions with endogenous entry. Journal of Econometrics, 180(2):198–216, June 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000426.

Florens:2015:IVE

[FV15] Jean-Pierre Florens and Sébastien Van Bellegem. Instrumental variable estimation in functional linear models. Journal of Econometrics, 186(2):465–476, June 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000445.

Fernandez-Villaverde:2015:EDE

[FVGQRR15] Jesús Fernández-Villaverde, Pablo Guerrón-Quintana, and Juan F. Rubio-Ramírez. Estimating dynamic equilibrium models with stochastic volatility. Journal of Econometrics, 185(1):216– 229, March 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614002310.

Fernandez-Val:2011:BCT

[FVV11] Iván Fernández-Val and Francis Vella. Bias corrections for two-step fixed effects panel data estimators. Journal of Econometrics, 163(2):144–162, August 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611000649.

Fernandez-Val:2016:ITE

[FVW16] Iván Fernández-Val and Martin Weidner. Individual and time effects in nonlinear panel models with large N, T. Journal of Econometrics, 192(1):291–312, May 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002997.

Francq:2013:GMP

[FWZ13] Christian Francq, Olivier Wintenberger, and Jean-Michel Zakoïan. GARCH models without positivity constraints: Exponential or log GARCH? Journal of Econometrics, 177(1):34–46, November 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613001267.

Fan:2019:RCE

[FWZ19] Jianqing Fan, Weichen Wang, and Yiqiao Zhong. Robust covariance estimation for approximate factor models. Journal of Econometrics, 208(1):5–22, January 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301684.

Fan:2017:SFU

[FXY17] Jianqing Fan, Lingzhou Xue, and Jiawei Yao. Sufficient forecasting using factor models. Journal of Econometrics, 201(2):292– 306, December 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617301616.

Francq:2010:IMI

[FZ10] Christian Francq and Jean-Michel Zakoïan. Inconsistency of the MLE and inference based on weighted LS for LARCH models. Journal of Econometrics, 159(1):151–165, November 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001284.

Francq:2015:RPE

[FZ15] Christian Francq and Jean-Michel Zakoïan. Risk-parameter estimation in volatility models. Journal of Econometrics, 184(1):158–173, January 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001766.

Francq:2018:ERV

[FZ18] Christian Francq and Jean-Michel Zakoïan. Estimation risk for the VaR of portfolios driven by semi-parametric multivariate models. Journal of Econometrics, 205(2):381–401, August 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300575.

Geweke:2011:OPP

[GA11] John Geweke and Gianni Amisano. Optimal prediction pools. Journal of Econometrics, 164(1):130–141, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000455.

Galvao:2011:QRD

[Gal11] Antonio F. Galvao. Quantile regression for dynamic panel data with fixed effects. Journal of Econometrics, 164(1):142–157, September 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611000443.

Galvao:2017:DRD

[Gal17] Ana Beatriz Galvão. Data revisions and DSGE models. Journal of Econometrics, 196(1):215–232, January 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407616301701.

Gayle:2013:ICE

[Gay13] Wayne-Roy Gayle. Identification and N-consistent estimation of a nonlinear panel data model with correlated unobserved effects. Journal of Econometrics, 175(2):71–83, August 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000584.

Gomez-Biscarri:2015:RBAb

[GBH15a] Javier Gomez-Biscarri and Javier Hualde. Regression-based analysis of cointegration systems. Journal of Econometrics, 186(1): 32–50, May 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614002978.

Gomez-Biscarri:2015:RBAa

[GBH15b] Javier Gomez-Biscarri and Javier Hualde. A residual-based ADF test for stationary cointegration in I(2) settings. Journal of Econometrics, 184(2):280–294, February 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001882.

Geweke:2019:SAB

[GD19] John Geweke and Garland Durham. Sequentially adaptive Bayesian learning algorithms for inference and optimization. Journal of Econometrics, 210(1):4–25, May 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618302021.

Georgiev:2010:MBA

[Geo10] Iliyan Georgiev. Model-based asymptotic inference on the effect of infrequent large shocks on cointegrated variables. Journal of Econometrics, 158(1):37–50, September 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000473.

Geweke:1996:BRR

[Gew96] John Geweke. Bayesian reduced rank regression in econometrics. Journal of Econometrics, 75(1):121–146, November 1996. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407695017739. See corrigendum [Kar17].

Geweke:2012:NBM

[Gew12] John Geweke. Nonparametric Bayesian modelling of monotone preferences for discrete choice experiments. Journal of Econometrics, 171(2):185–204, December 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001509.

Gagliardini:2017:DIV

[GG17] Patrick Gagliardini and Christian Gouriéroux. Double instrumental variable estimation of interaction models with big data. Journal of Econometrics, 201(2):176–197, December 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301495.

Gagliardini:2019:ILT

[GG19] Patrick Gagliardini and Christian Gouriéroux. Identification by Laplace transforms in nonlinear time series and panel models with unobserved stochastic dynamic effects. Journal of Econometrics, 208(2):613–637, February 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618302185.

Gan:2015:SMR

[GGHM15] Li Gan, Guan Gong, Michael Hurd, and Daniel McFadden. Subjective mortality risk and bequests. Journal of Econometrics, 188(2):514–525, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000846.

Golosnoy:2012:CAW

[GGL12] Vasyl Golosnoy, Bastian Gribisch, and Roman Liesenfeld. The conditional autoregressive Wishart model for multivariate stock market volatility. Journal of Econometrics, 167(1):211–223, March 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002582.

Guay:2013:RAR

[GGL13] Alain Guay, Emmanuel Guerre, and Stepána Lazarová. Robust adaptive rate-optimal testing for the white noise hypothesis. Journal of Econometrics, 176(2):134–145, October 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613001152.

Gallant:2017:BES

[GGR17] A. Ronald Gallant, Raffaella Giacomini, and Giuseppe Ragusa. Bayesian estimation of state space models using moment conditions. Journal of Econometrics, 201(2):198–211, December 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301501.

Galichon:2013:DB

[GH13] Alfred Galichon and Marc Henry. Dilation bootstrap. Journal of Econometrics, 177(1):109–115, November 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613001292.

Griffiths:2016:SMS

[GH16] William E. Griffiths and Gholamreza Hajargasht. Some models for stochastic frontiers with endogeneity. Journal of Econometrics, 190(2):341–348, February 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001815.

Giessing:2019:PRM

[GH19] Alexander Giessing and Xuming He. On the predictive risk in misspecified quantile regression. Journal of Econometrics, 213(1):235–260, November 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619300685.

Ghanem:2017:TIA

[Gha17] Dalia Ghanem. Testing identifying assumptions in nonseparable panel data models. Journal of Econometrics, 197(2):202– 217, April 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616302159.

Gallant:2018:BAE

[GHK18] A. Ronald Gallant, Han Hong, and Ahmed Khwaja. A Bayesian approach to estimation of dynamic models with small and large number of heterogeneous players and latent serially correlated states. Journal of Econometrics, 203(1):19–32, March 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617302336.

Georgiev:2018:TPI

[GHLT18] Iliyan Georgiev, David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Testing for parameter instability in predictive regression models. Journal of Econometrics, 204(1):101– 118, May 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618300095.

Granziera:2014:PTS

[GHM14] Eleonora Granziera, Kirstin Hubrich, and Hyungsik Roger Moon. A predictability test for a small number of nested models. Journal of Econometrics, 182(1):174–185, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761400075X.

Ghysels:2016:TGC

[GHM16] Eric Ghysels, Jonathan B. Hill, and Kaiji Motegi. Testing for Granger causality with mixed frequency data. Journal of Econometrics, 192(1):207–230, May 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615003024.

Graham:2018:QCR

[GHPP18] Bryan S. Graham, Jinyong Hahn, Alexandre Poirier, and James L. Powell. A quantile correlated random coefficients panel data model. Journal of Econometrics, 206(2):305–335, October 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300952.

Gotz:2016:TGC

[GHS16] Thomas B. Götz, Alain Hecq, and Stephan Smeekes. Testing for Granger causality in large mixed-frequency VARs. Journal of Econometrics, 193(2):418–432, August 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300768.

Gan:2014:MST

[GHX14] Li Gan, Cheng Hsiao, and Shu Xu. Model specification test with correlated but not cointegrated variables. Journal of Econometrics, 178(Part 1):80–85, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001589.

Ghysels:2016:MRM

[Ghy16] Eric Ghysels. Macroeconomics and the reality of mixed frequency data. Journal of Econometrics, 193(2):294–314, August 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300653.

Gilleskie:2010:WAD

[Gil10] Donna Gilleskie. Work absences and doctor visits during an illness episode: The differential role of preferences, production, and policies among men and women. Journal of Econometrics, 156(1):148–163, May 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002115.

Geweke:2011:IPM

[GJ11] John Geweke and Yu Jiang. Inference and prediction in a multiple-structural-break model. Journal of Econometrics, 163(2):172–185, August 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611000674.

Gourieroux:2017:NVA

[GJ17] Christian Gourieroux and Joann Jasiak. Noncausal vector autoregressive process: Representation, identification and semiparametric estimation. Journal of Econometrics, 200(1):118–134, September 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300799.

Gourieroux:2018:MNO

[GJ18] Christian Gourieroux and Joann Jasiak. Misspecification of noncausal order in autoregressive processes. Journal of Econometrics, 205(1):226–248, July 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618300514.

Gan:2015:NES

[GJZ15] Li Gan, Gaosheng Ju, and Xi Zhu. Nonparametric estimation of structural labor supply and exact welfare change under nonconvex piecewise-linear budget sets. Journal of Econometrics, 188(2):526–544, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000858.

Gilleskie:2010:SMO

[GK10] Donna B. Gilleskie and Ahmed Khwaja. Structural models of optimization behavior in labor, aging and health. Journal of Econometrics, 156(1):1–2, May 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609001936.

Goncalves:2015:BIL

[GK15] Sílvia Gonçalves and Maximilien Kaffo. Bootstrap inference for linear dynamic panel data models with individual fixed effects. Journal of Econometrics, 186(2):407–426, June 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761500041X.

Galvao:2016:SQR

[GK16] Antonio F. Galvao and Kengo Kato. Smoothed quantile regression for panel data. Journal of Econometrics, 193(1):92–112, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300239.

Goldman:2017:FOS

[GK17] Matt Goldman and David M. Kaplan. Fractional order statistic approximation for nonparametric conditional quantile inference. Journal of Econometrics, 196(2):331–346, February 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301944.

Goldman:2018:CDM

[GK18a] Matt Goldman and David M. Kaplan. Comparing distributions by multiple testing across quantiles or CDF values. Journal of Econometrics, 206(1):143–166, September 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300812.

Grammig:2018:TSI

[GK18b] Joachim Grammig and Eva-Maria Küchlin. A two-step indirect inference approach to estimate the long-run risk asset pricing model. Journal of Econometrics, 205(1):6–33, July 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300423.

Guo:2018:TEH

[GKCS18] Zijian Guo, Hyunseung Kang, T. Tony Cai, and Dylan S. Small. Testing endogeneity with high dimensional covariates. Journal of Econometrics, 207(1):175–187, November 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301325.

Grothe:2014:MME

[GKM14] Oliver Grothe, Volodymyr Korniichuk, and Hans Manner. Modeling multivariate extreme events using self-exciting point processes. Journal of Econometrics, 182(2):269–289, October 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761400089X.

Gospodinov:2017:SMD

[GKN17] Nikolay Gospodinov, Ivana Komunjer, and Serena Ng. Simulated minimum distance estimation of dynamic models with errorsin-variables. Journal of Econometrics, 200(2):181–193, October 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300854.

Geweke:2012:IAI

[GKP12a] John Geweke, Gary Koop, and Richard Paap. Introduction for the annals issue of the journal of econometrics on “Bayesian models, methods and applications”. Journal of Econometrics, 171(2):99–100, December 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001492.

Gundersen:2012:INS

[GKP12b] Craig Gundersen, Brent Kreider, and John Pepper. The impact of the national school lunch program on child health: a nonparametric bounds analysis. Journal of Econometrics, 166(1): 79–91, January 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001205.

Giraitis:2013:AFP

[GKP13] Liudas Giraitis, George Kapetanios, and Simon Price. Adaptive forecasting in the presence of recent and ongoing structural change. Journal of Econometrics, 177(2):153–170, December 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000699.

Gospodinov:2013:CST

[GKR13] Nikolay Gospodinov, Raymond Kan, and Cesare Robotti. Chisquared tests for evaluation and comparison of asset pricing models. Journal of Econometrics, 173(1):108–125, March 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612002485.

Gao:2015:MTM

[GKS15] Jiti Gao, Nam Hyun Kim, and Patrick W. Saart. A misspecification test for multiplicative error models of non-negative time series processes. Journal of Econometrics, 189(2):346–359, December 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001062.

Glass:2016:SAS

[GKS16] Anthony J. Glass, Karligash Kenjegalieva, and Robin C. Sickles. A spatial autoregressive stochastic frontier model for panel data with asymmetric efficiency spillovers. Journal of Econometrics, 190(2):289–300, February 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001803.

Giraitis:2014:IST

[GKY14] L. Giraitis, G. Kapetanios, and T. Yates. Inference on stochastic time-varying coefficient models. Journal of Econometrics, 179(1): 46–65, March 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613002248.

Gu:2014:TCR

[GL14] Jingping Gu and Zhongwen Liang. Testing cointegration relationship in a semiparametric varying coefficient model. Journal of Econometrics, 178(Part 1):57–70, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001565.

Gan:2016:ETT

[GL16] Li Gan and Qi Li. Efficiency of thin and thick markets. Journal of Econometrics, 192(1):40–54, May 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002742.

Green:2015:TES

[GLH15] Carl Green, Wei Long, and Cheng Hsiao. Testing error serial correlation in fixed effects nonparametric panel data models. Journal of Econometrics, 188(2):466–473, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000809.

Gao:2015:BRC

[GLL15] Yichen Gao, Cong Li, and Zhongwen Liang. Binary response correlated random coefficient panel data models. Journal of Econometrics, 188(2):421–434, October 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000779.

Guo:2019:SST

[GLL19] Shaojun Guo, Dong Li, and Muyi Li. Strict stationarity testing and GLAD estimation of double autoregressive models. Journal of Econometrics, 211(2):319–337, August 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300466.

Garcia:2011:EOR

[GLPR11] René Garcia, Marc-André Lewis, Sergio Pastorello, and Éric Renault. Estimation of objective and risk-neutral distributions based on moments of integrated volatility. Journal of Econometrics, 160(1):22–32, January 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000539.

Gilboa:2011:SBA

[GLS11] Itzhak Gilboa, Offer Lieberman, and David Schmeidler. A similarity-based approach to prediction. Journal of Econometrics, 162(1):124–131, May 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002681.

Goncalves:2011:BCT

[GM11] Sílvia Gonçalves and Nour Meddahi. Box–Cox transforms for realized volatility. Journal of Econometrics, 160(1):129–144, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000680.

Gourieroux:2015:PFD

[GM15] C. Gourieroux and A. Monfort. Pricing with finite dimensional dependence. Journal of Econometrics, 187(2):408–417, August 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000524.

Ghysels:2016:EAM

[GM16] Eric Ghysels and Massimiliano Marcellino. The econometric analysis of mixed frequency data sampling. Journal of Econometrics, 193(2):291–293, August 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300641.

Garcia:2017:STU

[GM17] Tanya P. Garcia and Yanyuan Ma. Simultaneous treatment of unspecified heteroskedastic model error distribution and mismeasured covariates for restricted moment models. Journal of Econometrics, 200(2):194–206, October 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300908.

Greenaway-McGrevy:2012:ADF

[GMHS12] Ryan Greenaway-McGrevy, Chirok Han, and Donggyu Sul. Asymptotic distribution of factor augmented estimators for panel regression. Journal of Econometrics, 169(1):48–53, July 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000048.

Gafarov:2018:DMI

[GMO18] Bulat Gafarov, Matthias Meier, and José Luis Montiel Olea. Delta-method inference for a class of set-identified SVARs. Journal of Econometrics, 203(2):316–327, April 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617302440.

Goncalves:2017:TEA

[GMP17] Sílvia Gonçalves, Michael W. McCracken, and Benoit Perron. Tests of equal accuracy for nested models with estimated factors. Journal of Econometrics, 198(2):231–252, June 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300180.

Gourieroux:2014:EPD

[GMR14a] C. Gouriéroux, A. Monfort, and J. P. Renne. Erratum to “Pricing default events: Surprise, exogeneity and contagion” [J. Econometrics 182(2) (2014) 397–411]. Journal of Econometrics, 183(2):150, December 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614002280. See [GMR14b].

Gourieroux:2014:PDE

[GMR14b] C. Gouriéroux, A. Monfort, and J. P. Renne. Pricing default events: Surprise, exogeneity and contagion. Journal of Econometrics, 182(2):397–411, October 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001080. See erratum [GMR14a].

Gourieroux:2017:SII

[GMR17] Christian Gouriéroux, Alain Monfort, and Jean-Paul Renne. Statistical inference for independent component analysis: Application to structural VAR models. Journal of Econometrics, 196(1):111–126, January 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407616301749.

Gao:2019:BST

[GMWY19] Zhaoxing Gao, Yingying Ma, Hansheng Wang, and Qiwei Yao. Banded spatio-temporal autoregressions. Journal of Econometrics, 208(1):211–230, January 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301775.

Gayle:2013:ENP

[GN13] Wayne-Roy Gayle and Soiliou Daw Namoro. Estimation of a nonlinear panel data model with semiparametric individual effects. Journal of Econometrics, 175(1):46–59, July 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000614.

Griffin:2011:CMP

[GO11] Jim E. Griffin and Roel C. A. Oomen. Covariance measurement in the presence of non-synchronous trading and market microstructure noise. Journal of Econometrics, 160(1):58–68, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000576.

Gospodinov:2012:LGE

[GO12] Nikolay Gospodinov and Taisuke Otsu. Local GMM estimation of time series models with conditional moment restrictions. Journal of Econometrics, 170(2):476–490, October 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001297.

Gagliardini:2019:DCA

[GOS19] Patrick Gagliardini, Elisa Ossola, and Olivier Scaillet. A diagnostic criterion for approximate factor structure. Journal of Econometrics, 212(2):503–521, October 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761930137X.

Giraitis:2012:MAF

[GP12] Liudas Giraitis and Peter C. B. Phillips. Mean and autocovariance function estimation near the boundary of stationarity. Journal of Econometrics, 169(2):166–178, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000309.

Gao:2013:SET

[GP13] Jiti Gao and Peter C. B. Phillips. Semiparametric estimation in triangular system equations with nonstationarity. Journal of Econometrics, 176(1):59–79, September 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761300095X.

Geweke:2014:LBI

[GP14a] John Geweke and Lea Petrella. Likelihood-based inference for regular functions with fractional polynomial approximations. Journal of Econometrics, 183(1):22–30, November 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001481.

Goncalves:2014:BFA

[GP14b] Sílvia Gonçalves and Benoit Perron. Bootstrapping factoraugmented regression models. Journal of Econometrics, 182 (1):156–173, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614000748.

Gourieroux:2010:IID

[GPY10] Christian Gouriéroux, Peter C. B. Phillips, and Jun Yu. Indirect inference for dynamic panel models. Journal of Econometrics, 157(1):68–77, July 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002772.

Guerron-Quintana:2017:IRM

[GQIK17] Pablo Guerron-Quintana, Atsushi Inoue, and Lutz Kilian. Impulse response matching estimators for DSGE models. Journal of Econometrics, 196(1):144–155, January 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407616301762.

Gagliardini:2013:SPE

[GR13] Patrick Gagliardini and Diego Ronchetti. Semi-parametric estimation of American option prices. Journal of Econometrics, 173(1):57–82, March 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612002345.

Giacomini:2014:TCF

[GR14] Raffaella Giacomini and Giuseppe Ragusa. Theory-coherent forecasting. Journal of Econometrics, 182(1):145–155, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000736.

Gupta:2015:IHO

[GR15] Abhimanyu Gupta and Peter M. Robinson. Inference on higher-order spatial autoregressive models with increasingly many parameters. Journal of Econometrics, 186(1):19–31, May 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761400298X.

Golombek:2018:EDS

[GR18a] Rolf Golombek and Arvid Raknerud. Exit dynamics of start-up firms: Structural estimation using indirect inference. Journal of Econometrics, 205(1):204–225, July 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300502.

Gupta:2018:PML

[GR18b] Abhimanyu Gupta and Peter M. Robinson. Pseudo maximum likelihood estimation of spatial autoregressive models with increasing dimension. Journal of Econometrics, 202(1):92–107, January 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407617301458.

Granger:2010:STD

[Gra10] Clive W. J. Granger. Some thoughts on the development of cointegration. Journal of Econometrics, 158(1):3–6, September 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000448.

Granger:2012:UCS

[Gra12] Clive W. J. Granger. Useful conclusions from surprising results. Journal of Econometrics, 169(2):142–146, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000413.

Guggenberger:2012:GSU

[GRS12] Patrik Guggenberger, Joaquim J. S. Ramalho, and Richard J. Smith. GEL statistics under weak identification. Journal of Econometrics, 170(2):331–349, October 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001212.

Garcia:2011:ESD

[GRV11] René Garcia, Eric Renault, and David Veredas. Estimation of stable distributions by indirect inference. Journal of Econometrics, 161(2):325–337, April 1, 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610002514.

Ghysels:2011:VFM

[GS11a] Eric Ghysels and Arthur Sinko. Volatility forecasting and microstructure noise. Journal of Econometrics, 160(1):257–271, January 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000771.

Griffin:2011:SBA

[GS11b] J. E. Griffin and M. F. J. Steel. Stick-breaking autoregressive processes. Journal of Econometrics, 162(2):383–396, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000613.

Gagliardini:2012:TRN

[GS12] Patrick Gagliardini and Olivier Scaillet. Tikhonov regularization for nonparametric instrumental variable estimators. Journal of Econometrics, 167(1):61–75, March 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611002375.

Gossner:2013:FSE

[GS13] Olivier Gossner and Karl H. Schlag. Finite-sample exact tests for linear regressions with bounded dependent variables. Journal of Econometrics, 177(1):75–84, November 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613001346.

Gencay:2015:MST

[GS15] Ramazan Gençay and Daniele Signori. Multi-scale tests for serial correlation. Journal of Econometrics, 184(1):62–80, January 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001754.

Giesecke:2018:FLP

[GS18] Kay Giesecke and Gustavo Schwenkler. Filtered likelihood for point processes. Journal of Econometrics, 204(1):33–53, May 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300058.

Giesecke:2019:SLE

[GS19] K. Giesecke and G. Schwenkler. Simulated likelihood estimators for discretely observed jump-diffusions. Journal of Econometrics, 213(2):297–320, December 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619301460.

Gallant:2018:EBM

[GT18] A. Ronald Gallant and George Tauchen. Exact Bayesian moment based inference for the distribution of the small-time movements of an Itô semimartingale. Journal of Econometrics, 205(1):140– 155, July 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618300472.

Gao:2013:ETA

[GTY13] Jiti Gao, Dag Tjøstheim, and Jiying Yin. Estimation in threshold autoregressive models with a stationary and a unit root regime. Journal of Econometrics, 172(1):1–13, January 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612002047.

Guggenberger:2010:IHP

[Gug10] Patrik Guggenberger. The impact of a Hausman pretest on the size of a hypothesis test: The panel data case.

Journal of Econometrics, 156(2):337–343, June 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002929.

Gupta:2018:ASS

[Gup18a] Abhimanyu Gupta. Autoregressive spatial spectral estimates. Journal of Econometrics, 203(1):80–95, March 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617302361.

Gupta:2018:NST

[Gup18b] Abhimanyu Gupta. Nonparametric specification testing via the trinity of tests. Journal of Econometrics, 203(1):169–185, March 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617302403.

Gutknecht:2016:TMU

[Gut16] Daniel Gutknecht. Testing for monotonicity under endogeneity: an application to the reservation wage function. Journal of Econometrics, 190(1):100–114, January 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002341.

Gu:2019:PDQ

[GV19] Jiaying Gu and Stanislav Volgushev. Panel data quantile regression with grouped fixed effects. Journal of Econometrics, 213(1):68–91, November 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619300612.

Golinski:2016:LMA

[GZ16] Adam Goliński and Paolo Zaffaroni. Long memory affine term structure models. Journal of Econometrics, 191(1):33–56, March 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002523.

Gao:2016:MAB

[GZWZ16] Yan Gao, Xinyu Zhang, Shouyang Wang, and Guohua Zou. Model averaging based on leave-subject-out cross-validation. Journal of Econometrics, 192(1):139–151, May 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615003012.

Hill:2013:MCT

[HA13] Jonathan B. Hill and Mike Aguilar. Moment condition tests for heavy tailed time series. Journal of Econometrics, 172(2):255– 274, February 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001972.

Hagemann:2012:STR

[Hag12] Andreas Hagemann. A simple test for regression specification with non-nested alternatives. Journal of Econometrics, 166(2):247–254, February 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611002181.

Hagemann:2019:PIT

[Hag19] Andreas Hagemann. Placebo inference on treatment effects when the number of clusters is small. Journal of Econometrics, 213(1):190–209, November 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619300661.

Hansen:2010:AEA

[Han10] Bruce E. Hansen. Averaging estimators for autoregressions with a near unit root. Journal of Econometrics, 158(1):142–155, September 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000643.

Hansen:2012:PLS

[Han12] Lars Peter Hansen. Proofs for large sample properties of generalized method of moments estimators. Journal of Econometrics, 170(2):325–330, October 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612001200.

Han:2015:TOR

[Han15] Xu Han. Tests for overidentifying restrictions in factoraugmented VAR models. Journal of Econometrics, 184(2):394– 419, February 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614002425.

Hansen:2016:ESP

[Han16] Bruce E. Hansen. Efficient shrinkage in parametric models. Journal of Econometrics, 190(1):115–132, January 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002365.

Han:2018:EID

[Han18] Xu Han. Estimation and inference of dynamic structural factor models with over-identifying restrictions. Journal of Econometrics, 202(2):125–147, February 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617301896.

Hassler:2011:EFI

[Has11] Uwe Hassler. Estimation of fractional integration under temporal aggregation. Journal of Econometrics, 162(2):240–247, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000145.

Hausman:2019:EL

[Hau19] Jerry Hausman. An econometric life. Journal of Econometrics, 211(1):4–10, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618302306.

Hayakawa:2010:EDF

[Hay10] Kazuhiko Hayakawa. The effects of dynamic feedbacks on LS and MM estimator accuracy in panel data models: Some additional results. Journal of Econometrics, 159(1):202–208, November 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001417.

He:2016:TSD

[HC16] Jing He and Song Xi Chen. Testing super-diagonal structure in high dimensional covariance matrices. Journal of Econometrics, 194(2):283–297, October 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616300975.

Hall:2012:IRM

[HHB12] Alastair R. Hall, Sanggohn Han, and Otilia Boldea. Inference regarding multiple structural changes in linear models with endogenous regressors. Journal of Econometrics, 170(2):281–302, October 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001182.

Hahn:2011:HTW

[HHM11] Jinyong Hahn, John C. Ham, and Hyungsik Roger Moon. The Hausman test and weak instruments. Journal of Econometrics, 160(2):289–299, February 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610001892.

Hoderlein:2017:TMR

[HHM17] Stefan Hoderlein, Hajo Holzmann, and Alexander Meister. The triangular model with random coefficients. Journal of Econometrics, 201(1):144–169, November 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300829.

Hicks:2012:SSC

[HHS12] Robert L. Hicks, William C. Horrace, and Kurt E. Schnier. Strategic substitutes or complements? The game of where to fish. Journal of Econometrics, 168(1):70–80, May 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611001709.

Heckman:2016:DTE

[HHV16] James J. Heckman, John Eric Humphries, and Gregory Veramendi. Dynamic treatment effects. Journal of Econometrics, 191(2):276–292, April 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615002778.

Hill:2016:LSA

[Hil16] Robert J. Hill. A least squares approach to imposing within-region fixity in the international comparisons program. Journal of Econometrics, 191(2):407–413, April 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002870.

Hall:2012:ICI

[HINR12] Alastair R. Hall, Atsushi Inoue, James M. Nason, and Barbara Rossi. Information criteria for impulse response function matching estimation of DSGE models. Journal of Econometrics, 170(2):499–518, October 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612001315.

See corrigendum [HINR14].

Hall:2014:CT

[HINR14] Alastair R. Hall, Atsushi Inoue, James M. Nason, and Barbara Rossi. Corrigendum to “Information criteria for impulse response function matching estimation of DSGE models” [J. Econom. 170 (2012) 499–518]. Journal of Econometrics, 178(Part 3):706, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613002029. See [HINR12].

Hsu:2011:ECM

[HK11] Shih-Hsun Hsu and Chung-Ming Kuan. Estimation of conditional moment restrictions without assuming parameter identifiability in the implied unconditional moments. Journal of Econometrics, 165(1):87–99, November 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761100100X.

Hansen:2014:IVE

[HK14a] Christian Hansen and Damian Kozbur. Instrumental variables estimation with many weak instruments using regularized JIVE. Journal of Econometrics, 182(2):290–308, October 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000918.

Hsu:2014:CST

[HK14b] Shih-Hsun Hsu and Chung-Ming Kuan. Constructing smooth tests without estimating the eigenpairs of the limiting process. Journal of Econometrics, 178(Part 1):71–79, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001577.

Henderson:2015:SCE

[HKLP15] Daniel J. Henderson, Subal C. Kumbhakar, Qi Li, and Christopher F. Parmeter. Smooth coefficient estimation of a seemingly unrelated regression. Journal of Econometrics, 189(1):148–162, November 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002043.

Horvath:2014:TSF

[HKR14] Lajos Horváth, Piotr Kokoszka, and Gregory Rice. Testing stationarity of functional time series. Journal of Econometrics, 179(1):66–82, March 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613002327.

Halbleib:2018:IAE

[HKRV18] Roxana Halbleib, Dennis Kristensen, Eric Renault, and David Veredas. Issue of the annals of econometrics on indirect estimation methods in finance and economics. Journal of Econometrics, 205(1):1–5, July 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618300411.

Hong:2012:EI

[HKW12] Han Hong, Chung-Ming Kuan, and Yoon-Jae Whang. Editors’ introduction. Journal of Econometrics, 167(2):295–296, April 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001965.

Huang:2019:IET

[HKY19] Liquan Huang, Umair Khalil, and Nese Yildiz. Identification and estimation of a triangular model with multiple endogenous variables and insufficiently many instrumental variables. Journal of Econometrics, 208(2):346–366, February 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301854.

Hafner:2010:EEM

[HL10] Christian M. Hafner and Oliver Linton. Efficient estimation of a multivariate multiplicative volatility model. Journal of Econometrics, 159(1):55–73, November 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001132.

Hallin:2011:DFP

[HL11a] Marc Hallin and Roman Liska. Dynamic factors in the presence of blocks. Journal of Econometrics, 163(1):29–41, July 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610002083.

Heiman:2011:EIA

[HL11b] Amir Heiman and Oded Lowengart. The effects of information about health hazards in food on consumers’ choice process. Journal of Econometrics, 162(1):140–147, May 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001466.

Horowitz:2012:UCB

[HL12] Joel L. Horowitz and Sokbae Lee. Uniform confidence bands for functions estimated nonparametrically with instrumental variables. Journal of Econometrics, 168(2):175–188, June 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002739.

Harding:2014:ETQ

[HL14a] Matthew Harding and Carlos Lamarche. Estimating and testing a quantile regression model with interactive effects. Journal of Econometrics, 178(Part 1):101–113, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001607.

Herwartz:2014:SVA

[HL14b] Helmut Herwartz and Helmut Lütkepohl. Structural vector autoregressions with Markov switching: Combining conventional with statistical identification of shocks. Journal of Econometrics, 183(1):104–116, November 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001535.

Harvey:2015:CSD

[HL15] David I. Harvey and Stephen J. Leybourne. Confidence sets for the date of a break in level and trend when the order of integration is unknown. Journal of Econometrics, 184(2):262–279, February 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001894.

Hallin:2017:RES

[HL17a] Marc Hallin and Davide La Vecchia. R-estimation in semiparametric dynamic location-scale models. Journal of Econometrics, 196(2):233–247, February 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616301725.

Hirschberg:2017:IIE

[HL17b] Joe Hirschberg and Jenny Lye. Inverting the indirect — the ellipse and the boomerang: Visualizing the confidence intervals of the structural coefficient from two-stage least squares. Journal of Econometrics, 199(2):173–183, August 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300702.

Horowitz:2017:NEI

[HL17c] Joel L. Horowitz and Sokbae Lee. Nonparametric estimation and inference under shape restrictions. Journal of Econometrics, 201(1):108–126, November 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617301057.

Hong:2018:NDM

[HL18] Han Hong and Jessie Li. The numerical delta method. Journal of Econometrics, 206(2):379–394, October 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300988.

Hale:2019:MBS

[HL19a] Galina Hale and Jose A. Lopez. Monitoring banking system connectedness with big data. Journal of Econometrics, 212(1):203–220, September 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761930082X.

Harding:2019:PQA

[HL19b] Matthew Harding and Carlos Lamarche. A panel quantile approach to attrition bias in Big Data: Evidence from a randomized experiment. Journal of Econometrics, 211(1):61–82, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302355.

Henderson:2012:EIN

[HLM^{+}12] Daniel J. Henderson, John A. List, Daniel L. Millimet, Christopher F. Parmeter, and Michael K. Price. Empirical implementation of nonparametric first-price auction models. Journal of Econometrics, 168(1):17–28, May 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611001710.

Hurn:2013:QML

[HLM13] A. S. Hurn, K. A. Lindsay, and A. J. McClelland. A quasimaximum likelihood method for estimating the parameters of multivariate diffusions. Journal of Econometrics, 172(1):106– 126, January 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612002187.

Hausman:2011:PCE

[HLMN11] Jerry Hausman, Randall Lewis, Konrad Menzel, and Whitney Newey. Properties of the CUE estimator and a modification with moments. Journal of Econometrics, 165(1):45–57, November 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000972.

Han:2016:CQM

[HLOW16] Heejoon Han, Oliver Linton, Tatsushi Oka, and Yoon-Jae Whang. The cross-quantilogram: Measuring quantile dependence and testing directional predictability between time series. Journal of Econometrics, 193(1):251–270, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300458.

Hubbard:2012:SEM

[HLP12] Timothy P. Hubbard, Tong Li, and Harry J. Paarsch. Semiparametric estimation in models of first-price, sealed-bid auctions with affiliation. Journal of Econometrics, 168(1):4–16, May 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001692.

Horrace:2016:ENP

[HLP16] William C. Horrace, Xiaodong Liu, and Eleonora Patacchini. Endogenous network production functions with selectivity. Journal of Econometrics, 190(2):222–232, February 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001748.

Henderson:2015:GBS

[HLPY15] Daniel J. Henderson, Qi Li, Christopher F. Parmeter, and Shuang Yao. Gradient-based smoothing parameter selection for nonparametric regression estimation. Journal of Econometrics, 184(2):233–241, February 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001924.

Ham:2011:MSP

[HLR11] John C. Ham, Xianghong Li, and Patricia B. Reagan. Matching and semi-parametric IV estimation, a distance-based measure of migration, and the wages of young men. Journal of Econometrics, 161(2):208–227, April 1, 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610002460.

Harvey:2006:MTC

[HLT06] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Modified tests for a change in persistence. Journal of Econometrics, 134(2):441–469, October 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407605001521. See corrigendum [HLT12a].

Harvey:2010:RMD

[HLT10] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Robust methods for detecting multiple level breaks in autocorrelated time series. Journal of Econometrics, 157(2):342– 358, August 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000424.

Harvey:2012:CMT

[HLT12a] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Corrigendum to “Modified tests for a change in persistence” [J. Econom. 134 (2006) 441–469]. Journal of Econometrics, 168(2): 407, June 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002570. See [HLT06].

Harvey:2012:TUR

[HLT12b] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Testing for unit roots in the presence of uncertainty over both the trend and initial condition. Journal of Econometrics, 169(2):188– 195, August 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000280.

Harvey:2012:URT

[HLT12c] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Unit root testing under a local break in trend. Journal of Econometrics, 167(1):140–167, March 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611002569.

Harvey:2013:TUR

[HLT13] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Testing for unit roots in the possible presence of multiple trend breaks using minimum Dickey–Fuller statistics. Journal of Econometrics, 177(2):265–284, December 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000894.

Harris:2016:TCI

[HLT16] David Harris, Stephen J. Leybourne, and A. M. Robert Taylor. Tests of the co-integration rank in VAR models in the presence of a possible break in trend at an unknown point. Journal of Econometrics, 192(2):451–467, June 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300136.

Huber:2013:PEB

[HLW13] Martin Huber, Michael Lechner, and Conny Wunsch. The performance of estimators based on the propensity score. Journal of Econometrics, 175(1):1–21, July 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613000390.

Hong:2015:EDD

[HLW15] Han Hong, Weiming Li, and Boyu Wang. Estimation of dynamic discrete models from time aggregated data. Journal of Econometrics, 188(2):435–446, October 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000780.

Hong:2019:CRM

[HLX19] Harrison Hong, Frank Weikai Li, and Jiangmin Xu. Climate risks and market efficiency. Journal of Econometrics, 208(1):265– 281, January 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301817.

Hendry:2014:UEA

[HM14] David F. Hendry and Grayham E. Mizon. Unpredictability in economic analysis, econometric modeling and forecasting. Journal of Econometrics, 182(1):186–195, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000761.

Hillier:2018:EHO

[HM18] Grant Hillier and Federico Martellosio. Exact and higher-order properties of the MLE in spatial autoregressive models, with applications to inference. Journal of Econometrics, 205(2):402– 422, August 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618300587.

Hong:2015:EEN

[HMN15] Han Hong, Aprajit Mahajan, and Denis Nekipelov. Extremum estimation and numerical derivatives. Journal of Econometrics, 188(1):250–263, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615001207.

Hallin:2011:MLD

[HMPV11] Marc Hallin, Charles Mathias, Hugues Pirotte, and David Veredas. Market liquidity as dynamic factors. Journal of Econometrics, 163(1):42–50, July 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610002095.

Holly:2011:FOP

[HMR11] Alberto Holly, Alain Monfort, and Michael Rockinger. Fourth order pseudo maximum likelihood methods. Journal of Econometrics, 162(2):278–293, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761100025X.

Handel:2013:RFP

[HMR13] Benjamin R. Handel, Kanishka Misra, and James W. Roberts. Robust firm pricing with panel data. Journal of Econometrics, 174(2):165–185, June 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613000420.

Hu:2013:IFP

[HMS13] Yingyao Hu, David McAdams, and Matthew Shum. Identification of first-price auctions with non-separable unobserved heterogeneity. Journal of Econometrics, 174(2):186–193, June 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000407.

Hnatkovska:2012:CMC

[HMT12] Viktoria Hnatkovska, Vadim Marmer, and Yao Tang. Comparison of misspecified calibrated models: The minimum distance approach. Journal of Econometrics, 169(1):131–138, July 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000085.

Hong:2012:ELI

[HN12] Han Hong and Denis Nekipelov. Efficient local IV estimation of an empirical auction model. Journal of Econometrics, 168(1): 60–69, May 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001722.

Hahn:2014:NHM

[HNS14] Jinyong Hahn, Whitney K. Newey, and Richard J. Smith. Neglected heterogeneity in moment condition models. Journal of Econometrics, 178(Part 1):86–100, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001590.

Halunga:2012:RBE

[HO12] Andreea G. Halunga and Denise R. Osborn. Ratio-based estimators for a change point in persistence. Journal of Econometrics, 171(1):24–31, November 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612001716.

Ho:2015:SQA

[Ho15] Hwai-Chung Ho. Sample quantile analysis for long-memory stochastic volatility models. Journal of Econometrics, 189 (2):360–370, December 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615001074.

Hoderlein:2011:HMC

[Hod11] Stefan Hoderlein. How many consumers are rational? Journal of Econometrics, 164(2):294–309, October 1, 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761100128X.

Hong:2017:ISC

[Hon17] Shengjie Hong. Inference in semiparametric conditional moment models with partial identification. Journal of Econometrics, 196(1):156–179, January 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407616301816.

Horowitz:2012:STN

[Hor12] Joel L. Horowitz. Specification testing in nonparametric instrumental variable estimation. Journal of Econometrics, 167(2): 383–396, April 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002041.

Horowitz:2014:ANI

[Hor14] Joel L. Horowitz. Adaptive nonparametric instrumental variables estimation: Empirical choice of the regularization parameter. Journal of Econometrics, 180(2):158–173, June 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000463.

Hounyo:2017:BIC

[Hou17] Ulrich Hounyo. Bootstrapping integrated covariance matrix estimators in noisy jump-diffusion models with non-synchronous trading. Journal of Econometrics, 197(1):130–152, March 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616302123.

Hoogerheide:2012:CAI

[HOvD12] Lennart Hoogerheide, Anne Opschoor, and Herman K. van Dijk. A class of adaptive importance sampling weighted EM algorithms for efficient and robust posterior and predictive simulation. Journal of Econometrics, 171(2):101–120, December 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001583.

Halunga:2017:HRB

[HOY17] Andreea G. Halunga, Chris D. Orme, and Takashi Yamagata. A heteroskedasticity robust Breusch–Pagan test for contemporaneous correlation in dynamic panel data models. Journal of Econometrics, 198(2):209–230, June 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300179.

Han:2012:AGP

[HP12a] Heejoon Han and Joon Y. Park. ARCH/GARCH with persistent covariate: Asymptotic theory of MLE. Journal of Econometrics, 167(1):95–112, March 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611002351.

Hong:2012:BAP

[HP12b] Han Hong and Bruce Preston. Bayesian averaging, prediction and nonnested model selection. Journal of Econometrics, 167(2): 358–369, April 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002028.

Han:2013:FDM

[HP13] Chirok Han and Peter C. B. Phillips. First difference maximum likelihood and dynamic panel estimation. Journal of Econometrics, 175(1):35–45, July 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613000572.

Haan:2014:LLC

[HP14a] Peter Haan and Victoria Prowse. Longevity, life-cycle behavior and pension reform. Journal of Econometrics, 178(Part 3):582– 601, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613002042.

Hou:2014:MLW

[HP14b] Jie Hou and Pierre Perron. Modified local Whittle estimator for long memory processes in the presence of low frequency (and other) contaminations. Journal of Econometrics, 182(2):309– 328, October 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001079.

Hayakawa:2015:RSE

[HP15] Kazuhiko Hayakawa and M. Hashem Pesaran. Robust standard errors in transformed likelihood estimation of dynamic panel data models with cross-sectional heteroskedasticity. Journal of Econometrics, 188(1):111–134, September 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001244.

Hill:2016:GEH

[HP16] Jonathan B. Hill and Artem Prokhorov. GEL estimation for heavy-tailed GARCH models with robust empirical likelihood inference. Journal of Econometrics, 190(1):18–45, January 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761500233X.

Hirukawa:2018:CEL

[HP18] Masayuki Hirukawa and Artem Prokhorov. Consistent estimation of linear regression models using matched data. Journal of Econometrics, 203(2):344–358, April 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617302464.

Hall:2014:BIN

[HPK14] Jamie Hall, Michael K. Pitt, and Robert Kohn. Bayesian inference for nonlinear structural time series models. Journal of Econometrics, 179(2):99–111, April 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407613002819.

Holly:2010:STM

[HPY10] Sean Holly, M. Hashem Pesaran, and Takashi Yamagata. A spatio-temporal model of house prices in the USA. Journal of Econometrics, 158(1):160–173, September 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000837.

Hualde:2010:SIM

[HR10] J. Hualde and P. M. Robinson. Semiparametric inference in multivariate fractionally cointegrated systems. Journal of Econometrics, 157(2):492–511, August 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001089.

Hansen:2012:JMA

[HR12a] Bruce E. Hansen and Jeffrey S. Racine. Jackknife model averaging. Journal of Econometrics, 167(1):38–46, March 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002405.

Hong:2012:LUP

[HR12b] Seung-Hyun Hong and Leonardo Rezende. Lock-in and unobserved preferences in server operating systems: a case of Linux vs. Windows. Journal of Econometrics, 167(2):494–503, April 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002120.

Hall:2015:IOC

[HR15a] Peter G. Hall and Jeffrey S. Racine. Infinite order cross-validated local polynomial regression. Journal of Econometrics, 185(2): 510–525, April 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001432.

Horvath:2015:TIB

[HR15b] Lajos Horváth and Gregory Rice. Testing for independence between functional time series. Journal of Econometrics, 189(2):371–382, December 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615001086.

Heckman:2016:ILT

[HR16] James J. Heckman and Lakshmi K. Raut. Intergenerational long-term effects of preschool-structural estimates from a discrete dynamic programming model. Journal of Econometrics, 191(1):164–175, March 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615002493.

Hahn:2017:IVE

[HR17] Jinyong Hahn and Geert Ridder. Instrumental variable estimation of nonlinear models with nonclassical measurement error using control variables. Journal of Econometrics, 200(2):238– 250, October 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300945.

Hahn:2019:TSS

[HR19] Jinyong Hahn and Geert Ridder. Three-stage semi-parametric inference: Control variables and differentiability. Journal of Econometrics, 211(1):262–293, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618302525.

Herbst:2012:EDM

[HS12a] Edward Herbst and Frank Schorfheide. Evaluating DSGE model forecasts of comovements. Journal of Econometrics, 171(2):152– 166, December 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001558.

Hu:2012:NID

[HS12b] Yingyao Hu and Matthew Shum. Nonparametric identification of dynamic models with unobserved state variables. Journal of Econometrics, 171(1):32–44, November 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001479.

Hidalgo:2013:TSS

[HS13a] Javier Hidalgo and Myung Hwan Seo. Testing for structural stability in the whole sample. Journal of Econometrics, 175(2): 84–93, August 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613000626.

Hill:2013:TCV

[HS13b] Jonathan B. Hill and Artyom Shneyerov. Are there common values in first-price auctions? A tail-index nonparametric test. Journal of Econometrics, 174(2):144–164, June 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000377.

Heckman:2014:IIC

[HS14] James J. Heckman and Apostolos Serletis. Introduction to internally consistent modeling, aggregation, inference, and policy. Journal of Econometrics, 183(1):1–4, November 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001468.

Hoderlein:2015:IEC

[HS15a] Stefan Hoderlein and Robert Sherman. Identification and estimation in a correlated random coefficients binary response model. Journal of Econometrics, 188(1):135–149, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761500127X.

Hu:2015:CFE

[HS15b] Yingyao Hu and Yuya Sasaki. Closed-form estimation of nonparametric models with non-classical measurement errors. Journal of Econometrics, 185(2):392–408, April 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614002796.

Hidalgo:2017:ITB

[HS17a] Javier Hidalgo and Marcia Schafgans. Inference and testing breaks in large dynamic panels with strong cross sectional dependence. Journal of Econometrics, 196(2):259–274, February 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301750.

Hwang:2017:AFT

[HS17b] Jungbin Hwang and Yixiao Sun. Asymptotic f and t tests in an efficient GMM setting. Journal of Econometrics, 198(2):277– 295, June 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300246.

Hwang:2018:TSS

[HS18a] Eunju Hwang and Dong Wan Shin. Two-stage stationary bootstrapping for bivariate average realized volatility matrix under market microstructure noise and asynchronicity. Journal of Econometrics, 202(2):178–195, February 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617302099.

Hwang:2018:SWG

[HS18b] Jungbin Hwang and Yixiao Sun. Should we go one step further? An accurate comparison of one-step and twostep procedures in a generalized method of moments framework. Journal of Econometrics, 207(2):381–405, December 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301556.

Herbst:2019:TPF

[HS19a] Edward Herbst and Frank Schorfheide. Tempered particle filtering. Journal of Econometrics, 210(1):26–44, May 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302033.

Hsu:2019:TTE

[HS19b]  Yu-Chin Hsu and Shu Shen.  Testing treatment effect

heterogeneity in regression discontinuity designs. Journal of Econometrics, 208(2):468–486, February 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301982.

Hsiao:2018:PMI

[Hsi18] Cheng Hsiao. Panel models with interactive effects. Journal of Econometrics, 206(2):645–673, October 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301131.

Hu:2017:ICI

[HSS17] Yingyao Hu, Susanne M. Schennach, and Ji-Liang Shiu. Injectivity of a class of integral operators with compactly supported kernels. Journal of Econometrics, 200(1):48–58, September 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300751.

Heckman:2010:TCR

[HSU10] James J. Heckman, Daniel Schmierer, and Sergio Urzua. Testing the correlated random coefficient model. Journal of Econometrics, 158(2):177–203, October 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000084.

Hallin:2013:OSE

[HSVV13] Marc Hallin, Yvik Swan, Thomas Verdebout, and David Veredas. One-step R-estimation in linear models with stable errors. Journal of Econometrics, 172(2):195–204, February 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761200200X.

Hoderlein:2016:TMU

[HSWY16] Stefan Hoderlein, Liangjun Su, Halbert White, and Thomas Tao Yang. Testing for monotonicity in unobservables under unconfoundedness. Journal of Econometrics, 193(1):183–202, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300252.

Horvath:2016:SIR

[HT16] Lajos Horváth and Lorenzo Trapani. Statistical inference in a random coefficient panel model. Journal of Econometrics, 193 (1):54–75, July 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300203.

Horvath:2019:TRR

[HT19] Lajos Horváth and Lorenzo Trapani. Testing for randomness in a random coefficient autoregression model. Journal of Econometrics, 209(2):338–352, April 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300090.

Heckman:2010:CIS

[HU10] James J. Heckman and Sergio Urzúa. Comparing IV with structural models: What simple IV can and cannot identify. Journal of Econometrics, 156(1):27–37, May 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609001973.

Hu:2017:EUA

[Hu17] Yingyao Hu. The econometrics of unobservables: Applications of measurement error models in empirical industrial organization and labor economics. Journal of Econometrics, 200(2):154– 168, October 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300830.

Hualde:2014:ELR

[Hua14] Javier Hualde. Estimation of long-run parameters in unbalanced cointegration. Journal of Econometrics, 178(2):761–778, February 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613002297.

Haldrup:2017:LMF

[HV17a] Niels Haldrup and J. Eduardo Vera Valdés. Long memory, fractional integration, and cross-sectional aggregation.

Journal of Econometrics, 199(1):1–11, July 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300428.

Han:2017:IGB

[HV17b] Sukjin Han and Edward J. Vytlacil. Identification in a generalization of bivariate probit models with dummy endogenous regressors. Journal of Econometrics, 199(1):63–73, July 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300465.

Hounyo:2017:LSB

[HV17c] Ulrich Hounyo and Rasmus T. Varneskov. A local stable bootstrap for power variations of pure-jump semimartingales and activity index estimation. Journal of Econometrics, 198(1):10–28, May 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300039.

Hautsch:2019:LSP

[HV19] Nikolaus Hautsch and Stefan Voigt. Large-scale portfolio allocation under transaction costs and model uncertainty. Journal of Econometrics, 212(1):221–240, September 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300831.

Hallin:2011:CSD

[HvdAW11] Marc Hallin, Ramon van den Akker, and Bas J. M. Werker. A class of simple distribution-free rank-based unit root tests. Journal of Econometrics, 163(2):200–214, August 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611000698.

Hallin:2016:SEC

[HvdAW16] Marc Hallin, Ramon van den Akker, and Bas J. M. Werker. Semiparametric error-correction models for cointegration with trends: Pseudo-Gaussian and optimal rank-based tests of the cointegration rank. Journal of Econometrics, 190(1):46–61, January 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002237.

Hoderlein:2010:SME

[HW10] Stefan Hoderlein and Joachim Winter. Structural measurement errors in nonseparable models. Journal of Econometrics, 157(2):432–440, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000965.

Hamilton:2012:IEG

[HW12a] James D. Hamilton and Jing Cynthia Wu. Identification and estimation of Gaussian affine term structure models. Journal of Econometrics, 168(2):315–331, June 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000450.

Hoderlein:2012:NIN

[HW12b] Stefan Hoderlein and Halbert White. Nonparametric identification in nonseparable panel data models with generalized fixed effects. Journal of Econometrics, 168(2):300–314, June 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000437.

Hamilton:2014:TIA

[HW14a] James D. Hamilton and Jing Cynthia Wu. Testable implications of affine term structure models. Journal of Econometrics, 178(Part 2):231–242, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001887.

Hausman:2014:ESP

[HW14b] Jerry A. Hausman and Tiemen Woutersen. Estimating a semi-parametric duration model without specifying heterogeneity. Journal of Econometrics, 178(Part 1):114–131, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001619.

Hsiao:2014:TOF

[HW14c] Cheng Hsiao and Shui Ki Wan. Is there an optimal forecast combination? Journal of Econometrics, 178(Part 2):294–309, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613002339.

Hu:2017:MEM

[HW17] Yingyao Hu and Tom Wansbeek. Measurement error models: Editors’ introduction. Journal of Econometrics, 200(2):151–153, October 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300817.

Hardle:2016:TTE

[HWY16] Wolfgang Karl Härdle, Weining Wang, and Lining Yu. TENET: Tail-event driven NETwork risk. Journal of Econometrics, 192(2):499–513, June 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616300161.

Hall:2010:NLS

[HY10] Peter Hall and Adonis Yatchew. Nonparametric least squares estimation in derivative families. Journal of Econometrics, 157(2):362–374, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000813.

Hsiao:2015:IGL

[HZ15a] Cheng Hsiao and Junwei Zhang. IV, GMM or likelihood approach to estimate dynamic panel models when either n or t or both are large. Journal of Econometrics, 187(1):312–322, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000962.

Hsiao:2015:SIP

[HZ15b] Cheng Hsiao and Qiankun Zhou. Statistical inference for panel dynamic simultaneous equations models. Journal of Econometrics, 189(2):383–396, December 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001098.

Hsiao:2018:IPI

[HZ18] Cheng Hsiao and Qiankun Zhou. Incidental parameters, initial conditions and sample size in statistical inference for dynamic panel data models. Journal of Econometrics, 207(1):114–128, November 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301209.

Ibragimov:2014:RLE

[Ibr14] Rustam Ibragimov. On the robustness of location estimators in models of firm growth under heavy-tailedness. Journal of Econometrics, 181(1):25–33, July 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000347.

Inklaar:2016:MIP

[ID16] Robert Inklaar and W. Erwin Diewert. Measuring industry productivity and cross-country convergence. Journal of Econometrics, 191(2):426–433, April 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002894.

Ivancic:2011:SDT

[IDF11] Lorraine Ivancic, W. Erwin Diewert, and Kevin J. Fox. Scanner data, time aggregation and the construction of price indexes. Journal of Econometrics, 161(1):24–35, March 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001831.

Inoue:2017:RWS

[IJR17] Atsushi Inoue, Lu Jin, and Barbara Rossi. Rolling window selection for out-of-sample forecasting with time-varying parameters. Journal of Econometrics, 196(1):55–67, January 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407616301713.

Inoue:2013:IIR

[IK13] Atsushi Inoue and Lutz Kilian. Inference on impulse response functions in structural VAR models. Journal of Econometrics, 177(1):1–13, November 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613001310.

See corrigendum [IK19].

Inoue:2016:JCS

[IK16] Atsushi Inoue and Lutz Kilian. Joint confidence sets for structural impulse responses. Journal of Econometrics, 192(2):421– 432, June 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300112.

Inoue:2019:CII

[IK19] Atsushi Inoue and Lutz Kilian. Corrigendum to “Inference on impulse response functions in structural VAR models” [J. Econometrics 177 (2013) 1–13]. Journal of Econometrics, 209(1):139– 143, March 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302288. See [IK13].

Ikeda:2016:BCE

[Ike16] Shin S. Ikeda. A bias-corrected estimator of the covariation matrix of multiple security prices when both microstructure effects and sampling durations are persistent and endogenous. Journal of Econometrics, 193(1):203–214, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300422.

ilKim:2016:EPF

[iKPS16] Kyoo il Kim, Amil Petrin, and Suyong Song. Estimating production functions with control functions when capital is measured with error. Journal of Econometrics, 190(2):267–279, February 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001852.

Ichimura:2010:CAD

[IL10] Hidehiko Ichimura and Sokbae Lee. Characterization of the asymptotic distribution of semiparametric M-estimators. Journal of Econometrics, 159(2):252–266, December 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001302. See corrigendum [IL18].

Ichimura:2018:CCA

[IL18] Hidehiko Ichimura and Sokbae Lee. Corrigendum to “Characterization of the asymptotic distribution of semiparametric M-estimators” [J. Econometrics 159 (2) (2010) 252– 266]. Journal of Econometrics, 202(2):306–307, February 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761730146X. See [IL10].

Isakov:2019:FTC

[ILM19] Leah Isakov, Andrew W. Lo, and Vahid Montazerhodjat. Is the FDA too conservative or too aggressive?: a Bayesian decision analysis of clinical trial design. Journal of Econometrics, 211(1): 117–136, July 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302380.

Issler:2011:AIF

[ILT11] João Victor Issler, Oliver Linton, and Allan Timmermann. Annals issue on forecasting — Guest Editors’ introduction. Journal of Econometrics, 164(1):1–3, September 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761100042X.

Iacone:2013:TBT

[ILT13] Fabrizio Iacone, Stephen J. Leybourne, and A. M. Robert Taylor. Testing for a break in trend when the order of integration is unknown. Journal of Econometrics, 176(1):30–45, September 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000663.

Inoue:2011:TWI

[IR11] Atsushi Inoue and Barbara Rossi. Testing for weak identification in possibly nonlinear models. Journal of Econometrics, 161(2):246–261, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610002575.

Jacho-Chavez:2010:INE

[JCLL10] David Jacho-Chávez, Arthur Lewbel, and Oliver Linton. Identification and nonparametric estimation of a transformed additively separable model. Journal of Econometrics, 156(2):392–407, June 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002978.

Jenish:2012:NSR

[Jen12] Nazgul Jenish. Nonparametric spatial regression under nearepoch dependence. Journal of Econometrics, 167(1):224–239, March 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002715.

Jin:2015:BMT

[JfL15] Fei Jin and Lung fei Lee. On the bootstrap for Moran’s i test for spatial dependence. Journal of Econometrics, 184(2):295– 314, February 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001900.

Jin:2018:INL

[JfL18] Fei Jin and Lung fei Lee. Irregular N2SLS and LASSO estimation of the matrix exponential spatial specification model. Journal of Econometrics, 206(2):336–358, October 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300964.

Jin:2019:GET

[JfL19] Fei Jin and Lung fei Lee. GEL estimation and tests of spatial autoregressive models. Journal of Econometrics, 208(2):585– 612, February 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302173.

Jin:2010:EMT

[JJ10] Hui Jin and Dale W. Jorgenson. Econometric modeling of technical change. Journal of Econometrics, 157(2):205–219, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609003005.

Just:2011:GIR

[JJ11] Richard E. Just and David R. Just. Global identification of risk preferences with revealed preference data. Journal of Econometrics, 162(1):6–17, May 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002577.

Johansen:2014:AIP

[JJ14] Søren Johansen and Katarina Juselius. An asymptotic invariance property of the common trends under linear transformations of the data. Journal of Econometrics, 178(Part 2):310–315, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001930.

Johansen:2010:THM

[JJFG10] Søren Johansen, Katarina Juselius, Roman Frydman, and Michael Goldberg. Testing hypotheses in an I(2) model with piecewise linear trends. An analysis of the persistent long swings in the Dmk/$ rate. Journal of Econometrics, 158(1):117–129, September 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000606.

Jarrow:2015:STC

[JK15] Robert Jarrow and Simon Sai Man Kwok. Specification tests of calibrated option pricing models. Journal of Econometrics, 189(2):397–414, December 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615001104.

Jing:2012:JAI

[JKLM12] Bing-Yi Jing, Xin-Bing Kong, Zhi Liu, and Per Mykland. On the jump activity index for semimartingales. Journal of Econometrics, 166(2):213–223, February 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761100217X.

Jacod:2017:TNC

[JKM17] Jean Jacod, Claudia Klüppelberg, and Gernot Müller. Testing for non-correlation between price and volatility jumps.

Journal of Econometrics, 197(2):284–297, April 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616302196.

Lee:2012:TES

[jL12] Myoung jae Lee. Treatment effects in sample selection models and their nonparametric estimation. Journal of Econometrics, 167(2):317–329, April 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611001990.

Johansen:2013:LSE

[JL13] Søren Johansen and Theis Lange. Least squares estimation in a simple random coefficient autoregressive model. Journal of Econometrics, 177(2):285–288, December 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613000900.

Jacod:2019:EIV

[JLZ19] Jean Jacod, Yingying Li, and Xinghua Zheng. Estimating the integrated volatility with tick observations. Journal of Econometrics, 208(1):80–100, January 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301714.

Jensen:2010:BSS

[JM10] Mark J. Jensen and John M. Maheu. Bayesian semiparametric stochastic volatility modeling. Journal of Econometrics, 157(2):306–316, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000278.

Jensen:2013:BSM

[JM13] Mark J. Jensen and John M. Maheu. Bayesian semiparametric multivariate GARCH modeling. Journal of Econometrics, 176(1):3–17, September 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613000808.

Jensen:2014:ESA

[JM14] Mark J. Jensen and John M. Maheu. Estimating a semiparametric asymmetric stochastic volatility model with a Dirichlet process mixture. Journal of Econometrics, 178(Part 3):523– 538, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613001681.

Jin:2016:BSM

[JM16a] Xin Jin and John M. Maheu. Bayesian semiparametric modeling of realized covariance matrices. Journal of Econometrics, 192(1): 19–39, May 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002729.

Jin:2016:MCB

[JM16b] Xin Jin and John M. Maheu. Modeling covariance breakdowns in multivariate GARCH. Journal of Econometrics, 194(1):1–23, September 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300562.

Jarocinski:2019:PAO

[JM19] Marek Jarociński and Albert Marcet. Priors about observables in vector autoregressions. Journal of Econometrics, 209(2):238– 255, April 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300065.

Johansen:2010:LIN

[JN10] Søren Johansen and Morten Ørregaard Nielsen. Likelihood inference for a nonstationary fractional autoregressive model. Journal of Econometrics, 158(1):51–66, September 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000485.

Johansen:2018:CVA

[JN18] Søren Johansen and Morten Ørregaard Nielsen. The cointegrated vector autoregressive model with general deterministic terms. Journal of Econometrics, 202(2):214–229, February 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617302117.

Jochmans:2015:MEM

[Joc15] Koen Jochmans. Multiplicative-error models with sample selection. Journal of Econometrics, 184(2):315–327, February 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614002152.

Johansen:2010:SIP

[Joh10] Søren Johansen. Some identification problems in the cointegrated vector autoregressive model. Journal of Econometrics, 158(2):262–273, October 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000102.

Joskow:2019:JH

[Jos19] Paul L. Joskow. Jerry Hausman. Journal of Econometrics, 211 (1):11–15, July 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302318.

Jenish:2012:SPA

[JP12] Nazgul Jenish and Ingmar R. Prucha. On spatial processes and asymptotic inference under near-epoch dependence. Journal of Econometrics, 170(1):178–190, September 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001340.

Jun:2010:CNT

[JPW10] Sung Jae Jun, Joris Pinkse, and Yuanyuan Wan. A consistent nonparametric test of affiliation in auction models. Journal of Econometrics, 159(1):46–54, November 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001120.

Jun:2015:CLE

[JPW15] Sung Jae Jun, Joris Pinkse, and Yuanyuan Wan. Classical Laplace estimation for ^{3} \sqrt{n}-consistent estimators: Improved convergence rates and rate-adaptive inference. Journal of Econometrics, 187(1):201–216, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000147.

Jun:2011:TBT

[JPX11] Sung Jae Jun, Joris Pinkse, and Haiqing Xu. Tighter bounds in triangular systems. Journal of Econometrics, 161(2):122– 128, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610002265.

Jentsch:2015:TSO

[JR15] Carsten Jentsch and Suhasini Subba Rao. A test for second order stationarity of a multivariate time series. Journal of Econometrics, 185(1):124–161, March 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761400195X.

Juhl:2014:THF

[JSE14] Ted Juhl and Walter Sosa-Escudero. Testing for heteroskedasticity in fixed effects models. Journal of Econometrics, 178 (Part 3):484–494, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407613001498.

Judge:2013:FOC

[Jud13] George Judge. Fellow’s opinion corner: Econometric information recovery. Journal of Econometrics, 176(1):1–2, September 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000638.

Just:2011:CWE

[Jus11] David R. Just. Calibrating the wealth effects of decoupled payments: Does decreasing absolute risk aversion matter? Journal of Econometrics, 162(1):25–34, May 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002590.

Jacobs:2011:MDR

[JvN11] Jan P. A. M. Jacobs and Simon van Norden. Modeling data revisions: Measurement error and dynamics of “true” values. Journal of Econometrics, 161(2):101–109, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610002526.

Jacobs:2010:CLR

[JW10] Jan P. A. M. Jacobs and Kenneth F. Wallis. Cointegration, longrun structural modelling and weak exogeneity: Two models of the UK economy. Journal of Econometrics, 158(1):108–116, September 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761000059X.

Jacobi:2016:BTE

[JWFS16] Liana Jacobi, Helga Wagner, and Sylvia Frühwirth-Schnatter. Bayesian treatment effects models with variable selection for panel outcomes with an application to earnings effects of maternity leave. Journal of Econometrics, 193(1):234–250, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300197.

Jiang:2018:NDT

[JWY18] Liang Jiang, Xiaohu Wang, and Jun Yu. New distribution theory for the estimation of structural break point in mean. Journal of Econometrics, 205(1):156–176, July 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300484.

Kaido:2016:DAI

[Kai16] Hiroaki Kaido. A dual approach to inference for partially identified econometric models. Journal of Econometrics, 192(1):269– 290, May 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615003164.

Kalnina:2011:SHF

[Kal11] Ilze Kalnina. Subsampling high frequency data. Journal of Econometrics, 161(2):262–283, April 1, 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610002563.

Kaplan:2015:IQI

[Kap15a] David M. Kaplan. Improved quantile inference via fixedsmoothing asymptotics and Edgeworth expansion. Journal of Econometrics, 185(1):20–32, March 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407614002401.

Kasparis:2015:NPR

[KAP15b] Ioannis Kasparis, Elena Andreou, and Peter C. B. Phillips. Nonparametric predictive regression. Journal of Econometrics, 185(2):468–494, April 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001377.

Karlsson:2017:CBR

[Kar17] Sune Karlsson. Corrigendum to “Bayesian reduced rank regression in econometrics” [J. Econometrics 75 (1996) 121– 146]. Journal of Econometrics, 201(1):170–171, November 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002754.

Kasparis:2010:BTC

[Kas10] Ioannis Kasparis. The bierens test for certain nonstationary models. Journal of Econometrics, 158(2):221–230, October 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000114.

Kastner:2019:SBT

[Kas19] Gregor Kastner. Sparse Bayesian time-varying covariance estimation in many dimensions. Journal of Econometrics, 210(1): 98–115, May 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302082.

Kaufmann:2015:SSM

[Kau15] Sylvia Kaufmann. K-state switching models with time-varying transition distributions — does loan growth signal stronger effects of variables on inflation? Journal of Econometrics, 187(1): 82–94, July 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761500024X.

Kawaguchi:2017:TRR

[Kaw17] Kohei Kawaguchi. Testing rationality without restricting heterogeneity. Journal of Econometrics, 197(1):153–171, March 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616302135.

Kock:2015:OIH

[KC15] Anders Bredahl Kock and Laurent Callot. Oracle inequalities for high dimensional vector autoregressions. Journal of Econometrics, 186(2):325–344, June 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000378.

Keane:2010:SVA

[Kea10] Michael P. Keane. Structural vs. atheoretic approaches to econometrics. Journal of Econometrics, 156(1):3–20, May 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-

6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609001948. See comments

[Rus10, Blu10].

Ketz:2018:SIW

[Ket18] Philipp Ketz. Subvector inference when the true parameter vector may be near or at the boundary. Journal of Econometrics, 207(2):285–306, December 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761830143X.

Ketz:2019:ASD

[Ket19] Philipp Ketz. On asymptotic size distortions in the random coefficients logit model. Journal of Econometrics, 212(2):413– 432, October 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619301010.

Kim:2019:FGI

[KF19] Donggyu Kim and Jianqing Fan. Factor GARCH-Itô models for high-frequency data with application to large volatility matrix prediction. Journal of Econometrics, 208(2):395–417, February 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301957.

Kaufmann:2019:EIC

[KFSvD19] Sylvia Kaufmann, Sylvia Frühwirth-Schnatter, and Herman K. van Dijk. Editorial introduction on complexity and big data in economics and finance: Recent developments from a Bayesian perspective. Journal of Econometrics, 210(1):1–3, May 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761830201X.

Kalli:2014:TVS

[KG14] Maria Kalli and Jim E. Griffin. Time-varying sparsity in dynamic regression models. Journal of Econometrics, 178(2):779– 793, February 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613002273.

Kalli:2018:BNV

[KG18] Maria Kalli and Jim E. Griffin. Bayesian nonparametric vector autoregressive models. Journal of Econometrics, 203(2):267– 282, April 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617302415.

Kato:2012:APQ

[KGMR12] Kengo Kato, Antonio F. Galvao, and Gabriel V. Montes-Rojas. Asymptotics for panel quantile regression models with individual effects. Journal of Econometrics, 170(1):76–91, September 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000760.

Khan:2013:DFE

[Kha13] Shakeeb Khan. Distribution free estimation of heteroskedastic binary response models using Probit/Logit criterion functions. Journal of Econometrics, 172(1):168–182, January 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001753.

Khwaja:2010:EWP

[Khw10] Ahmed Khwaja. Estimating willingness to pay for Medicare using a dynamic life-cycle model of demand for health insurance. Journal of Econometrics, 156(1):130–147, May 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002103.

Kim:2011:ECD

[Kim11] Dukpa Kim. Estimating a common deterministic time trend break in large panels with cross sectional dependence. Journal of Econometrics, 164(2):310–330, October 1, 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761100131X.

Kim:2012:MSP

[Kim12] Jae-Young Kim. Model selection in the presence of nonstationarity. Journal of Econometrics, 169(2):247–257, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000395.

Kim:2014:AQL

[Kim14] Jae-Young Kim. An alternative quasi likelihood approach, Bayesian analysis and data-based inference for model specification. Journal of Econometrics, 178(Part 1):132–145, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001620.

Kim:2012:PPE

[KK12] Chang Sik Kim and In-Moo Kim. Partial parametric estimation for nonstationary nonlinear regressions. Journal of Econometrics, 167(2):448–457, April 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611002089.

Koop:2013:LTV

[KK13] Gary Koop and Dimitris Korobilis. Large time-varying parameter VARs. Journal of Econometrics, 177(2):185–198, December 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000845.

Kurz-Kim:2014:PCD

[KKL14] Jeong-Ryeol Kurz-Kim and Mico Loretan. On the properties of the coefficient of determination in regression models with infinite variance variables. Journal of Econometrics, 181(1):15–24, July 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000335.

Kim:2018:ATL

[KKLW18] Donggyu Kim, Xin-Bing Kong, Cui-Xia Li, and Yazhen Wang. Adaptive thresholding for large volatility matrix estimation based on high-frequency financial data. Journal of Econometrics, 203(1):69–79, March 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617302270.

Koop:2019:BCV

[KKP19] Gary Koop, Dimitris Korobilis, and Davide Pettenuzzo. Bayesian compressed vector autoregressions. Journal of Econometrics, 210(1):135–154, May 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618302100.

Koo:2012:ESL

[KL12] Bonsoo Koo and Oliver Linton. Estimation of semiparametric locally stationary diffusion models. Journal of Econometrics, 170(1):210–233, September 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612001157.

Kong:2018:TAC

[KL18] Xin-Bing Kong and Cheng Liu. Testing against constant factor loading matrix with large panel high-frequency data. Journal of Econometrics, 204(2):301–319, June 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300393.

Klein:2010:HTE

[Kle10] Tobias J. Klein. Heterogeneous treatment effects: Instrumental variables without monotonicity? Journal of Econometrics, 155(2):99–116, April 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760900219X.

Koop:2011:BIT

[KLGS11] Gary Koop, Roberto Leon-Gonzalez, and Rodney W. Strachan. Bayesian inference in a time varying cointegration model. Journal of Econometrics, 165(2):210–220, December 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001588.

Koop:2012:BMA

[KLGS12] Gary Koop, Roberto Leon-Gonzalez, and Rodney Strachan. Bayesian model averaging in the instrumental variable regression model. Journal of Econometrics, 171(2):237–250, December 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001522.

Kline:2015:ICI

[Kli15] Brendan Kline. Identification of complete information games. Journal of Econometrics, 189(1):117–131, November 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001955.

Kim:2015:QOP

[KLMP15] Young Shin Kim, Jaesung Lee, Stefan Mittnik, and Jiho Park. Quanto option pricing in the presence of fat tails and asymmetric dependence. Journal of Econometrics, 187(2):512–520, August 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000603.

Koopman:2011:MFC

[KLS11] Siem Jan Koopman, André Lucas, and Bernd Schwaab. Modeling frailty-correlated defaults using many macroeconomic covariates. Journal of Econometrics, 162(2):312–325, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000303.

Kong:2019:RTN

[KLZ19] Xin-Bing Kong, Zhi Liu, and Wang Zhou. A rank test for the number of factors with high-frequency data. Journal of Econometrics, 211(2):439–460, August 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300533.

Kitagawa:2016:MAS

[KM16] Toru Kitagawa and Chris Muris. Model averaging in semiparametric estimation of treatment effects. Journal of Econometrics, 193(1):271–289, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761630046X.

Kunitomo:2011:MRB

[KMN11] Naoto Kunitomo, Michael McAleer, and Yoshihiko Nishiyama. Moment restriction-based econometric methods: an overview. Journal of Econometrics, 165(1):1–4, November 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611000935.

Kapetanios:2015:GDF

[KMPF15] G. Kapetanios, J. Mitchell, S. Price, and N. Fawcett. Generalised density forecast combinations. Journal of Econometrics, 188(1):150–165, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615001232.

Kapetanios:2014:NPD

[KMS14] George Kapetanios, James Mitchell, and Yongcheol Shin. A nonlinear panel data model of cross-sectional dependence.

Journal of Econometrics, 179(2):134–157, April 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000037.

Kalliovirta:2016:GMV

[KMS16] Leena Kalliovirta, Mika Meitz, and Pentti Saikkonen. Gaussian mixture vector autoregression. Journal of Econometrics, 192(2): 485–498, June 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761630015X.

Kock:2016:OIV

[Koc16] Anders Bredahl Kock. Oracle inequalities, variable selection and uniform inference in high-dimensional correlated random effects panel data models. Journal of Econometrics, 195(1):71–85, November 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616301221.

Kolesar:2018:MDA

[Kol18] Michal Kolesár. Minimum distance approach to inference with many instruments. Journal of Econometrics, 204(1):86–100, May 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300083.

Komarova:2013:BCM

[Kom13] Tatiana Komarova. Binary choice models with discrete regressors: Identification and misspecification. Journal of Econometrics, 177(1):14–33, November 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613001279.

Kelejian:2010:SES

[KP10a] Harry H. Kelejian and Ingmar R. Prucha. Specification and estimation of spatial autoregressive models with autoregressive and heteroskedastic disturbances. Journal of Econometrics, 157(1): 53–67, July 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609002784.

Koop:2010:FAP

[KP10b] Gary Koop and Simon Potter. A flexible approach to parametric inference in nonlinear and time varying time series models. Journal of Econometrics, 159(1):134–150, November 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001272.

Kasparis:2012:DMN

[KP12] Ioannis Kasparis and Peter C. B. Phillips. Dynamic misspecification in nonparametric cointegrating regression. Journal of Econometrics, 168(2):270–284, June 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000474.

Kuersteiner:2013:LTP

[KP13] Guido M. Kuersteiner and Ingmar R. Prucha. Limit theory for panel data models with cross sectional dependence and sequential exogeneity. Journal of Econometrics, 174(2):107–126, June 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000389.

Kelly:2015:TPR

[KP15] Bryan Kelly and Seth Pruitt. The three-pass regression filter: a new approach to forecasting using many predictors. Journal of Econometrics, 186(2):294–316, June 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000354.

Kim:2017:ARD

[KP17] Jihyun Kim and Joon Y. Park. Asymptotics for recurrent diffusions with application to high frequency regression. Journal of Econometrics, 196(1):37–54, January 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S030440761630166X.

Korobilis:2019:AHP

[KP19] Dimitris Korobilis and Davide Pettenuzzo. Adaptive hierarchical priors for high-dimensional vector autoregressions. Journal of Econometrics, 212(1):241–271, September 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300843.

Kong:2019:WCT

[KPS19] Jianning Kong, Peter C. B. Phillips, and Donggyu Sul. Weak σ-convergence: Theory and applications. Journal of Econometrics, 209(2):185–207, April 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619300041.

Kumbhakar:2012:BEA

[KPT12] Subal C. Kumbhakar, Christopher F. Parmeter, and Efthymios G. Tsionas. Bayesian estimation approaches to first-price auctions. Journal of Econometrics, 168(1):47–59, May 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611001680.

Kumbhakar:2013:ZIS

[KPT13] Subal C. Kumbhakar, Christopher F. Parmeter, and Efthymios G. Tsionas. A zero inefficiency stochastic frontier model. Journal of Econometrics, 172(1):66–76, January 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612002163.

Khan:2016:IPD

[KPT16] Shakeeb Khan, Maria Ponomareva, and Elie Tamer. Identification of panel data models with endogenous censoring. Journal of Econometrics, 194(1):57–75, September 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300446.

Kapetanios:2011:PNS

[KPY11] G. Kapetanios, M. Hashem Pesaran, and T. Yamagata. Panels with non-stationary multifactor error structures. Journal of Econometrics, 160(2):326–348, February 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610002022.

Kristensen:2010:LBI

[KR10] Dennis Kristensen and Anders Rahbek. Likelihood-based inference for cointegration with nonlinear error-correction. Journal of Econometrics, 158(1):78–94, September 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000527.

Krause:2013:CEL

[Kra13] Melanie Krause. Corrigendum to “Elliptical Lorenz Curves” [J. Econom. 40 (1989) 327–338]. Journal of Econometrics, 174(1): 44, May 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613000158. See [VA89].

Kristensen:2010:PML

[Kri10] Dennis Kristensen. Pseudo-maximum likelihood estimation in two classes of semiparametric diffusion models. Journal of Econometrics, 156(2):239–259, June 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440760900270X.

Kristensen:2011:SNE

[Kri11] Dennis Kristensen. Semi-nonparametric estimation and misspecification testing of diffusion models. Journal of Econometrics, 164(2):382–403, October 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611001412.

Krief:2017:DIN

[Kri17] Jerome M. Krief. Direct instrumental nonparametric estimation of inverse regression functions. Journal of Econometrics, 201(1):95–107, November 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617301173.

Kruiniger:2008:MLE

[Kru08] Hugo Kruiniger. Maximum likelihood estimation and inference methods for the covariance stationary panel AR(1)/unit root model. Journal of Econometrics, 144(2):447–464, June 2008. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407608000390. See corrigendum [Kru14].

Kruiniger:2013:QME

[Kru13] Hugo Kruiniger. Quasi ML estimation of the panel AR(1) model with arbitrary initial conditions. Journal of Econometrics, 173(2):175–188, April 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612002618.

Kruiniger:2014:CML

[Kru14] Hugo Kruiniger. Corrigendum to “Maximum likelihood estimation and inference methods for the covariance stationary panel AR(1) /unit root model” [J. Econom. 144 (2008) 447–464]. Journal of Econometrics, 178(2):824, February 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613002340. See [Kru08].

Karabiyik:2017:RRC

[KRW17] Hande Karabiyik, Simon Reese, and Joakim Westerlund. On the role of the rank condition in CCE estimation of factor-augmented panel regressions. Journal of Econometrics, 197(1):60–64, March 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616302007.

Kim:2011:SHA

[KS11] Min Seong Kim and Yixiao Sun. Spatial heteroskedasticity and autocorrelation consistent estimation of covariance matrix. Journal of Econometrics, 160(2):349–371, February 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610002034.

Kemp:2012:RTM

[KS12a] Gordon C. R. Kemp and J. M. C. Santos Silva. Regression towards the mode. Journal of Econometrics, 170(1):92–101, September 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000735.

Kim:2012:TSM

[KS12b] Don H. Kim and Kenneth J. Singleton. Term structure models and the zero bound: an empirical investigation of Japanese yields. Journal of Econometrics, 170(1):32–49, September 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001352.

Kline:2012:HOP

[KS12c] Patrick Kline and Andres Santos. Higher order properties of the wild bootstrap under misspecification. Journal of Econometrics, 171(1):54–70, November 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001480.

Kristensen:2012:EDM

[KS12d] Dennis Kristensen and Yongseok Shin. Estimation of dynamic models with nonparametric simulated maximum likelihood. Journal of Econometrics, 167(1):76–94, March 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611002363.

Kumbhakar:2012:EI

[KS12e] Subal C. Kumbhakar and Robin C. Sickles. Editors’ introduction. Journal of Econometrics, 168(1):1–3, May 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611001771.

Kutlu:2012:EMP

[KS12f] Levent Kutlu and Robin C. Sickles. Estimation of market power in the presence of firm level inefficiencies. Journal of Econometrics, 168(1):141–155, May 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611002442.

Kim:2013:HSD

[KS13] Min Seong Kim and Yixiao Sun. Heteroskedasticity and spatiotemporal dependence robust inference for linear panel models with fixed effects. Journal of Econometrics, 177(1):85–108, November 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613001309.

Kim:2014:FFM

[KS14] Hyun Hak Kim and Norman R. Swanson. Forecasting financial and macroeconomic variables using data reduction methods: New empirical evidence. Journal of Econometrics, 178 (Part 2):352–367, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407613001978.

Koo:2015:SBM

[KS15] Bonsoo Koo and Myung Hwan Seo. Structural-break models under mis-specification: Implications for forecasting. Journal of Econometrics, 188(1):166–181, September 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001384.

Keane:2016:ASM

[KS16a] Michael Keane and Olena Stavrunova. Adverse selection, moral hazard and the demand for medigap insurance. Journal of Econometrics, 190(1):62–78, January 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002225.

Kumbhakar:2016:EI

[KS16b] Subal C. Kumbhakar and Peter Schmidt. Editors’ introduction. Journal of Econometrics, 190(2):209–211, February 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001724.

Kristensen:2017:HOP

[KS17] Dennis Kristensen and Bernard Salanié. Higher-order properties of approximate estimators. Journal of Econometrics, 198(2):189– 208, June 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300155.

Kato:2018:UCB

[KS18a] Kengo Kato and Yuya Sasaki. Uniform confidence bands in deconvolution with unknown error distribution. Journal of Econometrics, 207(1):129–161, November 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301301.

Kim:2018:EPR

[KS18b] Soohun Kim and Georgios Skoulakis. Ex-post risk premia estimation and asset pricing tests using large cross sections: The regression-calibration approach. Journal of Econometrics, 204(2):159–188, June 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618300198.

Kasahara:2019:APM

[KS19a] Hiroyuki Kasahara and Katsumi Shimotsu. Asymptotic properties of the maximum likelihood estimator in regime switching econometric models. Journal of Econometrics, 208(2):442–467, February 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301970.

Kato:2019:UCB

[KS19b] Kengo Kato and Yuya Sasaki. Uniform confidence bands for nonparametric errors-in-variables regression. Journal of Econometrics, 213(2):516–555, December 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619301605.

Kaufmann:2019:BES

[KS19c] Sylvia Kaufmann and Christian Schumacher. Bayesian estimation of sparse dynamic factor models with order-independent and ex-post mode identification. Journal of Econometrics, 210(1): 116–134, May 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302094.

Klein:2015:EME

[KSV15a] Roger Klein, Chan Shen, and Francis Vella. Estimation of marginal effects in semiparametric selection models with binary outcomes. Journal of Econometrics, 185(1):82–94, March 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614002449.

Kneip:2015:FEP

[KSV15b] Alois Kneip, Léopold Simar, and Ingrid Van Keilegom. Frontier estimation in the presence of measurement error with unknown variance. Journal of Econometrics, 184(2):379–393, February 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614002279.

Kim:2017:FBV

[KSY17] Min Seong Kim, Yixiao Sun, and Jingjing Yang. A fixedbandwidth view of the pre-asymptotic inference for kernel smoothing with time series data. Journal of Econometrics, 197(2):298–322, April 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616302305.

Kurozumi:2011:MSC

[KT11] Eiji Kurozumi and Purevdorj Tuvaandorj. Model selection criteria in multivariate models with multiple structural changes. Journal of Econometrics, 164(2):218–238, October 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000820.

Kline:2012:BBR

[KT12] Brendan Kline and Elie Tamer. Bounds for best response functions in binary games. Journal of Econometrics, 166(1):92– 105, January 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001217.

Kuchler:2014:ESM

[KT14] Uwe Küchler and Stefan Tappe. Exponential stock models driven by tempered stable processes. Journal of Econometrics, 181(1): 53–63, July 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614000372.

Kumbhakar:2016:GBT

[KT16] Subal C. Kumbhakar and Efthymios G. Tsionas. The good, the bad and the technology: Endogeneity in environmental production models. Journal of Econometrics, 190(2):315–327, February 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001773.

Kutlu:2019:TVT

[KTT19] Levent Kutlu, Kien C. Tran, and Mike G. Tsionas. A time-varying true individual effects model with endogenous regressors. Journal of Econometrics, 211(2):539–559, August 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300922.

Khalaf:2014:IRI

[KU14] Lynda Khalaf and Giovanni Urga. Identification robust inference in cointegrating regressions. Journal of Econometrics, 182(2):385–396, October 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001419.

Kuersteiner:2012:KWG

[Kue12] Guido M. Kuersteiner. Kernel-weighted GMM estimators for linear time series models. Journal of Econometrics, 170(2):399– 421, October 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761200125X.

Kuersteiner:2019:IPD

[Kue19] Guido M. Kuersteiner. Invariance principles for dependent processes indexed by Besov classes with an application to a Hausman test for linearity. Journal of Econometrics, 211(1):243–261, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302483.

Klein:2010:ECT

[KV10a] Roger Klein and Francis Vella. Estimating a class of triangular simultaneous equations models without exclusion restrictions. Journal of Econometrics, 154(2):154–164, February 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609001675.

Komunjer:2010:EED

[KV10b] Ivana Komunjer and Quang Vuong. Efficient estimation in dynamic conditional quantile models. Journal of Econometrics, 157(2):272–285, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000035.

Kheifets:2017:NGF

[KV17] Igor Kheifets and Carlos Velasco. New goodness-of-fit diagnostics for conditional discrete response models. Journal of Econometrics, 200(1):135–149, September 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300805.

Kan:2010:DSA

[KW10a] Raymond Kan and Xiaolu Wang. On the distribution of the sample autocorrelation coefficients. Journal of Econometrics, 154(2):101–121, February 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609001638.

Kennan:2010:WWB

[KW10b] John Kennan and James R. Walker. Wages, welfare benefits and migration. Journal of Econometrics, 156(1):229–238, May 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002188.

Kaido:2014:TSP

[KW14] Hiroaki Kaido and Halbert White. A two-stage procedure for partially identified models. Journal of Econometrics, 182(1):5– 13, September 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614000633.

Kim:2016:UDT

[KW16] Donggyu Kim and Yazhen Wang. Unified discrete-time and continuous-time models and statistical inferences for merged low-frequency and high-frequency financial data. Journal of Econometrics, 194(2):220–230, October 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300914.

Kleppe:2014:MLE

[KYS14] Tore Selland Kleppe, Jun Yu, and Hans J. Skaug. Maximum likelihood estimation of partially observed diffusion models. Journal of Econometrics, 180(1):73–80, May 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000311.

Kleibergen:2015:UFT

[KZ15] Frank Kleibergen and Zhaoguo Zhan. Unexplained factors and their effects on second pass R-squared’s. Journal of Econometrics, 189(1):101–116, November 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001876.

Lamarche:2010:RPQ

[Lam10] Carlos Lamarche. Robust penalized quantile regression estimation for panel data. Journal of Econometrics, 157(2):396– 408, August 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000850.

Lamy:2012:EAA

[Lam12] Laurent Lamy. The econometrics of auctions with asymmetric anonymous bidders. Journal of Econometrics, 167(1):113– 132, March 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002703.

Large:2011:EQV

[Lar11] Jeremy Large. Estimating quadratic variation when quoted prices change by a constant increment. Journal of Econometrics, 160(1):2–11, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000497.

Lasak:2010:LBT

[Las10] Katarzyna Lasak. Likelihood based testing for no fractional cointegration. Journal of Econometrics, 158(1):67–77, September 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000503.

Lavergne:2014:MET

[Lav14] Pascal Lavergne. Model equivalence tests in a parametric framework. Journal of Econometrics, 178(Part 3):414–425, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001814.

Li:2017:EAT

[LB17] Kathleen T. Li and David R. Bell. Estimation of average treatment effects with panel data: Asymptotic theory and implementation. Journal of Econometrics, 197(1):65–75, March 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616302019.

Lee:2019:AWA

[LB19] Ying-Ying Lee and Debopam Bhattacharya. Applied welfare analysis for discrete choice with interval-data on income. Journal of Econometrics, 211(2):361–387, August 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300508.

Li:2016:EJD

[LC16] Chenxu Li and Dachuan Chen. Estimating jump-diffusions using closed-form likelihood expansions. Journal of Econometrics, 195(1):51–70, November 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616301348.

Lin:2015:OSN

[LCLS15] Wei Lin, Zongwu Cai, Zheng Li, and Li Su. Optimal smoothing in nonparametric conditional quantile derivative function estimation. Journal of Econometrics, 188(2):502–513, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000834.

Liu:2010:NTF

[LCY10] Jun M. Liu, Rong Chen, and Qiwei Yao. Nonparametric transfer function models. Journal of Econometrics, 157(1):151–164, July 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002826.

Leamer:2016:VCC

[Lea16] Edward E. Leamer. S-values: Conventional context-minimal measures of the sturdiness of regression coefficients. Journal of Econometrics, 193(1):147–161, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407616300185.

Lee:2012:BDP

[Lee12] Yoonseok Lee. Bias in dynamic panel models under time series misspecification. Journal of Econometrics, 169(1):54–60, July 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000103.

Lee:2014:ARM

[Lee14a] Seojeong Lee. Asymptotic refinements of a misspecificationrobust bootstrap for generalized method of moments estimators. Journal of Econometrics, 178(Part 3):398–413, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001838.

Lee:2014:TLD

[Lee14b] Yoon-Jin Lee. Testing a linear dynamic panel data model against nonlinear alternatives. Journal of Econometrics, 178 (Part 1):146–166, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407613001632.

Lee:2015:SSI

[Lee15] Jiyon Lee. A semiparametric single index model with heterogeneous impacts on an unobserved variable. Journal of Econometrics, 184(1):13–36, January 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001742.

Lee:2016:PQR

[Lee16a] Ji Hyung Lee. Predictive quantile regression with persistent covariates: IVX-QR approach. Journal of Econometrics, 192(1): 105–118, May 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615003000.

Lee:2016:ARM

[Lee16b] Seojeong Lee. Asymptotic refinements of a misspecificationrobust bootstrap for GEL estimators. Journal of Econometrics, 192(1):86–104, May 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615002900.

Lee:2018:EPS

[Lee18] Ying-Ying Lee. Efficient propensity score regression estimators of multivalued treatment effects for the treated. Journal of Econometrics, 204(2):207–222, June 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300290.

Leung:2015:TSE

[Leu15] Michael P. Leung. Two-step estimation of network-formation models with incomplete information. Journal of Econometrics, 188(1):182–195, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615001396.

Lam:2018:NER

[LF18] Clifford Lam and Phoenix Feng. A nonparametric eigenvalueregularized integrated covariance matrix estimator for asset return data. Journal of Econometrics, 206(1):226–257, September 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300927.

Lin:2010:GES

[LfL10a] Xu Lin and Lung fei Lee. GMM estimation of spatial autoregressive models with unknown heteroskedasticity. Journal of Econometrics, 157(1):34–52, July 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002887.

Liu:2010:GES

[LfL10b] Xiaodong Liu and Lung fei Lee. GMM estimation of social interaction models with centrality. Journal of Econometrics, 159(1):99–115, November 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610001259.

Liu:2019:LRT

[LfL19] Tuo Liu and Lung fei Lee. A likelihood ratio test for spatial model selection. Journal of Econometrics, 213(2):434–458, December 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301496.

Liu:2010:EGE

[LfLB10] Xiaodong Liu, Lung fei Lee, and Christopher R. Bollinger. An efficient GMM estimator of spatial autoregressive models. Journal of Econometrics, 159(2):303–319, December 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001715.

Li:2010:IIS

[Li10] Tong Li. Indirect inference in structural econometric models. Journal of Econometrics, 157(1):120–128, July 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002802.

Li:2017:FED

[Li17] Kunpeng Li. Fixed-effects dynamic spatial panel data models and impulse response analysis. Journal of Econometrics, 198(1): 102–121, May 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300167.

Liu:2015:DTL

[Liu15] Chu-An Liu. Distribution theory of the least squares averaging estimator. Journal of Econometrics, 186(1):142–159, May 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001596.

Lee:2014:TIR

[LKH14] Wei-Ming Lee, Chung-Ming Kuan, and Yu-Chin Hsu. Testing over-identifying restrictions without consistent estimation of the asymptotic covariance matrix. Journal of Econometrics, 181(2):181–193, August 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761400061X.

Liu:2018:EIV

[LKJ18] Zhi Liu, Xin-Bing Kong, and Bing-Yi Jing. Estimating the integrated volatility using high-frequency data with zero durations. Journal of Econometrics, 204(1):18–32, May 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300046.

Lengwiler:2010:IFY

[LL10a] Yvan Lengwiler and Carlos Lenz. Intelligible factors for the yield curve. Journal of Econometrics, 157(2):481–491, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001077.

Li:2010:NSE

[LL10b] Dong Li and Qi Li. Nonparametric/semiparametric estimation and testing of econometric models with data dependent smoothing parameters. Journal of Econometrics, 157(1):179–190, July 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002796.

Li:2012:LSE

[LL12a] Dong Li and Shiqing Ling. On the least squares estimation of multiple-regime threshold autoregressive models. Journal of Econometrics, 167(1):240–253, March 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611002685.

Liang:2012:FCR

[LL12b] Zhongwen Liang and Qi Li. Functional coefficient regression models with time trend. Journal of Econometrics, 170(1):15–31, September 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000784.

Li:2015:ANS

[LL15a] Cong Li and Zhongwen Liang. Asymptotics for nonparametric and semiparametric fixed effects panel models. Journal of Econometrics, 185(2):420–434, April 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614002942.

Li:2015:HPD

[LL15b] Qi Li and Tong Li. Heterogeneity in panel data and in nonparametric analysis. Journal of Econometrics, 188(2):313–315, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000706.

Li:2016:LCQ

[LL16] Degui Li and Runze Li. Local composite quantile regression smoothing for Harris recurrent Markov processes. Journal of Econometrics, 194(1):44–56, September 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300586.

Li:2015:FSF

[LLL15a] Degui Li, Oliver Linton, and Zudi Lu. A flexible semiparametric forecasting model for time series. Journal of Econometrics, 187(1):345–357, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000500.

Li:2015:NHG

[LLL15b] Muyi Li, Wai Keung Li, and Guodong Li. A new hyperbolic GARCH model. Journal of Econometrics, 189(2):428–436, December 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615001128.

Li:2016:CMS

[LLL16] Hongjun Li, Qi Li, and Ruixuan Liu. Consistent model specification tests based on k-nearest-neighbor estimation method. Journal of Econometrics, 194(1):187–202, September 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301075.

Li:2018:QML

[LLL18] Kunpeng Li, Qi Li, and Lina Lu. Quasi maximum likelihood analysis of high dimensional constrained factor models. Journal of Econometrics, 206(2):574–612, October 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301118.

Lin:2014:CNT

[LLS14] Zhongjian Lin, Qi Li, and Yiguo Sun. A consistent nonparametric test of parametric regression functional form in fixed effects panel data models. Journal of Econometrics, 178(Part 1):167–179, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001644.

Lewbel:2015:STT

[LLS15] Arthur Lewbel, Xun Lu, and Liangjun Su. Specification testing for transformation models with an application to generalized accelerated failure-time models. Journal of Econometrics, 184(1): 81–96, January 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614001936.

Li:2017:DNF

[LLS17] Hongjun Li, Qi Li, and Yutang Shi. Determining the number of factors when the number of factors can increase with sample size. Journal of Econometrics, 197(1):76–86, March 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616302020.

Li:2015:BCS

[LLY15] Yong Li, Xiao-Bin Liu, and Jun Yu. A Bayesian chi-squared test for hypothesis testing. Journal of Econometrics, 189(1):54–69, November 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761500192X.

Li:2015:AIM

[LLZ15] Dong Li, Shiqing Ling, and Jean-Michel Zakoïan. Asymptotic inference in multiple-threshold double autoregressive models. Journal of Econometrics, 189(2):415–427, December 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001116.

List:2011:CEU

[LM11] John A. List and Charles F. Mason. Are CEOs expected utility maximizers? Journal of Econometrics, 162(1):114–123, May 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760900267X.

Lee:2012:JEP

[LM12] Suzanne S. Lee and Per A. Mykland. Jumps in equilibrium prices and market microstructure noise. Journal of Econometrics, 168(2):396–406, June 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000711.

Lewbel:2011:EFD

[LML11] Arthur Lewbel, Daniel McFadden, and Oliver Linton. Estimating features of a distribution from binomial data. Journal of Econometrics, 162(2):170–188, June 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610002101.

Lanne:2017:IEN

[LMS17] Markku Lanne, Mika Meitz, and Pentti Saikkonen. Identification and estimation of non-Gaussian structural vector autoregressions. Journal of Econometrics, 196(2):288–304, February 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301828.

Li:2012:BAP

[LMT12] Mingliang Li, Kevin J. Mumford, and Justin L. Tobias. A Bayesian analysis of payday loans and their regulation. Journal of Econometrics, 171(2):205–216, December 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001571.

Ling:2015:FTS

[LMT15] Shiqing Ling, Michael McAleer, and Howell Tong. Frontiers in time series and financial econometrics: An overview. Journal of Econometrics, 189(2):245–250, December 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000974.

Liu:2010:MEM

[LMvdK10] Haiyong Liu, Thomas A. Mroz, and Wilbert van der Klaauw. Maternal employment, migration, and child development. Journal of Econometrics, 156(1):212–228, May 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002176.

Lee:2017:MIE

[LMZ17] Nayoung Lee, Hyungsik Roger Moon, and Qiankun Zhou. Many IVs estimation of dynamic panel regression models with measurement error. Journal of Econometrics, 200(2):251–259, October 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300957.

Lee:2012:HHT

[LO12] Yoonseok Lee and Ryo Okui. Hahn-hausman test as a specification test. Journal of Econometrics, 167(1):133–139, March 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002430.

Li:2015:SIC

[LO15] Tong Li and Tatsushi Oka. Set identification of the censored quantile regression model for short panels with fixed effects. Journal of Econometrics, 188(2):363–377, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000743.

Lee:2018:AID

[LOS18] Yoon-Jin Lee, Ryo Okui, and Mototsugu Shintani. Asymptotic inference for dynamic panel estimators of infinite order autoregressive processes. Journal of Econometrics, 204(2):147– 158, June 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618300186.

Lavergne:2013:SMD

[LP13] Pascal Lavergne and Valentin Patilea. Smooth minimum distance estimation and testing with conditional estimating equations: Uniform in bandwidth theory. Journal of Econometrics, 177(1):47–59, November 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407613001280.

Lee:2015:MSP

[LP15] Yoonseok Lee and Peter C. B. Phillips. Model selection in the presence of incidental parameters. Journal of Econometrics, 188(2):474–489, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000810.

Lieberman:2017:MSU

[LP17] Offer Lieberman and Peter C. B. Phillips. A multivariate stochastic unit root model with an application to derivative pricing. Journal of Econometrics, 196(1):99–110, January 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407616301695.

Li:2018:AIA

[LP18a] Jia Li and Andrew J. Patton. Asymptotic inference about predictive accuracy using high frequency data. Journal of Econometrics, 203(2):223–240, April 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761730235X.

Liu:2018:RTN

[LP18b] Xiaodong Liu and Ingmar R. Prucha. A robust test for network generated dependence. Journal of Econometrics, 207(1):92–113, November 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301192.

Lin:2018:EEC

[LPLZ18] Huazhen Lin, Lixian Pan, Shaogao Lv, and Wenyang Zhang. Efficient estimation and computation for the generalised additive models with unknown link function. Journal of Econometrics, 202(2):230–244, February 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617302208.

Liu:2015:DAB

[LPS15] Lily Y. Liu, Andrew J. Patton, and Kevin Sheppard. Does anything beat 5-minute RV? A comparison of realized measures across multiple asset classes. Journal of Econometrics, 187(1): 293–311, July 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000329.

Leeper:2010:DFF

[LPT10] Eric M. Leeper, Michael Plante, and Nora Traum. Dynamics of fiscal financing in the United States. Journal of Econometrics, 156(2):304–321, June 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002905.

Li:2019:BPM

[LPZ19] Chuhui Li, D. S. Poskitt, and Xueyan Zhao. The bivariate probit model, maximum likelihood estimation, pseudo true parameters and partial identification. Journal of Econometrics, 209(1):94– 113, March 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302562.

Liesenfeld:2010:DIM

[LR10a] Roman Liesenfeld and Jean-François Richard. The dynamic invariant multinomial probit model: Identification, pretesting and estimation. Journal of Econometrics, 155(2):117–127, April 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002346.

Liesenfeld:2010:EEP

[LR10b] Roman Liesenfeld and Jean-François Richard. Efficient estimation of probit models with correlated errors. Journal of Econometrics, 156(2):367–376, June 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002954.

Lee:2015:PNR

[LR15] Jungyoon Lee and Peter M. Robinson. Panel nonparametric regression with fixed effects. Journal of Econometrics, 188(2):346– 362, October 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000731.

Lee:2016:SEU

[LR16] Jungyoon Lee and Peter M. Robinson. Series estimation under cross-sectional dependence. Journal of Econometrics, 190(1):1– 17, January 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002213.

LaVecchia:2019:SAS

[LR19] Davide La Vecchia and Elvezio Ronchetti. Saddlepoint approximations for short and long memory time series: a frequency domain approach. Journal of Econometrics, 213(2):578–592, December 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619301629.

Laurent:2013:LFR

[LRV13] Sébastien Laurent, Jeroen V. K. Rombouts, and Francesco Violante. On loss functions and ranking forecasting performances of multivariate volatility models. Journal of Econometrics, 173(1): 1–10, March 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001777.

Ley:2012:MPB

[LS12] Eduardo Ley and Mark F. J. Steel. Mixtures of g-priors for Bayesian model averaging with economic applications. Journal of Econometrics, 171(2):251–266, December 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761200156X.

Lee:2015:SML

[LS15a] Donghoon Lee and Kyungchul Song. Simulated maximum likelihood estimation for discrete choices using transformed simulated frequencies. Journal of Econometrics, 187(1):131–153, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000238.

Li:2015:DAC

[LS15b] Xianghong Li and Barry Smith. Diagnostic analysis and computational strategies for estimating discrete time duration models — a Monte Carlo study. Journal of Econometrics, 187(1): 275–292, July 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000494.

Lu:2015:JMA

[LS15c] Xun Lu and Liangjun Su. Jackknife model averaging for quantile regressions. Journal of Econometrics, 188(1):40–58, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001256.

Lu:2016:SED

[LS16] Xun Lu and Liangjun Su. Shrinkage estimation of dynamic panel data models with interactive fixed effects. Journal of Econometrics, 190(1):148–175, January 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002389.

Liang:2019:DVE

[LS19a] Chong Liang and Melanie Schienle. Determination of vector error correction models in high dimensions. Journal of Econometrics, 208(2):418–441, February 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301969.

Liao:2019:BIP

[LS19b] Yuan Liao and Anna Simoni. Bayesian inference for partially identified smooth convex models. Journal of Econometrics, 211(2):338–360, August 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619300399.

Lee:2014:DET

[LSL14] Yoon Dong Lee, Seongjoo Song, and Eun-Kyung Lee. The delta expansion for the transition density of diffusion models. Journal of Econometrics, 178(Part 3):694–705, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613002212.

Linton:2010:IBT

[LSW10] Oliver Linton, Kyungchul Song, and Yoon-Jae Whang. An improved bootstrap test of stochastic dominance. Journal of Econometrics, 154(2):186–202, February 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609001882.

Lee:2013:TFI

[LSW13] Sokbae Lee, Kyungchul Song, and Yoon-Jae Whang. Testing functional inequalities. Journal of Econometrics, 172(1):14–32, January 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761200190X.

Li:2011:BIC

[LT11] Mingliang Li and Justin L. Tobias. Bayesian inference in a correlated random coefficients model: Modeling causal effect heterogeneity with an application to heterogeneous returns to schooling. Journal of Econometrics, 162(2):345–361, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000339.

Levy:2014:ACC

[LT14a] Joshua B. Levy and Murad S. Taqqu. The asymptotic codifference and covariation of log-fractional stable noise. Journal of Econometrics, 181(1):34–43, July 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000359.

Liu:2014:QML

[LT14b] Cheng Liu and Cheng Yong Tang. A quasi-maximum likelihood approach for integrated covariance matrix estimation with high frequency data. Journal of Econometrics, 180(2):217–232, June 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761400030X.

Lewbel:2015:IEG

[LT15a] Arthur Lewbel and Xun Tang. Identification and estimation of games with incomplete information using excluded regressors. Journal of Econometrics, 189(1):229–244, November 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001943.

Liu:2015:IVR

[LT15b] Shouwei Liu and Yiu-Kuen Tse. Intraday value-at-risk: an asymmetric autoregressive conditional duration approach. Journal of Econometrics, 189(2):437–446, December 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761500113X.

Larsen:2019:VNE

[LT19] Vegard H. Larsen and Leif A. Thorsrud. The value of news for economic developments. Journal of Econometrics, 210(1):203– 218, May 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302148.

Li:2016:ITV

[LTT16] Jia Li, Viktor Todorov, and George Tauchen. Inference theory for volatility functional dependencies. Journal of Econometrics, 193 (1):17–34, July 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300033.

Li:2017:AEC

[LTT17] Jia Li, Viktor Todorov, and George Tauchen. Adaptive estimation of continuous-time regression models using highfrequency data. Journal of Econometrics, 200(1):36–47, September 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300635.

Li:2017:MSJ

[LTTC17] Jia Li, Viktor Todorov, George Tauchen, and Rui Chen. Mixed-scale jump regressions with bootstrap inference. Journal of Econometrics, 201(2):417–432, December 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617301690.

Lee:2014:NSR

[LTU14] Tae-Hwy Lee, Yundong Tu, and Aman Ullah. Nonparametric and semiparametric regressions subject to monotonicity constraints: Estimation and forecasting. Journal of Econometrics, 182(1):196–210, September 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000773.

Liu:2017:RIB

[LVX17] Nianqing Liu, Quang Vuong, and Haiqing Xu. Rationalization and identification of binary games with correlated types. Journal of Econometrics, 201(2):249–268, December 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301574.

Lee:2010:AWE

[LW10a] Donghoon Lee and Kenneth I. Wolpin. Accounting for wage and employment changes in the US from 1968–2000: a dynamic model of labor market equilibrium. Journal of Econometrics, 156 (1):68–85, May 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609002073.

Lieli:2010:CEC

[LW10b] Robert P. Lieli and Halbert White. The construction of empirical credit scoring rules based on maximization principles. Journal of Econometrics, 157(1):110–119, July 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002814.

Lu:2014:RCR

[LW14a] Xun Lu and Halbert White. Robustness checks and robustness tests in applied economics. Journal of Econometrics, 178 (Part 1):194–206, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407613001668.

Lu:2014:TSS

[LW14b] Xun Lu and Halbert White. Testing for separability in structural equations. Journal of Econometrics, 182(1):14–26, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000645.

Linton:2016:NTS

[LWY16] Oliver Linton, Yoon-Jae Whang, and Yu-Min Yen. A nonparametric test of a strong leverage hypothesis. Journal of Econometrics, 194(1):153–186, September 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616301099.

Linton:2019:EEN

[LX19] Oliver Linton and Zhijie Xiao. Efficient estimation of nonparametric regression in the presence of dynamic heteroskedasticity. Journal of Econometrics, 213(2):608–631, December 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301666.

Liu:2016:CAM

[LXC16] Xialu Liu, Han Xiao, and Rong Chen. Convolutional autoregressive models for functional time series. Journal of Econometrics, 194(2):263–282, October 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300963.

Li:2016:EEI

[LXZ16] Yingying Li, Shangyu Xie, and Xinghua Zheng. Efficient estimation of integrated volatility incorporating trading information. Journal of Econometrics, 195(1):33–50, November 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301336.

Li:2012:BHT

[LY12] Yong Li and Jun Yu. Bayesian hypothesis testing in latent variable models. Journal of Econometrics, 166(2):237–246, February 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002211.

Lewbel:2016:IAT

[LY16] Arthur Lewbel and Thomas Tao Yang. Identifying the average treatment effect in ordered treatment models without unconfoundedness. Journal of Econometrics, 195(1):1–22, November 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301063.

Liu:2019:UTP

[LYCP19] Xiaohui Liu, Bingduo Yang, Zongwu Cai, and Liang Peng. A unified test for predictability of asset returns regardless of properties of predicting variables. Journal of Econometrics, 208(1):141– 159, January 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761830174X.

Li:2018:STB

[LYZ18] Yong Li, Jun Yu, and Tao Zeng. Specification tests based on MCMC output. Journal of Econometrics, 207(1):237–260, November 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301416.

Li:2012:IAB

[LZ12] Tong Li and Xiaoyong Zheng. Information acquisition and/or bid preparation: a structural analysis of entry and bidding in timber sale auctions. Journal of Econometrics, 168(1):29–46, May 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001679.

Linton:2019:ESI

[LZ19] Oliver Linton and Zhengjun Zhang. Editorial for the special issue on financial engineering and risk management for JoE. Journal of Econometrics, 208(1):1–4, January 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301672.

Li:2018:FMA

[LZK18] Jialiang Li, Wenyang Zhang, and Efang Kong. Factor models for asset returns based on transformed factors. Journal of Econometrics, 207(2):432–448, December 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301660.

Lan:2016:TSR

[LZL^{+}16] Wei Lan, Ping-Shou Zhong, Runze Li, Hansheng Wang, and Chih-Ling Tsai. Testing a single regression coefficient in high dimensional linear models. Journal of Econometrics, 195(1):154– 168, November 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616301087.

Li:2018:UAV

[LZL18] Yingying Li, Zhiyuan Zhang, and Yichu Li. A unified approach to volatility estimation in the presence of both rounding and random market microstructure noise. Journal of Econometrics, 203(2): 187–222, April 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617302294.

Lin:2018:REE

[LZW^{+}18] Huazhen Lin, Fanyin Zhou, Qiuxia Wang, Ling Zhou, and Jing Qin. Robust and efficient estimation for the treatment effect in causal inference and missing data problems. Journal of Econometrics, 205(2):363–380, August 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618300563.

Li:2014:NAB

[LZY14] Yong Li, Tao Zeng, and Jun Yu. A new approach to Bayesian hypothesis testing. Journal of Econometrics, 178(Part 3):602– 612, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613001991.

Li:2018:ZGM

[LZZL18] Dong Li, Xingfa Zhang, Ke Zhu, and Shiqing Ling. The ZD-GARCH model: a new way to study heteroscedasticity. Journal of Econometrics, 202(1):1–17, January 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407617301926.

Liao:2019:MAB

[LZZZ19] Jun Liao, Xianpeng Zong, Xinyu Zhang, and Guohua Zou. Model averaging based on leave-subject-out cross-validation for vector autoregressions. Journal of Econometrics, 209(1):35–60, March 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302598.

Macaro:2010:BNP

[Mac10] Christian Macaro. Bayesian non-parametric signal extraction for Gaussian time series. Journal of Econometrics, 157(2):381– 395, August 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000849.

Magdalinos:2012:MEA

[Mag12] Tassos Magdalinos. Mildly explosive autoregression under weak and strong dependence. Journal of Econometrics, 169(2):179– 187, August 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000346.

Manski:2016:CIE

[Man16] Charles F. Manski. Credible interval estimates for official statistics with survey nonresponse. Journal of Econometrics, 191(2): 293–301, April 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761500278X.

Maruyama:2014:EFS

[Mar14] Shiko Maruyama. Estimation of finite sequential games. Journal of Econometrics, 178(2):716–726, February 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613002261.

Massacci:2017:LSE

[Mas17] Daniele Massacci. Least squares estimation of large dimensional threshold factor models. Journal of Econometrics, 197(1):101– 129, March 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616302111.

Matzkin:2012:INL

[Mat12] Rosa L. Matzkin. Identification in nonparametric limited dependent variable models with simultaneity and unobserved heterogeneity. Journal of Econometrics, 166(1):106–115, January 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001229.

Matzkin:2016:ICN

[Mat16] Rosa L. Matzkin. On independence conditions in nonseparable models: Observable and unobservable instruments.

Journal of Econometrics, 191(2):302–311, April 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002791.

Matzkin:2019:CIS

[Mat19] Rosa L. Matzkin. Constructive identification in some nonseparable discrete choice models. Journal of Econometrics, 211(1): 83–103, July 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302367.

Majewski:2015:SPG

[MBC15] Adam A. Majewski, Giacomo Bormetti, and Fulvio Corsi. Smile from the past: a general option pricing framework with multiple volatility and leverage components. Journal of Econometrics, 187(2):521–531, August 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000615.

McCausland:2012:HMH

[McC12] William J. McCausland. The HESSIAN method: Highly efficient simulation smoothing, in a nutshell. Journal of Econometrics, 168(2):189–206, June 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611002752.

McCloskey:2017:BBS

[McC17] Adam McCloskey. Bonferroni-based size-correction for nonstandard testing problems. Journal of Econometrics, 200(1):17–35, September 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300556.

McFadden:2012:EJP

[McF12] Daniel McFadden. Economic juries and public project provision. Journal of Econometrics, 166(1):116–126, January 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001230.

Mikosch:2013:HTO

[MdV13] Thomas Mikosch and Casper G. de Vries. Heavy tails of OLS. Journal of Econometrics, 172(2):205–221, February 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001996.

Menzel:2014:CEM

[Men14] Konrad Menzel. Consistent estimation with many moment inequalities. Journal of Econometrics, 182(2):329–350, October 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001389.

Mikkelsen:2019:CET

[MHU19] Jakob Guldbæk Mikkelsen, Eric Hillebrand, and Giovanni Urga. Consistent estimation of time-varying loadings in highdimensional factor models. Journal of Econometrics, 208(2):535– 562, February 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761830215X.

Mikusheva:2010:RCS

[Mik10] Anna Mikusheva. Robust confidence sets in the presence of weak instruments. Journal of Econometrics, 157(2):236–247, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609003029.

Miyauchi:2016:SEP

[Miy16] Yuhei Miyauchi. Structural estimation of pairwise stable networks with nonnegative externality. Journal of Econometrics, 195(2):224–235, December 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616301592.

Mittelhammer:2011:FEL

[MJ11] Ron C. Mittelhammer and George Judge. A family of empirical likelihood functions and estimators for the binary response model. Journal of Econometrics, 164(2):207–217, October 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000819.

Mesters:2014:GDP

[MK14] G. Mesters and S. J. Koopman. Generalized dynamic panel data models with random effects for cross-section and time. Journal of Econometrics, 180(2):127–140, June 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761400044X.

Malikov:2016:VCP

[MKS16] Emir Malikov, Subal C. Kumbhakar, and Yiguo Sun. Varying coefficient panel data model in the presence of endogenous selectivity and fixed effects. Journal of Econometrics, 190(2):233– 251, February 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615001761.

Maheu:2011:DHF

[MM11] John M. Maheu and Thomas H. McCurdy. Do high-frequency measures of volatility improve forecasts of return distributions? Journal of Econometrics, 160(1):69–76, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000588.

Medeiros:2016:RHD

[MM16] Marcelo C. Medeiros and Eduardo F. Mendes. l_{1}-regularization of high-dimensional time-series models with non-Gaussian and heteroskedastic errors. Journal of Econometrics, 191(1):255– 271, March 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002638.

McDonough:2017:MDI

[MM17a] Ian K. McDonough and Daniel L. Millimet. Missing data, imputation, and endogeneity. Journal of Econometrics, 199(2):141– 155, August 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300684.

Meyer:2017:MBC

[MM17b] Bruce D. Meyer and Nikolas Mittag. Misclassification in binary choice models. Journal of Econometrics, 200(2):295–311, October 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300982.

Moreira:2019:OTS

[MM19] Humberto Moreira and Marcelo J. Moreira. Optimal two-sided tests for instrumental variables regression with heteroskedastic and autocorrelated errors. Journal of Econometrics, 213(2):398– 433, December 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619301393.

Maneesoonthorn:2012:PFV

[MMFG12] Worapree Maneesoonthorn, Gael M. Martin, Catherine S. Forbes, and Simone D. Grose. Probabilistic forecasts of volatility and its risk premia. Journal of Econometrics, 171(2):217–236, December 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001534.

Meddahi:2011:RV

[MMS11] Nour Meddahi, Per Mykland, and Neil Shephard. Realized volatility. Journal of Econometrics, 160(1):1, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001570.

Ma:2019:IFP

[MMS19] Jun Ma, Vadim Marmer, and Artyom Shneyerov. Inference for first-price auctions with Guerre, Perrigne, and Vuong’s estimator. Journal of Econometrics, 211(2):507–538, August 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300478.

Mills:2014:TBS

[MMV14] Benjamin Mills, Marcelo J. Moreira, and Lucas P. Vilela. Tests based on t-statistics for IV regression with weak instruments. Journal of Econometrics, 182(2):351–363, October 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001067.

Marmer:2012:OCM

[MO12] Vadim Marmer and Taisuke Otsu. Optimal comparison of misspecified moment restriction models under a chosen measure of fit. Journal of Econometrics, 170(2):538–550, October 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001339.

Mourifie:2015:SBT

[Mou15] Ismael Mourifié. Sharp bounds on treatment effects in a binary triangular system. Journal of Econometrics, 187(1):74–81, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000251.

Miller:2010:NNT

[MP10] J. Isaac Miller and Joon Y. Park. Nonlinearity, nonstationarity, and thick tails: How they interact to generate persistence in memory. Journal of Econometrics, 155(1):83–89, March 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002139.

Malik:2011:PFC

[MP11] Sheheryar Malik and Michael K. Pitt. Particle filters for continuous likelihood evaluation and maximisation. Journal of Econometrics, 165(2):190–209, December 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611001473.

Mariano:2012:STM

[MP12a] Roberto S. Mariano and Daniel Preve. Statistical tests for multiple forecast comparison. Journal of Econometrics, 169(1):123– 130, July 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000152.

Moon:2012:BPU

[MP12b] H. R. Moon and B. Perron. Beyond panel unit root tests: Using multiple testing to determine the nonstationarity properties of individual series in a panel. Journal of Econometrics, 169(1):

29–33, July 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000097.

McCulloch:2013:ENS

[MP13a] J. Huston McCulloch and E. Richard Percy. Extended Neyman smooth goodness-of-fit tests, applied to competing heavy-tailed distributions. Journal of Econometrics, 172(2):275–282, February 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612002023.

McElroy:2013:DTS

[MP13b] Tucker McElroy and Dimitris N. Politis. Distribution theory for the Studentized mean for long, short, and negative memory time series. Journal of Econometrics, 177(1):60–74, November 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613001334.

McElroy:2014:SDS

[MP14] Tucker S. McElroy and Dimitris N. Politis. Spectral density and spectral distribution inference for long memory time series via fixed-b asymptotics. Journal of Econometrics, 182(1):211–225, September 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614000785.

Mosconi:2017:ICS

[MP17] Rocco Mosconi and Paolo Paruolo. Identification conditions in simultaneous systems of cointegrating equations with integrated variables of higher order. Journal of Econometrics, 198(2):271– 276, June 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407617300222.

Magnus:2010:CTM

[MPP10] Jan R. Magnus, Owen Powell, and Patricia Prüfer. A comparison of two model averaging techniques with an application to growth empirics. Journal of Econometrics, 154(2):139–153, February 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609001663.

Monfort:2017:SZA

[MPRR17] Alain Monfort, Fulvio Pegoraro, Jean-Paul Renne, and Guillaume Roussellet. Staying at zero with affine processes: an application to term structure modelling. Journal of Econometrics, 201(2):348–366, December 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617301653.

Mancini:2011:TEM

[MR11] Cecilia Mancini and Roberto Renò. Threshold estimation of Markov models with jumps and interest rate modeling. Journal of Econometrics, 160(1):77–92, January 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000618.

Maasoumi:2016:SAA

[MR16] Esfandiar Maasoumi and Jeffrey S. Racine. A solution to aggregation and an application to multidimensional ‘well-being’ frontiers. Journal of Econometrics, 191(2):374–383, April 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002845.

Monfort:2015:QKF

[MRR15] Alain Monfort, Jean-Paul Renne, and Guillaume Roussellet. A quadratic Kalman filter. Journal of Econometrics, 187(1):43–56, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000123.

Montes-Rojas:2011:RTH

[MRSE11] Gabriel Montes-Rojas and Walter Sosa-Escudero. Robust tests for heteroskedasticity in the one-way error components model. Journal of Econometrics, 160(2):300–310, February 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001909.

Maller:2016:LSD

[MRT16] Ross Maller, Steven Roberts, and Rabee Tourky. The largesample distribution of the maximum Sharpe ratio with and without short sales. Journal of Econometrics, 194(1):138–152, September 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300598.

Marmer:2012:QBN

[MS12] Vadim Marmer and Artyom Shneyerov. Quantile-based nonparametric inference for first-price auctions. Journal of Econometrics, 167(2):345–357, April 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611002016.

Marcellino:2016:MFO

[MS16] Massimiliano Marcellino and Vasja Sivec. Monetary, fiscal and oil shocks: Evidence based on mixed frequency structural FAVARs. Journal of Econometrics, 193(2):335–348, August 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300689.

Malikov:2017:SET

[MS17] Emir Malikov and Yiguo Sun. Semiparametric estimation and testing of smooth coefficient spatial autoregressive models. Journal of Econometrics, 199(1):12–34, July 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761730043X.

Ma:2018:ELD

[MS18] Shujie Ma and Liangjun Su. Estimation of large dimensional factor models with an unknown number of breaks. Journal of Econometrics, 207(1):1–29, November 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301155.

Machado:2019:QM

[MS19] José A. F. Machado and J. M. C. Santos Silva. Quantiles via moments. Journal of Econometrics, 213(1):145–173, November 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300648.

Muller:2011:FDA

[MSS11] Hans-Georg Müller, Rituparna Sen, and Ulrich Stadtmüller. Functional data analysis for volatility. Journal of Econometrics, 165(2):233–245, December 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611001606.

Machado:2019:IVS

[MSV19] Cecilia Machado, Azeem M. Shaikh, and Edward J. Vytlacil. Instrumental variables and the sign of the average treatment effect. Journal of Econometrics, 212(2):522–555, October 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301381.

Mavroeidis:2015:EHA

[MSW15] Sophocles Mavroeidis, Yuya Sasaki, and Ivo Welch. Estimation of heterogeneous autoregressive parameters with short panel data. Journal of Econometrics, 188(1):219–235, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001517.

Meijer:2017:CEL

[MSW17] Erik Meijer, Laura Spierdijk, and Tom Wansbeek. Consistent estimation of linear panel data models with measurement error. Journal of Econometrics, 200(2):169–180, October 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300842.

Moon:2018:ERC

[MSW18] Hyungsik Roger Moon, Matthew Shum, and Martin Weidner. Estimation of random coefficients logit demand models with interactive fixed effects. Journal of Econometrics, 206(2):613– 644, October 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761830112X.

Manner:2019:TSB

[MSW19] Hans Manner, Florian Stark, and Dominik Wied. Testing for structural breaks in factor copula models. Journal of Econometrics, 208(2):324–345, February 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301842.

Marmer:2013:WME

[MSX13] Vadim Marmer, Artyom Shneyerov, and Pai Xu. What model for entry in first-price auctions? A nonparametric approach. Journal of Econometrics, 176(1):46–58, September 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000821.

Molinari:2012:AII

[MT12] Francesca Molinari and Elie Tamer. Annals issue on identification and decisions. Journal of Econometrics, 166(1):1–2, January 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002545.

Merlo:2019:NRI

[MT19] Antonio Merlo and Xun Tang. New results on the identification of stochastic bargaining models. Journal of Econometrics, 209(1): 79–93, March 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302550.

Moon:2013:TDB

[MV13] Seongman Moon and Carlos Velasco. Tests for m-dependence based on sample splitting methods. Journal of Econometrics, 173(2):143–159, April 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612002679.

Muller:2013:LFR

[MW13] Ulrich K. Müller and Mark W. Watson. Low-frequency robust cointegration testing. Journal of Econometrics, 174(2): 66–81, June 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761300033X.

Murtazashvili:2016:CFA

[MW16] Irina Murtazashvili and Jeffrey M. Wooldridge. A control function approach to estimating switching regression models with endogenous explanatory variables and endogenous switching. Journal of Econometrics, 190(2):252–266, February 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001839.

Maasoumi:2017:WCW

[MW17] Esfandiar Maasoumi and Le Wang. What can we learn about the racial gap in the presence of sample selection? Journal of Econometrics, 199(2):117–130, August 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300660.

McAlinn:2019:DBP

[MW19a] Kenichiro McAlinn and Mike West. Dynamic Bayesian predictive synthesis in time series forecasting. Journal of Econometrics, 210(1):155–169, May 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618302112.

Muller:2019:NWR

[MW19b] Ulrich K. Müller and Yulong Wang. Nearly weighted risk minimal unbiased estimation. Journal of Econometrics, 209(1):18– 34, March 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618302586.

Mariano:2012:RAN

[MXY12a] Roberto S. Mariano, Zhijie Xiao, and Jun Yu. Recent advances in nonstationary time series: a festschrift in honor of Peter C. B. Phillips. Journal of Econometrics, 169(2):139–141, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000255.

Mariano:2012:RAP

[MXY12b] Roberto S. Mariano, Zhijie Xiao, and Jun Yu. Recent advances in panel data, nonlinear and nonparametric models: a festschrift in honor of Peter C. B. Phillips. Journal of Econometrics, 169 (1):1–3, July 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000036.

Mykland:2019:CSI

[Myk19] Per Aslak Mykland. Combining statistical intervals and market prices: the worst case state price distribution. Journal of Econometrics, 212(1):272–285, September 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300855.

Mykland:2016:BDC

[MZ16] Per A. Mykland and Lan Zhang. Between data cleaning and inference: Pre-averaging and robust estimators of the efficient price. Journal of Econometrics, 194(2):242–262, October 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300951.

Mao:2018:STI

[MZ18] Guangyu Mao and Zhengjun Zhang. Stochastic tail index model for high frequency financial data with Bayesian analysis. Journal of Econometrics, 205(2):470–487, August 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300678.

Mykland:2019:ATS

[MZC19] Per A. Mykland, Lan Zhang, and Dachuan Chen. The algebra of two scales estimation, and the S-TSRV: High frequency estimation that is robust to sampling times. Journal of Econometrics, 208(1):101–119, January 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301726.

Nguimkeu:2019:ETE

[NDT19] Pierre Nguimkeu, Augustine Denteh, and Rusty Tchernis. On the estimation of treatment effects with endogenous misreporting. Journal of Econometrics, 208(2):487–506, February 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301994.

Nishiyama:2011:CNT

[NHKJ11] Yoshihiko Nishiyama, Kohtaro Hitomi, Yoshinori Kawasaki, and Kiho Jeong. A consistent nonparametric test for nonlinear causality — specification in time series regression. Journal of Econometrics, 165(1):112–127, November 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611001023.

Nielsen:2010:NCA

[Nie10] Morten Ørregaard Nielsen. Nonparametric cointegration analysis of fractional systems with unknown integration orders. Journal of Econometrics, 155(2):170–187, April 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002383.

Nakamura:2011:PDR

[NNN11] Alice O. Nakamura, Emi Nakamura, and Leonard I. Nakamura. Price dynamics, retail chains and inflation measurement. Journal of Econometrics, 161(1):47–55, March 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001855.

Nolan:2013:LNR

[NOR13] John P. Nolan and Diana Ojeda-Revah. Linear and nonlinear regression with stable errors. Journal of Econometrics, 172(2):186– 194, February 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001923.

Norets:2015:BRN

[Nor15] Andriy Norets. Bayesian regression with nonparametric heteroskedasticity. Journal of Econometrics, 185(2):409–419, April 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614002966.

Norets:2012:BMJ

[NP12] Andriy Norets and Justinas Pelenis. Bayesian modeling of joint and conditional distributions. Journal of Econometrics, 168(2): 332–346, June 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000577.

Nicoletti:2010:MSD

[NR10] Cheti Nicoletti and Concetta Rondinelli. The (mis)specification of discrete duration models with unobserved heterogeneity: a Monte Carlo study. Journal of Econometrics, 159(1):1–13, November 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610001090.

Noureldin:2014:MRA

[NSS14] Diaa Noureldin, Neil Shephard, and Kevin Sheppard. Multivariate rotated ARCH models. Journal of Econometrics, 179(1): 16–30, March 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407613002078.

Orea:2019:NSF

[OÁ19] Luis Orea and Inmaculada C. Álvarez. A new stochastic frontier model with cross-sectional effects in both noise and inefficiency terms. Journal of Econometrics, 213(2):556–577, December 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301599.

ODonnell:2016:UIA

[O’D16] C. J. O’Donnell. Using information about technologies, markets and firm behaviour to decompose a proper productivity index. Journal of Econometrics, 190(2):328–340, February 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001785.

Ogata:2013:EMS

[Oga13] Hiroaki Ogata. Estimation for multivariate stable distributions with generalized empirical likelihood. Journal of Econometrics, 172(2):248–254, February 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612002011.

Okui:2011:IVE

[Oku11] Ryo Okui. Instrumental variable estimation in the presence of many moment conditions. Journal of Econometrics, 165(1):70– 86, November 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611000996.

Onatski:2012:APC

[Ona12] Alexei Onatski. Asymptotics of the principal components estimator of large factor models with weakly influential factors. Journal of Econometrics, 168(2):244–258, June 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000449.

Onatski:2015:AAS

[Ona15] Alexei Onatski. Asymptotic analysis of the squared estimation error in misspecified factor models. Journal of Econometrics, 186(2):388–406, June 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000408.

Okhrin:2013:SEH

[OOS13] Ostap Okhrin, Yarema Okhrin, and Wolfgang Schmid. On the structure and estimation of hierarchical Archimedean copulas. Journal of Econometrics, 173(2):189–204, April 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612002667.

Ouyang:2015:TEE

[OP15] Min Ouyang and Yulei Peng. The treatment-effect estimation: a case study of the 2008 economic stimulus package of China. Journal of Econometrics, 188(2):545–557, October 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761500086X.

Oh:2016:HDC

[OP16] Dong Hwan Oh and Andrew J. Patton. High-dimensional copula-based distributions with mixed frequency data. Journal of Econometrics, 193(2):349–366, August 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300707.

Oka:2018:TCB

[OP18] Tatsushi Oka and Pierre Perron. Testing for common breaks in a multiple equations system. Journal of Econometrics, 204(1): 66–85, May 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618300071.

Oka:2011:ESC

[OQ11] Tatsushi Oka and Zhongjun Qu. Estimating structural changes in regression quantiles. Journal of Econometrics, 162(2):248– 267, June 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611000261.

Otsu:2012:TNN

[OSW12] Taisuke Otsu, Myung Hwan Seo, and Yoon-Jae Whang. Testing for non-nested conditional moment restrictions using unconditional empirical likelihood. Journal of Econometrics, 167(2): 370–382, April 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761100203X.

Otsu:2010:BEE

[Ots10] Taisuke Otsu. On Bahadur efficiency of empirical likelihood. Journal of Econometrics, 157(2):248–256, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609003030.

Onatski:2019:ECC

[OW19] Alexei Onatski and Chen Wang. Extreme canonical correlations and high-dimensional cointegration analysis. Journal of Econometrics, 212(1):307–322, September 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300879.

Otsu:2015:ELR

[OXM15] Taisuke Otsu, Ke-Li Xu, and Yukitoshi Matsushita. Empirical likelihood for regression discontinuity design. Journal of Econometrics, 186(1):94–112, May 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001444.

Okui:2019:PDA

[OY19] Ryo Okui and Takahide Yanagi. Panel data analysis with heterogeneous dynamics. Journal of Econometrics, 212(2):451–475, October 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619301022.

Pakel:2019:BRN

[Pak19] Cavit Pakel. Bias reduction in nonlinear and dynamic panels in the presence of cross-section dependence. Journal of Econometrics, 213(2):459–492, December 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619301484.

Park:2016:EAV

[Par16] Yang-Ho Park. The effects of asymmetric volatility and jumps on the pricing of VIX derivatives. Journal of Econometrics, 192(1): 313–328, May 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616000026.

Parker:2019:AIC

[Par19] Thomas Parker. Asymptotic inference for the constrained quantile regression process. Journal of Econometrics, 213(1):174–189, November 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761930065X.

Patton:2011:DBR

[Pat11a] Andrew J. Patton. Data-based ranking of realised volatility estimators. Journal of Econometrics, 161(2):284–303, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610002551.

Patton:2011:VFC

[Pat11b] Andrew J. Patton. Volatility forecast comparison using imperfect volatility proxies. Journal of Econometrics, 160(1):246–256, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761000076X.

Pesaran:2014:ALD

[PC14] M. Hashem Pesaran and Alexander Chudik. Aggregation in large dynamic panels. Journal of Econometrics, 178(Part 2):273– 285, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613001917.

Pitt:2012:SPM

[PdSSGK12] Michael K. Pitt, Ralph dos Santos Silva, Paolo Giordani, and Robert Kohn. On some properties of Markov chain Monte Carlo simulation methods based on the particle filter. Journal of Econometrics, 171(2):134–151, December 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001510.

Pedersen:2017:ITB

[Ped17] Rasmus Søndergaard Pedersen. Inference and testing on the boundary in extended constant conditional correlation GARCH models. Journal of Econometrics, 196(1):23–36, January 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407616301658.

Pelenis:2014:BRH

[Pel14] Justinas Pelenis. Bayesian regression with heteroscedastic error density and parametric mean function. Journal of Econometrics, 178(Part 3):624–638, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613002194.

Pelger:2019:LDF

[Pel19] Markus Pelger. Large-dimensional factor modeling based on high-frequency observations. Journal of Econometrics, 208(1): 23–42, January 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301696.

Petrova:2019:QBL

[Pet19] Katerina Petrova. A quasi-Bayesian local likelihood approach to time varying parameter VAR models. Journal of Econometrics, 212(1):286–306, September 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300867.

Papanicolaou:2016:VBT

[PG16] Alex Papanicolaou and Kay Giesecke. Variation-based tests for volatility misspecification. Journal of Econometrics, 191(1):217– 230, March 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002602.

Phillips:2017:SIR

[PG17] Peter C. B. Phillips and Wayne Yuan Gao. Structural inference from reduced forms with many instruments. Journal of Econometrics, 199(2):96–116, August 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300659.

Poskitt:2015:HOI

[PGM15] D. S. Poskitt, Simone D. Grose, and Gael M. Martin. Higherorder improvements of the sieve bootstrap for fractionally integrated processes. Journal of Econometrics, 188(1):94–110, September 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615001372.

Phillips:1991:CML

[Phi91] Robert F. Phillips. A constrained maximum-likelihood approach to estimating switching regressions. Journal of Econometrics, 48(1–2):241–262, April/May 1991. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/030440769190040K.

See note [XTZ10].

Phillips:2010:BD

[Phi10] Peter C. B. Phillips. Bootstrapping I(1) data. Journal of Econometrics, 158(2):280–284, October 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000138.

Phillips:2014:OEC

[Phi14] Peter C. B. Phillips. Optimal estimation of cointegrated systems with irrelevant instruments. Journal of Econometrics, 178(Part 2):210–224, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407613001863.

Park:2016:EQC

[PHL16] Sujin Park, Seok Young Hong, and Oliver Linton. Estimating the quadratic covariation matrix for asynchronously observed high frequency stock returns corrupted by additive measurement error. Journal of Econometrics, 191(2):325–347, April 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761500281X.

Pei:2018:NFE

[PHY18] Youquan Pei, Tao Huang, and Jinhong You. Nonparametric fixed effects model for panel data with locally stationary regressors. Journal of Econometrics, 202(2):286–305, February 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617302233.

Perera:2017:FTP

[PK17] Indeewara Perera and Hira L. Koul. Fitting a two phase threshold multiplicative error model. Journal of Econometrics, 197(2):348– 367, April 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616302329.

Post:2018:POB

[PKA18] Thierry Post, Selçuk Karabati, and Stelios Arvanitis. Portfolio optimization based on stochastic dominance and empirical likelihood. Journal of Econometrics, 206(1):167–186, September 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300824.

Phillips:2013:PRU

[PL13] Peter C. B. Phillips and Ji Hyung Lee. Predictive regression under various degrees of persistence and robust long-horizon regression. Journal of Econometrics, 177(2):250–264, December 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000882.

Proietti:2015:GAF

[PL15] Tommaso Proietti and Alessandra Luati. The generalised autocovariance function. Journal of Econometrics, 186(1):245–257, May 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001730.

Phillips:2016:REI

[PL16] Peter C. B. Phillips and Ji Hyung Lee. Robust econometric inference with mixed integrated and mildly explosive regressors. Journal of Econometrics, 192(2):433–450, June 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300124.

Phillips:2017:ESS

[PLG17] Peter C. B. Phillips, Degui Li, and Jiti Gao. Estimating smooth structural change in cointegration models. Journal of Econometrics, 196(1):180–195, January 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407616301804.

Pope:2011:AAR

[PLJ11] Rulon D. Pope, Jeffrey T. LaFrance, and Richard E. Just. Agricultural arbitrage and risk preferences. Journal of Econometrics, 162(1):35–43, May 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002607.

Preve:2011:LPB

[PM11] Daniel Preve and Marcelo C. Medeiros. Linear programmingbased estimators in simple linear regression. Journal of Econometrics, 165(1):128–136, November 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407611001035.

Potiron:2017:EIQ

[PM17] Yoann Potiron and Per A. Mykland. Estimation of integrated quadratic covariation with endogenous sampling times. Journal of Econometrics, 197(1):20–41, March 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616301981.

Phillips:2010:SLM

[PMG10] Peter C. B. Phillips, Tassos Magdalinos, and Liudas Giraitis. Smoothing local-to-moderate unit root theory. Journal of Econometrics, 158(2):274–279, October 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000126.

Peluso:2015:RUP

[PMM15] Stefano Peluso, Antonietta Mira, and Pietro Muliere. Reinforced urn processes for credit risk models. Journal of Econometrics, 184(1):1–12, January 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001791.

Poirier:2012:WSY

[Poi12] Dale J. Poirier. What is sensible for your agents should be sensible for yourself. Journal of Econometrics, 170(1):249–250, September 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000747.

Poirier:2017:EEM

[Poi17] Alexandre Poirier. Efficient estimation in models with independence restrictions. Journal of Econometrics, 196(1):1–22, January 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407616301646.

Portnoy:2019:ETS

[Por19] Stephen Portnoy. Edgeworth’s time series model: Not AR(1) but same covariance structure. Journal of Econometrics, 213(1):281– 288, November 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619300703.

Poskitt:2016:VAM

[Pos16] D. S. Poskitt. Vector autoregressive moving average identification for macroeconomic modeling: a new methodology. Journal of Econometrics, 192(2):468–484, June 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300148.

Ploberger:2012:OEU

[PP12] Werner Ploberger and Peter C. B. Phillips. Optimal estimation under nonstandard conditions. Journal of Econometrics, 169(2):258–265, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612000358.

Paolella:2015:CCM

[PP15] Marc S. Paolella and Paweł Polak. COMFORT: a common market factor non-Gaussian returns model. Journal of Econometrics, 187(2):593–605, August 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000664.

Potscher:2018:CSA

[PP18] Benedikt M. Pötscher and David Preinerstorfer. Controlling the size of autocorrelation robust tests. Journal of Econometrics, 207(2):406–431, December 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618301568.

Pesaran:2013:OFP

[PPP13] M. Hashem Pesaran, Andreas Pick, and Mikhail Pranovich. Optimal forecasts in the presence of structural breaks. Journal of Econometrics, 177(2):134–152, December 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613000687.

Pesaran:2011:VSE

[PPT11] M. Hashem Pesaran, Andreas Pick, and Allan Timmermann. Variable selection, estimation and inference for multi-period forecasting problems. Journal of Econometrics, 164(1):173–187, September 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611000467.

Paolella:2019:RSD

[PPW19] Marc S. Paolella, Paweł Polak, and Patrick S. Walker. Regime switching dynamic correlations for asymmetric and fat-tailed conditional returns. Journal of Econometrics, 213(2):493–515, December 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407619301563.

Park:2012:FRC

[PQ12] Joon Y. Park and Junhui Qian. Functional regression of continuous state distributions. Journal of Econometrics, 167(2):397– 412, April 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002053.

Paolella:2013:LDH

[PRSV13] Marc Paolella, Eric Renault, Gennady Samorodnitsky, and David Veredas. Latest developments on heavy-tailed distributions. Journal of Econometrics, 172(2):183–185, February 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612001911.

Penaranda:2012:STR

[PS12] Francisco Peñaranda and Enrique Sentana. Spanning tests in return and stochastic discount factor mean-variance frontiers: a unifying approach. Journal of Econometrics, 170(2):303–324, October 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001194.

Pelagatti:2013:RTS

[PS13] Matteo M. Pelagatti and Pranab K. Sen. Rank tests for short memory stationarity. Journal of Econometrics, 172(1):90–105, January 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612002151.

Parente:2017:TAC

[PS17] Paulo M. D. C. Parente and Richard J. Smith. Tests of additional conditional moment restrictions. Journal of Econometrics, 200(1):1–16, September 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617300349.

Patra:2018:CBP

[PSS18] Rohit Kumar Patra, Emilio Seijo, and Bodhisattva Sen. A consistent bootstrap procedure for the maximum score estimator. Journal of Econometrics, 205(2):488–507, August 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761830068X.

Pettenuzzo:2022:CPS

[PST22] Davide Pettenuzzo, Yong Song, and Allan Timmermann. Corrigendum to “Predictability of stock returns and asset allocation under structural breaks” [j. econometrics 164 (2011) 60–78]. Journal of Econometrics, 227(2):513–517, April 2022. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407621000476.

Palm:2011:CSD

[PSU11] Franz C. Palm, Stephan Smeekes, and Jean-Pierre Urbain. Cross-sectional dependence robust block bootstrap panel unit root tests. Journal of Econometrics, 163(1):85–104, July 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610002149.

Park:2010:SCR

[PSW10] Joon Y. Park, Kwanho Shin, and Yoon-Jae Whang. A semiparametric cointegrating regression: Investigating the effects of age distributions on consumption and saving. Journal of Econometrics, 157(1):165–178, July 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002851.

Pesaran:2013:PUR

[PSY13] M. Hashem Pesaran, L. Vanessa Smith, and Takashi Yamagata. Panel unit root tests in the presence of a multifactor error structure. Journal of Econometrics, 175(2):94–115, August 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000353.

Pawasutipaisit:2011:WAF

[PT11a] Anan Pawasutipaisit and Robert M. Townsend. Wealth accumulation and factors accounting for success. Journal of Econometrics, 161(1):56–81, March 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610001879.

Pesaran:2011:LPC

[PT11b] M. Hashem Pesaran and Elisa Tosetti. Large panels with common factors and spatial correlation. Journal of Econometrics, 161(2):182–202, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610002459.

Pettenuzzo:2011:PSR

[PT11c] Davide Pettenuzzo and Allan Timmermann. Predictability of stock returns and asset allocation under structural breaks. Journal of Econometrics, 164(1):60–78, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000479. See corrigendum [PST22].

Pettenuzzo:2016:MAM

[PTV16] Davide Pettenuzzo, Allan Timmermann, and Rossen Valkanov. A MIDAS approach to modeling first and second moment dynamics. Journal of Econometrics, 193(2):315–334, August 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300665.

Palm:2011:FSP

[PU11] Franz C. Palm and Jean-Pierre Urbain. Factor structures for panel and multivariate time series data. Journal of Econometrics, 163(1):1–3, July 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610002058.

Pedroni:2015:NRT

[PVWW15] Peter L. Pedroni, Timothy J. Vogelsang, Martin Wagner, and Joakim Westerlund. Nonparametric rank tests for nonstationary panels. Journal of Econometrics, 185(2):378–391, April 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407614002802.

Park:2012:RWC

[PW12] Joon Y. Park and Yoon-Jae Whang. Random walk or chaos: a formal test on the Lyapunov exponent. Journal of Econometrics, 169(1):61–74, July 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612000139.

Papay:2011:ERD

[PWM11] John P. Papay, John B. Willett, and Richard J. Murnane. Extending the regression-discontinuity approach to multiple assignment variables. Journal of Econometrics, 161(2):203–207, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610002538.

Porter:2015:RDD

[PY15] Jack Porter and Ping Yu. Regression discontinuity designs with unknown discontinuity points: Testing and estimation. Journal of Econometrics, 189(1):132–147, November 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001712.

Patton:2019:DSM

[PZC19] Andrew J. Patton, Johanna F. Ziegel, and Rui Chen. Dynamic semiparametric models for expected shortfall (and Valueat-Risk). Journal of Econometrics, 211(2):388–413, August 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761930048X.

Qu:2015:ESA

[QfL15] Xi Qu and Lung fei Lee. Estimating a spatial autoregressive model with an endogenous spatial weight matrix. Journal of Econometrics, 184(2):209–232, February 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001870.

Qu:2017:QES

[QfLY17] Xi Qu, Lung fei Lee, and Jihai Yu. QML estimation of spatial dynamic panel data models with endogenous time varying spatial weights matrices. Journal of Econometrics, 197(2):173– 201, April 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616302147.

Qian:2016:CEM

[Qia16] Hang Qian. A computationally efficient method for vector autoregression with mixed frequency data. Journal of Econometrics, 193(2):433–437, August 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616300781.

Qian:2016:SEC

[QS16] Junhui Qian and Liangjun Su. Shrinkage estimation of common breaks in panel data models via adaptive group fused lasso. Journal of Econometrics, 191(1):86–109, March 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002377.

Qian:2012:ESP

[QW12] Junhui Qian and Le Wang. Estimating semiparametric panel data models by marginal integration. Journal of Econometrics, 167(2):483–493, April 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611002119.

Qu:2015:NEI

[QY15] Zhongjun Qu and Jungmo Yoon. Nonparametric estimation and inference on conditional quantile processes. Journal of Econometrics, 185(1):1–19, March 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614002462.

Rao:2016:SAC

[RH16] D. S. Prasada Rao and Gholamreza Hajargasht. Stochastic approach to computation of purchasing power parities in the international comparison program (ICP). Journal of Econometrics, 191(2):414–425, April 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002882.

Richard:2019:RBT

[Ric19] Patrick Richard. Residual bootstrap tests in linear models with many regressors. Journal of Econometrics, 208(2):367–394, February 2019. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618301921.

Racine:2017:NCQ

[RL17] Jeffrey S. Racine and Kevin Li. Nonparametric conditional quantile estimation: a locally weighted quantile kernel approach. Journal of Econometrics, 201(1):72–94, November 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301185.

Robinson:2010:EES

[Rob10] P. M. Robinson. Efficient estimation of the semiparametric spatial autoregressive model. Journal of Econometrics, 157(1):6–17, July 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440760900284X.

Robinson:2011:ATN

[Rob11] P. M. Robinson. Asymptotic theory for nonparametric regression with spatial data. Journal of Econometrics, 165(1):5–19, November 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000947.

Robinson:2012:NTR

[Rob12] Peter M. Robinson. Nonparametric trending regression with cross-sectional dependence. Journal of Econometrics, 169(1): 4–14, July 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000061.

Robinson:2014:EML

[Rob14] Peter M. Robinson. The estimation of misspecified long memory models. Journal of Econometrics, 178(Part 2):225–230, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001875.

Rosen:2012:SIQ

[Ros12] Adam M. Rosen. Set identification via quantile restrictions in short panels. Journal of Econometrics, 166(1):127–137, January 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001242.

Rothe:2010:NED

[Rot10] Christoph Rothe. Nonparametric estimation of distributional policy effects. Journal of Econometrics, 155(1):56–70, March 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609001924.

Robinson:2015:RML

[RR15] Peter M. Robinson and Francesca Rossi. Refinements in maximum likelihood inference on spatial autocorrelation in panel data. Journal of Econometrics, 189(2):447–456, December 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001141.

Rossi:2011:UMF

[RS11] Barbara Rossi and Tatevik Sekhposyan. Understanding models’ forecasting performance. Journal of Econometrics, 164 (1):158–172, September 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407611000480.

Rossi:2013:CPD

[RS13] Barbara Rossi and Tatevik Sekhposyan. Conditional predictive density evaluation in the presence of instabilities. Journal of Econometrics, 177(2):199–212, December 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613000857.

Robertson:2015:IEP

[RS15] Donald Robertson and Vasilis Sarafidis. IV estimation of panels with factor residuals. Journal of Econometrics, 185(2):526– 541, April 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304407614002899. See comment

[Ahn15].

Ryu:2017:MEE

[RS17] Hang K. Ryu and Daniel J. Slottje. Maximum entropy estimation of income distributions from Basmann’s weighted geometric mean measure. Journal of Econometrics, 199(2):221–231, August 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761730074X.

Rossi:2019:ATC

[RS19] Barbara Rossi and Tatevik Sekhposyan. Alternative tests for correct specification of conditional predictive densities. Journal of Econometrics, 208(2):638–657, February 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618302197.

Robinson:2012:SIR

[RT12] Peter M. Robinson and Supachoke Thawornkaiwong. Statistical inference on regression with spatial dependence. Journal of Econometrics, 167(2):521–542, April 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611002144.

Rust:2010:CSV

[Rus10] John Rust. Comments on: “Structural vs. atheoretic approaches to econometrics” by Michael Keane. Journal of Econometrics, 156(1):21–24, May 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440760900195X.

See [Kea10].

Reynaert:2014:IPR

[RV14] Mathias Reynaert and Frank Verboven. Improving the performance of random coefficients demand models: The role of optimal instruments. Journal of Econometrics, 179(1):83–98, March 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613002649.

Robinson:2015:EIF

[RV15] Peter M. Robinson and Carlos Velasco. Efficient inference on fractionally integrated panel data models with fixed effects. Journal of Econometrics, 185(2):435–452, April 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614002930.

Robinson:2018:ITP

[RV18] Peter M. Robinson and Carlos Velasco. Inference on trending panel data. Journal of Econometrics, 206(2):282–304, October 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618300940.

Renault:2014:DMH

[RvdHW14] Eric Renault, Thijs van der Heijden, and Bas J. M. Werker. The dynamic mixed hitting-time model for multiple transaction prices and times. Journal of Econometrics, 180(2):233–250, June 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000396.

Renault:2011:CER

[RW11] Eric Renault and Bas J. M. Werker. Causality effects in return volatility measures with random times. Journal of Econometrics, 160(1):272–279, January 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000783.

Romano:2017:RWL

[RW17] Joseph P. Romano and Michael Wolf. Resurrecting weighted least squares. Journal of Econometrics, 197(1):1–19, March 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761630197X.

Russo:2011:DIS

[RY11] J. E. Russo and Kevyn Yong. The distortion of information to support an emerging evaluation of risk. Journal of Econometrics, 162(1):132–139, May 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610001478.

Saijo:2013:EDM

[Sai13] Hikaru Saijo. Estimating DSGE models using seasonally adjusted and unadjusted data. Journal of Econometrics, 173(1): 22–35, March 2013. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612002461.

Salimans:2012:VSF

[Sal12] Tim Salimans. Variable selection and functional form uncertainty in cross-country growth regressions. Journal of Econometrics, 171(2):267–280, December 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612001546.

Santos:2011:IVM

[San11] Andres Santos. Instrumental variable methods for recovering continuous linear functionals. Journal of Econometrics, 161(2):129– 146, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610002253.

Sasaki:2015:HSD

[Sas15] Yuya Sasaki. Heterogeneity and selection in dynamic panel data. Journal of Econometrics, 188(1):236–249, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001542.

So:2015:SIC

[SC15a] Mike K. P. So and Ray S. W. Chung. Statistical inference for conditional quantiles in nonlinear time series models. Journal of Econometrics, 189(2):457–472, December 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001153.

Su:2015:QLE

[SC15b] Fei Su and Kung-Sik Chan. Quasi-likelihood estimation of a threshold diffusion process. Journal of Econometrics, 189(2):473– 484, December 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615001165.

Choi:2016:ITM

[sC16] Hwan sik Choi. Information theory for maximum likelihood estimation of diffusion models. Journal of Econometrics, 191(1):110– 128, March 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615002511.

Schumann:2011:SNT

[Sch11] Keith D. Schumann. Semi-nonparametric test of second degree stochastic dominance with respect to a function. Journal of Econometrics, 162(1):71–78, May 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002620.

Schennach:2019:CI

[Sch19] Susanne M. Schennach. Convolution without independence. Journal of Econometrics, 211(1):308–318, July 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618302549.

Choi:2014:AAL

[sCJP14] Hwan sik Choi, Minsoo Jeong, and Joon Y. Park. An asymptotic analysis of likelihood-based diffusion model selection using high frequency data. Journal of Econometrics, 178(Part 3):539– 557, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613002005.

Seo:2018:TSM

[Seo18] Juwon Seo. Tests of stochastic monotonicity with improved power. Journal of Econometrics, 207(1):53–70, November 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301179.

Shi:2017:SDP

[SfL17] Wei Shi and Lung fei Lee. Spatial dynamic panel data models with interactive fixed effects. Journal of Econometrics, 197(2): 323–347, April 2017. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616302317.

Salish:2019:MBN

[SG19] Nazarii Salish and Alexander Gleim. A moment-based notion of time dependence for functional time series. Journal of Econometrics, 212(2):377–392, October 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300910.

Serra:2011:RBP

[SGF11] Teresa Serra, Barry K. Goodwin, and Allen M. Featherstone. Risk behavior in the presence of government programs. Journal of Econometrics, 162(1):18–24, May 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002589.

Shiu:2013:IEN

[SH13] Ji-Liang Shiu and Yingyao Hu. Identification and estimation of nonlinear dynamic panel data models with unobserved covariates. Journal of Econometrics, 175(2):116–131, August 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000559.

Su:2016:SIV

[SH16] Liangjun Su and Tadao Hoshino. Sieve instrumental variable quantile regression estimation of functional coefficient models. Journal of Econometrics, 191(1):231–254, March 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002572.

Shao:2011:BAS

[Sha11] Xiaofeng Shao. A bootstrap-assisted spectral test of white noise under unknown dependence. Journal of Econometrics, 162(2): 213–224, June 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611000030.

Shaliastovich:2015:LCO

[Sha15] Ivan Shaliastovich. Learning, confidence, and option prices. Journal of Econometrics, 187(1):18–42, July 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615000317.

Ho:2017:BMR

[sHDW17] Chi san Ho, Paul Damien, and Stephen Walker. Bayesian mode regression using mixtures of triangular densities. Journal of Econometrics, 197(2):273–283, April 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616302184.

Sun:2018:TAM

[SHHW18] Yuying Sun, Ai Han, Yongmiao Hong, and Shouyang Wang. Threshold autoregressive models for interval-valued time series data. Journal of Econometrics, 206(2):414–446, October 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301039.

Shimotsu:2012:ELW

[Shi12] Katsumi Shimotsu. Exact local Whittle estimation of fractionally cointegrated systems. Journal of Econometrics, 169(2):266– 278, August 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000383.

Shi:2015:MST

[Shi15] Xiaoxia Shi. Model selection tests for moment inequality models. Journal of Econometrics, 187(1):1–17, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000135.

Shi:2016:EEH

[Shi16] Zhentao Shi. Econometric estimation with high-dimensional moment equalities. Journal of Econometrics, 195(1):104–119, November 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616301373.

Sun:2011:MCI

[SHL11] Yiguo Sun, Cheng Hsiao, and Qi Li. Measuring correlations of integrated but not cointegrated variables: a semiparametric approach. Journal of Econometrics, 164(2):252–267, October 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001138.

Shonkwiler:2016:VTN

[Sho16] J. S. Shonkwiler. Variance of the truncated negative binomial distribution. Journal of Econometrics, 195(2):209–210, December 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616301610.

Sun:2019:CEF

[SI19] Yutec Sun and Masakazu Ishihara. A computationally efficient fixed point approach to dynamic structural demand estimation. Journal of Econometrics, 208(2):563–584, February 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302161.

Sianesi:2017:ERB

[Sia17] Barbara Sianesi. Evidence of randomisation bias in a largescale social experiment: The case of ERA. Journal of Econometrics, 198(1):41–64, May 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761730012X.

Sizova:2011:IVF

[Siz11] Natalia Sizova. Integrated variance forecasting: Model based vs. reduced form. Journal of Econometrics, 162(2):294–311, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000315.

Su:2010:PQM

[SJ10] Liangjun Su and Sainan Jin. Profile quasi-maximum likelihood estimation of partially linear spatial autoregressive models. Journal of Econometrics, 157(1):18–33, July 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002863.

Su:2012:SEP

[SJ12] Liangjun Su and Sainan Jin. Sieve estimation of panel data models with cross section dependence. Journal of Econometrics, 169 (1):34–47, July 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612000073.

Su:2018:ILG

[SJ18] Liangjun Su and Gaosheng Ju. Identifying latent grouped patterns in panel data models with interactive fixed effects. Journal of Econometrics, 206(2):554–573, October 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301106.

Su:2015:STP

[SJZ15] Liangjun Su, Sainan Jin, and Yonghui Zhang. Specification test for panel data models with interactive fixed effects. Journal of Econometrics, 186(1):222–244, May 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001729.

Sun:2012:SPG

[SK12] Yixiao Sun and Min Seong Kim. Simple and powerful GMM over-identification tests with accurate size. Journal of Econometrics, 166(2):267–281, February 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761100220X.

Scharth:2016:PEI

[SK16] Marcel Scharth and Robert Kohn. Particle efficient importance sampling. Journal of Econometrics, 190(1):133–147, January 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615002201.

Srisuma:2012:SEM

[SL12] Sorawoot Srisuma and Oliver Linton. Semiparametric estimation of Markov decision processes with continuous state space. Journal of Econometrics, 166(2):320–341, February 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761100234X.

Su:2013:NDP

[SL13] Liangjun Su and Xun Lu. Nonparametric dynamic panel data models: Kernel estimation and specification testing. Journal of Econometrics, 176(2):112–133, October 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613001140.

Sibbertsen:2018:MTA

[SLB18] Philipp Sibbertsen, Christian Leschinski, and Marie Busch. A multivariate test against spurious long memory. Journal of Econometrics, 203(1):33–49, March 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407617302324.

Slottje:2017:CME

[Slo17] Dan Slottje. The creative mind in econometrics: Studies in celebration of Robert Basmann’s 90th year on causation, identification and structural equation estimation. Journal of Econometrics, 199(2):93–95, August 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300647.

Sun:2018:EIF

[SM18] Yiguo Sun and Emir Malikov. Estimation and inference in functional-coefficient spatial autoregressive panel data models with fixed effects. Journal of Econometrics, 203(2):359–378, April 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407618300010.

Song:2010:TSC

[Son10] Kyungchul Song. Testing semiparametric conditional moment restrictions using conditional martingale transforms. Journal of Econometrics, 154(1):74–84, January 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609001511.

Song:2011:MAT

[Son11] Zhaogang Song. A martingale approach for testing diffusion models based on infinitesimal operator. Journal of Econometrics, 162(2):189–212, June 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610002472.

Song:2014:SMS

[Son14] Kyungchul Song. Semiparametric models with single-index nuisance parameters. Journal of Econometrics, 178(Part 3):471– 483, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613001486.

Song:2015:SEM

[Son15] Suyong Song. Semiparametric estimation of models with conditional moment restrictions in the presence of nonclassical measurement errors. Journal of Econometrics, 185(1):95–109, March 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614002450.

Spanos:2010:ATC

[Spa10] Aris Spanos. Akaike-type criteria and the reliability of inference: Model selection versus statistical model specification. Journal of Econometrics, 158(2):204–220, October 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761000014X.

Seo:2016:DPT

[SS16] Myung Hwan Seo and Yongcheol Shin. Dynamic panels with threshold effect and endogeneity. Journal of Econometrics, 195(2):169–186, December 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616301506.

Severini:2012:EBE

[ST12] Thomas A. Severini and Gautam Tripathi. Efficiency bounds for estimating linear functionals of nonparametric regression models with endogenous regressors. Journal of Econometrics, 170(2):491–498, October 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407612001303.

Sojli:2015:DGF

[ST15] Elvira Sojli and Wing Wah Tham. Divided governments and futures prices. Journal of Econometrics, 187(2):622–633, August 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000688.

Stoye:2012:MRT

[Sto12] Jörg Stoye. Minimax regret treatment choice with covariates or with limited validity of experiments. Journal of Econometrics, 166(1):138–156, January 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611001254.

Smith:2019:VSP

[STZ19] Simon C. Smith, Allan Timmermann, and Yinchu Zhu. Variable selection in panel models with breaks. Journal of Econometrics, 212(1):323–344, September 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300880.

Su:2012:SGE

[Su12] Liangjun Su. Semiparametric GMM estimation of spatial autoregressive models. Journal of Econometrics, 167(2):543–560, April 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002156.

Sun:2011:RTI

[Sun11] Yixiao Sun. Robust trend inference with series variance estimator and testing-optimal smoothing parameter. Journal of Econometrics, 164(2):345–366, October 1, 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611001308.

Sun:2014:LFI

[Sun14] Yixiao Sun. Let’s fix it: Fixed-b asymptotics versus small-b asymptotics in heteroskedasticity and autocorrelation robust inference. Journal of Econometrics, 178(Part 3):659–677, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613002054.

Sun:2016:FCS

[Sun16] Yiguo Sun. Functional-coefficient spatial autoregressive models with nonparametric spatial weights. Journal of Econometrics, 195(1):134–153, November 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761630149X.

Su:2019:NSM

[SUZ19] Liangjun Su, Takuya Ura, and Yichong Zhang. Non-separable models with high-dimensional data. Journal of Econometrics, 212(2):646–677, October 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407619301447.

Simar:2012:PCD

[SV12] Léopold Simar and Anne Vanhems. Probabilistic characterization of directional distances and their robust versions. Journal of Econometrics, 166(2):342–354, February 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002338.

Simar:2016:UHE

[SVV16] Léopold Simar, Anne Vanhems, and Ingrid Van Keilegom. Unobserved heterogeneity and endogeneity in nonparametric frontier estimation. Journal of Econometrics, 190(2):360–373, February 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001840.

Semykina:2010:EPD

[SW10] Anastasia Semykina and Jeffrey M. Wooldridge. Estimating panel data models in the presence of endogeneity and selection. Journal of Econometrics, 157(2):375–380, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000825.

Stock:2014:ETP

[SW14a] James H. Stock and Mark W. Watson. Estimating turning points using large data sets. Journal of Econometrics, 178(Part 2):368– 381, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S030440761300198X.

Su:2014:TCI

[SW14b] Liangjun Su and Halbert White. Testing conditional independence via empirical likelihood. Journal of Econometrics, 182(1):27–44, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614000657.

Sanderson:2016:WIT

[SW16] Eleanor Sanderson and Frank Windmeijer. A weak instrument F -test in linear IV models with multiple endogenous variables. Journal of Econometrics, 190(2):212–221, February 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001736.

Su:2017:TVF

[SW17] Liangjun Su and Xia Wang. On time-varying factor models: Estimation and testing. Journal of Econometrics, 198(1):84–101, May 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617300143.

Schennach:2012:LIL

[SWC12] Susanne Schennach, Halbert White, and Karim Chalak. Local indirect least squares and average marginal effects in nonseparable structural systems. Journal of Econometrics, 166(2):282– 302, February 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611002223.

Shively:2011:NFE

[SWD11] Thomas S. Shively, Stephen G. Walker, and Paul Damien. Nonparametric function estimation subject to monotonicity, convexity and other shape constraints. Journal of Econometrics, 161(2):166–181, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610002435.

Swensen:2011:BAT

[Swe11] Anders Rygh Swensen. A bootstrap algorithm for testing cointegration rank in VAR models in the presence of stationary variables. Journal of Econometrics, 165(2):152–162, December 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001436.

Song:2016:TTO

[SX16] Zhaogang Song and Dacheng Xiu. A tale of two option markets: Pricing kernels and volatility risk. Journal of Econometrics, 190(1):176–196, January 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615002328.

Sasaki:2017:USD

[SX17a] Yuya Sasaki and Yi Xin. Unequal spacing in dynamic panel data: Identification and estimation. Journal of Econometrics, 196(2):320–330, February 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616301932.

Shephard:2017:EAM

[SX17b] Neil Shephard and Dacheng Xiu. Econometric analysis of multivariate realised QML: Estimation of the covariation of equity prices under asynchronous trading. Journal of Econometrics, 201(1):19–42, November 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617301434.

Su:2015:QED

[SY15] Liangjun Su and Zhenlin Yang. QML estimation of dynamic panel data models with spatial errors. Journal of Econometrics, 185(1):230–258, March 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614002668.

Sun:2019:IDD

[SY19] Yu Sun and Karen X. Yan. Inference on Difference-in-Differences average treatment effects: a fixed-b approach. Journal of Econometrics, 211(2):560–588, August 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300545.

Shintani:2012:SRT

[SYN12] Mototsugu Shintani, Tomoyoshi Yabu, and Daisuke Nagakura. Spurious regressions in technical trading. Journal of Econometrics, 169(2):301–309, August 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000292.

Todorov:2010:JBN

[TB10] Viktor Todorov and Tim Bollerslev. Jumps and betas: a new framework for disentangling and estimating systematic risks. Journal of Econometrics, 157(2):220–235, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609003017.

Taamouti:2014:NEI

[TBE14] Abderrahim Taamouti, Taoufik Bouezmarni, and Anouar El Ghouch. Nonparametric estimation and inference for conditional density based Granger causality measures. Journal of Econometrics, 180(2):251–264, June 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000402.

Tetenov:2012:STC

[Tet12] Aleksey Tetenov. Statistical treatment choice based on asymmetric minimax regret criteria. Journal of Econometrics, 166(1):157– 165, January 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407611001266.

Tjostheim:2013:LGC

[TH13] Dag Tjøstheim and Karl Ove Hufthammer. Local Gaussian correlation: a new measure of dependence. Journal of Econometrics, 172(1):33–48, January 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001741.

Thornton:2014:ADR

[Tho14] Michael A. Thornton. The aggregation of dynamic relationships caused by incomplete information. Journal of Econometrics, 178(Part 2):342–351, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001966.

Todorov:2011:EAJ

[Tod11] Viktor Todorov. Econometric analysis of jump-driven stochastic volatility models. Journal of Econometrics, 160(1):12–21, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000515.

Tong:2015:TMT

[Ton15] Howell Tong. Threshold models in time series analysis — some reflections. Journal of Econometrics, 189(2):485–491, December 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001177.

Torgovitsky:2017:MDI

[Tor17] Alexander Torgovitsky. Minimum distance from independence estimation of nonseparable instrumental variables models. Journal of Econometrics, 199(1):35–48, July 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300441.

Tao:2019:RCC

[TPY19] Yubo Tao, Peter C. B. Phillips, and Jun Yu. Random coefficient continuous systems: Testing for extreme sample path behavior. Journal of Econometrics, 209(2):208–237, April 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300053.

Trapani:2013:BPF

[Tra13] Lorenzo Trapani. On bootstrapping panel factor series. Journal of Econometrics, 172(1):127–141, January 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612002175.

Trapani:2016:TFM

[Tra16] Lorenzo Trapani. Testing for (in)finite moments. Journal of Econometrics, 191(1):57–68, March 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407615002596.

Tripathi:2011:GMM

[Tri11] Gautam Tripathi. Generalized method of moments (GMM) based inference with stratified samples when the aggregate shares are known. Journal of Econometrics, 165(2):258–265, December 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761100162X.

Tsionas:2012:MLE

[Tsi12] Efthymios G. Tsionas. Maximum likelihood estimation of stochastic frontier models by the Fourier transform. Journal of Econometrics, 170(1):234–248, September 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612000796.

Todorov:2010:ASF

[TT10] Viktor Todorov and George Tauchen. Activity signature functions for high-frequency data analysis. Journal of Econometrics, 154(2):125–138, February 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609001651.

Todorov:2011:RLT

[TTG11] Viktor Todorov, George Tauchen, and Iaryna Grynkiv. Realized Laplace transforms for estimation of jump diffusive volatility models. Journal of Econometrics, 164(2):367–381, October 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001291.

Todorov:2014:VAS

[TTG14] Viktor Todorov, George Tauchen, and Iaryna Grynkiv. Volatility activity: Specification and estimation. Journal of Econometrics, 178(Part 1):180–193, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0304407613001656.

Trapani:2010:MVM

[TU10] Lorenzo Trapani and Giovanni Urga. Micro versus macro cointegration in heterogeneous panels. Journal of Econometrics, 155(1):1–18, March 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609001626.

Timmermann:2013:DEM

[TvD13] Allan Timmermann and Herman K. van Dijk. Dynamic econometric modeling and forecasting in the presence of instability. Journal of Econometrics, 177(2):131–133, December 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000675.

Tu:2017:FCN

[TY17] Yundong Tu and Yanping Yi. Forecasting cointegrated nonstationary time series with time-varying variance. Journal of Econometrics, 196(1):83–98, January 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407616301798.

Tang:2018:ETL

[TYZ18] Niansheng Tang, Xiaodong Yan, and Puying Zhao. Exponentially tilted likelihood inference on growing dimensional unconditional moment models. Journal of Econometrics, 202(1):57– 74, January 2018. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407617301719.

Tauchen:2011:RJF

[TZ11] George Tauchen and Hao Zhou. Realized jumps on financial markets and predicting credit spreads. Journal of Econometrics, 160(1):102–118, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407610000655.

Villasenor:1989:ELC

[VA89] José A. Villaseñor and Barry C. Arnold. Elliptical Lorenz curves. Journal of Econometrics, 40(2):327–338, February 1989. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304407689900894. See corrigendum [Kra13].

vonEije:2014:DIC

[vEGM14] Henk von Eije, Abhinav Goyal, and Cal B. Muckley. Does the information content of payout initiations and omissions influence firm risks? Journal of Econometrics, 183(2):222–229, December 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001158.

vanHasselt:2011:BIS

[vH11] Martijn van Hasselt. Bayesian inference in a sample selection model. Journal of Econometrics, 165(2):221–232, December 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611001618.

Villani:2012:GSF

[VKN12] Mattias Villani, Robert Kohn, and David J. Nott. Generalized smooth finite mixtures. Journal of Econometrics, 171(2):121– 133, December 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001595.

Vogelsang:2012:HAS

[Vog12] Timothy J. Vogelsang. Heteroskedasticity, autocorrelation, and spatial correlation robust inference in linear panel models with fixed-effects. Journal of Econometrics, 166(2):303–319, February 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002326.

Vikstrom:2018:BTE

[VRW18] Johan Vikström, Geert Ridder, and Martin Weidner. Bounds on treatment effects on transitions. Journal of Econometrics, 205(2):448–469, August 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618300605.

Vogelsang:2014:IMO

[VW14a] Timothy J. Vogelsang and Martin Wagner. Integrated modified OLS estimation and fixed-b inference for cointegrating regressions. Journal of Econometrics, 178(2):741–760, February 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613002303.

Voss:2014:STM

[VW14b] Sebastian Voß and Rafael Weißbach. A score-test on measurement errors in rating transition times. Journal of Econometrics, 180(1):16–29, May 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614000165.

Wang:2013:FLM

[WBH13] Cindy Shin-Huei Wang, Luc Bauwens, and Cheng Hsiao. Forecasting a long memory process subject to structural breaks. Journal of Econometrics, 177(2):171–184, December 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407613000833.

West:2012:EAP

[Wes12] Kenneth D. West. Econometric analysis of present value models when the discount factor is near one. Journal of Econometrics, 171(1):86–97, November 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761200173X.

Westerlund:2015:ERD

[Wes15a] Joakim Westerlund. The effect of recursive detrending on panel unit root tests. Journal of Econometrics, 185(2):453–467, April 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001560.

Westerlund:2015:PP

[Wes15b] Joakim Westerlund. The power of PANIC. Journal of Econometrics, 185(2):495–509, April 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001572.

Wang:2010:EFE

[WH10] Hung-Jen Wang and Chia-Wen Ho. Estimating fixed-effect panel stochastic frontier models by model transformation. Journal of Econometrics, 157(2):286–296, August 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610000047.

Wang:2011:MME

[WH11] Liqun Wang and Cheng Hsiao. Method of moments estimation and identifiability of semiparametric nonlinear errors-in-variables models. Journal of Econometrics, 165(1):30–44, November 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611000960.

Woutersen:2019:IPS

[WH19] Tiemen Woutersen and Jerry A. Hausman. Increasing the power of specification tests. Journal of Econometrics, 211(1):166–175, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302458.

Wilcox:2011:SMR

[Wil11] Nathaniel T. Wilcox. ‘stochastically more risk averse:’ a contextual theory of stochastic discrete choice under risk. Journal of Econometrics, 162(1):89–104, May 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002656.

Wang:2013:PML

[WIW13] Honglin Wang, Emma M. Iglesias, and Jeffrey M. Wooldridge. Partial maximum likelihood estimation of spatial probit models. Journal of Econometrics, 172(1):77–89, January 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612001893.

Wang:2016:BII

[WK16] Wenjie Wang and Maximilien Kaffo. Bootstrap inference for instrumental variable models with many weak instruments. Journal of Econometrics, 192(1):231–268, May 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615003152.

White:2015:VVM

[WKM15] Halbert White, Tae-Hwan Kim, and Simone Manganelli. VAR for VaR: Measuring tail dependence using multivariate regression quantiles. Journal of Econometrics, 187(1):169–188, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000287.

Westerlund:2012:TUR

[WL12] Joakim Westerlund and Rolf Larsson. Testing for a unit root in a random coefficient panel data model. Journal of Econometrics, 167(1):254–273, March 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611002727.

Wu:2014:MBT

[WL14] Jianhong Wu and Guodong Li. Moment-based tests for individual and time effects in panel data models. Journal of Econometrics, 178(Part 3):569–581, January 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001796.

Westerlund:2015:NTU

[WL15] Joakim Westerlund and Rolf Larsson. New tools for understanding the local asymptotic power of panel unit root tests. Journal of Econometrics, 188(1):59–93, September 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001268.

Wang:2019:FMM

[WLC19] Dong Wang, Xialu Liu, and Rong Chen. Factor models for matrix-valued high-dimensional time series. Journal of Econometrics, 208(1):231–248, January 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301787.

Wolter:2016:KEH

[Wol16] James Lewis Wolter. Kernel estimation of hazard functions when observations have dependent and common covariates. Journal of Econometrics, 193(1):1–16, July 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761630001X.

Wooldridge:2014:QML

[Woo14] Jeffrey M. Wooldridge. Quasi-maximum likelihood estimation and testing for nonlinear models with endogenous explanatory variables. Journal of Econometrics, 182(1):226–234, September 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000797.

Wooldridge:2019:CRE

[Woo19] Jeffrey M. Wooldridge. Correlated random effects models with unbalanced panels. Journal of Econometrics, 211(1):137–150, July 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618302392.

White:2014:GCE

[WP14] Halbert White and Davide Pettenuzzo. Granger causality, exogeneity, cointegration, and economic policy analysis. Journal of Econometrics, 178(Part 2):316–330, January 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407613001942.

Wang:2011:BEM

[WPY11] Xiaohu Wang, Peter C. B. Phillips, and Jun Yu. Bias in estimating multivariate and univariate diffusions. Journal of Econometrics, 161(2):228–245, April 1, 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610002484.

Wang:2018:BIB

[WT18] Wenjie Wang and Firmin Doko Tchatoka. On bootstrap inconsistency and Bonferroni-based size-correction for the subset Anderson–Rubin test under conditional homoskedasticity. Journal of Econometrics, 207(1):188–211, November 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301337.

Wu:2010:ESE

[Wu10] Ximing Wu. Exponential series estimator of multivariate densities. Journal of Econometrics, 156(2):354–366, June 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002942.

Wu:2011:VDJ

[Wu11] Liuren Wu. Variance dynamics: Joint evidence from options and high-frequency returns. Journal of Econometrics, 160(1):280– 287, January 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610000795.

Westerlund:2015:CSA

[WU15] Joakim Westerlund and Jean-Pierre Urbain. Cross-sectional averages versus principal components. Journal of Econometrics, 185(2):372–377, April 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614002784.

Weissbach:2010:LRT

[WW10] Rafael Weißbach and Ronja Walter. A likelihood ratio test for stationarity of rating transitions. Journal of Econometrics, 155(2):188–194, April 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002693.

Wachter:2015:WCE

[WW15] Jessica A. Wachter and Missaka Warusawitharana. What is the chance that the equity premium varies over time? Evidence from regressions on the dividend–price ratio. Journal of Econometrics, 186(1):74–93, May 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001407.

Wang:2010:GNI

[WWYL10] Shaoping Wang, Peng Wang, Jisheng Yang, and Zinai Li. A generalized nonlinear IV unit root test for panel data with crosssectional dependence. Journal of Econometrics, 157(1):101–109, July 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407609002875.

Waggoner:2016:SMH

[WWZ16] Daniel F. Waggoner, Hongwei Wu, and Tao Zha. Striated Metropolis–Hastings sampler for high-dimensional models. Journal of Econometrics, 192(2):406–420, June 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300100.

Wang:2018:MCN

[WWZ18] Qiying Wang, Dongsheng Wu, and Ke Zhu. Model checks for nonlinear cointegrating regression. Journal of Econometrics, 207(2):261–284, December 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618301428.

Wan:2014:SIB

[WX14] Yuanyuan Wan and Haiqing Xu. Semiparametric identification of binary decision games of incomplete information with correlated private signals. Journal of Econometrics, 182(2):235–246, October 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614001043.

Wan:2015:ISB

[WX15] Yuanyuan Wan and Haiqing Xu. Inference in semiparametric binary response models with interval data. Journal of Econometrics, 184(2):347–360, February 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614001948.

White:2014:CDG

[WXC14] Halbert White, Haiqing Xu, and Karim Chalak. Causal discourse in a game of incomplete information. Journal of Econometrics, 182(1):45–58, September 2014. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614000669.

Wei:2012:RFC

[WY12] Xiaoqiao Wei and Yuhong Yang. Robust forecast combinations. Journal of Econometrics, 166(2):224–236, February 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002168.

Wang:2016:DAE

[WY16] Xiaohu Wang and Jun Yu. Double asymptotics for explosive continuous time models. Journal of Econometrics, 193(1):35–53, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407616300173.

Waggoner:2012:CMM

[WZ12] Daniel F. Waggoner and Tao Zha. Confronting model misspecification in macroeconomics. Journal of Econometrics, 171(2):167– 184, December 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407612001601.

Wang:2016:CVR

[WZ16] Chuan-Sheng Wang and Zhibiao Zhao. Conditional valueat-risk: Semiparametric estimation and inference. Journal of Econometrics, 195(1):86–103, November 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761630135X.

Wan:2010:LSM

[WZZ10] Alan T. K. Wan, Xinyu Zhang, and Guohua Zou. Least squares model averaging by Mallows criterion. Journal of Econometrics, 156(2):277–283, June 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002838.

Xu:2015:MLE

[XfL15a] Xingbai Xu and Lung fei Lee. Maximum likelihood estimation of a spatial autoregressive Tobit model. Journal of Econometrics, 188(1):264–280, September 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001657.

Xu:2015:SAM

[XfL15b] Xingbai Xu and Lung fei Lee. A spatial autoregressive model with a nonlinear transformation of the dependent variable. Journal of Econometrics, 186(1):1–18, May 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407614002954.

Xu:2018:SML

[XfL18] Xingbai Xu and Lung fei Lee. Sieve maximum likelihood estimation of the spatial autoregressive Tobit model. Journal of Econometrics, 203(1):96–112, March 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617302385.

Xiao:2012:RIN

[Xia12] Zhijie Xiao. Robust inference in nonstationary time series models. Journal of Econometrics, 169(2):211–223, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000371.

Xiao:2018:IEI

[Xia18] Ruli Xiao. Identification and estimation of incomplete information games with multiple equilibria. Journal of Econometrics, 203(2):328–343, April 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407617302452.

Xie:1988:SWC

[Xie88] Wen Zhi Xie. A simple way of computing the inverse moments of a non-central chi-square random variable. Journal of Econometrics, 37(3):389–393, March 1988. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0304407688900139. See corrigendum [Xie11].

Xie:2011:CSW

[Xie11] Wen Zhi Xie. Corrigendum to “A simple way of computing the inverse moments of a non-central chi-square random variable” [J. Econom. 37 (1988) 389–393]. Journal of Econometrics, 161(2):

338, April 1, 2011. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610002241. See [Xie88].

Xiu:2010:QML

[Xiu10] Dacheng Xiu. Quasi-maximum likelihood estimation of volatility with high frequency data. Journal of Econometrics, 159(1):235– 250, November 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407610001454.

Xiu:2014:HPB

[Xiu14] Dacheng Xiu. Hermite polynomial based expansion of European option prices. Journal of Econometrics, 179(2):158–177, April 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614000153.

Xu:2010:NPC

[XTZ10] Jianjun Xu, Xianming Tan, and Runchu Zhang. A note on Phillips (1991): “A constrained maximum likelihood approach to estimating switching regressions”. Journal of Econometrics, 154(1):35–41, January 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609001481.

See [Phi91].

Xu:2012:RMT

[Xu12] Ke-Li Xu. Robustifying multivariate trend tests to nonstationary volatility. Journal of Econometrics, 169(2):147–154, August 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407612000267.

Xu:2013:PTS

[Xu13] Ke-Li Xu. Powerful tests for structural changes in volatility. Journal of Econometrics, 173(1):126–142, March 2013. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407612002473.

Xu:2017:RDC

[Xu17] Ke-Li Xu. Regression discontinuity with categorical outcomes. Journal of Econometrics, 201(1):1–18, November 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617301471.

Xu:2018:SNE

[Xu18] Ke-Li Xu. A semi-nonparametric estimator of regression discontinuity design with discrete duration outcomes. Journal of Econometrics, 206(1):258–278, September 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301143.

Xiao:2019:WDM

[XX19] Zhijie Xiao and Lan Xu. What do mean impacts miss? Distributional effects of corporate diversification. Journal of Econometrics, 213(1):92–120, November 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300624.

Yang:2015:GMT

[Yan15a] Zhenlin Yang. A general method for third-order bias and variance corrections on a nonlinear estimator. Journal of Econometrics, 186(1):178–200, May 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614001705.

Yang:2015:LTS

[Yan15b] Zhenlin Yang. LM tests of spatial dependence based on bootstrap critical values. Journal of Econometrics, 185(1):33–59, March 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407614002413.

Yang:2018:UEF

[Yan18] Zhenlin Yang. Unified M-estimation of fixed-effects spatial dynamic models with short panels. Journal of Econometrics, 205(2):423–447, August 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618300599.

Yang:2019:NDE

[YCW19] Nian Yang, Nan Chen, and Xiangwei Wan. A new delta expansion for multivariate diffusions via the Itô–Taylor expansion. Journal of Econometrics, 209(2):256–288, April 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619300077.

Yu:2012:ESD

[YdJfL12] Jihai Yu, Robert de Jong, and Lung fei Lee. Estimation for spatial dynamic panel data with fixed effects: The case of spatial cointegration. Journal of Econometrics, 167(1):16–37, March 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002417.

Yang:2017:SIU

[YfL17a] Chao Yang and Lung fei Lee. Social interactions under incomplete information with heterogeneous expectations. Journal of Econometrics, 198(1):65–83, May 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300131.

Yang:2017:IQE

[YfL17b] Kai Yang and Lung fei Lee. Identification and QML estimation of multivariate and simultaneous equations spatial autoregressive models. Journal of Econometrics, 196(1):196–214, January 2017. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0304407616301683.

Yang:2017:SWL

[YL17] Yaxing Yang and Shiqing Ling. Self-weighted LAD-based inference for heavy-tailed threshold autoregressive models. Journal of Econometrics, 197(2):368–381, April 2017. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617300015.

Yu:2018:TRE

[YP18] Ping Yu and Peter C. B. Phillips. Threshold regression with endogeneity. Journal of Econometrics, 203(1):50–68, March 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407617302348.

Yamamoto:2015:TFL

[YT15] Yohei Yamamoto and Shinya Tanaka. Testing for factor loading structural change under common breaks. Journal of Econometrics, 189(1):187–206, November 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S030440761500189X.

Yu:2012:BEM

[Yu12a] Jun Yu. Bias in the estimation of the mean reversion parameter in continuous time models. Journal of Econometrics, 169(1): 114–122, July 2012. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030440761200005X.

Yu:2012:SSV

[Yu12b] Jun Yu. A semiparametric stochastic volatility model. Journal of Econometrics, 167(2):473–482, April 2012. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611002107.

Yu:2012:LEI

[Yu12c] Ping Yu. Likelihood estimation and inference in threshold regression. Journal of Econometrics, 167(1):274–294, March 2012. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407611002740.

Yu:2015:AET

[Yu15] Ping Yu. Adaptive estimation of the threshold point in threshold regression. Journal of Econometrics, 189(1):83–100, November 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615001888.

Yuan:2014:ELS

[YXZ14] Ao Yuan, Jinfeng Xu, and Gang Zheng. On empirical likelihood statistical functions. Journal of Econometrics, 178(Part 3):613– 623, January 2014. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0304407613002017.

Yan:2019:SER

[YY19] Jin Yan and Hong Il Yoo. Semiparametric estimation of the random utility model with rank-ordered choice data. Journal of Econometrics, 211(2):414–438, August 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300521.

Zellner:2010:DMC

[ZA10] Arnold Zellner and Tomohiro Ando. A direct Monte Carlo approach for Bayesian analysis of the seemingly unrelated regression model. Journal of Econometrics, 159(1):33–45, November 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610001119.

Zadrozny:2016:EYW

[Zad16] Peter A. Zadrozny. Extended Yule–Walker identification of VARMA models with single- or mixed-frequency data. Journal of Econometrics, 193(2):438–446, August 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300793.

Zamarro:2010:AHR

[Zam10] Gema Zamarro. Accounting for heterogeneous returns in sequential schooling decisions. Journal of Econometrics, 156(2):260– 276, June 2010. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407609002711.

Zu:2014:ESV

[ZB14] Yang Zu and H. Peter Boswijk. Estimating spot volatility with high-frequency financial data. Journal of Econometrics, 181(2):117–135, August 2014. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407614000608.

Zhang:2018:PTT

[ZC18] Rongmao Zhang and Ngai Hang Chan. Portmanteau-type tests for unit-root and cointegration. Journal of Econometrics, 207(2):307–324, December 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407618301519.

Zhu:2019:PNS

[ZCLW19] Xuening Zhu, Xiangyu Chang, Runze Li, and Hansheng Wang. Portal nodes screening for large scale social networks. Journal of Econometrics, 209(2):145–157, April 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618302689.

Zhu:2010:GAS

[ZG10] Dongming Zhu and John W. Galbraith. A generalized asymmetric Student-t distribution with application to financial econometrics. Journal of Econometrics, 157(2):297–305, August 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000266.

Zhao:2010:DEN

[Zha10] Zhibiao Zhao. Density estimation for nonlinear parametric models with conditional heteroscedasticity. Journal of Econometrics, 155(1):71–82, March 2010. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407609002127.

Zhang:2011:ECE

[Zha11a] Lan Zhang. Estimating covariation: Epps effect, microstructure noise. Journal of Econometrics, 160(1):33–47, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000540.

Zhao:2011:NMV

[Zha11b] Zhibiao Zhao. Nonparametric model validations for hidden Markov models with applications in financial econometrics. Journal of Econometrics, 162(2):225–239, June 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407611000042.

Zhang:2015:SMB

[Zha15] Ting Zhang. Semiparametric model building for regression models with time-varying parameters. Journal of Econometrics, 187(1):189–200, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615000469.

Zhang:2016:FSA

[Zha16a] Xianyang Zhang. Fixed-smoothing asymptotics in the generalized empirical likelihood estimation framework. Journal of Econometrics, 193(1):123–146, July 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0304407616300240.

Zhang:2016:WNT

[Zha16b] Xianyang Zhang. White noise testing and model diagnostic checking for functional time series. Journal of Econometrics, 194(1):76–95, September 2016. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407616300604.

Zhu:2018:SLM

[Zhu18] Ying Zhu. Sparse linear models and l_{1}-regularized 2SLS with high-dimensional endogenous regressors and instruments. Journal of Econometrics, 202(2):196–213, February 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617302105.

Zincenko:2018:NEF

[Zin18] Federico Zincenko. Nonparametric estimation of first-price auctions with risk-averse bidders. Journal of Econometrics, 205(2):303–335, August 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030440761830054X.

Zhu:2015:BST

[ZL15a] Ke Zhu and Wai Keung Li. A bootstrapped spectral test for adequacy in weak ARMA models. Journal of Econometrics, 187(1): 113–130, July 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000299.

Zhu:2015:MBP

[ZL15b] Ke Zhu and Shiqing Ling. Model-based pricing for financial derivatives. Journal of Econometrics, 187(2):447–457, August 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S030440761500055X.

Zhou:2019:EEC

[ZLCL19] Ling Zhou, Huazhen Lin, Kani Chen, and Hua Liang. Efficient estimation and computation of parameters and nonparametric functions in generalized semi/non-parametric regression models. Journal of Econometrics, 213(2):593–607, December 2019. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407619301642.

Zhang:2015:EGV

[ZLX15] Wenyang Zhang, Degui Li, and Yingcun Xia. Estimation in generalised varying-coefficient models with unspecified link functions. Journal of Econometrics, 187(1):238–255, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407615000470.

Zhang:2013:ACF

[ZLZ13] Xinyu Zhang, Zudi Lu, and Guohua Zou. Adaptively combined forecasting for discrete response time series. Journal of Econometrics, 176(1):80–91, September 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613001048.

Zhang:2011:EER

[ZMAS11] Lan Zhang, Per A. Mykland, and Yacine Aït-Sahalia. Edgeworth expansions for realized volatility and related estimators. Journal of Econometrics, 160(1):190–203, January 2011. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407610000722.

Zhang:2016:GFT

[ZOZS16] Shulin Zhang, Ostap Okhrin, Qian M. Zhou, and Peter X.-K. Song. Goodness-of-fit test for specification of semiparametric copula dependence models. Journal of Econometrics, 193(1): 215–233, July 2016. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407616300434.

Zu:2015:NST

[Zu15] Yang Zu. Nonparametric specification tests for stochastic volatility models based on volatility density. Journal of Econometrics, 187(1):323–344, July 2015. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304407615001190.

Zhu:2019:NQA

[ZWWH19] Xuening Zhu, Weining Wang, Hansheng Wang, and Wolfgang Karl Härdle. Network quantile autoregression. Journal of Econometrics, 212(1):345–358, September 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300892.

Zhou:2010:WAC

[ZWXW10] Yong Zhou, Alan T. K. Wan, Shangyu Xie, and Xiaojing Wang. Wavelet analysis of change-points in a nonparametric regression with heteroscedastic variance. Journal of Econometrics, 159(1):183–201, November 2010. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407610001405.

Zhang:2013:MAJ

[ZWZ13] Xinyu Zhang, Alan T. K. Wan, and Guohua Zou. Model averaging by jackknife criterion in models with dependent data. Journal of Econometrics, 174(2):82–94, June 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407613000183.

Zhang:2019:QRB

[ZWZ19] Yingying Zhang, Huixia Judy Wang, and Zhongyi Zhu. Quantile-regression-based clustering for panel data. Journal of Econometrics, 213(1):54–67, November 2019. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407619300600.

Zheng:2015:GAM

[ZXC15] Tingguo Zheng, Han Xiao, and Rong Chen. Generalized ARMA models with martingale difference errors. Journal of Econometrics, 189(2):492–506, December 2015. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407615001189.

Zhang:2015:CVS

[ZY15] Yongli Zhang and Yuhong Yang. Cross-validation for selecting a model selection procedure. Journal of Econometrics, 187(1): 95–112, July 2015. CODEN JECMB6. ISSN 0304-4076 (print),

1872-6895 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304407615000305.

Zhang:2018:SWM

[ZY18] Xinyu Zhang and Jihai Yu. Spatial weights matrix selection and model averaging for spatial autoregressive models. Journal of Econometrics, 203(1):1–18, March 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407617302245.

Zellner:2011:EER

[ZZ11] Arnold Zellner and David Zilberman. The economics and econometrics of risk: an introduction to the special issue. Journal of Econometrics, 162(1):1–5, May 2011. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407609002474.

Zhang:2016:CSM

[ZZ16] Zhengjun Zhang and Bin Zhu. Copula structured M4 processes with application to high-frequency financial data. Journal of Econometrics, 194(2):231–241, October 2016. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407616300938.

Zhao:2018:MMA

[ZZC18] Zifeng Zhao, Zhengjun Zhang, and Rong Chen. Modeling maxima with autoregressive conditional Fréchet model. Journal of Econometrics, 207(2):325–351, December 2018. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0304407618301520.

Zhu:2018:LDA

[ZZL18] Qianqian Zhu, Yao Zheng, and Guodong Li. Linear double autoregression. Journal of Econometrics, 207(1):162–174, November 2018. CODEN JECMB6. ISSN 0304-4076 (print), 1872-6895

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304407618301313.